A Bibliography of Books about the Internet and Networking: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 August 2024 Version 2.97

Title word cross-reference

$99 [Kro00].

.com [Wei01a]. .NET [BFS^{+}02, SS02].

’01 [Ano00a, Ano01b, USE01]. ’03 [Ano03a]. ’05 [Ano05a].

1.1 [BS02c]. 10 [Loc03]. 10-Gigabit [cFJH^{+}03]. 101 [Dav01b]. 10th [IEE01b]. 11th [ACK00, IEE01b]. 12 [Ame00, BMR01, Cra00, SNB^{+}00]. 14th [FHM00]. 1st [IEE01e].

2 [FM05, FK04]. 2.0 [MD02, Ten05]. 2000 [AD01, Cal00, Dre00, IEE01a, Kam00, Kro00, LL01, New01a, NB00, Nor00a, Tre01, Wil02b]. 2001 [BU01, IEE01b, Pal01, Sta01b, Sta01c]. 2002 [B^{+}02, IEE02b, IEE02a]. 2003 [ACM03, FLA^{+}03, Mos05]. 2004 [ACM04]. 21st [DL01]. 24 [BS01a]. 24/7 [Foc01]. 26 [BS00]. 29th [FLA^{+}03]. 2nd [BBP00]. 2U [Ano00h].

3.0 [ER02, Wil04]. 39th [IEE01d]. 3G [Bra03, SC07, Vac01]. 3rd [Ano00a, Ano01b, Cal00, Dwi02, USE01].

4 [Cas00b, PFAHA03, SC05]. 400-2U [Ano00h]. 4G [Gli06]. 4th [Ano03a, Cal01, Fra01c, IEE01c, KK00, ASP01a, ASP01b].

5 [GBR05]. 5.0 [Tre01]. 505 [MD02]. 5th [AH02, BMM07, Cal02, HL00, KZ01, RS00a, WL04].

6 [CdBvdM01, Smi00b]. 640 [MD02]. 640-505 [MD02].

7 [Wan06a].

8 [Ber00]. 8-10 [AD01]. 802.11 [OP04]. 802.11b [Ros03]. 802.15.4 [GCB04]. 802.16 [Coo04].

’99 [HL00].

= [CSS00, Kur00a].

A-Interface [LS00]. AAA [LKP02]. aandachtspunten [PG00]. Abbas [Puz03b]. ABR [KJF^{+}00]. Abschluss [Wey01]. abuse [Ana02, Arn01]. abusive [LSC^{+}04]. Academic [LM00]. Académie [dB01]. Academy [dB01]. accelerate [BG00b, Pau03]. Accelerates [Ano03c]. Accept [Coc03a]. Acceptable [SNT02]. Access [Bar00c, Che02, IAI02, MD02, O’L00, Uni01, All09, Ano02, EP06, Gre00c, IEE01a, LFD01, LSC^{+}04, Nal02, Nua07, RB02, VN03, Wis00]. Accesses [DL02]. Accessible [SKS^{+}02, MWD00]. Accessing [GDC^{+}02, Gil00]. accountability [JSK02, Ros01]. Accountant [Coh00a]. accounting [Cho01b]. accredited [BB00a]. accuracy [ZFdRD05]. Accurate [KOS^{+}02]. ACLU [Bar01b]. ACM [ACM04, IEE02c, Bar00a]. acquisitions [Tim00]. Acronyms [Bar01a]. Across [KSSK02, BS01b, But02, Koh02]. Act [Kam00]. active [Ras05a, Gre02a]. activism [HM02a]. activities [Bra01, BS01b, Flo01]. actor [Koz00]. Ad [ABC^{+}07, Toh02, AM07, Cha05, Cra07, Lab07]. ad-hoc [AM07, Cra07]. Adaptation [Abi02]. Adapting [PKP03]. Adaptive [Abi02, KJL02, LLZ03, MWD00, SHK02, KD05].

Address [Lea00, SPP^{+}07]. Address-Space [SPP^{+}07]. Addressing [Gra01b]. Adi [Coc03b]. Adleman [Coc03b]. Administering [ST00]. Administration [Dan02, Kre03, Zad01]. administrator [Ano00i]. Administrators [ES03, Hal00b, Mau03, Sta03b, VE05]. admission [BFOBR02, LIB04]. adult [Ano02]. Advance [CZB^{+}01, Lea00]. Advanced [Cha02, Gli06, SZGbC04]. advances [CPM01]. advantage [FN03, Har00d, Moh02]. adventure [Bar00d]. advertising [KM01a]. advice [Pew00]. adviser [Bor00]. Advocacy [HM02a]. Africa [CL00]. Afrique [CL00]. After [Mar00]. afterwards [Man00a, Man01]. against [SX03]. Age [Sin01a, AD01, Bie01, Bla02a, Foc01, Gol00a, Gra01a, GM01, Hir00, JF00, LS01b, Lev01b, Lip00, LW00, Log00, McB02, McC00b, MB02, Mol02, MM00b, Moo00, Nat01, RM02, Shu01, Cap00b]. agencies [Hil01]. agenda [Fre01]. Agent [CLZ00]. Agents [Huh03, RHN02, SP02, KK00, KZ01, Omi01, RB01]. aggregation [CTYS03]. Aggressive [Wyl05]. Agile [THB05]. agreements [BG00a]. Ahead [Wes02]. AI [Bar00c]. aided [LSS01]. AILA [BeM00]. air [Cur00]. Airport [Ano05b]. AISA [Cha02]. Aix [TL02]. Aix-en-Provence [TL02]. AJAX [JWC08, JF06, Mor08]. Alan [Puz03a]. Algebra [Sob02]. Algorithm [Bar01b, GT02, Kim02, KJF^{+}00, LCH^{+}08, WZZ04]. algorithmics [Goo02a, Var05]. Algorithms [SM02b, Bou06, MR07, Sob02]. alien [Mar00]. Aligning [Sim02a]. all-in-one [ISI00a, ISI00b, ISI01]. All-to-one [LJ01]. aller [DF00]. alliances [EL01]. allied [Edw02]. Allyn [OG00]. Alternatives [Swa01]. Altertumswissenschaften [AR01]. altertumswissenschaftlichen [AR01]. always [Fra01b]. amazing [Bar00d]. America [Gar00c, SP00, Tob01]. American [Bro02, FHM00, ABM00, Har00d].

Americans [Llo01d]. among [Kay00]. amorphous [AAC^{+}00]. Analysis [BCG06, CNI^{+}04, JY02, Ket02, McC03, EL01, GT02, Mai06, MS00b, Nor01c, SGD^{+}02, STPM05, ZLL01]. analyst [NNM01]. analysts [Ano01e]. analytical [KMK04]. analyze [Cur00]. Analyzer [Ano03c]. Analyzing [JRSW04]. Anarchy [Bar00a]. ancestors [Fli00]. Anders [Bar01a]. angleren [GS00]. angleren-ruseren-hayeren [GS00]. Anguilla [Fra01c]. Anguraintanetto [Kur00a]. Animal [Kis00]. animals [BG01a]. animation [Dud08]. Announcement [Ano00a]. annoyances [Tyn05]. annoying [Tyn05]. Annual [Dav00, WL04, BBL00a, Cal00, Cal01, Cal02, FHM00]. Anonymity [Bar01a, Dav02a]. answers [VC00]. ANTARCTICA [GIM^{+}02]. anti [CN00]. anti-globalisation [CN00]. antichita [CSS00]. Anticipating [DR03]. Antike [CSS00]. Antipasto [Swa00a]. antiquity [AR01, AR01]. Apache [Goo02b, Ris05, TSC02]. APh [Soc01]. APIs [KKT04]. Apple [Owa05]. Appliance [Lev01a]. Appliances [LKK02, Ris01]. Application [LKP02, Man02, McF03, Sin01b, SM00b, BvS00, EST00, GPR^{+}05, KKT04, Mat02, Neu00, SW01]. Applications [Abu01, Ano03c, BFS^{+}02, CLZ00, CS01, Cri03, Dwi02, FB05, Gra05, Gre02a, Hel00, HW00, IEE01e, IEE01f, IEE02b, IEE02a, Ins03, Kro00, Leo02, Mes00, PJN^{+}02, Ros02, WL04, WZ01a, Abe03, Bea02, Bur02, Cha02, Gre00b, HCP05, HL00, JWC08, LMV02, Mor08, Ngu01, Gal02a, Omi01, RG02a, Ric01a, SSS06, Sha01, S^{+}06, Sin01b, SMLN^{+}03, SM02b, Tra00, Var03, WA06]. Applied [SS02]. applying [BL00a]. Approach [DYN03, Far03, KR03, KR05, PK02, PD03, Zha03, CSMW02, Har05, HL02b, KMK04, KR01b, KR08, MS00c, PG01, Pan04, Pau03, PD07, TS00, Var05, Ver01, ZG04]. Approaches [MB00a, Hak00, Ora02]. April [Hoh01, IEE01a, IEE01c]. Arab [AS02]. Arachne [CSS00]. Arachnes [CSS00]. Architecture [Bla00a, Day08, EPR03, FB05, Gil07, Jep03, McC03, NDK^{+}03, SP02, SKS^{+}02, Bea02, Bow00, Bur02, CS02, DDPP00, FB04, HB05, SKW^{+}07, WCB01]. Architectures [BW02, CZ04, Com00b, Puz03b, Sch06, BDFG02, CML00a, Dot01, Hal00a, MR07, Sun01, Wan01b]. archive [Ano00d]. Area [Fib01, SI01, Slo00, BvS00, CRW01, GCB04]. areas [Hak00]. ARF [Adv00b, Adv00a]. arguments [BC01]. Ariel [Ive00]. Armies [Bar00b]. Art [MSM02, BF01, Fra00, Zdz05]. arts [Spo01]. Artsguide [Spo01]. AS-level [CGJ^{+}04, SFFF03]. Asian [Ros01]. ASIP [SKW^{+}07]. aspecten [PG00]. aspects [BK00a, WHdM00]. Aspekte [WHdM00]. Assembly [Sno01]. Assessing [MTK03, Wes00a, Wes00b]. Assessment [Dur02, Lan06]. assessments [Lan06]. Asset [Kro00]. Associated [Gol00c]. associations [Hil01]. AT&T [Bes00]. ATM [Ahm02, IEE01c, KJF^{+}00]. attack [Mir05]. Attacks [Gar00a, SX03, Dwi09, Sch00a, SL06, Zal05]. Attention [Nis01]. Attitudes [ET03]. attract [Sin01a]. auction [DV02]. auctions [Ell00, RO00]. audience [Pew00]. audiences [CR01b]. Audio [Bai01, ABGR01, FF00, HSW07]. Audit [Mai06]. August [B^{+}02, Cha02, IEE01d, JC01, Koh02, MS00b, vdMdB01, Mou01, Smi01c, New01a]. Austin [ACK00]. Australia [CN00]. Austria [Juk02]. Author [dB01]. Automated [HL02a, Rie02]. automatic [Pal02]. Automation [Str03]. Automotive [CHB02]. Automounter [Zad01]. autonomous [BNHGV04, Gao01]. av [Her01]. Availability [NKB^{+}03, Yu00a, Yu00b]. aviation [Cur00].

Aviel [ET03]. Avoid [VM02, Joh03, Tyn05]. Avoiding [LBJ^{+}02]. Award [Coc03b]. aware [YFB02]. awareness [Gur01]. Away [Coc03b, KM01a]. AXIS [BI02].

B [Puz03a]. B2B [SW01]. Bachelor [BB00a]. Back [CZB^{+}01, Owa05]. backbone [BSS02, RZMD05]. backbones [MTK03]. Backplane [BBM^{+}03]. Bacon [OG00]. Bad [McC00a]. Bahia [MFdSeS01]. balancing [Moo02]. Ballistic [Pau03]. Ballots [Coc03a]. Baltimore [IEE02c]. ban [Cen00]. Bandwidth [LKKS05, Uhl02, BK00b]. bang [LR02]. Banners [SZ01]. bararan [GS00]. Barbara [AH02, IEE01d]. Bark [Bar00d]. barrier [LFD01]. BASCOM [Ano00g, Ano00h]. Base [Abi02]. Based [Bar00c, CZ04, CWZ^{+}05, DWX02, Kur02, LPK02, NKB^{+}03, PFAHA03, PcC02, SP02, San01, Str04, SX03, TKF^{+}02, Wan08, ZY02, Ano00g, Ber00, BKY^{+}03, CKK01, CTYS03, Dot01, DBS00, GBCH00, Joc01, JKH^{+}00, Kro00, LA02, Law04, LSC^{+}04, MV03, Ngu01, Pan04, Rya00, Sun01, SDB00, STYC02, WAF00, WL04, ZFdRD05, GY01]. Bases [B^{+}02, FLA^{+}03, AR01]. basic [RWS00, Kap00, Kur00b]. basica [Ano01c]. Basics [Bey00, Fli00]. bats’atrakan [GS00]. battle [Mod00]. Bay [BL00b]. Bayesian [Zdz05]. BCRAN [MD02]. be [Dav02a, Gar00b, Man00a, Man01, Uni01]. become [HNV04]. bed [ABC^{+}07]. BEEP [Ros02]. Before [Lig03, Six00]. begin [Pau03]. beginner [Cli01, Tre01]. Beginning [BCK^{+}00, Ten05]. behavior [BMR00, Che01, CML00b, RZMD05]. Behaviour [RG02b]. Behind [RD05, Zal03, BL00b, BB00b, YA04]. Beit [Soc00]. Beitrage [CSS00]. Bell [Bes00]. Bellovin [ET03]. Ben [Smi00d]. Benchmark [Bar01b]. Benchmarking [LRJ^{+}03]. ber [DF00]. Bergen [RS00a]. Berlin [CRS00, FLA^{+}03]. Best [Kai06a, Ano00f, BF01, McB02]. Beta [Ano00h]. betrayal [MSCG01]. better [Smi00c]. Between [TT03, Dre00, Ros00]. Beyond [Bie01, Cer01, Jen01, LR01a, Rob00, Sti03, Bej05, CG01, Dot01, DSM00, KV00, Loc03, Man00b, Nai02, Sun01]. BGP [BNHGV04]. Bhubaneswar [GM04]. Biblical [Wor02]. Bibliografieren [RS00b]. bibliographic [ISI00a, ISI00b, ISI01]. Bibliographie [Fra00]. bibliographies [ISI00a, ISI00b, ISI01]. bibliography [Kis00, Lin00, Nor00b, Nor01a, RS00b]. bidding [RO00]. Big [Los00c, Los00b, Los00a, Ano00e, LR02, RR03a]. bigwigs [Ano01e]. Bildbearbeitung [DF00]. Bilder [DF00]. Bill [And00a]. BIND [AL01]. biocomputing [Bro00b]. Bioinformatics [Bro00b]. biologist [Bro00b]. Bioterrorism [Bar02]. birth [Ste00a]. bis [Hoh01]. Blind [GBM02]. bloopers [Joh03]. blue [And00b]. Blueprint [FK03, FK04, Mur00a]. Bluetooth [BS02c, KKT04]. body [Gab01b]. Book [Dwi02, ET03, KE01, Law04, Puz03a, Puz03b, Sti03, Zal03, Com00a, Jak00, Los00c, Los00b, Los00a, Ros03, Ros08, Wan01a, Wan06a, Wan06b]. Bookshelf [LS01a, Sta00b, Tay00, Woe02]. boom [Man00a, Man00b, Man01]. booms [Fra01b]. boost [Won00]. boosts [Pau03]. Boston [Ber00, CdBvdM01, KK00, Smi00b]. Botnet [LWD07]. boundaries [Eve00, Lip00]. Box [Wil03]. boys [Kat00, Kat01]. Brain [Bar00b, Goe02]. Brand [LA00c, BL00a, MM00b, Tur06]. branding [BL00a, RR00]. Brave [LR02, BC01]. Brazil [MFdSeS01]. Break [HM04, Pri00, WA06]. breaking [LFD01, YA04]. breakneck [Car00a]. breakthrough [Dot01, Sun01]. breast [Abe01]. Bridges [Per00]. Bridging [ACM04]. Brief [LCC^{+}09, Nau00, Sha01]. briefing [Gol00c]. Briefs [Lea00, Pau03]. bringing [FHLW03]. British [Fra01c, GW00]. Broadband [Bey00, EP06, IEE01a, IEE01b, Tru00, Ale00, AGM07, Nua07, Ros08, Vac01, WZ01b]. Broadcast [Mea00, Pew00]. Broadcasting [DSM00, LKMK02, TKF^{+}02, Gar00d, SL01, TPB01]. broke [Bes00]. brokerage [Bak00b]. Browsing [McF05]. Brutus [CJM00]. BSD [DF00]. bubble [PP01a]. budget [Ano00e, Dif02]. Build [Kro00, AKY01, Ano00e, Gre00b]. Building [Dea00, Gre02a, Gre00c, HW00, Kai06a, KE01, McG06, MD02, VM02, ZCC00, Ano00d, ABM00, Bea02, BG00b, Car00a, Foc01, JF06, JWC08, Lim06, LMV02, LA00c, Moh02, MM00b, Mor08, Mur00a, Ste00b, Tob01]. Bulgaria [AD01]. Bull [Pau03]. Bulletproof [Cha05]. burning [Wan06a]. burst [PP01a]. Business [OFA00, Pan02, Sil00, AT01, Ano00e, BSW00, Bod00, Bon00, Bra00, BG00b, BR00, BL00b, Car00a, Cas01, CPM01, Dei01a, Dif02, EHT02, Eng00, Flo01, Foc01, Gla00, GM00a, HM02b, Kam01, KM00, KI00, KW02, LCL04, Lou01, McC00b, MVK01, MM00a, MR00, Mer01, Mor00, PB00, Pow02, Pri00, RBB01, Sau01b, SW01, Sim02a, Til05, VS01, Was00, WHdM00, Wei01a, WB02, McB02]. Business-to-business [Sil00, SW01]. bust [Man00a, Man00b, Man01]. busts [Fra01b]. Buy [PLW07]. bytecode [Qia00].

C [DF00, DC01, Don03, Mic00]. C# [Bar01a]. CA [AH02, Koh02, MS00b]. Cable [Lig03, Llo00, CFLA04, DSM00, LFD01]. Cables [BS02c]. Cabling [Cri03, GM00b]. Cache [cCP00, GcC02, Wie02, PB03, WWD04]. Caching [CWZ^{+}05]. Caesar [Chu02]. calculus [LT01b]. Caldera [Ano00g]. Calif [IEE01e]. California [BS01a, BS02a, IEE01d, NB00, USE01, Ada00, Ada01, BS00, IEE01f, Lev01a]. call [Bla01, Hus04, LF00]. Caltech [Bar00a, Coc03a]. Campaigning [MSM02]. campus [Haw01]. Can [CZB^{+}01, PP01a]. Canaan [EDGW00]. Canada [Gei00]. Canadian [JT00]. cancer [Abe01]. Capable [PcC02]. Capacity [CW00, MA02, Pau03, PM04]. capital [BK00a, Joc01]. Capron [Cap00a]. capturing [CGJ^{+}04, Gre01a]. car [MZSR00, SZSR00]. Cards [Ano00h]. care [Gol00b, MCH01, Woo00]. career [BK05, Har01]. CareZone [Hun02]. Caribbean [Dur02]. caring [PG01]. caring-centered [PG01]. Carnegie [FHM00]. Carrier [CAN^{+}04, Wil01]. Case [Abi02, cFJH^{+}03, Gla00, MG01, SKW^{+}07]. cases [AT01, EHT02, MSS01, RB01]. Castle [KS05]. Catalyst [Ano03c]. category [Wil06]. Caterer [GM00a]. CCNP [MD02]. CCSA [Dan02]. CD [DF00, AOB00, Gil00, Har00b, Ste03, Wan06a]. CD-quality [Gil00]. CD-Rom [AOB00]. CD-ROMs [Har00b]. CDMA [Bra03, LS00, ML02]. CEH [Gra07, KV08]. Cells [RS02b]. Cellular [Byr03]. centered [PG01]. Centers [Meh02, Hil01]. Central [Ano00g]. Centric [RFP02, Foc01]. centuries [BG01b]. century [DL01]. CEO [Ano01e]. Certified [Opp07, Gra07, KV08]. CGI [HL02b]. Chabad [Soc00]. chain [Cho01b, KR01a, PB00]. challenged [Kro00]. Challenges [Bar01b, MB00a, Tir03, Chi01, Das01]. Chancen [LSS01]. change [Amo01, Fri01, Mod00, Sur01, Moo02]. changed [McC00b]. changing [BSW00, LT01a, Mur00b, PHCMS05, SP00, WN01a]. channel [LLL03]. chaos [Spa01]. Characteristics [IAI02]. Characterizing [BvS00]. Charge [LA00a, LA00b]. Charge-sensitive [LA00a, LA00b]. charging [Sti02]. Cheap [Bar00b]. cheating [LF00]. Check [Dan02]. Checklists [Kai06a, Kai06b]. Cher [Lej00]. Chess [BCG06]. Cheswick [ET03]. Chicago [ASP01a, ASP01b]. Child [Ano01a, Arn01, Jen01]. children [Ian01, WM00]. China [B^{+}02, HL00, Jef00, WN01b]. choices [Mur00b]. Chord [SMLN^{+}03]. CIA [KK00, KZ01]. ciphers [Chu02]. circuits [Ini07]. Cisco [BB00b, Hel02, MD02]. citation [RS00b]. Citrix [Wil04]. City [IEE02b, IEE02a, Adv00b, Adv00a]. civic [Nor01b]. civil [Red00, War01]. Claims [Bar00c, WS01]. Class [Sch02a, Nor01a]. classica [CSS00]. classification [LCH^{+}08, Zdz05]. classroom [And01, Dud00, FKC00, Hir00]. clear [Sut00b]. clearinghouses [Hil01]. Clearly [Los03b, Pfa01]. click [KM01a]. Clicks [PP00, PP01b]. Client [CS01]. Client-Server [CS01]. Clock [WZZ04]. Closed [RD05]. Club [BCG06]. Cluetrain [Swa00a]. Cluster [Ano00g, ER02, KJL02, NKB^{+}03, PcC02, CTYS03, GBCH00, JKH^{+}00, STYC02]. Cluster-Based [NKB^{+}03, PcC02, CTYS03, GBCH00, JKH^{+}00, STYC02]. clustering [WZZ04]. Clusters [cFJH^{+}03, SZGbC04]. coastal [Hak00]. Code [CAN^{+}04, HM04, HM00, CP01, HL03, Hus04, Lav06, OC07, PC01]. Codes [Chu02]. codice [Tos00]. Coding [Bur01, FPM05]. Coimbra [BMM07, Smi01a]. Coli [Coc03a]. Collaboration [RHN02, SEGS03]. collapse [LBJ^{+}02]. collapses [Met00]. collections [Fra00]. College [BB00a]. Color [Bar00c, LH00]. com [AA01, Car01b, Dot01, Sun01]. Come [Bar00a]. Comer [Dwi02]. comercio [MRM00]. coming [Man00a, Man01, Nat01]. comments [MSS01]. Commerce [Cam01, KE01, Sta00c, Bak00b, Bal01, BMP00, BMP01, Bor00, BK00d, BG00b, Cal00, Cal01, Cal02, Cam00, MRM00, EW00, Fit00, Gla00, Gol00b, JSK02, Kam00, KM01b, KR01a, Lev01b, Lou01, Luc02, MM00a, Mor00, RB01, Sin01a, Soo01, Sta00e, Var03, Wey01, WN01b]. commercial [ILS00]. commercializing [Gre00c]. commercio [Dra01]. Common [CAN^{+}04, Joh03]. Communication [CNI^{+}04, Coh00b, KMK04, Leo02, Str03, WV00, WZ01a, CL00, Coo04, GG05, HM00, Gon00, MS00a, OC07, PM04, RK01b, RM02, Sam02, TV05, HM00]. Communications [Bey00, Bur01, CG01, Cri03, Gib01, Hub02, IEE01b, Mea00, Pal01, Puz03a, Rap02, Sta00a, BMM07, BCKT05, CFLA04, HSW07, Gon00, LT02, MTK03, Odl00, Sch05, SJ01, Ste00a, VC00]. CommuniGate [Ano00g]. Communities [BG00b, ACM04, FN03]. community [BL01, Gat01, HM02a, O’L00]. companies [Bur00, Fra01b, Mod00]. companies- [Bur00]. companion [Ben01, Kis00, OP04, Owe02a, Pet01]. Companionship [CLWR00]. Company [Sta02b, Ano00e, BL00b, Car01b, KR01a, MB02, VS01]. Comparative [CJL02, Far03]. Comparing [ABGR01]. compass [Spa01]. compelling [Dud08, Wei01a]. competition [Gar00d, Llo00]. competitive [BSS02, BG00c, FN03, GB00, Ste00b]. complete [Bay00, Ben01, Bow00, Cli01, GM00b, Kad00, Koz01, Lan06, Smi00c, UG00]. compliance [Lay06]. components [JF06]. Compositional [Abi02]. Comprehension [CAN^{+}04]. comprehensive [Dav01a, How01, Koz05, KV08]. Compression [Say03, HSW07]. computation [BW00b, Sob02]. Computational [Kim02, LD03]. Computer [AS02, Ano00g, Bar01a, BH08, Coc03b, Dwi02, Haw01, HL00, Joh06, Kay00, KR01b, KR03, KR05, KR08, Mil06, PD03, PD07, STHK02a, Soo01, Sta00a, Sta03a, Tan03, Tyn05, Vac09, WS01, Ano00c, Ano01d, Cas00a, Cli01, Col00, Com00a, DCC00, JBR05, Lej00, LSS01, MSS01, PM04, SL06, Wan01a, Wan06b]. computer-aided [LSS01].

Computer-Science [Coc03b]. computergestutzter [LSS01]. Computers [Ano00h, Cap00b, Coc03b, Nic01, Cas00a, Kam00, New01a]. Computing [ACM04, Bir03, CP03, CSK00, Cla03a, DF03, Fra03a, GIM^{+}02, HH03, Huh03, JZ02, Kro00, LV03, Meh02, Met03, MRK^{+}03, PKP03, PN03, PHM03, Rao02, RFP02, San02, San03, TT03, Vin02, Vin03c, Vin03b, Vin03a, ZN05, AAC^{+}00, Amo01, BU01, Bou06, Bur02, Chi01, FK03, FK04, Fur00, GM04, Goe02, OLLY02, RR02, Sto02, SZGbC04, Ten00, UBM02, Wal01b, ZLL01]. comunicacion [Gon00]. comunicaciones [Gon00]. comunitaria [Tos00]. concept [Ler00b]. conceptos [Gar00b]. Concepts [ET03, WL04, CSMW02, Gar00b, Lav06]. concerns [Nor00b]. conclusione [Gio00]. conditioned [WCB01]. Condor [FMM00]. Conference [ACM04, AD01, ASP01a, ASP01b, AH02, Ano01b, Ano03a, Ano05a, B^{+}02, DWX02, FLA^{+}03, HL00, IEE01c, IEE01d, IEE02c, Juk02, Kim01, WL04, Abe03, BMP00, BMP01, BMM07, BCKT05, FHM00, Fra01c, GM04, TL02]. Configuration [CSK00, RWS00]. configure [Ros03]. Configuring [Mos05]. conflict [BK00a, LLL03]. confronting [Bie01]. confusion [Bur00]. Congestion [KCC02, PT00, JT01, LBJ^{+}02]. Congestion-dependent [PT00]. Congress [MFdSeS01]. Connect [BS02c, Ano00h]. Connected [DHL^{+}04]. Connection [PJ00a, BB00b, HPS01]. connections [Bar01c, Hil00, LFD01]. Connectivity [CHB02, Wil03]. Connexion [PJ00a]. consciousness [Goe02]. consequences [AN00]. Consider [CAN^{+}04]. Considerations [Rub01]. consistency [Yu00a, Yu00b]. consolidation [Mai06]. constant [Moo02]. constitutionality [Ano02]. construction [Mor08]. Constructive [Ora02]. Consumer [Ano00g, Dav00, Hil01, KC00]. consumers [Mod00, Mur00b, Woo00]. contacts [Hil01]. containment [CCC^{+}05]. contemporain [Fra00]. contemporary [SV01]. Content [HB05, KJL02, PKP03, San01, Sim02b, Ver01, Wie02, FB04, RB02, SGD^{+}02, Zdz05]. Content-Adaptive [KJL02]. Content-Based [San01]. Contest [Bar00a]. context [HJ04]. continuous [Moo02]. contract [Imp00]. contratto [Gio00]. contributi [CSS00]. Control [LKK02, LPK02, Mas06, SHK02, deC00, BFOBR02, Bie01, CdBvdM01, JT01, LA00a, LA00b, LA02, LIB04, LSC^{+}04, vdMdB01, YCY03]. Controlling [Kro00]. controversy [Smi01d]. convenience [BMR00]. converge [Ale00]. Convergence [EPR03, JY02, Lla00, LABJ01]. Conversation [Hut01]. cookbook [OBB05]. CoolKeyboards [Ano00h]. Cooperation [Leh05]. Cooperative [KK00, KZ01, CR01a]. Coordination [CLZ00, Omi01, ZY02, GR01]. Coping [LLL03]. Copper [Cri03]. copyright [Lit01, Sha00, ZB01]. CORC [CR01a]. Core [Dud08, Hal00b, Muc02]. corporate [CKK01, Col00, LW00]. Corporation [Ano00g, Ano00h]. Correction [CTBA^{+}01]. correlation [GG05]. COST [CRS00, vdB00, Lim06, Smi01a, Sti02]. Cost-Effective [vdB00]. Costs [BC02, Bak00b]. could [Pau03]. Counter [SL06]. counterculture [Tur06]. countering [LWD07]. countermeasures [Dwi09]. Countries [Leh05]. Course [CAN^{+}04, Mes00, GM03, KCF02, Lou01]. Courses [GDC^{+}02]. Cover [LRJ^{+}03, MRK^{+}03, STHK02b, TWW03]. coverage [RGE00]. CPXe [TWW03]. Crackers [SEK01b]. crash [KCF02, Lou01, Man00a, Man01]. crawlies [Llo01a]. create [ISI00a, ISI00b, ISI01, Wei01a]. Creating [Dud08, FN03, Goe02, Jan01, And01, Com01, HP02, KM00]. creation [Owa05]. Creative [MVK01]. creepy [Llo01a]. Crime [Joh06, Wal01a, Cas00a, Kip07, KB00]. criminalité [PJ00b]. Criminality [PJ00b]. crisis [Com01]. Critical [Jac00, Gla01, GM03, PN01, SE00, SV01]. critique [Fra00]. CRM [Gre01a]. Cryptographers [Coc03b]. Cryptography [Gal02a, Fra01c, GG05]. CS [CAN^{+}04]. CSV [Ano00h]. cueing [BS02b]. cult [Kah05]. cultural [Gat01]. culture [WM00]. cultures [Llo01b]. Cure [CLWR00]. curiosity [Smi00d]. Curl [Müf01]. current [Van01a]. curricula [Cra00]. curriculum [BS01b, Lew01]. Customer [Ste00b, AKY01, BG00b, Bro00a, Foc01, KM00, Koh02, LCL04, MB02, New00a]. customer-centric [Foc01]. customer-oriented [LCL04]. Customers [Pic03, Gre01a, PB00, Sin01a]. Cuts [Pau03]. cutting [Col00]. Cyber [FNRC05, Car09, DF01, Hir00]. cyber-savvy [Hir00]. cyber-space [DF01]. Cybercrime [Col00]. cyberculture [Tur06]. Cyberdroit [FS00]. CyberEducator [BMR01]. Cyberlaw [FS00]. cyberlaws [Kam00]. Cyberliteracy [Gur01]. Cybermarketing [BBP00]. Cybernet [Ano00h]. Cybersecurity [PLW07]. CyberShock [Sch00b]. cyberspace [Ano01d, Bie01, Fra02, KK00, Sam02, Wer00]. cycles [Spa01]. Cyndi [How01].

D [ET03]. d’abondance [LD00]. Dagstuhl [KS05]. danger [FK01, Ger00]. Dangerous [Lea00]. Dante [Wer00]. Dark [Vin02]. Data [Bar01b, B^{+}02, Bey00, But02, CIA^{+}03, Cla03b, Coc03a, Cri03, EPR03, FTH00, FLA^{+}03, GIM^{+}02, Ken02a, Ken02b, Leo02, Mac03, ML02, Meh02, Sta00a, TPB01, ZLL01, Ahm05, BK00b, CTYS03, CFLA04, Col02, Cur00, GBCH00, Koh02, LT02, Pol00, AR01]. Database [HW00, Leo02, Gal02a, Ric01a]. Databases [San01, Abe03, Hil01, ISI00a, ISI00b, ISI01, RG02a]. Datacenter [Jol03]. date [Smi00c]. Dateien [DF00]. Datenbanken [AR01]. Datensatz [Pol00]. d’auteur [dB01]. d’autore [ZB01]. days [Kuo01, Nau00]. DDoS [SX03]. deadly [HLV05]. Dear [Lej00]. Death [CN00]. debate [Got04]. Debian [DF00]. Debugger [Ano00g]. decade [FKC00]. Decaf [Bar01b]. December [ACK00, Ano05b, GM04, HL00]. Deep [BL00a]. defamation [Col01]. Defending [SX03, YA04]. defense [Mir05, WN06, Wyl05]. defenses [SL06, WS01]. definitions [Soo01]. Definitive [Hal00b, Ros02]. degrees [BB00a]. Degunking [BD05, PP05]. del [Car01a, CSS00, Gio00, Tos00]. Delay [PJN^{+}02, Rao02]. Delayed [LABJ01]. delays [JT01]. Deliberative [Wes02]. Delivered [Kur02]. Delivering [GBM02, LFD01, Puz03a, Gre00c, SJ01]. Delivery [CWZ^{+}05, PFAHA03, Wie02, Ive00, KD05, SGD^{+}02, Wes00a, Wes00b]. dell’informatica [Tos00]. Democracies [CZB^{+}01]. Democracy [CTBA^{+}01, SM02a, Wes02, Ano01g, Bro02, Cai01, Cli00, Fer00, GW00, Wes00a, Wes00b]. democratization [Fer00]. Demography [Coc03b]. Denial [Gar00a, Mir05]. Denial-of-Service [Gar00a]. denied [Uni01]. Density [Ano00h]. Denver [JC01, vdMdB01, Mou01, Smi01c]. dependence [GLMT05, PHCMS05]. dependent [PT00]. Deploying [HB03, LS00, Wil04, Ras05a]. Deployment [Meh02, ES01, Ler00b]. depression [Man00a, Man00b, Man01]. Derecho [Ano01c, Gar00b]. desarrollo [MRM00]. description [CR01a]. Design [Ano03c, Ano05b, Bra03, BKY^{+}03, BF01, CRW01, cCP00, Cla03b, Cla03a, CFHO02a, CFHO02b, DWX02, EPR03, Fra03b, Hel00, Ken02b, LKP02, Mac03, McC03, Sut00a, TKF^{+}02, Wie02, And01, BC01, Bod00, Coh00b, Dil00, GG05, GT02, Joh03, Juk02, Lan05b, Lay06, Ler00b, OC07, PM04, YFB02, YCY03]. designer [OP04]. Designing [Abi02, Bar00a, DE02, ER02, Gra01b, HP02, Sch02a, Wel02, Wil02b, Zel07, Fly01, JG08, Var05, Wil01]. desk [Sco01]. Desktop [DHL^{+}04]. destruction [MVK01]. detecting [NLiMK04]. Detection [Pro00, Wan08, Bej05, Bej06, LWD07, NNM01, NN02]. deterministic [LT01b]. deterritorialisation [CN00]. Deutsch [Hoh01]. Deutsch-Schwedisches [Hoh01]. developer [Joh00]. Developers [CLWR00, Sta03b]. Developing [Gra05, Lim06, LS01a, MV03, Ros02, GMP00, HL02b, Gal02a, Par05]. Development [Bra03, KOS^{+}02, Kro00, LKK02, McF03, SM00b, BvS00, MRM00, EST00, Fra02, Har01, Lai01, LCL04, MM07, Sta00e, THB05, Tim00, VW07]. Developments [AS02, Cal00, Cal01, Cal02]. Device [Ano03c, SKS^{+}02, Kha02]. Devices [DE02, Kro00, STHK02b, STHK02a, Ini07, Var05]. d’expression [Fra00]. DHTML [Lan05b]. diagnostic [STPM05]. diary [Lej00]. Diccionario [dAA00]. dictionaries [LSS01]. Dictionary [Ano01d, DCC00, Pho01, Gil05, Kur00a, New00b, New01b]. Did [Gre03]. didattica [CSS00]. Diderot [BBL00b]. Die.com [Lev01b]. Diego [IEE01e, MS00b]. Difficulties [FP01]. diffusion [DR03]. digest [IEE01b]. Digital [AA01, Bai01, Ber03, BW02, Cas00a, CZB^{+}01, Kro00, Lit01, Nor01b, Pal01, AR01, Bar00d, Cho01a, Com01, Fit00, FF00, Gil00, Goe02, GM01, HKLS03, JvEB^{+}04, JBR05, KV00, KW06, LSS01, LR02, Gon00, Tur06, TPB01]. digitala [Her01]. digitale [ZB01]. digitaler [LSS01]. digitales [Gon00]. digitally [AR01]. d’images [BG01b]. Dimensioning [BK00b]. direcciones [Ano00f, Ano01c]. Direct [Mea00, Bro00a, Sil00]. Directions [Gib01, Ano00f, Pur02, S^{+}06]. Director [Neu00]. directories [Gre00b]. Directory [HB03, LH00, Was00, Gre02a]. diritto [Dra01, Tos00, ZB01]. dis [Cai01]. disabilities [Kay00]. discipline [Ber03]. disclosure [LW00]. discovery [Spa01]. Discussion [Bla02d, Col00]. Displays [Bar00c]. disruption [Sch00b]. Distance [DMP00, CN00, FJJ^{+}01]. Distributed [ACK00, DWX02, Jep03, LPK02, MRK^{+}03, Rao02, Tso02, CML00a, Goe02, GBCH00, KW06, RR02, GM04]. Distributing [KSSK02]. Distribution [KJL02, SA00, Ver01]. diverse [Gat01]. divide [Cho01a, Com01, HKLS03, JvEB^{+}04, Nor01b]. Diving [LWPM08]. DivX [ZB01]. Dix [BG01b]. DNA [Ano03c]. DNS [AL01, Gre00b, LL01]. Do [Car01b, VS01, Mat02, McC00b, PP01a, Sau01a, Smi00c]. Doctorates [BB00a]. document [Ive00]. Documents [GBM02, Pau03, Sim02b]. Does [BG00c, CML00b, GB00, Owe02b, UJ02]. dogs [Ems00]. doing [BSW00, Bon00, KW02, Pow02]. DOM [Lan05b]. Domain [And03, Bur00, CDT02, Dre01b, FB04, Six00]. domaine [Dre01b]. domaines [Fra00]. Domanennamen [Six00]. données [FTH00]. Don’t [Bar00c]. doorman [Pal02]. Doors [RD05]. Dorado [Gar00c]. DoS-Resistant [ABB^{+}02, Bla02d]. Dot [Dot01, Sun01, AA01, Car01b]. Dot-com [Dot01, Sun01, AA01, Car01b]. Dot.bomb [Kuo01]. dot.coms [LR01a]. Douglas [Dwi02]. Down [KR03, KR05, KR01b, KR08]. downloading [Gil00, GPR^{+}05]. downloads [Pau03]. Drafting [BG00a]. DRI [Sta00d]. drive [Flo01]. Driven [Hou01, Fri01, PP00, PP01b]. drivers [BDFG02, Kha02]. driving [Par05, PB00]. droit [AOB00, FS00, dB01]. Drupal [VW07]. DSOM [ACK00]. Dulles [Ano05b]. dummies [Bri08, LPM01]. Dumpster [LWPM08]. d’un [Ano00c]. d’usages [BG01b]. DVD [Wan06a]. dwarf [Ano00g]. Dynamic [Ano00g, Har00c, Nis01, RB02, Dud08, HP02, Mur00a, WWD04]. Dynamics [OMM02].

e-auction [DV02]. e-Bay [BL00b]. e-business [Mor00, Dif02, WB02, McB02]. e-care [Gol00b]. e-commerce [Bor00, Fit00, Gla00, Gol00b, Kam00, Lev01b, Sin01a, WN01b, KE01, Sta00c, Bal01, KR01a, MM00a]. e-Democracy [Wes02]. E-Empire [KE01]. e-factor [Foc01]. e-Government [Dur02, Wes00a, Wes00b]. E-healthcare [Gol00b]. E-learning [Lai01]. e-mail [Fly01, Swa00b, KM01c]. E-patent [Gla00]. e-Pharmacy [Hun02]. e-PhilanthropyFoundation.org [WHA02]. e-policy [Fly01]. E-politics [Hay00]. e-privacy [MB02]. e-resources [AR01]. e-Ressourcen [AR01]. E-Services [KSSK02]. E-shock [KS01]. E-supply [PB00]. e-Video [Ale00]. E-Volve-or-Die.com [Lev01b]. E. [Coc03a]. E.U. [Gar00d]. E1 [Ano00h]. Earth [Tur06]. easier [Won00]. easily [Sur00]. East [Ros01]. Easy [Ano00h, Lev00, Wan06a]. eBay [Wil06]. eBizBench [LCL04]. eBrands [Car00a]. ebXML [KW02, PN03]. EC [BMP00, BMP01]. EC-WEB [BMP01, BMP00]. economic [GG03, Kim01, LS01b, LR01a, LR01b, Nor00b]. Economically [Hed02]. Economics [Cla03a, Gre02c, Gre03, KV00, SV01, SMKR04]. economy [Ano01e, BK00a, Bra00, Fer02, Gou00, KM01b, LJ01, MVK01, Moh02, Wis00, WN01b, ZW00]. ecosystem [Hak00]. écran [Lej00]. ed [CSS00]. ed.} [Dwi02]. edge [Col00, KM00, KM01b, Mar01]. éditeurs [Fra00]. Edition [BBP00, ET03, Per00, Ada01, Mar01]. Editor [Loc03, Meh02, Met03, San02]. editorial [CS02]. Editors [CP03, DF03, HH03, LV03]. EDP [Red00]. eds [Sti03]. Education [DMP00, Goo00, SNB^{+}00, BMR01, CPM01, MG00, Mer00, PG01]. educational [Gre00a]. educators [Pet01]. EDV [CSS00, Red00]. EDV-Prozess [Red00]. effect [FW00]. effectenrechtelijk [Fei01]. Effectenverkeer [Fei01]. Effective [CSK00, RHN02, vdB00, Fly01, SL06]. Effectiveness [SPP^{+}07]. effects [BSS02]. Efficient [ABB^{+}02, BKB04, Bla02d, Sut00a, SKW^{+}07, SM02b]. eficientes [Gar00c]. EGOV [TL02]. eHRM [Gow01]. Eighth [B^{+}02]. EIGRP [RWS00]. EIS [Ano00h]. EJB [Mar01]. elastic [BK00b, LA02]. elections [PvS01]. Electrical [Str03]. Electronic [Bak00b, Bar01b, BMP00, BMP01, Bro02, Cal00, Cal01, Cal02, CPM01, LPSW04, MSM02, Rub01, Smi00c, SA00, Ste03, TL02, BK00d, CR01a, MRM00, DSM00, EW00, EDGW00, Foc01, Luc02, MWD00, MR00, MB02, RK01a, RS00b, Soo01, UJ02, Wey01]. electronico [MRM00]. électronique [BK00d]. Elegant [CAN^{+}04]. elektronische [PG00]. elettronico [Dra01]. ELT [TG00]. Email [Gag00, Joh00]. Emarketing [SF01]. Embedded [Ano00g, Ben00, DE02, Ins03, Nis01, Woe02, BW00b, Kha02]. Embedding [AAC^{+}00, BW00b, CLWR00, EGH00, PK00, SM00c, Ten00]. Embedix [Ano00g]. emergence [Fra01a]. Emerging [IEE01a, IEE01b, Ini07, Goe02, WN01b]. Emitting [Bar00c]. EMJ [Ano00g]. Empire [KE01, And00a]. empirical [EL01]. employee [DF01, UJ02]. employees [Col00]. Emulation [Ano03c]. Emulex [CZB^{+}01, CTBA^{+}01]. Enabled [Che02, Gre02a, LIB04]. enablers [BDFG02]. enabling [GCB04]. encryption [SKW^{+}07]. Encyclopedia [Ste03, KS00]. End [Ano00g, JT01, Rao02, BC01, CCC^{+}05, Got04, KD05, Wey01]. End-To-End [Rao02, Ano00g, JT01, BC01, CCC^{+}05, Got04, KD05]. Ending [Zdz05, Arn01, RO00]. EndNote [ISI00a, ISI00b, ISI01]. Endpoint [Kad07]. Enforcement [ZY02]. Enforcing [Cen00]. eng [Bro00a]. engagement [Nor01b]. Engineering [Coc03a, DYN03, LWPM08, AJ02, Flo01, FPM05, MFdSeS01, Ver01]. Engines [KRZ02, Nai02, Par05]. English [Jak00, CK00, WSM00]. Enhanced [ML02]. Enigma [Chu02]. Enjeux [CL00]. Enlighten [Ano00h]. EnlightenDSM [Ano00h]. enlightened [Dif02]. Enough [Bar00a, San03]. ensure [BL00a, MB02]. entenderse [Gar00b]. Entera [Ano00h]. Enterprise [ER02, NCC^{+}02, Pic03, GY01, Lla00, OFA00, Sha01, WB02, Ano03c, JWC08]. entire [PB00]. entrepreneurs [Joc01, Pri00]. entropia [Chi01]. Environment [CSK00, HB03, Hel02, Müf01, Phi01, ZY02, BSW00, KB00, LCL04, NLiMK04, Nal02, Sin01a, Var03, WN01a]. Environments [CJL02, CHB02, Jol03, BhHSS00]. epic [Tob01]. épreuve [FS00]. eProcess [KM00]. equality [Agr03]. Era [IEE01b, Bru00, GM03, JSK02, KM00, KM01b, LF00, MFdSeS01, New00a]. Erfurt [WL04]. Ergotracer} [RG02b]. ERICA [KJF^{+}00]. eServer [Ano00g]. ESP [Ano00h]. español [Mil01a]. Essential [Gre02b, Kai06b, O’D00, Wei02, Ano00i, Che00, HB01, HM02b, KM01c, Wal01b]. essentials [Pan04]. ESSIR [ACP01]. establishing [Eng00]. estado [Ano01c]. estate [yJC01]. esthétique [Fra00]. Estimation [DLOM02, Kro00, PJN^{+}02, FJJ^{+}01, STPM05]. Ethernet [cFJH^{+}03, Rey03]. ethical [Gra07, KV08, Til05]. Ethics [Ved01, Lan00a, Lan00b]. ethnographic [MS00c]. Etnus [Ano00g]. EU [Eur00].

EURO [KS05]. EURO-NGI [KS05]. Europe [LPSW04, MSM02]. European [ACP01]. evaluate [Jac00]. Evaluating [BSS02, CML00a]. Evaluation [ABC^{+}07, CJL02, DWX02, PcC02, CRW01, RC05, YFB02]. event [Adv00b, Adv00a, CRW01]. events [Van01a]. every [Dif02, Wil06]. Everyone [Her00, Jon00b]. Everything [CTBA^{+}01, Com00a, Pow02]. evidence [BG00c, Cas00a, CML00b, GLMT05, GB00, RO00]. Evolutionary [Pal01]. evoluzione [Gio00, ZB01]. Evolving [Kle02]. Exam [MD02]. examples [CSMW02, GT02]. exams [Har01]. excellence [KS05]. Exchange [ABB^{+}02, Bla02d]. exchanges [SW01]. Executable [OKB02]. Exhibition [IEE01d]. Expansion [OTPK02]. expect [Fra01b]. expectations [RK01b]. experience [RK01b]. Experiences [Kur02]. experiment [RO00]. Experimental [ABC^{+}07, BhHSS00]. Experiments [Abo03, OTPK02, Bir00]. experts [Adv00b, Adv00a, SP00]. Explained [Gal02b, Los03b, MC00, Pfa01, Rho02]. Explicit [KCC02]. exploit [MM07]. Exploiting [HM04]. exploits [FPM05]. exploration [SKW^{+}07]. Exploratory [San01]. exploring [HPS01]. explosive [LT02]. exposed [HD01, MSK03, SSS06]. expositions [Fra00]. exposure [Ian01]. Expression [CNI^{+}04, Lip00]. Extensible [Coc03a]. extremism [FK01]. Extrusion [Bej06]. Eye [GDC^{+}02]. EZTerm [Ano00h].

f [DF00]. faces [Bar00d, Got04].

Fachbegriffe [KS00]. Facial [CNI^{+}04]. facing [Com01]. factor [Foc01]. Fail [Bar00c]. fairy [GT01]. FAM [KKYK04]. family [FNRC05, Hil00]. Fast [Bar00a, PJ05, SP00, Var05, Pau03]. Faster [Lea00]. Fastest [Bar00c]. Fault [DHL^{+}04, Moo00]. Fault-Tolerant [DHL^{+}04]. FAX [Ano00h]. FC [Fra01c].

Feature [LRJ^{+}03, MRK^{+}03, STHK02b, TWW03, CIA^{+}03, NLiMK04]. Featuring [Abi02, KR03, KR05, KR01b]. February [Fra01c, IEE02b, IEE02a, Juk02, Lev01a]. federal [And00b, Hil01]. federated [Abe03]. Feel [Kro00]. Feet [Bar00c]. Fellow [Dre00]. Fellowship [STT00]. Femtoseconds [Bar00a]. Fewer [Bar00a]. Fi [Ros03, Ros08, Bar03]. Fiber [And03]. Fibre [Fib01]. Field [Bar00a, Mau03, Car01b, Zal05]. Fifth [Juk02]. file [Los00a]. filled [Swe01]. Filtering [SX03, Zdz05]. finance [JL01]. financed [Wei01a]. Financial [Fra01c, Mau01, FSW00, Gla00]. find [Joc01, KC00]. Finding [B^{+}01, Fli00, Kad00, Koz01]. finish [B^{+}01]. Firebrands [MM00b]. Firefox [Che05, HKH05, McF05]. Firewall [Dan02]. Firewall-1 [Dan02]. Firewalls [CBR03, ET03, VE05, ZCC00]. Firmen [Bur00]. Firmen- [Bur00]. firms [Bak00b]. First [Kim01, Mes00, Abe03, BMP00, Cha02, CRS00, GM04, KS05, TL02]. Five [Leh05, Bar00d, EST00, Fra02]. five-module [EST00]. fix [HLV05]. Fixed [ABC^{+}07, Bla05, EP06]. fixpoint [Qia00]. flaws [HLV05]. Flexible [SP02]. Flow [SHK02, BFOBR02, PM04]. Focal [Ste03]. focus [PB00]. folk [Llo01c]. Foo [VGM04]. Food [Smi01b]. force [Jon00a]. forces [Mer01]. forecast [Ano01e]. Forecasting [LT02]. forecasts [Gla00]. Forensic [Joh06, Cas00a, Kip07]. forensics [Ber03, JBR05]. forest [GB04]. Forever [CAN^{+}04]. forging [FHM00]. Formal [OKB02, ABGR01]. Forms [JG08, Bal01]. Foro [Car01a]. Forschung [AR01]. Forum [CLWR00, CZB^{+}01, CTBA^{+}01, CAN^{+}04]. Forward [Bar00c, SP00]. Fotos [DF00]. foundations [Arm00, GT02]. Fourth [BBL00a]. Frame [Sta00b]. Framework [ABC^{+}07, GIM^{+}02, Kro00, LPK02, Tso02, EST00, EW00, KW06, Til05]. France [TL02]. Francisco [Koh02, USE01]. Franklin [Smi00d]. fraud [Col00]. Free [Ano03c, Bar00a, CC01, HH00, Lip00, Haw01, ZB01]. Freedom [Bar01b, KM01b]. Freenet [Bar00a]. French [Ano00c, AOB00, BBL00b, BG01b, BK00d, LD00, CL00, Dre01b, FTH00, FS00, HM00, dL00, Lej00, PJ00b, PJ00a, Soc01, Thi00, WJ00, dB01]. Frequency [DLOM02]. frequent [Sta02b]. frontier [Sam02]. fulfilling [GM01]. full [Bak00b, FHLW03]. fully [Gol00c]. functional [WA06]. Functions [Mit02b, RG02a]. Fundamentals [AGM07, BFS^{+}02, Day08, Dwi02, TV05]. funding [B^{+}01, Pet01]. Fundraising [WHA02]. Future [All03, Met03, PKN^{+}04, Sti02, Amo01, Ano01e, BDFG02, Bod00, Cho01a, CRS00, Dre00, FSW00, FHM00, KK00, Mol02, Nau00, RHC05, S^{+}06, Smi01a, VN03, Ano05b].

G [Coc03b]. G.lite [Ano00g]. GAAP [LW00]. gai [Zhu00]. gaining [KM01b]. Gains [Lea00]. galaxy [Cas01]. gambling [Cen00]. Game [NCC^{+}02, CSMW02]. games [Par05]. gap [LW00]. Garage [HKH05, Owa05]. garden [Gol00a]. gardeners [HH00]. gates [Wer00, And00a]. Gateway [LKP02, deC00]. Gauging [PvS01]. Gazetteers [Rie02]. ged [Bro00a]. Geeks [Kat00, Kat01]. Gefahr [FK01]. Gefahrenabwehr [Ger00]. gender [Nor01a]. Genealogy [Fli00, Cli01, How01]. Generalized [LKKS05]. Generation [Bla02c, CZ04, ML02, MSC00, McF05, RFP02, Str04, Wil02a, ZN05, BI02, DDPP00, DRH09, JC01, KS05, Lou01, RS00a, Wei01b]. generations [Dre00]. généraux [Fra00]. Genomes [RG02a]. gentle [RR03a]. Geographic [Tso02, HL02b, ZFdRD05]. geography [LS01b, SE00, Van01a]. Geospatial [But02]. German [Hoh01, AR01, Bor00, Bur00, Dre00, FK01, Ger00, Hoh01, ILS00, KS00, LSS01, Mer00, Pol00, Red00, RS00b, Six00, SF00, WHdM00, Wey01]. German-Swedish [Hoh01]. Germany [BMP01, BU01, CRS00, FLA^{+}03, KS05, UBM02, WL04]. Geschaftsplattform [WHdM00]. gestalten [DF00]. Get [Lea00, Wei01a]. Getting [Coc03a, Ell00, GM00a, Lev00]. gift [GM01]. Gigabit [cFJH^{+}03]. GIMP [DF00, DF00]. GIMP- [DF00]. Give [CAN^{+}04]. Global [Ano00g, Ano00h, Meh02, BK00a, BK00c, FJJ^{+}01, GR01, Goe02, Gou00, Har00d, KW02, KB00, MVK01, War01, HKT04]. globalisation [CN00]. Globalization [Che02, BK00a]. Globus [SC05]. glossary [Gol00c]. Glucksman [STT00]. GMPLS [FB05]. GNU [DF00, ZB01]. Gnu/Linux [ZB01]. Go [CZB^{+}01, Gre03, Bra01, GPR^{+}05, Man00a, Man01, Pew00]. Goals [PHM03]. Goes [Wes02]. Goethe [Dre00]. Goethes [Dre00]. Going [Fit00, Fra03a]. gold [Ash00, Gou00]. Goldberg [Sti03]. Goliath [Kuo01]. Good [McC00a, Mit02b, San03, GG05, Pri00]. Goods [BBP00]. Governance [ZW00, Ros01]. Governing [Fra01a]. Government [Dur02, GN01, Not00, And00b, Hay00, Hil01, SV01, TL02, Wes00a, Wes00b]. grab [O’R00]. Graduates [CAN^{+}04]. Grafiken [DF00]. Grail [Toi03]. grants [Pet01]. grantwriter [Pet01]. grantwriting [B^{+}01]. graphic [BF01]. Graphs [VHE02]. grassroots [Hil01]. Great [RD05, Tob01]. Greece [BCKT05]. GRID [Kim02, DF03, FK03, FK04]. Grids [BDFG02, DHL^{+}04, cFJH^{+}03, LD03, TT03]. Group [FHM00, dB01]. groupe [dB01]. groupware [Smi00a]. Grows [DF03]. Growth [Rob00, BG00b, LT02, Mer01, PP00, PP01b]. Grundlagen [DF00]. GSM [Cha05]. GTK [DF00]. GTK-Versionen [DF00]. Guest [CS02, CP03, DF03, HH03, LV03, Loc03, Meh02, Met03, San02]. guia [Ano01c].

Guide

[BM02, BK00d, Cal03, Dan02, Dav00, Don03, GM00a, Gre02b, Hal00b, Hel01, LWPM08, Mau03, O’D00, Poo03, Ros02, Til01, Vac06, Wei02, Abe01, AH01, Ada00, Ada01, And01, And00b, Ano00b, Ano00j, Ano01d, Bar02, Bay00, BeM00, BC00, Bla05, Bon00, Bro00b, BSP00, CK00, Car01b, Che00, Cli01, Coh00a, Cri00, Dav01a, DL01, Doc02, DC01, Eng00, Eur00, FF00, GR02, Gil04, Gil00, Gol00c, Gow01, Gra07, Gri00, GM00b, HB01, Haw01, Hil01, HM02b, Hun02, Hus00, Jef00, Joc01, Joh00, Kad00, Kam00, Kam01, Kis00, Koz05, Koz00, Koz01, KV08, Lim06, Lan06, Law00, Lev00, MA03, MG00, MP01, MCH01, Nic01, OG00, Pew01, PHT00, Pow01, Pow02, PJ05, RG02a, Rho00, RG05, Ros08, RS00b, SM00a, SL06, Smi00c, SK00, Spo01]. guide [SR00, Sut00b, THB05, Tre01, TM00, Uni01, UG00, Wal01b, WHA02, Wil06, Woo00, Wor02, Zal05]. Guidebook [SEK01b, ES01]. Gutenberg [BB02].

Hack [KLG02, Syn00, SL06, Til05]. Hacker [CBR03, Gra07, YA04, ET03]. Hackers [SEK01b, Sch00b]. Hacking [DRH09, Dwi09, LWPM08, MSK03, SSS06, VGM04, BK05, KV08]. Hacks [McF05, FW06]. Hall [SR00]. Hand [Lea00]. Handbook [Bou06, Bov00, Fur00, Lus00, Pro00, Say03, Slo00, Sto02, Ano00i, Bra00, BR00, Cal00, Cal01, Cal02, Doc02, Ear06, Far01, Fly01, FF00, JT00, KB00, Lan06, MS00a, McG00, NNM01, OP04, Sta01a, Sta01b, Sta01c, SB01, TK06, Vac01, Vac09, YA04]. handel [PG00]. Handheld [Lea00]. Hands [CL02]. Hands-on [CL02]. happened [Wei01b]. Harcourt [Pow01]. hardware [Ano01d]. harness [Gol00b]. Hash [Mit02b]. hasn’t [McC00b]. hayeren [GS00]. Head [GDC^{+}02]. Health [Ano00b, Dav00, Dav01a, Dav01b, Woo00, Bar02, Edw02, MCH01, OG00, RK01b, WC00]. healthcare [Gol00b, KC00]. heat [BL00a]. Hedy [Bar00b]. held [AD01]. Helix [Ano03c]. Help [Hal00b]. Henry [Puz03a]. herbal [Owe02a]. herbs [Owe02a]. Here [Bar00a, GPR^{+}05]. Herndon [Ano05b]. Herriot [Coc03b]. Heuristically [FKP02]. Hewlett [Ano00h]. Hides [Zal03]. High [ACM04, Ano00h, Bru00, JKH^{+}00, Ken02b, KOS^{+}02, KB00, LCH^{+}08, NKB^{+}03, Sta02a, WV00, BK00b, Dav02a, JWC08, Man00a, Man01, You02, IEE01c]. High-Accurate [KOS^{+}02]. High-Density [Ano00h]. high-intensity [You02]. High-Performance [NKB^{+}03, WV00, JKH^{+}00, LCH^{+}08]. High-Speed [Sta02a, BK00b]. High-tech [Bru00, Man00a, Man01]. High-technology-crime [KB00]. highway [EDGW00, TM00]. Hilton [Ano05b]. Histoire [BBL00b]. historical [ZW00]. History [Leh05, LCC^{+}09, SR00, BB02, BV01, Nau00, Odl00, SK00, TM00, Wer00, BBL00b]. Hoax [CZB^{+}01, CTBA^{+}01]. Hoc [ABC^{+}07, Toh02, AM07, Cha05, Cra07, Lab07]. Hogging [Bar01a]. Holy [Toi03]. Home [Cro00, LKK02, Low05, New01a, O’D00, PKN^{+}04, SEK01b, BB00a, Bri08]. Homeland [RR03b]. Hong [B^{+}02, HL00]. hop [Sob02]. hop-by-hop [Sob02]. host [FJJ^{+}01, Los00b, Ras05a]. Hosted [Abu01]. hosts [ZFdRD05]. Hot [Loc03]. Hotel [IEE01b, Lev01a]. Hotelkeeper [GM00a]. hottest [BL00b, Kad00]. how-to-do-it [Mat02]. HTML [AF02, Cas00b, Jan01, Knu01, Nie06]. HTTP [Zal03]. hu [Zhu00]. Hubs [Tel00]. humaines [Ano00c]. Human [Bar00b, HHH00, Kim01, Gow01]. Human.Society [Kim01]. hundredth [JF00]. hunts [Lei00]. Hurst [STPM05]. hut [Pol00]. hyperlinks [EL01]. Hypermedia [Abi02]. HyperNietzsche [Ano00c]. Hypertext [LSS01, Ano00c, LSS01]. hypertexte [Ano00c]. hypervideo [FB04].

I.T. [KLG02]. IBM [Bar00c, Bon00]. ICANN [Bar01b]. ICATM’01 [IEE01c]. ICDCIT [GM04]. ICMP [McC00a]. ICQT [Sti02]. ICs [Bar00a]. ICSC [HL00]. Idaho [Kat00, Kat01]. IDE [Ano03c]. idea [Pol00, RR03a]. idealists [And00a]. Idee [Pol00]. Identification [OMM02]. Identifiers [Coc03a]. identify [Sin01a]. identity [AML05, Sch00b]. IDEs [Sor01]. idiot [UG00]. IDMaps [FJJ^{+}01]. IEE [IEE01c]. IEEE [ACM04, ASP01a, ASP01b, IEE01a, AH02, ACK00, Bir03, CP03, Cla03a, Coo04, DF03, Fra03a, Gil04, Gil05, GCB04, HH03, Huh03, LV03, Met03, OP04, PKP03, PN03, PHM03, San02, San03, TT03, Vin03c, Vin03b, Vin03a]. IEEE/ACM [ACM04]. IEM [RC05]. IFIP [ASP01a, ASP01b, AH02, ACK00, Juk02, RS00a]. IFIP/IEEE [ASP01a, ASP01b, AH02, ACK00]. Igniting [ACM03]. II [vdMdB01, Nor01a, Smi01c, BS01a, Kut01]. IICS [BU01, UBM02]. III [BS02a]. IIS [ST00, Tre01]. I’ll [PLW07]. illustrated [Koz05]. Ilmenau [BU01]. im [CSS00, FK01, Ger00, Mer00, Pol00, RS00b, Six00, Wey01, RR05b]. iMacs [LPM01]. Image [Bov00, MS01b, San01, ZLL01, RBS02]. image/pattern [ZLL01]. images [BG01b]. ImageStream [Ano00h]. Imaging [TWW03, BS00, BS01a, BS02a]. IMAP [Rho00]. iMigo [Ano00h]. immigration [BeM00]. immutable [RR00]. impact [Cho01b, Hir00, JvEB^{+}04, Liu01, MG02, Ros01]. Impaired [AS02]. imperative [MB02]. impersonal [Ros01]. Implementation [ABC^{+}07, Bra03, BKY^{+}03, Gil07, PcC02, TKF^{+}02, Cha00, CP01, Lay06, PC01, YCY03]. Implementing [Ole03, Sch02a, Bla05, Fly01, SS02]. Implications [CAN^{+}04, Ket02, Bor00, GLMT05, Har00d, Odl00]. Improve [Owe02b]. improved [KKYK04]. Improvements [MB00b]. Improving [GcC02, NKB^{+}03, SHK02, Wes02, ZFdRD05]. inappropriate [Ian01]. Inc. [Ano00g, Ano00h]. incident [JBR05]. including [B^{+}01, Smi00c]. increase [BG00b, CML00b, Swe01]. Increasing [All03, Che01, Sil00, Ste00b]. indexes [Lin00]. Indexing [Tso02, TS00]. India [GM04, Mil01b]. Indicators [KH02]. Indies [Fra01c]. indirect [Zal05]. Indirection [SAR^{+}02]. indispensable [HNV04]. Industrial [CAN^{+}04, Mac03, WB02]. Industries [Gre02c]. Industry [Bar01a, PKP03, Rey03, Bak00b, BG00c, GB00, LT01a, Mil01b, Smi01b]. Infantile [Bar00b]. Inference [KKYK04]. Inferno [SM00b]. inferring [Gao01]. Infidelity [MSCG01]. infinite [Gil00, LR02]. informacion [Gon00]. informatica [dAA00]. Informatika [GS00]. Information [Ano03b, KT00, Kro00, Law04, Lay06, LKK02, Rie02, Sno01, TK06, Tso02, Zha03, AH01, ACP01, BK00c, CML00a, yJC01, Dav01a, Eur00, Gri00, GN01, Hil01, KV00, KKYK04, KK00, KZ01, KB00, KC00, Koz01, Liu01, Gon00, Mil01b, Nor01b, Not00, PL05, PHT00, RS00a, Shu01, Vac09, Wal01b, WN01a, WWD04, ZG04, Cap00b, Kam00]. informatiques [Ano00c]. Informix [Har00c]. InfoSec [BK05]. InfoWorld [Met00]. Infrastructure [Bra03, EKS02, SAR^{+}02, FK03, FK04, Lim06, Ler00b, Ben02]. Infrastructures [PHM03, Sch02a, CPM01]. Initial [Dur02]. Initiation [PJ00a, Puz03a, SJ01, PJ00a]. Innocent [Hus04]. Innovation [ACM03]. Innovations [Gib01]. Innovative [BU01, UBM02]. ins [DF00]. Insight [BU03, IEE02c]. insolvency [BK00a].

Instability [JY02]. install [Ros03]. Installation [Mac03, Ras05b]. Installations [Kro00]. Instant [RR05b, Bay00]. Institute [BBL00a]. insurance [BG00c, GB00]. Integrated [STYC02, WZP00, PK00]. integrating [Lla00, TSC02]. Integration [Vin02, Vin03c, Vin03b, Vin03a, YCY03, GY01]. integrity [Bur00]. intellectual [Haw01, KV00, Lit01]. intelligence [Ano00e, Ber00, Goe02, Mar00]. Intelligent [IEE01c, SX03, ACK00]. IntelliJ [Ano03c]. intensity [You02]. Inter [CDT02]. Inter-domain [CDT02]. interact [Hay00]. Interaction [STHK02b, STHK02a]. interactions [NLiMK04]. Interactive [DADAPW01, SL02, Knu01, Müf01]. Interconnecting [ABC^{+}07]. Interconnection [DT03, DYN03]. Interconnections [Per00]. interconnectivity [BNHGV04]. Interconnects [Loc03]. interdisciplinary [Var05]. Interest [FHM00]. Interface [Wil00, Dud08, SL01, LS00]. Interfaces [Jan01]. Interim [Lea00]. Internal [Ano00g, Bej06]. International [AD01, ASP01a, ASP01b, AH02, ACK00, Ano00g, B^{+}02, EDGW00, FLA^{+}03, HL00, IEE01c, IEE01d, Kim01, KK00, MFdSeS01, Pal01, Smi01a, Sti02, WL04, vdB00, Abe03, BMP00, BMP01, BMM07, BCKT05, Cha02, yJC01, CRS00, Fra01a, Fra01c, FW00, Fri01, GM04, ILS00, JL01, KZ01, Koh02, KS05, MS00b, RS00a, TL02, UBM02, BU01]. internationale [ILS00]. Internationalization [Kap00]. internazionale [Dra01]. Internet [Adv00b, Adv00a, Abe03, AD01, ASP01a, AH02, AR01, Ano00a, Ano00g, Ano00h, Ano01b, Ano03a, Ano05a, Ano05b, BMM07, BCKT05, Bur00, Cha02, CWZ^{+}05, CSS00, CRS00, Dre00, DF00, Dwi02, ET03, Fre01, FK01, GM04, Hoh01, IEE01a, IEE01b, IEE01c, IEE01e, IEE02b, IEE02a, JC01, Juk02, Ker03, Kim01, KS05, Law04, LSS01, MFdSeS01, Puz03a, Puz03b, RS00b, Smi01a, Sti02, SMKR04, SX03, UBM02, WHdM00, WL04, AOB00, BK00d, Dre01b, dB01, Hed02, Abe01, AAC^{+}00, Abi02, AH01, Ada00, Ada01, AT01, Agr03, ABB^{+}02, AWW00, AS02, ASP01b, ABGR01, Ale00, Alf05, All01, dAA00, AN00, AR01, Ame00, Amo01, Ana02, ABC^{+}07, And01, And05, And00a, And00b, Ano00c, Ano00b, Ano00d, Ano00e, Ano00f, Ano00i, Ano00j, Ano01a, Ano01d, Ano01c, Ano01e]. Internet [Ano01g, Ano02, Ano03c, Arm00, Arn01, Ash00, AA01, AOB00, AJ02, Bak00a, Bak00b, BvS00, BBL00a, Bal01, BKB04, BW00a, BC02, BD00, Bar00b, Bar00c, Bar00a, Bar01a, Bar01b, BBL00b, BNHGV04, Bar00d, Bar01c, Bar02, BK00a, BBM^{+}03, BG00a, BDFG02, Bay00, BeM00, BB00a, Bea02, BSW00, BCK06, BFOBR02, Ben02, Ben01, BS00, BS01a, BS02a, Ber03, Ber00, BSS02, Bes00, Bha05, BMR00, BBP00, BC00, Bir03, Bir00, BMR01, Bla00a, Bla00b, Bla01, Bla02a, BCG06, Bla02d, BC01, Bod00, BU01, Bon00, Bor00, BW00b, BG01a, BL01, Bou01, Bow00, BG01b, Bra00, BU03, BS02b, Bra01, BL00a, BK00d, BG00b, B^{+}01, BB02, BR00, Bro00a, Bro00b, BG00c, Bro01, BSP00, Bro02, Bru00, BKY^{+}03, BL00b, BB00b, BGE00, BV01, BS01b, BF01, Bur00, Bur02]. Internet [But02, Cal00, Cal01, Cal02, Cen00, CLZ00, Cai01, Cam00, Cam01, CK00, CSMW02, CN00, Cap00a, Car01a, Car01b, CC01, CIA^{+}03, Car00a, CS02, Cas00a, Cas01, Cer01, Cha00, LD00, CNI^{+}04, CGJ^{+}04, CTLL07, CHCC07, yJC01, Che00, Che01, CDT02, CP03, CL00, Che02, CBR03, CR01b, Chi01, CPM01, CKK01, CdBvdM01, cCP00, CML00b, CJL02, Cho01a, Cho01b, CTYS03, CSK00, Chu02, CHB02, Cla03b, Cla03a, Cli01, Cli00, Coc03b, Coc03a, CW00, Coh00a, Coh00b, Col02, Col01, Col00, Com00a, CL02, CG01, Con02, MRM00, CCC^{+}05, Cot02, CH00, Cra00, CLWR00, CAN^{+}04, Cri00, CP01, Cry01, Cur00, Dot01, DWX02, Das01, Dav02a, Dav00, Dav01a, Dav01b, Dav02b, DE02, Dea00, DDN00, Dei01a, DDN02, DM02, DLOM02, Den02, DL01, DF01, DV02, Dif02, Dil00, DBS00, DHL^{+}04]. Internet [Doc02, DSM00, DF03, DCC00, Dra01, Dre00, Dre01a, Dre01b, Dud00, DL02, Dur02, DR03, EST00, EW00, Edw02, EDGW00, EHT02, EL01, Ell00, ES01, ES03, Ems00, Eng00, EKS02, ET03, EGH00, Eur00, Eve00, Fra00, FKP02, FM05, Fal00, FSW00, Far01, Far03, Fei01, FKC00, FTH00, FS00, Fer00, Fer02, FMM00, Fie01, FB04, Fit00, FM00, Fli00, Flo01, FP01, Fly01, Foc01, FJJ^{+}01, Fra03a, Fra01a, Fra02, Fra01b, FW00, Fri01, FF00, FK01, Fur00, GT01, Gab01b, Gab01a, Gab01c, Gab01d, Gao01, GR01, Gar00a, GSS03, Gar00c, Gar00d, GB04, Gat01, Gei00, Ger00, GS00, GPR^{+}05, GW00, Gil00, Gio00, GDC^{+}02, Gla01, Gla00, Goe02, Gol00a, Gol00c, Gol00b, GLMT05, GIM^{+}02, GT02, Goo02a, GB00, Goo00, Gor01, GG03]. Internet [Got04, Gou00, Gow01, Gra01a, GMP00, GM03, Gra02, GM00a, Gre00a, Gre01a, Gre00b, Gre00c, Gre02c, Gre03, Gre01b, GBCH00, Gri00, GM01, GN01, GBM02, GY01, Gur01, Hak00, Hal00a, Hal00b, Han00, HB01, Har01, Har00a, Har00b, HD01, Har00d, Haw01, Hay00, HC04, HH03, HH00, Her00, Her01, HHH00, HM02a, Hil01, HC02, Hil00, HL02a, Hir00, Hof01, HC01, HNV04, HKT04, HCP05, Hoh01, HM02b, Hoo01, Hou01, HJ04, How01, HL02b, HKLS03, Hub02, HHL^{+}03, Huh03, HL00, Hui00, Hun02, HPS01, Hus00, Hut01, IEE01f, ISI00a, ISI00b, ISI01, Ian01, ILS00, Imp00, Ins03, IAI02, Ive00, JT00, JvEB^{+}04, JF06, Jac00, JSK02, Jak00, Jan00, JRSW04, Jef00, Jen01, JF00, JY02, JLHD01, JL01, Joc01, JT01, Joh00, Jon00b, Jon00a, JKH^{+}00, Kad00]. Internet [KV00, Kam00, Kam01, KS01, KRZ02, Kat00, Kat01, Kay00, KM01a, KM00, KM01b, KSSK02, Ket02, KM01c, Key00, Kha02, KI00, KCF02, KJL02, KCC02, Kim02, KKYK04, Kis00, KOS^{+}02, Kle02, KH02, Knu01, KW02, KC00, Koz05, Koz00, Koz01, Kro00, KR01a, KS00, Kuo01, Kur02, Kur00a, Kur00b, KR01b, KR03, KR05, KD05, Kut01, LA00a, LA00b, LA02, LABJ01, Gar00b, Lai01, LT01a, Lan05a, Lan00a, Lan00b, LF00, LFD01, dL00, LM00, Law00, LH00, LT01b, LPM01, LV03, LS01b, Lea00, Lee01, LKP02, LBJ^{+}02, LKK02, LKMK02, LLL03, LCL04, LMC01, LCC^{+}09, Lei00, Lej00, LSS01, Lem00, Ler00b, Lev01a, Lev00, Lev01b, Lew01, LIB04, Lig03, LR02, LSC^{+}04, LA00c, Lin00, Lip00, LW00, LR01a, LR01b, Lit01, Liu01, LLZ03, Lla00, Gon00, Llo00, Llo01a, Llo01b]. Internet [Llo01c, Llo01d, Loc03, Log00, Lon00, LT02, Los00b, Los00a, Lou01, Luc02, LJ01, Lug00, LRJ^{+}03, MSS01, MSCG01, Mai00, MG00, Man00a, Man00b, Man01, MS00a, MTK03, Mar00, MP01, MCH01, MWD00, Mat02, Mau01, MS01a, Max00, McB02, MSC00, McC00b, MG02, McG00, MB00a, MB00b, MVK01, MM00a, MR00, Meh02, vdMdB01, MB02, Mer01, Mer00, Met00, Met03, Mew01, MRK^{+}03, Mil01a, MS00c, Mil01b, Mir05, Mod00, Moh02, MG01, Mol02, MSM02, MM00b, Moo00, Mor00, MZSR00, Müf01, Mur00b, Nat00, Nat01, NKB^{+}03, Nai02, NLiMK04, Nal02, NCC^{+}02, NDK^{+}03, Nau00, Neu00, New01a, New00a, New00b, New01b, Ngu01, Gal02a, Nic01, NB00, Nor00a, Nor00b, Nor01a, Nor01b, Not00, O’L00, O’R00, Odl00, OMM02, OTPK02, OP01, OG00, Omi01, OLLY02, Opp01a, Opp02]. Internet [Opp07, Ora02, OFA00, OKB02, Owe02a, Owe02b, Pew01, Pew00, Pal02, Pal01, PJ00b, PJ00a, Pap02, PKN^{+}04, PJN^{+}02, PHCMS05, PN01, PKP03, PN03, Pau03, PP01a, Smi02b, Pet01, PLW07, PvS01, PC01, Pho01, PB00, Pol00, PHM03, PK00, PP00, PP01b, PHT00, PFAHA03, Pow01, Pow02, Pri00, Pri02, PG00, Pro02, Pur02, Puz03b, Sta01a, RB01, RBS02, RM00, RG02a, Rao02, RHC05, RK01a, Red00, Ree00, RBB01, RHN02, Rho00, Rho02, RZMD05, Ric01a, RK01b, Ric00, Ric01b, Rie02, RR00, RS02b, Ris01, Rit00, Rob00, RB02, RG02b, Roe01, RFP02, RGE00, Ros01, Ros02, Ros00, RO00, RM02, Rub01, RS00b, Rya00, Sta01b, Sta01c, Soc00, Syn00, Sun01, SP02, Sam02, San02, San03, SGD^{+}02, Sau01a, Sau01b, Sch01, STHK02b, STHK02a, Sch06, SZ01]. Internet [Sch02b, Sch00b, SDB00, Sco00, Sco01, SZSR00, SW01, SHK02, SE00, Sha01, Sha00, SM02a, STYC02, SEK01a, SL02, SKS^{+}02, SEGS03, SV01, Shu01, SNT02, SFFF03, SP00, Sil00, Sim02a, SM00a, Sin01a, Sin01b, SJ01, Six00, Sle00, SL01, Smi02a, Smi00d, Smi00c, Smi00b, Smi00a, Smi01b, Smi01c, Smi01d, SC02, SK00, Sob02, SNB^{+}00, Soo01, Spa01, SEK01b, SF00, Spo01, Sta03a, SB01, Sta00e, Sta02b, Ste00a, Ste01, Ste00b, Sti03, STPM05, SAR^{+}02, SMLN^{+}03, Sto03, SG01, SR00, SM00c, SZGbC04, SM02b, LKKS05, Sur00, Sur01, Sut00b, Swe01, TT03, TT00, TG00, Tel00, Ten00, TD00, Thi00, TS00, TWW03, Tim00, Tir03, Tob01, TKF^{+}02, TMVTdM02, Tos00, TAG00, Tra00, TM00, Tso02, TPB01, USE01, Uni01, Uhl02, UG00, UJ02, Vac01, Van01a]. Internet [Van01b, Var03, VN03, Ved01, Vin02, Vin03c, Vin03b, Vin03a, VS01, VC00, VHE02, Wal01a, WM00, Wan01b, Wan01a, WZZ04, Wan06b, War01, WSM00, WHA02, Was00, WZP00, WHdM00, Weh01, Wei01a, WN00, Wei00, Wei01b, Wel02, WCB01, Moo02, Wer00, Wes00a, Wes00b, Wes02, Wey01, Wie02, Wil00, WS01, WN01a, Wil03, Wil01, WHE00, Wis00, WZ01b, WJ00, Won00, WN01b, Woo00, WC00, Wor02, WB02, WWD04, YFB02, YCY03, You02, Yu00a, Yu00b, Zal03, ZLL01, ZY02, ZR01, ZB01, Zim00, ZFdRD05, ZCC00, ZW00, dB01]. Internet-age [Hir00]. Internet-based [WL04, GY01, DWX02, Ber00, CKK01, Dot01, DBS00, Joc01, Ngu01, Sun01, SDB00].

Internet-challenged [Kro00]. Internet-driven [Fri01, PP00, PP01b]. Internet-HTTP [Zal03]. Internet-Provider [SF00]. Internet-ready [O’L00]. Internet-related [Kim01]. Internet-scale [NDK^{+}03, Chi01]. Internet2 [FHM00]. Internets [Dwi02, Sta02a]. internetworked [FN03, LPK02]. Internetworking [Cis00, Com00b, CS01, Ead03, LS00, Per00, Ahm02]. Interoperability [BM02]. interpersonal [Bar01c]. interplanetary [Hoo01, Bha05]. InterQoS [NCC^{+}02]. INTERWORKING [RS00a]. IntraLinux [Ano00g]. Intranet [Ano03c, Opp01a, Opp02, Ter00, Das01, Gri00, Jan00]. Introducing [Bra01, Ker03]. Introduction [Ana02, Ben02, Bla00a, Byr03, CP03, Col02, Cri03, DF03, Har02, HH03, Knu01, LV03, Loc03, Meh02, Met03, Mit02a, Neu00, San02, Sti03, DSM00, EGH00, Fli00, HC01, RR03a]. Intrusion [Nor01c, Pro00, Ras05a, Wan08, Bej05, NNM01, NN02]. intrusions [Bej06]. iNUX [Ano00g]. Invasion [Sto03]. invasions [Tyn05]. invention [Tob01]. Investigation [Joh06, Kip07]. investigator [KB00]. Investing [Gou00, Fra01b, Nai02]. Investment [Ket02, JL01]. investor [CML00b]. investors [Pew00]. Invocation [Vin03a]. IP [Mou01, Ahm02, Arm00, Ben00, Bla00a, Bla00b, Bla02b, Cal03, CZ04, Cla03b, Com00b, CS01, CP01, DC01, Don03, Gra01b, Hel01, Ins03, Koz05, Los03b, MV03, OP01, Pan04, PC01, PcC02, Puz03a, Räi01, RWS00, Rey03, SX03, VPD04, Wil01]. IP-Based [CZ04, MV03]. iPhone [Dud08]. iPod [Kah05]. IPS [Ras05a]. IPSec [SKW^{+}07, Til01, Los00c]. IPTV [Alf05]. IPv6 [And05, BCK06, Bla05, Los03a, Los04, Met03, Owe02b, VN03]. IQ2 [Ano00g]. IrisNet [NDK^{+}03]. Irresistible [Mer01]. ISAPI [Man02]. Issues [CFHO02a, CFHO02b, Fra03b, Sam02, Shu01, Fre01, Fit00, Gat01, Haw01, Kim01, Kut01, MR00, Nor01a, SV01]. Italy [ACP01, KZ01, Pal01]. ITC [MFdSeS01]. ITC-17 [MFdSeS01]. iteration [Qia00]. ITP [RBS02]. ITV [FB04]. IV [KK00]. IWDC [Pal01]. IXP [LCH^{+}08].

J2ME [Muc02]. Jack [Coc03b]. Jakarta [Goo02b]. Jakarta-Tomcat [Goo02b]. Jamalipour [Puz03b]. January [BS00, BS01a, BS02a, Cal02, IEE02b, IEE02a, Cal00]. Japan [IEE02b, IEE02a]. Japanese [Ano00g]. JASMINE [SEGS03]. Java [Ano03c, AF02, Bar01a, BI02, BKY^{+}03, Cal03, CJ02, CSK00, Eng02, Har00c, HL02b, Kan02, Kro00, KKT04, Mar01, Gal02a, Qia00, RR02, Ric01a, SEGS03, SC05, Swa01]. Java-based [BKY^{+}03]. Java/CGI [HL02b]. JavaScript [AF02, Har00c, HP02, Knu01, Lan05b, Mur00a, NS01]. JAWS [Ano00g]. JetBrains [Ano03c]. JetForm [Ano00g]. jeunes [Fra00]. Jini [YCY03]. jiten [Kur00a]. job [Smi00c]. jobs [Kad00]. John [Coc03b]. Johnston [Puz03a]. Joint [IEE01c]. Jones [Jon00a]. Jose [BS00, BS01a, BS02a, IEE01f, Lev01a]. journ [RK01a]. journal [Ano01a, Lej00]. journalists [McG00]. journey [ABM00, Ros00]. JSF [JF06]. JSP [Mar01]. judaica [Soc00]. judicial [Soo01]. jukebox [Gil00]. Julius [Chu02]. July [IEE01d, Kim01, KK00, IEE01f]. July-3 [IEE01d]. Jumpstart [ES03, Sta03b, VE05]. June [BU01, FHM00, KS05, UBM02]. junk [Pau03]. juridique [BK00d, PJ00a]. juridiques [Ano00c]. juridische [PG00]. jurisprudence [Gar00b]. jurisprudencia [Gar00b]. Juristentreffen [Hoh01]. Just [Bar01a, KM01a, Vin03b]. JWAVE [Ano00g].

Keeping [SEK01b, Gre01a]. Ken [Sti03]. Kernel [CTLL07, HB06]. Key [ABB^{+}02, Bar00a, Ben02, Bla02d, PHM03, Gal02a]. kids [Hir00]. Killer [Bar01a, SW01]. Kit [Bla02b, Mor08]. know [Com00a, Mew01, Pow02]. Knowledge [All03, Rie02, Smi00a, ZR01, DBS00, KM01c, LMV02, SDB00]. knowledge-sharing [LMV02]. Kong [B^{+}02, HL00]. Korea [Cha02, Kim01, IEE01c]. Kuhlungsborn [UBM02]. Kurzweil [Bar00b]. kyokasho [ASY02].

l [CLWR00]. lab [Pan04]. lab-based [Pan04]. Label [Har02, AJ02]. LabVIEW [Tra00]. lakes [Hak00]. Lamarr [Bar00b]. lan [Zhu00, Bar03, MC00]. land [O’R00]. Language [Cry01, Dud00, Log00, WAF00, Zdz05]. language-based [WAF00]. Languages [Dav02b, IEE01d, PN01]. L’année [Soc01]. LANs [Mau03]. Lantronix [Ano00g]. Large [B^{+}02, FLA^{+}03]. largest [LWD07]. Lassen [Pol00]. Last [RO00, WZ01b]. lasting [BL00a]. later [FKC00]. LaTeX2HTML [Yua01]. Latina [Gar00c]. Launching [Fra02]. Lauterkeitsrecht [Bur00]. Lava [Ano00g]. Law [BBL00a, BD00, Con02, DM02, EW00, FS00, HB01, Kam00, Law00, Rob00, Sch01, SG01, AWW00, AOB00, Bal01, BB00a, BSW00, Bla02a, Bor00, BR00, Cal00, Cal01, Cal02, Col01, Gar00d, Gei00, Gol00c, Har01, HC02, ILS00, KM01c, Gar00b, Lem00, Llo00, MSS01, Ree00, RBB01, Sam02, Sau01b, Sco01, Six00, SC02, Sur00, Hoh01]. Lawlinks [Car00b]. Laws [FKP02, BK00a, RR00, SFFF03]. lawyer [BC00, Hoh01, RBB01]. lawyers [Ada00, Ada01, HB01, JT00]. Layer [ABC^{+}07, CJL02]. Layer-2 [ABC^{+}07]. LDAP [Gre00b, Gre02a, HB03, Rho00]. LDAP-enabled [Gre02a]. lead [Wei01a]. leaders [PB00]. leading [Ano01e, SP00, Sur00]. Leads [Bar00a]. League [Bir03]. Lean [Ben00]. leap [Ash00]. Leapfrogging [Mil01b]. learners [Bra01]. Learning [Hir00, Kur02, Pap02, Sco00, FKC00, Gra01a, Har01, Lai01, Log00, Rya00]. lecteurs [Fra00]. Lecture [Dre00]. Lectures [ACP01]. legacy [Mer01]. Legal [BK00a, Bro01, Lon00, Ano00c, BeM00, Bie01, Bor00, Fit00, HB01, Haw01, Kam01, Koz01, Lip00, Gon00, MR00, PJ00a, Pow02, RM00, WHdM00, BK00d]. Legally [TAG00]. legislative [RGE00, Soo01]. Leisure [TD00]. Leonard [Coc03b]. lesson [Bra01]. lessons [Har00a]. level [BFOBR02, CGJ^{+}04, SFFF03]. lexicography [LSS01]. Lexikographie [LSS01]. Lexikon [KS00]. liability [WS01]. lian [Zhu00]. libel [TAG00]. librarian [Cri00, NB00]. librarians [Ben01, Mat02, Sau01a, Woo00]. libraries [Gor01, LM00, Shu01, AD01]. library [Gri00, Liu01, PHT00, Ste01, Ano00d, Jak00]. Licensing [CLWR00]. lies [Thi00]. life [BG00c, Car01b, GB00, SP00]. lifetime [Nau00]. Lifts [Bar00c]. Light [Bar00c, Gre01a]. Light-Emitting [Bar00c]. Like [Bar00b]. likely [AN00]. Limbo [SM00b]. Limited [DLOM02]. limits [Bie01]. Line [Sno01, Moo00]. Lineo [Ano00g, Ano00h]. Lingo [Neu00]. Links [EL01]. L’Internet [Ano01f]. Linux [Ano00g, Ano00h, DF00, BM02, CS01, Cro00, CP01, DMP00, Gag00, Kha02, Kro00, MSC00, Pfa01, PC01, Sor01, TSC02, Zad01, ZB01]. Linux-based [Ano00g]. Linux/Posix [CS01]. Linux/Unix [Sor01]. list [How01, RK01a]. Literacy [CAN^{+}04]. Literature [BSP00]. Little [Got04, WJ00]. Live [LKMK02]. Live-Broadcasting [LKMK02]. lives [FNRC05]. Living [Moo00, Raj00, Log00]. LMDS [Vac01]. Local [SI01, Slo00, Hil01]. Locality [IAI02]. Locating [Sim02b]. Location [Var03, Wie02, BKB04, ZFdRD05]. Log [OTPK02, SNB^{+}00, Koh02, Mai06]. Logic [Dav02b]. logistics [Cho01b]. London [BMP00]. Long [CAN^{+}04, PHCMS05, GLMT05]. Long-range [PHCMS05]. Long-Term [CAN^{+}04]. Look [Kro00, Sti03, GLMT05]. Lookup [GcC02, SMLN^{+}03]. loss [CTLL07, CHCC07, Pol00]. Lossless [Say03]. Lost [CLWR00, Kat00, Kat01]. love [Ash00]. Low [GCB04, Lim06]. low-cost [Lim06]. Low-rate [GCB04]. Loyalty [Sin01a, BG00b, MM00b, Ste00b]. Loyalty.com [New00a]. Ltd [Ano00g, Ano00h]. Lund [Hoh01, Hoh01].

M [ET03, KM01b, Lou01]. M-commerce [KM01b, Lou01]. MA [KK00, Bes00]. Mac [Dud08]. machine [CHCC07]. machines [LRJ^{+}03]. made [Wan06a, Won00]. madness [Ash00]. Magnetic [Bar00a]. Mail [Bar01b, Dil00, Opp07, Swa00b, BB00a, Fly01, KM01c, Pau03, Rho00]. Maintenance [Ole03]. Make [Bar00a, BG00c, GB00, Mur00b, TT00]. Making [BB00b, Hil00, Müf01, MWD00, O’L00]. Mamas [Ems00]. mammals [Gab01a]. man [Dre00]. manage [Gre00b]. Management [ASP01a, ASP01b, AH02, Jep03, LSC^{+}04, Nad03, RD05, Sch02a, Str04, Ter00, Toi03, vdB00, ACK00, B^{+}01, DBS00, Erb03, Gow01, GY01, Jan00, KJF^{+}00, Key00, Lev01b, Mai06, MM01, New00a, Pur02, SDB00, STYC02, Smi00b, Smi01c, TK06, Var03, ZLL01]. Manager [Kro00, Lus00]. Managers [ES03, Mai00]. Managing [Das01, Fri01, Gri00, Ras05b, Sch02a, Smi01d, You02, Moo00]. Manual [Abo03, DF00, Low05, Mat02, Sau01a, WJ00]. manuel [WJ00]. manufacturers [KS01]. manufacturing [Ber00, WB02]. Mapping [Vin03b]. Maps [HL02a]. March [Hoh01, Sta01a, USE01, Sta01b, Sta01c]. Marietta [Pol00, Pol00]. Maris [EDGW00]. Marken [Bur00]. Marken- [Bur00]. Market [BBP00, Lea00, MS01a, PB00, Wis00]. marketers [Mol02]. Marketing [MG02, Zim00, AKY01, Bay00, Bro00a, Cha00, Han00, Hof01, Lee01, LJ01, Moh02, Mol02, New00a, Ric01b, SEK01a, Sil00, Sin01a, Ste00a, Swe01, WZP00]. Marketplace [LPSW04, PLW07]. markets [BK00a, BG00c, DV02, FSW00, GB00, Nai02, TT00]. Marking [KCC02]. marknadsföring [Her01]. marks [Bur00]. marks- [Bur00]. Marques [Dre01b]. Marz [Hoh01]. mass [Sch00b]. Master [BB00a, Kur02]. material [Ian01]. Materialien [Dre00]. materials [Dre00, Ree00]. Math [CAN^{+}04, Van01b]. Mathematics [Ame00, DV02, Gla01]. Matter [Vin02]. Maximal [CAN^{+}04]. Maximizing [CKK01]. May [Bar00a, BMM07, BCKT05, Dav02a]. Mayfield [CK00]. MBone [DMP00]. MD [IEE02c]. Me [Bar00c, CAN^{+}04]. Meaningful [Abe03]. measurement [Lay06, ZFdRD05]. measurement-based [ZFdRD05]. Measurements [KOS^{+}02, WZZ04]. mechanism [WAF00]. mechanisms [ABGR01, Mir05, Wan01b]. Media [CWZ^{+}05, Sch02b, Ste03, BBL00b, BS02b, BB02, DSM00, Gol00c, Hil01, KS00, MR00, deC00]. medias [BBL00b]. Mediator [KRZ02]. medical [Dav01a, KC00]. Medicine [Dav00, Dav01a, Owe02a]. Medien [KS00]. Medien-Lexikon [KS00]. Meeting [Ano05b, BKY^{+}03, Hoh01]. Meets [PKN^{+}04, SL02, BW00b]. mejores [Ano00f]. Mellon [FHM00]. memoir [Ash00]. Memory [cCP00, DBS00, SDB00]. mensonges [Thi00]. mental [BU03]. mergers [BSS02]. merging [TPB01]. Messaging [Pic03, Opp01b, RR05b]. Metadata [Gre00a, Tso02]. MetaFrame [Wil04]. Metasploit [MM07]. Method [CSK00, Dil00]. methodologies [Dot01, Sun01]. Methods [MA02, ABGR01, Gla00]. Metrics [MA02, JSK02]. mezzi [ZB01]. MIBs [Nad03]. Micro [Gre02c, Gre03]. Microcontrollers [Ead03]. Microsoft [And00a, Bar01b, LMV02, Mos05, PP05]. Microsoft.NET [Sta03b]. Microsoft’s [Gre02a]. Middle [Got04]. Middleware [Kro00, LV03, Ler00b, Smi02a, Vin02]. MIDP [Muc02]. Migrating [Bla05]. mile [WZ01b]. MIME [Opp01b]. MIMO [OC07]. Mind [Bar01b]. Minimization [Rao02]. minimize [BG00b]. Minimum [BC02]. Mining [OTPK02, Koh02]. minute [RO00]. Misinformation [CZB^{+}01]. missing [Low05]. missions [RHC05]. Mississippi [Tob01]. Mistakes [Bar00a, Joh03]. Mitwelt [Dre00]. mix [PHCMS05]. MMNS [ASP01a, ASP01b, AH02]. mo [Ems00]. mo-mo [Ems00]. Mobile [CLZ00, CP03, CJL02, Hub02, Pas05, Puz03b, SP02, Sch05, S^{+}06, Toh02, VN03, ZN05, Ahm05, Bla05, Bou06, Bur02, Cra07, Dei01a, OP01, Sin01b, Sto02, Var03]. Mobile-Agent [CLZ00]. mobility [KS05, OP01]. Model [Ano00c, Uhl02, ZY02, LJ01]. Modèle [Ano00c]. Modeling [OMM02]. modellers [Hak00]. Modelling [Hak00, Juk02]. Models [CLZ00, MA02, AT01, BU03, Cra00, EHT02, JSK02, Omi01, SM02b]. Modem [Ano00g, Ano03c, Bar00c]. Modena [KZ01]. Modern [CFLA04, Wan08, DSM00, Har00d, Lan05b]. Modular [RS02b]. Modulation [Bur01]. module [EST00]. Molecular [Byr03]. Molecules [Byr03]. money [Joc01, TT00]. Monitoring [Man02, Bej05, Bej06, UJ02]. Monterey [NB00]. Moore [Rob00]. Moral [dB01]. morales [dB01]. mortar [PP00, PP01b]. most [Tyn05, VS01]. Motorola [Ano03c]. Move [Abe03, GPR^{+}05, Nai02]. Movement [GDC^{+}02, CN00]. Moves [Bar00c, MR00]. Moving [Met03]. Mozilla [McF03]. MP3 [FF00, Gil00, UG00, ZB01]. MPEG [FM05, PFAHA03]. MPEG-2 [FM05]. MPEG-4 [PFAHA03]. MPLS [AJ02, Har02, Nad03, VPD04]. Multi [Raj00, Arm00, AJ02, Hus00, Nal02]. Multi-Operating [Raj00]. multi-operator [Nal02]. multi-protocol [AJ02]. multi-service [Arm00, Hus00]. Multiagent [GIM^{+}02]. Multicast [LLZ03, MA03, TKF^{+}02, WZ01a, MS01a, YFB02]. Multicasting [BC02]. Multicultural [PG01]. Multimedia [Gib01, JZ02, Kur02, Mit02a, PJN^{+}02, PFAHA03, Puz03a, Sch02b, SEGS03, SL01, ASP01a, ASP01b, AH02, AOB00, Gil04, Hil01, Neu00, SJ01, Smi00b, Smi01c, AOB00]. Multiple [Mit02b]. Multiprotocol [Har02]. multithreaded [LCH^{+}08]. Munich [BMP01]. music [FF00, GPR^{+}05, Gil00, LT01a, Mew01, UG00]. musikaliska [Her01]. Must [Kai06a, Mod00]. my [Alf05, BL01, Gab01b, Kuo01]. MySQL [TSC02]. myth [Com01].

nach [Bur00]. Nachwelt [Dre00]. NADIA [SKS^{+}02]. Name [Six00]. Namens [Bur00]. Namens- [Bur00]. Namensschutz [Six00]. Names [Bur00, Dre01b, Six00]. names- [Bur00]. Nanoengineering [Bar00a]. Nanoseconds [Bar00a]. Nanotechnology [RR03a, RR03b]. Napster [Mer01]. Nara [IEE02b, IEE02a]. nation [ABM00, Eve00, GM01]. national [Hil01, Fre01, LH00]. Native [Llo01d]. natural [RO00]. Navigating [RM00, Gur01]. Navigation [TMVTdM02]. nazionale [Tos00]. NDS [Gre02a]. Neal [Ben01]. Neal-Schuman [Ben01]. need [Com00a, Mew01, Pow02]. Needs [CZB^{+}01]. negli [Lug00]. Negotiation [NCC^{+}02]. nel [Dra01]. nell’ambito [CSS00]. nell’era [ZB01]. NEOS [FMM00]. Nero [Wan06a]. Net [Lea00, Mau01, Ano00b, Ano00f]. Net.ObjectDays [WL04].

NetCONNECT [Ano00g]. Netlets [Rao02]. NetMax [Ano00h]. Netspionage [BK00c]. NetWare [Raj00]. Network [Abo03, Ano05b, Bai01, Bar03, Ber00, Bra03, Cro00, CFHO02a, CFHO02b, Day08, EPR03, FKC00, Fra03b, Gil07, GcC02, Hal00b, Har05, Hel00, Ken02b, Kre03, Kro00, LT01b, Leo02, Lus00, McC03, MR07, Nad03, NNM01, NN02, Poo03, RC05, SKS^{+}02, Slo00, Ste03, Str04, Str03, Toi03, Var05, VPD04, Wan08, Wil02a, AML05, Bej05, CdBvdM01, GM00b, Juk02, KS05, KR01a, LCH^{+}08, MSK03, vdMdB01, Mil06, MM01, PT00, PK00, Ras05a, RR02, Syn00, Sch00a, T^{+}07, Thi00, Tur06, VE05, Vac06, WN06, Wyl05, Mil06]. Network-based [Kro00]. Networked [Mes00, WL04, Moh02, Var05, Wel02]. Networking [ACM04, Ano03b, CL02, Cri03, Dwi02, Ead03, Gal02b, Hel02, Ini03, JZ02, KR03, KR05, Lim06, Leh05, Nat00, O’D00, Raj00, Rey03, Sta03a, Tru00, AGM07, Bou06, Bri08, Com00a, HB05, KMK04, KR01b, KR08, Low05, New01b, OFA00, Pfa01, Ros03, Smi00c]. Networks [ASP01a, ASP01b, All03, AH02, ABC^{+}07, And03, Bla02c, BW02, Byr03, CZ04, Cla03b, DT03, DYN03, Dwi02, Fib01, cFJH^{+}03, Gre02b, Jep03, Ken02a, Mac03, MD02, Ole03, PK02, PD03, RS02a, SI01, Sta02a, Sut00a, Til01, Toh02, WV00, Wie02, Wil02b, Zha03, vdB00, Ahm05, ACK00, AM07, Arm00, BDFG02, BK00b, Bla05, Cra07, FM05, Gli06, GCB04, HC04, Hus00, KJF^{+}00, Lab07, Ler00b, Lla00, PL05, PD07, PM04, RS00a, S^{+}06, SC07, Sto02, Tan03, Vac01, Ver01, Wil01, YA04, ZG04]. Networth [Fra01b]. Netz [CSS00]. neue [FK01]. Neuroscience [Byr03]. never [Mod00]. News [Bar00b, Bar00c, Bar00a, Bar01a, Bar01b, Coc03b, Coc03a, Gar00a, Lea00, Pau03, Sto03, VN03, Pew00]. Newton [GY01, New00b, New01b]. Next [And05, CZ04, MSC00, McF05, RS00a, RFP02, Str04, Wil02a, ZN05, BI02, DDPP00, DRH09, JC01, KS05, Lou01, RR03a, Wei01b, WZ01b]. Next-Generation [CZ04, McF05, ZN05, DDPP00, JC01, Wei01b]. NFS [Zad01]. NGI [KS05]. NGOs [War01]. nights [Kuo01]. nine [AA01]. No [LWPM08]. Noble [Kim02]. Nodes [DHL^{+}04]. noms [Dre01b]. Non [CZB^{+}01]. Non-OOSD [CZB^{+}01]. Nonlinear [MS01b]. Nontechnical [Ian01]. Nordic [Leh05]. NORDUnet [Leh05]. normativa [Gio00, Tos00]. North [FHM00]. Norway [RS00a]. Note [CDT02]. Notification [KCC02, CRW01]. Novel [SX03, WWD04]. Novell’s [Gre02a]. November [ACM03, ACM04, AD01, ASP01a, ASP01b, Ber00, CdBvdM01, GY01, IEE02c, NB00, Smi00b]. novice [BU03, Ten05]. NSA [RC05]. NT [Nor00a, Raj00]. NT/2000 [Nor00a]. Numerics [Ano00g]. nuove [CSS00]. nurse [Nic01]. nurses [Edw02, FM00, MCH01]. nursing [Nic01]. nutshell [Hud06]. NUTSS [Fra03a].

OBGP [JY02]. Object [KRZ02, Lav06, PFAHA03, WL04, KW06, IEE01d, WL04]. Object-Based [PFAHA03]. Object-Oriented [KRZ02, WL04, Lav06, IEE01d, WL04]. obligados [Gar00b]. observability [Che01]. occitan [BG01b]. oceans [Gab01c]. October [Adv00b, Adv00a, ASP01a, ASP01b, AH02, GY01, RS00a, Sti02]. ODBASE [Abe03]. Off [Bar00c]. offer [BB00a]. Offers [Ano03c, Lea00]. Office [PP05]. offices [Hil01]. official [Gra07, New00b, New01b]. officials [HC01]. offline [Flo01]. Offs [FKP02]. ok [Got04]. OmniMark [Bak00a]. ONDM [Juk02]. ONE [Pic03, AKY01, Dwi02, ISI00a, ISI00b, ISI01, LJ01]. One-Stop [Dwi02]. One-to-one [AKY01]. ones [Fra01b]. Online [Bar01c, CAN^{+}04, Kam01, Sti03, TT00, BL00a, Dav01b, FNRC05, Fli00, Gar00c, GM00a, Har00b, HJ04, Joc01, KC00, MS00a, Owe02a, Pew00, Pow02, RK01a, Ric01b, Smi00c, Tyn05, WHA02]. Ontologies [Kle02, Abe03]. OOSCD [CZB^{+}01]. OOSD [CZB^{+}01]. Open [BGE00, Cla03a, Kre03, OKB02, Sor01, ZB01, LS00]. OpenLinux [Ano00g]. OpenMail [Ano00h]. OpenMerchant [Ano00h]. OpenML [Bar01a]. openness [Ros01]. OpenPhone [Her00]. OpenSales [Ano00h]. Operating [HB03, Raj00, Wil01]. Operation [Cla03b, RWS00]. Operational [Tso02]. Operations [ACK00]. operator [Kam01, Nal02]. oportunidad [MRM00]. opportunities [Das01, RB01, LSS01]. opportunity [MRM00]. opposing [Cot02]. Optical [And03, Bla02c, Gre02b, Ini03, JY02, Lea00, RS02a, HC04, Juk02, VPD04]. Optimal [JLHD01]. Optimization [Ken02a, LKKS05, FMM00]. Optimized [FKP02]. Optimizing [CTYS03, cFJH^{+}03, Gra01b]. options [LM00]. Opts [Bar01b]. ordinary [ZW00]. ordinateur [Lej00]. Organising [HL02a]. organization [DF01, HM02a, PB00]. organizational [DBS00, Gat01, SDB00]. Organizations [LH00, Das01, Hil01]. organize [ISI00a, ISI00b, ISI01]. organizing [Gre00a]. Oriented [KRZ02, WL04, Lav06, LCL04, IEE01d, WL04]. origines [LD00]. origins [LD00, GLMT05]. OS/2 [DF00]. Other [CLWR00, Cur00, Gre00b, JvEB^{+}04, Met00, RK01a]. our [Bie01, WM00]. outils [Ano00c]. outlines [Jon00a]. Outlook [Mos05]. outsourcing [Mat02]. ouvrages [Fra00]. overlays [Cra07]. Overview [Cis00]. Own [Dea00, AML05].

P2P [Agr03, TT03]. PA [ACM04]. Packard [Ano00h]. Packet [DLOM02, KCC02, SX03, HC04, LCH^{+}08]. packet-switched [HC04]. Packeteer [Ano03c]. Page [OTPK02]. Pages [Kan02, And00b, HP02, Mur00a]. painless [Ros08]. PalmOS [Lea00]. papers [AD01, Har00b, Koh02, MS00b, KS05]. para [MRM00]. Paradigm [FKP02]. Paradigms [Swa00a, Swa01]. Parallel [Ano00g, Chi01, RB02]. parameter [STPM05]. parameters [Jon00a]. part [Abe03]. participate [Hay00]. Participatory [CTBA^{+}01]. Passes [Coc03b]. passion [PP00, PP01b]. passion-driven [PP00, PP01b]. passive [Zal05]. Patent [Sha00, Gla00, O’R00]. path [Sob02]. Paths [Rao02]. patient [Abe01]. pattern [Fer02, ZLL01]. Patterns [Day08, IAI02]. payoff [GG03, LR01b]. PC [Cer01]. PDF [DF00]. PDF-Dateien [DF00]. pearly [Wer00]. Peer [Bir03, Cla03a, Fra03a, TT03, BKB04, Cra07, SMLN^{+}03]. peer-to-peer [Cra07, SMLN^{+}03]. Penetration [WN06, All09, BH08, KLG02, MM07, Til05]. Pennsylvania [FHM00]. pénurie [LD00]. People [AS02, Kay00]. Performance [ACM04, CJL02, GcC02, Ken02a, Ken02b, Man02, NKB^{+}03, SHK02, Sta02a, Ter00, Vin03c, WV00, BvS00, CML00a, CdBvdM01, HC04, Hus00, JWC08, JKH^{+}00, LCH^{+}08, vdMdB01]. performing [Lan06]. périodiques [Fra00]. Perl [DF00, AF02, Sno01]. Persistent [GBCH00]. Personal [Ano00g, FTH00, LKMK02, Lej00, Mau01, SEK01b, GCB04, Tyn05]. Personalisation [HL02a]. Personalization [Abi02, SP02]. personnel [Lej00]. personnelles [FTH00]. perspectief [Fei01]. Perspective [RS02a, EP06]. perspectives [EDGW00, LSS01]. Perspektiven [LSS01]. Pervasive [Bur02, Amo01]. pesquisas [Gar00c]. petit [WJ00]. PGP [Opp01b]. Ph.D.s [Coc03a]. Pharmacy [Hun02]. phenomenon [BL00b]. PhilanthropyFoundation.org [WHA02]. philological [Soc01]. philologique [Soc01]. Philosophical [CAN^{+}04, Ano00c]. philosophie [Fra00]. philosophiques [Ano00c]. Phoenix [ACM03]. Phone [ZN05]. photographer [Far01]. photography [Eng00]. PHP [Abu01, GBR05, Hud06, Lav06, Ler00a]. PHP3 [HW00]. PHP4 [BCK^{+}00]. phreakers [Sch00b]. physical [All09]. physicians [Smi02b]. Phyton [DF00]. PIER [HHL^{+}03]. Pioneer [Bar00b, Coc03b]. Pittsburgh [ACM04, FHM00]. Placement [JLHD01]. plagiarism [LF00]. plan [Bay00, Wei01a]. Planning [Bra03, CW00, Mau01, MA02, Ole03, ES01, Lim06, Ngu01]. plans [Bra01, CML00b]. plants [Bou01]. plastiques [Fra00]. platform [WHdM00]. play [Mor08]. plenty [LD00]. Plug [DF00, Mor08]. plug-and-play [Mor08]. Plug-ins [DF00]. plugins [DDPP00]. Plymouth [Smi00d]. Pocket [Nie06, Cap00a, Ler00a]. Point [Dan02]. points [Koh02]. Policies [Wes02, Fly01, Kut01, SMKR04]. Policy [Str04, ZY02, CG01, Dav01b, Fly01, HM02a, HC02, HM00, Imp00, SNT02, VC00]. Policy-Based [Str04]. Political [dB01, AN00, Gat01, Nor00b]. politics [Bro02, Fra02, GW00, Gra01a, Hay00, Max00, War01, ZW00]. politiques [dB01]. Polymers [Bar00c]. POP [Rho00]. POP3 [Mic00]. popular [VS01]. pornography [Ano02, Jen01]. Portal [Kro00, EL01, LMV02]. Portals [Abi02]. porting [FPM05]. Portugal [BMM07, Smi01a]. Posix [CS01]. possibilities [CR01a]. Possible [SM02a]. post [LJ01, Mol02, Cer01]. post-Internet [LJ01, Mol02]. post-PC [Cer01]. poster [Yu00b]. potential [FHLW03]. poteri [Lug00]. pour [Ano00c, Fra00]. poverty [LD00, Nor01b]. Power [FKP02, Mau03, SFFF03, Gol00b, GBR05, HPS01]. practica [Gar00c]. Practical [Cal03, DJK02, Don03, GSS03, Gra05, Kha02, Mac03, Poo03, Pro00, RS02a, Rey03, Sau01b, Str03, Bla05, Doc02, DC01, Har05, Lim06, Law00, Lev00, MA03, Mol02, RG02a, RG05, SM00a, WHdM00]. Practice [Kai06a, Rap02, Cal00, Cal01, Cal02, Cha00, Fie01, HB05, Lee01, Los03a, Los04, Opp09, Sta01a, Sta01b, Sta01c]. Practices [CFHO02a, CFHO02b, DT03, Fra03b, Pic03, McB02]. practitioners [HB01]. Pragmatic [Mas06, THB05]. praktische [WHdM00]. prassi [Gio00]. predicted [Alf05]. Predicts [Bar00b]. Prentice [SR00]. prep [KV08]. Prepare [CAN^{+}04]. preparing [KR01a]. prescriptions [Nat00]. presence [Bay00]. Presentation [Wil04]. Press [dL00, Gol00c]. presse [dL00]. pressure [BL00a]. Presumption [Vin03c]. prevail [Fra01b]. Prevention [EKS02, Ras05a]. pricay [HD01]. Price [NCC^{+}02, Dav02a, RO00, Wil06]. Pricing [LIB04, MB00a, MB00b, CSMW02, PT00, SMKR04]. primary [Gla01, PN01, SE00, SV01]. Primer [Har02, AV04]. principle [Got04]. Principles [Com00b, DT03, Han00, Hel00, PK02, Rap02, Ric01a]. Print [Ano00h, KS00, KC00]. priorities [Sim02a]. priority [LSC^{+}04]. Privaatrechtelijke [PG00]. Privacy [EKS02, KH02, Con02, Haw01, Imp00, JF00, Lug00, MB02, Smi00d, Tyn05]. Private [Til01, Six00]. privati [Dra01]. privatrechtliche [Six00]. Pro [Ano00g, Ano00h, JF06, VW07]. proactive [Ten00]. Problem [Vin03b]. Probleme [Red00]. problèmes [Ano00c]. Problems [VM02, Ano00c, FMM00, Red00]. Proceedings [ACM04, FLA^{+}03, IEE01d, IEE01e, IEE02c, USE01, AH02, ACK00, BMP00, BMP01, B^{+}02, BU01, Cha02, FHM00, Fra01c, IEE01f, IEE02b, IEE02a, KK00, KZ01, Lev01a, MFdSeS01, Pal01, RS00a, RGE00, Smi01a, Sti02, TL02, UBM02, BCKT05, GM04, ASP01a, ASP01b, BMM07, CRS00, HL00, Kim01, NB00, WL04]. proceedings-2000 [NB00]. process [Red00]. Processes [ZR01]. Processing [Bov00, CJL02, Gra01a, MS01b, Zha03, Bla01, ZG04]. Processor [CFHO02a, CFHO02b, Fra03b, GcC02, LKKS05, LCH^{+}08]. Processors [cCP00, Gil07]. Produces [CAN^{+}04]. producing [Ale00]. Production [Pri02, LSS01]. Products [Ano00g, Ano00h, Ano03c, Kro00]. Produktion [LSS01]. Professional [Bir01, Hel01, Bea02, Gri00, KC00, Lai01, PHT00, Ten05]. professionals [Edw02, FPM05, MCH01]. professionellen [DF00]. Professor [Bar00d]. profile [Sta02b]. Profiling [CHB02, MS00b]. profit [PP01a]. profitable [Mor00]. profits [Sil00]. progetto [CSS00]. Program [Kur02, Cal00, Cal01, Cal02, Cli01, DDN00, Dei01a, Dei01b, DDN02, Sta01a, Sta01b, Sta01c, STT00]. programmable [NLiMK04]. Programmer [LS01a, Rho00, Sta00b, Tay00, Woe02]. Programmers [Cal03, Don03, DC01, Hus04, MA03]. Programming [Bar00a, Coc03b, CS01, CLWR00, Dav02b, Gil07, Müf01, RS02b, STK02, SC05, Swa00a, Swa01, WZ01a, Woe02, AF02, Bak00a, Chi01, GBR05, Har00c, HLV05, Kha02, Knu01, KKT04, Kur00b, RR02, SM02b, TSC02]. Programs [Jan01]. Project [Her00, Kro00, B^{+}01, Mai00, BF01]. projects [RG05, Tob01, You02, LH00]. Projekts [CSS00]. promise [Agr03, DF01, GM01, LR01a]. Promises [Lea00, CH00]. promote [Swe01]. promotion [Won00]. proofing [Syn00]. Propagation [BCK06, OC07]. Propellerhead [Jon05, PJ05]. Properties [Bar00a]. property [Haw01, KV00, Lit01]. prophets [Dif02]. Proposed [CIA^{+}03, MB00b]. prosecution [Ger00]. prospects [CH00]. prosper [DF01, Man00a, Man01]. Prosperity [All03]. protect [MB02]. Protecting [Bla00b, WM00, JF00, Lit01]. Protection [EKS02, Wan08, FTH00, PJ00a, Six00, VPD04]. protezione [ZB01]. Protocol [And05, BGE00, LKP02, Puz03a, Zal03, AJ02, CP01, Los03a, Los04, PC01, SMLN^{+}03, Alf05, BBM^{+}03, Ker03, Mil06, OKB02, RBS02, Sin01b, SJ01, deC00]. Protocols [ABB^{+}02, Bla00a, Bla00b, Bla02d, CZ04, CJL02, Cla03b, Com00b, Far03, Hal00b, Jep03, Per00, Puz03b, SZ01, Toh02, WZ01a, Bla01, BS02b, CJM00, Dwi09, HB05, JC01, Joh00, Koz05, MR07, HCP05, Sta03a]. proven [Sil00, Swe01]. Provence [TL02]. provider [Mat02, SF00, SF00]. providers [BSS02, MV03, PG00]. Providing [Opp07]. provisioning [Sti02]. Proxies [JLHD01]. Proxy [CWZ^{+}05]. Prozess [Red00]. Psychological [Bir00, Nor01a]. Pty [Ano00g]. pubblici [Lug00]. Public [Imp00, PHM03, Bar02, Cri00, HC01, Gal02a, Ben02]. publication [LSS01]. publications [Hil01]. Publikation [LSS01]. Publish [Yua01]. Publishing [ZR01, KV00, MR00, RS00b]. Publizieren [RS00b]. punditry [Met00]. putting [WB02]. Python [TSC02].

QofIS [CRS00, Smi01a, Sti02]. QoS

[ABGR01, Hus00, KD05, LIB04, NCC^{+}02, PcC02, Sob02, Sti02, TKF^{+}02, Wan01b, YFB02]. QoS-aware [YFB02]. QoS-Capable [PcC02]. QoS-enabled [LIB04]. QoSMIC [YFB02]. qualitative [MS00a]. Quality [Arm00, BFOBR02, Hel02, Räi01, Smi01a, Sta02a, Sti02, CdBvdM01, MTK03, SL01, Wan01b, CRS00, Gil00]. Quantifying [NKB^{+}03]. Quantum [Coc03b]. Quelltext [DF00]. Queries [KRZ02]. Query [OTPK02]. Querying [HHL^{+}03]. Questions [Ano00c, MSM02, MSS01, VC00]. Quests [BG01a, BV01, GT01, Gab01c, Llo01a, Llo01b, Llo01c, Van01a, BL01, Bou01, BS01b, Gab01b, Gab01a, Gab01d, Llo01d, Van01b]. queuing [LT01b]. Quick [MG00, CK00, OG00, Ano00h]. Quick-Connect [Ano00h]. quickly [Sur00]. quota [LSC^{+}04]. quota-based [LSC^{+}04]. quotation [RS00b]. quotes [Pew00].

R [ET03, DF00]. race [AA01, Nor01a]. racing [Par05]. radar [GG05]. Radio [Har00a, KS00, Pri02, Nau00]. radiocesium [Hak00]. Radisson [IEE01b]. Rahmen [CSS00]. railroads [Nai02]. rails [THB05]. raising [VC00]. Randomization [SPP^{+}07]. range [GLMT05, PHCMS05]. Rapid [McF03]. rapidly [BSW00]. rapport [dB01]. rate [GCB04, LA00a, LA00b, LA02]. RBAC [ZY02]. RDCT [WWD04]. reader [Tel00]. Readings [JZ02, Ric01b]. ready [Bra01, O’L00]. ready-to-go [Bra01]. Real [CNI^{+}04, CAN^{+}04, DWX02, JBR05, BB00b, yJC01, Gre01a, MSCG01, MM01, OC07]. Real-Time [CNI^{+}04, DWX02, MM01]. real-world [OC07]. Realistic [LD03]. Reality [CAN^{+}04, HL02b]. Really [CZB^{+}01]. RealNetworks [Ano03c]. Reason [Jon05, PJ05]. Reasoning [Abi02]. Rebol [Swa00b]. Receive [Coc03b]. reception [Dre00]. recherche [Ano00c]. Recht [Hoh01]. rechtliche [WHdM00]. Rechtsextremismus [FK01]. recommendation [KKYK04]. recommended [DL01]. reconfigurable [WWD04]. reconnaissance [Zal05]. record [Pol00]. recovery [CHCC07, VPD04]. recreation [Ora02]. Red [Ano00f]. reduce [Ian01]. Reducing [Ano05a]. Reference [Ano03c, Dwi02, Hel01, Nie06, FB04, Koz05, Ler00a, Sau01a, Sco01]. references [ISI00a, ISI00b, ISI01]. referencia [Soc00]. reflections [Cas01]. reflective [PG01]. Reflectometers [And03]. regime [Fra01a, Gon00]. regimen [Gon00]. region [AS02]. regional [MRM00, Hil01]. regions [Fra02]. Regularity [DL02]. regulated [Uni01]. Regulation [Roe01, Gar00d, SC02].

Regulations [vdB00]. Reinvigorating [GW00]. related [Hil01, Kim01]. relating [Kam00]. relationship [AKY01, Hil00, New00a, OFA00]. relationships [Bar01c, Gao01, MSCG01, Ste00b]. Relay [Sta00b]. Released [Bar01a, Bar01b]. Releases [Ano03c]. reliable [GPR^{+}05]. Relief [Bar01a]. Religion [CH00]. Religionsunterricht [Mer00]. religious [Mer00]. reloaded [SL06]. Remaining [MSM02]. remedies [Col00]. Remote [MD02, Pal02, Rub01, Ras05b]. Remotely [LKK02]. Repelling [CBR03, ET03]. replacement [PB03]. replica [Owa05]. Replicated [JLHD01, RB02, Yu00a, Yu00b]. Report [Abe03, Ano01b, Ano03a, Ano05a, Fre01, PKP03, Adv00b, dB01]. reporters [Pew01]. reports [STT00]. representation [AN00, SM02b]. Representative [CTBA^{+}01, CGJ^{+}04]. Representing [GB04]. Request [KJL02]. requirements [Fer02]. Research [CIA^{+}03, CDT02, AR01, Ano00c, Ano01e, BDFG02, BeM00, BC00, Bro01, CH00, Doc02, Fre01, Fie01, Har00b, LM00, Lee01, Lon00, MS00a, MM07, MG02, RM00, STT00]. researcher [Koz01]. Researchers [Bar00b, Doc02, McG00]. researching [AH01, MS00a, Weh01]. réseau [Thi00]. reservation [CKK01]. Reshaping [War01]. Resistant [ABB^{+}02, Bla02d]. Resource [Bla02b, Coc03a, MM01, Ame00, Bar02, BBM^{+}03, Bra00, CR01a, Gow01, HB01, Koz01, MCH01, Pew01, Pet01, Rya00, STYC02, Smi00c, Uni01, Wor02]. resource-based [Rya00]. Resources [TD00, Abe01, AR01, yJC01, FM00, Gre00a, Hil01, Jac00, JL01, MWD00, RK01a, Ric01b, Swe01, Tim00, TM00, Was00, Woo00]. resources.com [Was00]. response [JBR05, Ras05a]. responsibility [WM00]. responsible [Erb03]. Ressourcen [AR01]. restart [CTLL07]. restoration [VPD04]. restricting [Ano02]. resumes [Smi00c]. retail [Bak00b]. retailers [KS01]. retailing [MZSR00, SZSR00]. retain [Sin01a]. rete [CSS00]. Rethinking [BC01]. Retrieval [Rie02, San01, ACP01, CML00a, KT00, WWD04]. Retrieving [Tso02]. Return [Day08]. retuschieren [DF00]. reveal [SP00]. reverse [FPM05]. Review [Dwi02, ET03, Law04, Puz03a, Puz03b, Sti03, Zal03, Ano01a, Gra07]. Revised [KS05, ACP01, Gol00c, Koh02, MS00b]. revolution [FF00, KR01a, LR01b, Mew01, SW01, WZ01b]. revolutionize [PB00]. Rezeption [Dre00]. RF [Wei02]. RFCs [Los00c, Los00a]. rich [JF06]. Richardson [IEE01a, IEE01b]. riches [AA01]. richtig [DF00]. riding [CC01]. Right [VM02, Bur00, FK01]. Right-wing [FK01]. rights [Col00, HHH00, Kis00, SF00, dB01]. Rip [Gar00a]. rise [Tur06]. Risk [Cam01, BMR00, BG00b, Cam00, Cur00, Lan06]. risks [Jan00, PvS01, Sch00a, WN00, Wei00]. Rivest [Coc03b]. RMI [PKN^{+}04]. Roaming [LKP02]. Robocopter [Bar00c]. robot [Gol00a]. Robotic [Bar00a]. Robots [Bar00b, Sti03, SM00c]. Robust [CIA^{+}03, PJN^{+}02]. Rock [Smi00d]. rode [Kat00, Kat01]. Rodelu [Car01a]. Roland [Sti03]. role [BhHSS00, Liu01]. ROM [Ste03, AOB00]. ROMs [Har00b]. Ronald [Coc03b]. Rootkits [HB06]. Roots [CAN^{+}04]. Round [OMM02]. Round-Trip [OMM02]. Route [GcC02]. Router [DDPP00, Lea00, PcC02]. Routers [Ano00h, Per00, TKF^{+}02, DDPP00]. Routing [CDT02, Hui00, JY02, Ken02a, Kim02, PM04, BNHGV04, GR01, Hal00a, LABJ01, MR07, Sob02, YFB02]. RPC [DHL^{+}04]. RPC-V [DHL^{+}04]. RTEL [Ano00g]. RTP [Ano00h]. RTSP [Ano00h]. RTSP/RTP [Ano00h]. Rubin [ET03]. rules [Bro00a, KS01, Mur00b, Pri00, Ros01, RO00]. Ruling [Spa01]. running [KI00]. ruseren [GS00]. rush [Ash00, AA01, Gou00].

S [Opp01b]. S/MIME [Opp01b].

Safeguarding [Bar03]. safer [KM01c]. safety [Cur00]. SAFKASI [WAF00]. SAINT [IEE02b, IEE02a]. sales [CC01]. Salesman [Bar01b]. Salvador [MFdSeS01]. San [BS00, BS01a, BS02a, IEE01e, IEE01f, Koh02, Lev01a, MS00b, USE01]. Santa [AH02, IEE01d]. sapping [Pew00]. SAR [B^{+}02]. SAS [Ano00g]. Satellite [Mea00, TMVTdM02]. Satisfy [PHM03]. savant [Ano00c]. savvy [Hir00]. saw [GM03]. Says [Bar01a]. SC2002 [IEE02c]. SC2003 [ACM03]. Scalability [Owe02b]. Scalable [Hed02, Sch06, SMLN^{+}03, WCB01]. scale [Chi01, Wil01, NDK^{+}03]. scaling [RZMD05]. Scam [Ems00]. Scarcer [Coc03a]. scavenger [Lei00]. Scheme [DF00, Gra01b, SHK02]. scholars [BSP00, LM00]. school [Ano00i, Gre01b, ACP01]. schools [BB00a, SM00a]. Schreiben [Pol00]. Schreiben-Lassen [Pol00]. Schuman [Ben01]. Schwedisches [Hoh01]. schweizerischen [Bur00]. Science [Bar01a, Coc03b, Coc03a, CAN^{+}04, HL00, Yua01, AR01, Cas00a, FKC00, Shu01]. sciences [AR01, Ano00c, Doc02, dB01, dB01]. SCO [DF00]. screen [Lej00]. Screenwriting [Weh01]. script [Weh01]. SCTP [CIA^{+}03]. SDH [VPD04]. Search [KRZ02, Tir03, Uhl02, ISI00a, ISI00b, ISI01, Smi00c]. searcher [AH01]. searchers [BU03]. Searching [Tso02, AH01, Sha00, TS00]. Second [ET03, IEE01f, Per00, Smi01a, Tay00, BMP01, GLMT05, RO00, Sti02, UBM02]. second-price [RO00]. secret [FNRC05, Gal02a]. Secrets [RD05, VGM04, BL00b, Che05, Col00, MSK03, VS01]. Secure [ABB^{+}02, Bar03, Bla02d, GBM02, LPSW04, LKP02, McB02, Opp01b, Sto03, VM02, HL03, Jon00a, HB03]. Securing [Nor00a, McB02]. Securities [BSW00, Roe01]. Security [AM07, Ano00g, Bar00b, Bla00b, BCG06, CBR03, ES03, ET03, GSS03, Ken02a, Lan05a, McG06, Opp00, Opp03, Pho01, PHM03, Poo03, RR03b, RR05b, Rub01, San03, SEK01b, VM02, Wan08, All09, BH08, Bej05, Bej06, Cha05, CJM00, Con02, Ear06, ES01, FPM05, Har05, HLV05, Hus04, Jan00, JRSW04, JF00, JBR05, Kad07, KLG02, Lan06, Lay06, LWD07, MSK03, Nor01a, Opp01a, Opp02, RR05a, Ris05, RC05, TK06, Vac06, Vac09, WAF00, WA06]. SEDA [WCB01]. seeking [Pet01, VC00]. Segment [CWZ^{+}05]. Segment-Based [CWZ^{+}05]. selected [KS05]. Self [GLMT05, HL02a, Wyl05]. self-defense [Wyl05]. Self-Organising [HL02a]. Self-similarity [GLMT05]. sell [BK05]. selling [TT00, Weh01]. Sellphotos.com [Eng00]. sells [Wil06]. Semantic [Law04, Sch00a, AV04, FHLW03]. semantics [Abe03]. SEMPER [LPSW04]. sending [Pau03]. Sensing [NDK^{+}03]. sensitive [LA00a, LA00b]. Sensor [Zha03, Lab07, S^{+}06, ZG04]. sensors [GCB04, PK00]. Sentient [Huh03]. Seoul [Cha02, IEE01c, Kim01]. September [ACP01, BMP00, BMP01, CRS00, FLA^{+}03, IEE01b, KZ01, MFdSeS01, Pal01, Smi01a, TL02, WL04]. serials [FHM00]. Series [Ano00h, Ano03b]. serve [Sin01a]. Server [Ano00h, Ano03c, Bar01b, CS01, Gag00, Jan01, Kan02, KJL02, KOS^{+}02, LKK02, Wil04, LMV02, Ano03c, Pic03]. Servers [Ben00, DADAPW01, JLHD01, LKKS05, JKH^{+}00, LRJ^{+}03, Nor00a]. Service [CAN^{+}04, Gar00a, GBM02, Hel02, LKP02, Räi01, RFP02, Sta02a, vdB00, Arm00, Bak00b, BFOBR02, Cai01, CRW01, CTLL07, CHCC07, FJJ^{+}01, Hus00, Mat02, Mir05, MV03, NLiMK04, PG00, Ste00b, Wan01b, Wes00a, Wes00b]. ServiceGlobe [KSSK02]. Services [ASP01a, ASP01b, AH02, ACK00, Ano00g, Ano00h, BBP00, CW00, GIM^{+}02, Jol03, Kan02, KSSK02, LV03, ML02, MB00a, MB00b, MA02, NKB^{+}03, NDK^{+}03, OTPK02, Opp07, PKP03, PN03, STK02, SC05, Sti02, Swa00b, TWW03, Tso02, Wil02a, Amo01, Cha02, yJC01, CJ02, CTYS03, CRS00, Gla00, Gre00c, GBCH00, Gri00, Har00b, JC01, JL01, KW06, Liu01, Lou01, MV03, PT00, Pas05, RS00a, Ras05b, SS02, STYC02, SJ01, Smi01a, SZGbC04, WCB01, WA06, Yu00a, Yu00b, HB03, Puz03b, Wil04]. Session [Puz03a, Yu00b, SJ01]. Set [Leo02, Sta00c]. seven [Sil00]. Sex [Thi00]. Sexe [Thi00]. SGML [LSS01, LSS01]. SGML/XML [LSS01, LSS01]. Shamir [Coc03b]. shaping [OFA00]. shared [BhHSS00]. shareholder [Moo00]. SharePoint [LMV02]. Sharing [Ano03b, LKKS05, BBM^{+}03, LMV02, Nal02]. shellcode [FPM05]. Ship [RG05]. shock [KS01]. Shopping [EKS02, BMR00]. shorten [WWD04]. Should [Bar00c, Uni01]. Shoulder [LWPM08]. sic [HD01, IEE01c]. Side [CLWR00, JvEB^{+}04]. siécles [BG01b]. Siegwart [Sti03]. Signal [GG05]. Signatures [CZB^{+}01, Nor01c]. Silence [Zal05, Arn01]. similarity [GLMT05]. simple [ISI00a, ISI00b, ISI01]. simplify [GBCH00]. simulating [FP01]. Simulation [Abo03, SI01]. Sinnreich [Puz03a]. sins [HLV05]. SIP} [Puz03a, SJ01]. Site [Dea00, Pal02, Fra00, Smi00d, Swe01]. Sites [KH02, And00b, Ano00i, Cap00a, DL01, How01, Sur00]. Sixth [Hoh01, Log00]. size [WWD04]. skills [Gla01, PN01, SE00, SV01]. skillz [BK05]. Skip [CAN^{+}04]. skydd [Her01]. slander [TAG00]. Small [Kro00, Pan02, RZMD05, STHK02b, STHK02a, Ano00e]. Small-time [RZMD05]. Smart [ZN05, McC00b]. SMTP [Rho00]. snap [ISI00a, ISI00b, ISI01]. SNMP [Cis00]. Snort [OBB05]. SOAP [BI02, Eng02, LPK02, SS02, STK02]. SOAP-Based [LPK02]. Social [LWPM08, Ano00c, BB02, Doc02, HM02a, Lip00, MP01, Nor00b, Sur01]. Socially [Erb03]. Society [HJ04, Kim01, AWW00, Cas01, RS00a, Sam02, Sle00, War01, Wel02]. socio [Kim01]. socio-economic [Kim01]. Sockets [Cal03, CS01, Don03, FPM05, DC01, MA03]. Sofia [AD01]. Soft [Ano03c]. Software [Ano00g, Ano00h, Ano03c, Bar00a, CLWR00, HM04, Lem00, McG06, Pau03, Red00, SA00, VM02, Ano01d, DDPP00, Fit00, Fly01, Gla00, HLV05, Ive00, RB01, RG05, WA06, ZB01]. Solaris [DF00, HB03]. Solution [Ano00g, Lea00, ISI00a, ISI00b, ISI01]. Solutions [Ano00h, Che02, ER02, Str04, Jan00, Mai00, MSK03, MV03]. solving [FMM00]. sometimes [FNRC05]. SONET [BW02, VPD04]. SONET-SDH [VPD04]. Sophisticated [Kro00]. soul [BK05]. Source [Ano03c, Kre03, Sor01, ZB01]. Sourcebook [Ahm02, Hil01]. Sources [Rie02, Cur00, Eur00, Gla01, PN01, SE00, SV01, Tim00]. Space [BKY^{+}03, DLOM02, Nis01, SPP^{+}07, DF01, HCP05, OC07, RHC05, Wer00]. space-time [OC07]. spam [Zdz05]. Spammers [Pau03]. Spanish [Ano00f, MRM00, Gar00b, Gon00, Mil01a]. Spatial [IAI02]. speaking [TAG00]. Spec [Bar01a]. special [MG00]. Specification [OKB02]. Spectrum [Bar00b, VHE02, STPM05]. Speech [Bar01b, Haw01, Sch01]. Speed [Bar00c, IEE01c, Sta02a, BK00b, Car00a, Gre01a, KR01a]. Spinning [FHLW03]. spokes [Tel00]. Spotlight [PN03]. Spread [Bar00b]. Springer [Pho01]. spying [FNRC05]. SRUTI [Ano05a]. SSL [Opp09]. stability [JT01]. Stabilizing [CDT02]. Stable [GR01]. Stakes [CL00]. Stalker [Ano00g]. Standard [CIA^{+}03, Lea00, Qia00, Gil04, HCP05, KW02, Sta00c]. Standards [Bar00c, Coo04, Los00b, Zel07]. Start [Par05, AA01, B^{+}01]. start-ups [AA01]. started [Ell00, Lev00]. Starting [KI00]. State [Bar00b, MSM02, Eve00, Hil01]. Statehouse [Pew01]. states [Eve00, New01a]. Stati [Lug00]. Statistical [Wan08, Zdz05]. Statistically [Wan08]. Statistically-Based [Wan08]. Staying [McC00b, Ste00b]. Steal [Wan01a, Wan06b]. Stealing [AML05]. step [And01, Fli00, Gil00, ISI00a, ISI00b, ISI01, SL06]. step-by-step [And01, Fli00, Gil00, SL06]. Steps [Ano05a]. Steven [ET03]. Steward [Tur06]. stock [Col00, Ems00, Eng00, MS01a]. Stocks [Ket02, Gou00]. Stop [Dwi02]. Storage [ACM04, Coc03a, Fib01, Jep03, Toi03, Pau03]. Store [Bar01b, CTBA^{+}01]. stories [Jak00]. Storm [Ano00h]. story [BB00b, PP01a, Tob01, Wal01b]. Strafverfolgung [Ger00]. Strategies [BCK06, Luc02, vdB00, AT01, Amo01, Bru00, CC01, Cra00, Dif02, Flo01, Gla00, Hus00, Ian01, JWC08, KS01, MVK01, Mol02, PB03, Sil00, WS01]. Strategy [Abi02, KJL02, KCC02, NCC^{+}02, AKY01, Cha00, Gri00, YA04]. straw [Pol00]. Streaming [CWZ^{+}05, KD05, Pri02, Sch02b]. Streams [DLOM02, BS02b]. street [McC00b]. Streetwise [Wei01a]. strengthen [Lew01]. Strohhut [Pol00]. Stromix [Ano00h]. Structure [VHE02, Wis00]. Structured [Cra07]. structures [GBCH00]. Structuring [BS02b]. STS [Ano00g]. Student [LF00, STT00]. students [BSP00, CK00, Doc02, Hir00]. studies [Gla00, Hil00, Wor02]. Study [BNHGV04, Dan02, DL02, cFJH^{+}03, BhHSS00, BBM^{+}03, BB00a, Coo04, Har01, SKW^{+}07]. Studying [Har01]. stuff [HH00]. style [RS00b]. stylebook [Gol00c]. Styles [Vin03a]. sua [Soc00]. Sub [AS02]. Sub-region [AS02]. subject [Hil01, Kis00, TS00]. Subnet [Sut00a]. Subversion [Mas06]. subverting [HB06]. succeeding [Fer02]. Success [Mor00, WHA02, Wei01a]. successful [Eng00, Fra01b, RG05]. succinct [GB04]. summary [Ian01]. Summer [ACP01]. Summer-School [ACP01]. Sun [ER02, Pic03]. Super [Ano00g, Ano03b, Bar00b]. Super-Symmetric [Ano00g]. Supercomputer [Bar00b, Bar00c]. Supercomputing [ACM04]. SuperNets [Bir03]. superpower [BB00b]. superscalar [LRJ^{+}03]. supplement [LH00]. supply [Cho01b, KR01a, PB00]. support [BvS00, CTLL07, CHCC07, Hak00]. Supporting [DADAPW01, Kle02, FB04]. Supports [Ano03c]. Surfing [LWPM08, KM01c]. Surgeons [Bar00a, All01]. Survey [AS02, PB03]. surveys [Dil00]. survie [WJ00]. survival [Hus00, MVK01, MB02, WJ00]. surviving [Adv00b, Adv00a, Sch00b]. sustainable [FN03]. Swedish [Hoh01]. swindles [Col00]. Swiss [Bur00, Coc03a]. switch [KJF^{+}00]. switched [HC04]. Switches [Per00]. Switching [Har02, AJ02]. Switzerland [Sti02]. Symmetric [Ano00g]. Symposium [Ano00a, Ano01b, Ano01g, Ano03a, IEE01a, IEE01b, IEE01c, IEE01e, IEE02b, IEE02a, USE01, RS00a]. synchronization [Mil06, WZZ04]. Synergy [TT03]. Synthesis [CNI^{+}04]. Synthesizing [LD03]. System [DWX02, GIM^{+}02, LKMK02, OMM02, PKN^{+}04, Raj00, TKF^{+}02, Ano00e, BNHGV04, Bie01, EP06, Gao01, MS01a, YCY03, Bow00]. Systems [Abe03, ACK00, Ano00a, Ano00g, Ano00h, Ano01b, Ano03a, Ben00, Bla02c, CZ04, Cla03a, ES03, IEE01d, LS01a, LPK02, Leo02, ML02, Mit02a, PD03, Sch02a, Str03, Toh02, USE01, Woe02, BU01, CKK01, CdBvdM01, CP01, Ini07, Kha02, KS05, LT01b, McB02, vdMdB01, MM01, Ngu01, PD07, PC01, Ric01a, SGD^{+}02, Smi00b, Smi01c, Sta03b, UBM02, VE05, WAF00].

T1 [Ano00h, BW02]. Tables [Mit02b, BNHGV04]. TACT [Yu00a, Yu00b]. tailored [Dil00]. Talent [Bar01a]. tales [GT01, Llo01c]. tall [Llo01c]. Talmud [Ros00]. taming [Tob01]. Tao [Bej05]. Taormina [Pal01]. Targeting [CR01b]. task [Jon00a]. tax [Bor00, Kam01]. taxes [Imp00, Wis00]. TC6 [Juk02, RS00a]. tcl [DF00]. TCP [Ben00, Cal03, Com00b, CS01, CP01, DC01, Don03, Gra01b, HC04, Hel01, Ins03, Koz05, LA00a, LA00b, LBJ^{+}02, Los03b, Pan04, PC01, Rey03, SHK02]. TCP/IP [Ben00, Cal03, Com00b, CS01, CP01, DC01, Don03, Gra01b, Hel01, Ins03, Koz05, Los03b, Pan04, PC01, Rey03]. teach [Gla01, Lei00, PN01, SE00, SV01]. teacher [BMR01, DL01]. teachers [Ame00, And01, BSP00, Lev00, Pro02, Uni01]. Teaching [Cra00, Gor01, Kur02, GMP00, Hir00, HPS01, Jac00, Lai01, WSM00]. teams [All09]. Tech [LWPM08, Bru00, Man00a, Man01, Sha01]. Technical [KS00, RM02, Til01, Soo01]. Technique [SX03, Pau03]. Techniques [Ken02b, Kur02, Leo02, Nad03, Pic03, Wan08, Lav06, Gal02a, Swe01, TSC02]. Technologies [Hed02, Ano00a, Ano00g, Ano00h, Ano01b, Ano03a, CL02, IEE01a, IEE01b, Lea00, O’D00, Pan02, SL02, Sti02, Tir03, USE01, Ale00, BMP00, BMP01, CL00, Das01, Gli06, Ini07, Gon00, MC00, Omi01, Opp00, Opp03, RB01, Sut00b, WL04, JC01]. Technology [AS02, Bai01, Cis00, Gar00a, IEE01d, Kam00, Ker03, Ket02, Muc02, San02, VN03, Bur02, CFLA04, GM04, HM02b, Hut01, KB00, Gar00b, Lla00, Lou01, MWD00, Mil01b, Mou01, Nai02, Nua07, RHC05, Sam02, Sco01, S^{+}06, SL01, T^{+}07, Wal01b, YCY03]. tecnologia [Gar00b]. tecnologias [Gon00]. tecnologica [ZB01]. tecnologie [CSS00].

Telecom [vdB00, Gla00, New00b, New01b]. Telecommunications [Bla02a, Cri03, Gar00d, Pan02, Col02, Fri01, Lim06, New00b, New01b]. Telemachos [CSS00]. Telemaco [CSS00]. telephone [CL00, Hut01, CL00]. Telephony [Bla02b, Her00, Bla01, LMC01, Llo00, NLiMK04]. telepistemology [Gol00a]. telerobotics [Gol00a]. Teletraffic [MFdSeS01]. television [Alf05, LD00, CFLA04, Gra01a, Llo00, LD00, LD00, LD00]. tell [Wan01a, Wan06b]. Temporal [IAI02]. Ten [BG01b]. teoria [Gar00c]. Terabytes [IEE02c]. Term [CAN^{+}04, Har00b]. Terminal [Wil04]. terms [DCC00, KS00, Soo01]. terrorists [Sch00b]. tertulia [Car01a]. Test [ABC^{+}07, FS00, Ngu01]. Test-bed [ABC^{+}07]. Testing [Ngu01, All09, BH08, KLG02, MM07, Til05, WN06, WA06]. Text [DF00, AT01, Ree00]. textbook [Hak00]. texts [EHT02]. Theater [PKN^{+}04]. theft [Col00]. their [Pau03, PB00]. them [HLV05, Joh03, Pri00]. themes [Har00d]. theoretical [CSMW02]. theory [Fie01, LT01b, Lee01, Los03a, Los04, Opp09, RO00]. therapist [SB01]. There [Lig03, GG03]. Thesauri [Rie02]. they’re [Fra01b]. thieves [Sch00b]. Thin [Ano00h]. think [Man00a, Man01]. Thinking [Loc03, Gla01, GM03, Jac00, PN01, SE00, SV01]. Third [ACP01, Bla02c, ML02, Ada01, BCKT05, Koh02, Sti02]. Thorough [Sti03]. threat [BK00c, LWD07]. Three [Leo02]. Three-Volume [Leo02]. thriving [Lev01b]. Thunderbird [HKH05]. Thyrrhenian [Pal01]. Time [And03, CNI^{+}04, DWX02, KOS^{+}02, OMM02, Tay00, AKY01, Den02, Gre01a, Mil06, MM01, OC07, RZMD05, WWD04]. Time-Domain [And03]. TINI [Wil00]. Tips [McF05, Jon05, Swe01]. titres [Fra00]. TLS [Opp09]. to-do-it [Sau01a]. today [Fal00, Mod00]. tolerance [Jen01, Wey01].

Tolerant [DHL^{+}04]. Tomcat [Goo02b]. too [Dav02a]. Tool [SEGS03, Wes02, FPM05, HPS01, Sau01a]. Toolbox [Rit00]. Toolkit [SC05, HL02b, MM07, Ste01]. Tools [Ken02b, McF05, Nad03, Ano00c, CR01a, Cap00b, Swe01, IEE01d]. Top [KR03, KR05, KR01b, KR08]. Top-Down [KR03, KR05, KR01b, KR08]. topics [Lei00]. topologies [CGJ^{+}04]. Topology [BC02, VHE02, SFFF03]. TotalView [Ano00g]. touch [BhHSS00, Koh02]. Tourism [TD00]. Tournament [Coc03a]. tra [Gio00]. trace [Mai06]. Traceback [SX03]. Traceback-Based [SX03]. Tracer [RG02b]. Track [PHM03, San03, Max00]. Tracking [EKS02, FNRC05]. Trade [FKP02, Col00, FW00]. Trade-Offs [FKP02]. tradecraft [Syn00]. trademark [Sha00]. Trademarks [Dre01b]. tradeoffs [Yu00a, Yu00b]. Trading [BGE00, OKB02, Bak00b, CML00b, Ems00]. traditional [MR00]. Traffic [Ano05a, Bla00b, LKKS05, AJ02, BK00b, KJF^{+}00, LA02, Nal02, PHCMS05, RZMD05, STPM05, Swe01, Won00]. trainer [Ste01]. Transaction [CJL02]. Transcript [Bla02d]. transfer [Los00a]. transform [Bro02]. Transformation [ZR01, ZW00]. transformations [GR02]. transformed [And00a]. transition [CG01]. Translation [Lea00]. transmission [FM05, Ive00]. Transmitting [But02]. transparence [Thi00]. transparency [Thi00]. Transport [Bla02c, BW02, CIA^{+}03, CJL02, RBS02]. Trash [Bar01b]. travail [dB01]. travel [CKK01, Cur00]. Traveling [Bar01b]. treatise [Bal01]. Treaty [SF00]. Trends [AS02, Rob00, Pal01]. tricks [Jon05]. Trip [OMM02]. troubadors [BG01b]. troubadours [BG01b]. Troubleshooting [And03, Mac03, Ole03]. true [GM03]. Trust [Cam00, Cam01, HKLS03, San02]. truths [Moo02]. tunable [Yu00a, Yu00b]. tuner [WWD04]. tunnels [LBJ^{+}02]. Turing [Coc03a, Coc03b]. tussenstand [Fei01]. Tutorial [Sch02b, Smi02a, KC00]. TV [KS00, KS00, SL02]. Twenty [B^{+}02, Leh05]. Twenty-Eighth [B^{+}02]. Twenty-Five [Leh05]. Two [Rao02, Kat00, Kat01, Gar00b]. Two-Paths [Rao02]. TX [IEE01a, ACK00, IEE01b]. tyranny [CN00].

U.S. [Bar00a]. Ubiquitous [CP03]. UDP [HC04]. Ugly [McC00a]. UK [BMP00]. ultra [GPR^{+}05]. Umgang [DF00]. UMTS [Bra03, Cha05]. Unauthorised [All09]. underground [Kur00a, Mer01]. Understanding [And03, Cur00, Mes00, Sut00b, AGM07, Car01b, LT02]. understood [Gar00b]. unfair [LSC^{+}04]. Unicorn [OLLY02]. Unified [CZB^{+}01, PK02]. unionism [DF01]. United [New01a]. Uniti [Lug00]. University [FHM00, Rya00]. Unix [DF00, Raj00, GSS03, Sor01]. unlimited [Mar01]. Untangled [GMP00]. Unwanted [Ano05a]. unwired [LR02]. up-to-date [Smi00c]. updated [Gol00c]. upgrade [CHCC07]. Upgrades [Ano03c]. upheaval [VC00]. Uploads [Bar00b]. Upper [Lea00]. Upphovsrättsligt [Her01]. ups [AA01]. ura [Kur00a]. Uruguay [Car01a]. USA [ASP01a, ASP01b, AH02, ACK00, Ano05b, BS01a, BS02a, Ber00, CdBvdM01, GY01, IEE01a, IEE01d, IEE02c, JC01, KK00, Koh02, MS00b, vdMdB01, Mou01, Smi00b, Smi01c, USE01]. usability [JG08]. usage [MS00b, Sim02a, UJ02]. USB [Ano03c]. Use [BBP00, Gre02c, BG01b, B^{+}01, Har01, Haw01, JvEB^{+}04, Kay00, New01a, PJ00a, Ros03, SNT02, Smi00c, TG00]. Usefulness [AS02]. USENIX [Ano00a, Ano01b, Ano03a]. User [DL02, DF00, EKS02, Jan01, RG02b, BF01, Dud08, MS00b, Sta02b]. Users [HL02a, Mau03, Jac00]. Using [And03, BFS^{+}02, CNI^{+}04, DJK02, DMP00, Gla01, GcC02, Har00b, HW00, HL02a, HCP05, Lew01, Mau01, OMM02, OTPK02, PN01, Puz03a, Rao02, Rie02, Sau01a, STHK02b, STHK02a, SE00, SV01, AJ02, Bro02, CML00a, Cur00, Flo01, Har00c, Hay00, KKYK04, Knu01, Lan05b, LBJ^{+}02, Lon00, Mas06, Mat02, PB00, RK01a, RC05, SJ01, YCY03]. USITS [Ano00a, Ano01b, Ano03a, USE01]. utilisation [PJ00a]. Utility [LA02]. Utility-based [LA02]. utopia [Lan05b]. utopianism [Tur06].

V [Ano00h, DHL^{+}04, KZ01]. V-Systems [Ano00h]. v.5.7 [Ano00g]. VA [Ano05b]. Vain [Tir03]. Valuation [Ket02]. Value [All03, CKK01, KM00, Moo00, PB00, Til05]. vanuit [Fei01]. Varenna [ACP01]. variety [CML00a]. Vector [But02]. vendors [MV03]. verification [Qia00]. Verifying [CJM00]. verk [Her01]. Version [Ano00g, CS01, Mas06, Ste03, Ano00h]. Versionen [DF00]. versus [Bak00b]. VerteX [Ano00h]. Vertragen [Wey01]. Vertragsrecht [SF00]. Verwechselbarkeit [Bur00]. Very [B^{+}02, FLA^{+}03, STHK02b, STHK02a]. Veterinary [Ano00j]. via [ABGR01, Bro01, Col00, Fei01, Gio00, Pal02, EDGW00]. vid [Her01]. Video [Bov00, EPR03, LLZ03, Ale00, CFLA04, KD05, Llo00, Ale00]. Vienna [Juk02]. view [CK00]. Viewpoint [Cli00, Dav02a, O’R00]. viewpoints [Cot02]. Views [Bar00b, Bar00c, Bar00a, Bar01a, Bar01b, Coc03b, Coc03a]. Vigilante [CCC^{+}05]. village [Ber03]. Violation [KH02]. Virtual [CHCC07, Eve00, Kur02, Til01, And01, BhHSS00, Car01a, Gar00c, Gre00c, HL02b, MSCG01, Rya00]. Viruses [Lea00, Pau03]. visions [Juk02]. Visual [Ano00g, Mur00a, Kap00, Kur00b]. Visualization [Law04]. Visualizing [Law04]. Visually [AS02]. VLDB [FLA^{+}03]. VLDP [B^{+}02]. VM [Ano03c]. VMware [Ano03c]. vocabulary [ZW00]. VOCAL [DJK02]. VoD [SL02]. Voice [Che02, EPR03, HSW07, Hou01, Phi01, Puz03a, Räi01, CFLA04, Col02, MTK03, Ker03, Mou01]. Voice-Driven [Hou01]. Voice-Enabled [Che02]. Voice-Over-IP [Räi01]. Voice/Web [Phi01]. VoiceXML [Hou01, Phi01]. VoIP [Mou01, DJK02, Dwi09, EPR03, RR05a]. Vol [CS01]. Volatile [DHL^{+}04]. VolP [SJ01]. Volume [Com00b, CFHO02a, CFHO02b, Leo02, Kut01]. voluntary [OLLY02]. Volve [Lev01b]. vom [Hoh01]. voor [PG00]. vor [Six00]. vorbereiten [DF00]. Vortrage [Dre00]. Vote [MSM02]. Voting [Rub01, AN00, Ano01g, HC01, JRSW04, Jon00a, MG01, Wei00, Fre01]. Vs [CTBA^{+}01, BC01]. VSI [Ano00h]. VSI-FAX [Ano00h]. vulnerability [MM07].

W3 [AH01]. W3C [Bar00c]. wake [Hus04, LF00]. wake-up [Hus04, LF00]. WAN [Ano00h]. wang [Zhu00]. want [Alf05]. WAP [Sin01b]. warfare [Car09]. warning [Ger00]. Wars [RR03b]. Was [Lig03, And00a]. Washington [Ano05b]. Watch [MB00a, MB00b, Rob00]. wavelet [STPM05]. waves [Spa01]. Way [VM02, McC00b, Mur00b, VS01]. ways [Swe01, Won00]. WCDMA [OP01]. wealth [KM00, Spa01]. Weapons [RR03b, Sch00b]. Weather [Bar00b, Gab01d, Moo02]. Weaving [AF02]. Web [BMP01, Law04, MS00b, SS02, KT00, BMP01, Hed02, AKY01, And01, And00a, And00b, Ano00j, AV04, BMP00, Bay00, BFS^{+}02, Ben00, BMR01, BW00b, BS01b, BF01, Cas00b, CJ02, CZB^{+}01, DADAPW01, Dav01a, DDN00, DDN02, FHLW03, Flo01, GMP00, Har00c, HH03, HP02, HW00, Huh03, Hus04, HPS01, Jac00, JG08, JLHD01, Joh03, JWC08, Jol03, Kan02, Kro00, Lan05b, LH00, LV03, Man02, McF05, MA02, Mur00a, Mur00b, NS01, Ngu01, OTPK02, Opp00, Opp03, PKP03, Pas05, PN03, Phi01, PB03, Pri02, Pro02, RM00, SSS06, STHK02b, STHK02a, SL02, Smi00d, STK02, Spo01, Sto03, Swe01, TSC02, THB05, TWW03, Tir03, TKF^{+}02, Weh01, WA06, Won00, Yua01, Zel07]. Web-Based [TKF^{+}02]. Webcams [Sti03]. WEBKDD [Koh02]. WEBKDD’99 [MS00b]. Websights [Bod00]. Website [AF02, Kai06b]. Websites [Ano02, Kai06a]. Webstract [DRI00, Sta00d]. week [Adv00b, Adv00a]. welfare [Ano01a]. well [WCB01]. well-conditioned [WCB01]. wellness [Ano00b]. West [Fra01c]. Western [CZB^{+}01]. Whatever [Wei01b]. Where [Gre03]. Which [Gre02c]. white [Ano00g]. Who [Bes00, CZB^{+}01]. Whole [Tur06]. Wi [Ros03, VGM04, Ros08, Bar03]. Wi-Fi [Ros03, Ros08, Bar03]. Wi-Foo [VGM04]. WIAPP [IEE01f]. Wide [BMR01, BW00b, Cas00b, Dav01a, DDN00, DDN02, FHLW03, NS01, Opp00, Opp03, Pro02, RM00, BvS00, CRW01]. wide-area [BvS00, CRW01]. wild [Ems00, Goe02]. Wiley [Sha01]. Will [DF01, Amo01, Fra01b, Man00a, Man01, Wei01a]. William [ET03]. Wily [CBR03, ET03]. WiMAX [AGM07, Nua07]. win [Mod00]. Window [Bow00, JF00]. Windows [DF00, BD05, BM02, HB06, Kro00, LL01, Nor00a, Tre01, Wil02b, Wil04]. wing [FK01]. winner [Alf05]. Winners [Bar01a]. winning [LJ01]. wire [Zal05]. wired [Jef00, BMM07, BCKT05]. Wired/wireless [BMM07, BCKT05]. Wireless [Ahm05, Bar03, Bea02, Bri08, Bur01, CZ04, CJL02, Coo04, Dei01a, Ear06, EPR03, FW06, Gil04, GIM^{+}02, Gra05, KCF02, Kip07, Lim06, Lab07, LS01a, LKP02, Mau03, Ole03, PK02, PL05, PKP03, Puz03b, Rap02, SP02, STHK02a, SHK02, Sha01, T^{+}07, Toh02, Vac01, VGM04, Wei02, ZG04, AGM07, AM07, Bou06, Cha05, Cra07, EP06, GPR^{+}05, Gil05, Gli06, GG05, GCB04, HSW07, IEE01a, Ini07, KM01b, Lla00, Nua07, OC07, PK00, Rho02, Ros03, Ros08, Sch05, S^{+}06, SC07, Sto02, TV05, Vac06, Var03, VN03, KS05, Sin01b, BMM07, BCKT05]. wiretapping [Lan05a]. wiring [GM00b]. Wirtschaftsrecht [ILS00]. without [BS02c, GR01]. Women [GG03, LH00, WC00]. won [And00a]. won’t [Wan01a, Wan06b]. woolly [Ems00]. WordMage [Ano00g]. words [Ano01d]. work [JG08, RS00b, UJ02, WB02]. Workbook [DADAPW01]. worker [MP01]. Working [Juk02, KB00, dB01]. workloads [CML00a]. workplace [Ana02, Ora02]. works [Com00a, Gra02]. Workshop [ACK00, Ano05a, BU01, Fre01, IEE01f, KK00, Lev01a, Pal01, Smi01a, Sti02, Adv00b, Cha02, CRS00, Ian01, KZ01, Koh02, KS05, MS00b, UBM02]. Workshops [IEE02b, Adv00a]. Workstations [cFJH^{+}03]. World [Bar00c, Lim06, Puz03b, Sch02a, Sur00, Was00, WL04, Amo01, BC01, BL00b, BV01, Ems00, Fra02, Gre00c, Har00d, Kur00a, LR02, Llo01b, OC07, PN01, PP00, PP01b, SV01, Spo01, Van01a, VS01, War01, BMR01, BW00b, Cas00b, Dav01a, DDN00, DDN02, FHLW03, NS01, Opp00, Opp03, Pro02, RM00, Spo01]. World-Class [Sch02a]. worlds [Ros00]. worldwide [Bea02, Nor01b]. Worm [BCK06]. worms [CCC^{+}05]. worse [Man00a, Man01]. Worterbucher [LSS01]. Worth [Mau01]. Wrap [Sta00b]. Wrap-up [Sta00b]. writers [McG00, TT00]. Writing [HL03, Pol00, Har00b, Sin01b, TT00, Weh01]. WWIC [BMM07, BCKT05]. Wyndham [Lev01a].

X [Bow00, Dud08]. Xanthi [BCKT05]. XFree86 [DF00]. XFree86/23.3.6 [DF00]. XHTML [Nie06]. XLink [Bar00c]. XML [Law04, LSS01, ASY02, BFS^{+}02, Bir01, Dei01b, GR02, Kro00, LSS01, SS02, Sim02b, SM02b]. XML-based [Law04, Kro00]. XPath [GR02, Sim02b]. XPointer [Sim02b]. Xpresso [Ano00h]. XSLT [GR02, Ten05].

Yahoo [VS01]. year [Soc01]. Years [Leh05, Nau00]. Yellow [OTPK02]. yogo [Kur00a]. York [Adv00b, Adv00a]. young [Bra01]. youth [Uni01].

Z [KS00, Ano01d, KS00]. Zen [HH03]. zero [CTLL07, CHCC07]. zero-loss [CTLL07, CHCC07]. Zhongguo [Zhu00]. Zitat [RS00b]. Zitieren [RS00b]. zivilrechtliche [Red00]. zum [RS00b]. zur [Dre00]. Zurich [Sti02]. zwischen [Dre00].

References

Aspatore:2001:DRN

[AA01] Jonathan Reed Aspatore and Alicia Abell. Digital rush: nine Internet start-ups in the race for dot-com riches. AMA-COM, New York, NY, USA, 2001. ISBN 0-8144-0567-3. x + 290 pp. LCCN HF5548.32 .A836 2001.

Abelson:2000:EIA

[AAC^{+}00] Harold Abelson, Don Allen, Daniel Coore, Chris Hanson, George Homsy, Thomas F. Knight, Radhika Nagpal, Erik Rauch, Gerald Jay Sussman, and Ron Weiss. Embedding the Internet: amorphous computing. Communications of the ACM, 43(5):74, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                                     http://www.acm.org/pubs/

citations/journals/cacm/
2000-43-5/p74-abelson/.

Aiello:2002:ERS

[ABB^{+}02] William Aiello, Steven M.

Bellovin, Matt Blaze, Ran Canetti, John Ioannidis, Angelos D. Keromytis, and Omer Reingold. Efficient, DoS-resistant, secure key exchange for Internet protocols. Lecture Notes in Computer Science, 2467:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2467/24670027.htm;
http://link.springer.de/
link/service/series/0558/
papers/2467/24670027.pdf.

Ancillotti:2007:LFI

[ABC^{+}07] E. Ancillotti, R. Bruno, M. Conti, E. Gregori, and A. Pinizzotto. A layer-2 framework for interconnecting ad hoc networks to fixed Internet: Test-bed implementation and experimental evaluation. The Computer Journal, 50(4):478–499, July 2007. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
50/4/478;   http://comjnl.
oxfordjournals.org/cgi/content/
full/50/4/478;   http://
comjnl.oxfordjournals.org/
cgi/reprint/50/4/478.

Abeloff:2001:PGB

[Abe01] Martin D. Abeloff. A patient guide to breast cancer resources on the Internet. eMedguides.com, Princeton, NJ, USA, 2001. ISBN 0-9700525-7-X. 61 pp. LCCN WP 39 P298 2001.

Aberer:2003:RFI

[Abe03]  Karl Aberer.  Report on

the first international conference on ontologies, databases and applications of semantics (ODBASE): part of the federated conference “On the Move to Meaningful Internet Systems 2002”. SIGMOD, 32(1): 69–72, March 2003.

Aldini:2001:CQI

[ABGR01] Alessandro Aldini, Marco Bernardo, Roberto Gorrieri, and Marco Roccetti. Comparing the QoS of Internet audio mechanisms via formal methods. ACM Transactions on Modeling and Computer Simulation, 11(1):1–42, January 2001. CODEN ATMCEZ. ISSN 1049-3301.

Abidi:2002:DAH

[Abi02] Syed Sibte Raza Abidi. Designing adaptive hypermedia for Internet portals: A personalization strategy featuring case base reasoning with compositional adaptation. Lecture Notes in Computer Science, 2527:60–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2527/25270060.htm;
http://link.springer.de/
link/service/series/0558/
papers/2527/25270060.pdf.

Appleby:2000:AJB

[ABM00] Joyce Oldham Appleby, Alan Brinkley, and James M. McPherson. The American journey: building a nation. Glencoe/McGraw-Hill, New York, NY, USA, teacher’s wraparound ed edition, 2000. ISBN 0-02-821877-9. ???? pp. LCCN H311 .G50.32.

Aboelela:2003:NSE

[Abo03]  Emad Aboelela.  Network

Simulation Experiments Manual. Academic Press, New York, NY, USA, fourth edition, 2003. ISBN 0-12-042171-2. ???? pp. LCCN ???? US$19.95.

Abualsamid:2001:PHA

[Abu01] Ahmad Abualsamid. PHP and hosted applications. Dr. Dobbs Journal, 26(1):56, 58, 60–63, January 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/2001/2001_01/php.
txt.

Ambler:2000:SMI

[ACK00] Anthony Ambler, Seraphin B.

Calo, and Gautam Kar, editors. Services management in intelligent networks:

11th IFIP/IEEE International Workshop on Distributed Systems: Operations and Management, DSOM 2000, Austin, TX, USA, December 4–6, 2000: proceedings, volume 1960 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-41427-4 (softcover). LCCN TK5105.5 .I334 2000. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1960.htm.

ACM:2003:SII

[ACM03] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

ACM:2004:SHP

[ACM04] ACM, editor. SC 2004: High Performance Computing, Networking and Storage: Bridging communities: Proceedings of the IEEE/ACM Supercomputing 2004 Conference, Pittsburgh, PA, November 6– 12, 2004. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2153-3. LCCN ????

Agosti:2001:LIR

[ACP01] Maristella Agosti, Fabio Crestani, and Gabriella Pasi, editors. Lectures on information retrieval: Third European Summer-School, ESSIR 2000 Varenna, Italy, September 11– 15, 2000, revised lectures, volume 1980 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-41933-0. LCCN QA267.A1 L43 no.1980.

Achleitner:2001:LAI

[AD01] Herbert K. Achleitner and Aleksandur Georgiev Dimchev, editors. Libraries in the age of the Internet: papers from the International Conference held in Sofia, Bulgaria 8-10 November 2000. Union of Librarians and Information Services Officers (ULISO) Bulgaria, Sofia, Bulgaria, 2001. ISBN 0-934068-13-5. LCCN Z674.75.I58 L53 2001.

Adams:2000:IGC

[Ada00] Jean Murrell Adams, editor.

The Internet guide for California lawyers 2001. Continuing Education of the Bar, Berkeley, CA, USA, second edition, 2000. ISBN 0-7626-0476-X. xvii + 348 pp. LCCN KFC74 .I57 2000; KFC74.5 .I58 2000.

Adams:2001:IGC

[Ada01] Jean Murrwell Adams. The Internet guide for California lawyers third edition. Continuing Education of the Bar, Berkeley, CA, USA, third edition, 2001. xvii + 369 pp. LCCN KFC74 .I57 2001. “MI-33212”–T.p. verso. Disk contains all of the URLs listed in the Appendix in HTML format. Treatise Collection copy 2 lacks computer disk. C.

ARF:2000:ESI

[Adv00a] Advertising Research Foundation, editor. The experts on surviving the Internet: an ARF week of workshops event, October 18, 2000, New York City. Advertising Research Foundation, New York, NY, USA, 2000. LCCN HF5415.2 .E987 2000.

ARF:2000:ERS

[Adv00b] Advertising Research Foundation, editor. The experts report on surviving the Internet: an ARF week of workshop event, October 17, 2000, New York City. Advertising Research Foundation, New York, NY, USA, 2000. LCCN HF5415.2 .E985 2000.

Anderson-Freed:2002:WWP

[AF02] Susan Anderson-Freed. Weaving a Website: programming in HTML, JavaScript, Perl and Java. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-028220-0. 712 pp. LCCN QA76.625 .A64 2002. BUY-A-BOOK gift of Michael E. Hayden thanks to Professor Frank Deremer.

Andrews:2007:FWU

[AGM07] Jeffrey G. Andrews, Arunabha Ghosh, and Rias Muhamed. Fundamentals of WiMAX: understanding broadband wireless networking. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2007. ISBN 0-13-222552-2 (hardcover). 449 (est.) pp. LCCN TK5103.2 .A56 2007. URL http:/

/www.loc.gov/catdir/toc/
ecip075/2006038505.html.

Agre:2003:PPI

[Agr03] Philip E. Agre. P2P and the promise of Internet equality. Communications of the ACM, 46(2):39–42, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ackermann:2001:ISG

[AH01] Ernest C. Ackermann and Karen Hartman. The information searcher’s guide to searching + researching on the Internet + W3. ABF Content, Wilsonville, OR, USA, second edition, 2001. ISBN 1-887902-58-9. x + 438 pp. LCCN ZA4201 .A25 2001. Rev. ed. of: Information specialist’s guide to searching and researching on the Internet and World Wide Web. 1999.

Almeroth:2002:MMI

[AH02] Kevin C. Almeroth and Masum Hasan, editors. Management of multimedia on the Internet: 5th IFIP/IEEE International Conference on Management of Multimedia Networks and Services, MMNS 2002, Santa Barbara, CA, USA, October 2002: proceedings, volume 2496 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-44271-5 (softcover). LCCN TK5103.75 .I58 2002.

Ahmad:2002:SAI

[Ahm02] Khalid Ahmad. Sourcebook of ATM and IP internetworking. IEEE Press series on network management. Wiley-Interscience, New York, NY, USA, 2002. ISBN 0-471-20815-9. xiii + 470 pp. LCCN TK5105.35 .A36 2002. US$99.95.

Ahmad:2005:WMD

[Ahm05] Aftab Ahmad. Wireless and mobile data networks. Wiley-Interscience, New York, NY, USA, 2005. ISBN 0-471-67075-8 (hardcover). xxix + 346 pp. LCCN TK5103.2 .A43

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0617/2004025911-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0617/2004025911-
d.html;   http://www.loc.
gov/catdir/toc/ecip053/2004025911.
html.

Awduche:2002:ITE

[AJ02] Daniel O. Awduche and Bijan Jabbari. Internet traffic engineering using multi-protocol label switching (MPLS). Computer Networks (Amsterdam, Netherlands: 1999), 40(1):111–129, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/97/27/34/
abstract.html.

Allen:2001:OOW

[AKY01] Cliff Allen, Deborah Kania, and Beth Yaeckel. One-to-one Web marketing: build a relationship marketing strategy one customer at a time. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 2001. ISBN 0-471-40400-4 (paperback). xv + 351 pp. LCCN HF5415.1265 .A419 2001. “Wiley computer publishing”. Previously published as: Internet world guide to one-to-one Web marketing. New York: Wiley Computer Pub., c1998.

Albitz:2001:DB

[AL01] Paul Albitz and Cricket Liu.

DNS and BIND. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2001. ISBN 0-596-00158-4. xviii + 601 pp. LCCN TK5105.875.I57 A44 2001. Covers BIND 9.

Alesso:2000:VPI

[Ale00]  H. P. Alesso.  e-Video:

producing Internet video as broadband technologies converge. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-70314-9 (paperback). xiv + 289 pp. LCCN TK5105.887 .A43 2000. CD-ROM contains code and video demos of all the examples in the book.

Alfonsi:2005:WMI

[Alf05]  B. Alfonsi.  I want my

IPTV: Internet Protocol television predicted a winner. IEEE Distributed Systems Online, 6(2):??, February 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

                                        ieeexplore.ieee.org/iel5/
                                        8968/30522/01407761.pdf?
                                        isnumber=30522&prod=JNL&
                                        arnumber=1407761&arSt=&ared=
                                        &arAuthor=Alfonsi%2C+B.;
                                         http://ieeexplore.ieee.
                                        org/xpls/abs_all.jsp?isnumber=
                                        30522&arnumber=1407761&count=
                                        5&index=2.

Allen:2001:IS

[All01] Jeff W. (Jeff Wayne) Allen. The Internet for surgeons. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95319-1 (softcover). xvi + 152 pp. LCCN W 26.5 A427i 2001.

Allee:2003:FKI

[All03]  Verna Allee.  The Fu-

ture of Knowledge: Increasing Prosperity through Value Networks. Butterworth-Heinemann, Boston, MA, USA, 2003. ISBN 0-7506-7591-8. xx + 294 pp. LCCN HD58.82 .A367 2003. US$24.99.

Allsopp:2009:UAP

[All09] Wil Allsopp. Unauthorised access: physical penetration testing for IT security teams. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2009. ISBN 0-470-74761-7. xx + 287 pp. LCCN QA76.9.A25 A456

2009. URL http://www.loc.
gov/catdir/enhancements/
fy1013/2010487083-d.html;
 http://www.loc.gov/catdir/
enhancements/fy1013/2010487083-
t.html.

Anjum:2007:SWA

[AM07] Farooq Anjum and Petros Mouchtaris. Security for wireless ad-hoc networks. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2007. ISBN 0-471-75688-1. ???? pp. LCCN TK5105.59 .A54

2007. URL http://www.loc.
gov/catdir/toc/ecip0620/
2006029342.html.

Ameis:2000:MIR

[Ame00] Jerry A. Ameis. Mathematics on the Internet: a resource for K–12 teachers. Merrill, Upper Saddle River, NJ, USA, 2000. ISBN 0-13-011061-2. 129 pp. LCCN QA41.6 .A64 2000.

Alder:2005:SNH

[AML05] Raven Alder, Timothy M. Mullen, and Johnny Long, editors. Stealing the network: how to own an identity. Syngress Publishing, Inc., Rockland, MA, USA, 2005. ISBN 1-59749-006-7 (paperback). xxiv + 333 pp. LCCN PS648.C65 S742 2005.

Amor:2001:IFS

[Amo01] Daniel Amor. Internet future strategies: how pervasive computing services will change the world. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-041803-X. xxii + 294 pp. LCCN HF5548.32 .A463 2002.

Alvarez:2000:LCI

[AN00] R. Michael Alvarez and Jonathan Nagler. The likely consequences of Internet voting for political representation. Technical report, Loyola Law School, Los Angeles, CA, USA, September 19, 2000. 37 pp.

                                        URL http://www.lls.edu/
                                        internetvoting/ivote3c.pdf.

Anandarajan:2002:IAW

[Ana02] Murugan Anandarajan. Internet abuse in the workplace: Introduction. Communications of the ACM, 45(1):53– 54, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Andrews:2000:HWW

[And00a] Paul Andrews. How the Web was won: how Bill Gates and his Internet idealists transformed the Microsoft Empire. Broadway Books, New York, NY, USA, 2000. ISBN 0-7679-0049-9 (paperback). 352 pp. LCCN HD9696.63.U64 M532 2000. US$15.95.

Andriot:2000:IBP

[And00b] Laurie Andriot. Internet blue pages: the guide to federal government Web sites. CyberAge Books, Medford, NJ, USA, 2001-2002 ed edition, 2000. ISBN 0-910965-43-9. xi + 447 pp. LCCN ZA5075 .A53 2000. URL http://www.

fedweb.com.

Anderson:2001:IWD

[And01] Dennis Anderson. The Internet and Web design for teachers: a step-by-step guide to creating a virtual classroom. Longman, New York, NY, USA, 2001. ISBN 0-321-07751-2. x + 112 pp. LCCN LB1044.87 .A53 2001.

Anderson:2003:TOF

[And03] Duwayne Anderson. Troubleshooting Optical Fiber Networks: Understanding and Using Optical Time-Domain Reflectometers. Academic Press, New York, NY, USA, 2003. ISBN 0-12-058661-4. ???? pp. LCCN ???? US$79.95.

Andress:2005:INI

[And05] Jason G. Andress. IPv6: The next Internet protocol. ;login: the USENIX Association newsletter, 30 (2):??, April 2005. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/
login/2005-04/pdfs/andress0504.
pdf.

Anonymous:2000:AUS

[Ano00a] Anonymous. Announcement: 3rd USENIX Symposium on Internet Technologies and Systems (USITS ’01). ;login: the USENIX Association newsletter, 25(3):??, June 2000. CODEN LOGNEM. ISSN 1044-6397. URL http:

//www.usenix.org/events/
usits01.

Anonymous:2000:HNH

[Ano00b] Anonymous. Health net: a health and wellness guide to the Internet. McGraw-Hill, New York, NY, USA, third edition, 2000. ISBN 0-07-234591-8. 55 pp. LCCN RA773.6 .H44 2000.

Anonymous:2000:HMD

[Ano00c]  Anonymous.  HyperNiet-

zsche: Modèle d’un hypertexte savant sur Internet pour le recherche en sciences humaines. Questions philosophiques, problèmes juridiques, outils informatiques. (French) [HyperNietzsche: Model of a hypertext savant on the Internet for research in social sciences. Philosophical questions, legal problems, computer tools]. Écritures electroniques. Presses universitaires de France, Paris, France, 2000. ISBN 2-13-051156-2. 200 pp. LCCN B3315.A12 H96 2000.

Anonymous:2000:IAB

[Ano00d] Anonymous. The Internet archive: building an “Internet Library”. Internet Archive, San Francisco, CA, USA, 2000. LCCN ZA4201.

Anonymous:2000:IBI

[Ano00e] Anonymous. Internet business intelligence: how to build a big company system on a small company budget. CyberAge Books, Medford, NJ, USA, 2000. ISBN 0-910965-35-8. xx + 438 pp. LCCN HD38.7 .V56 2000.

Anonymous:2000:IMD

[Ano00f] Anonymous. Internet: las mejores direcciones en la Red. (Spanish) [Internet: The best directions in the net]. Endesa, Madrid, Spain, 2000. 207 pp. LCCN TK5105.875.I57 I54 2000.

Anonymous:2000:NPI

[Ano00g] Anonymous. New products:

IQ2 NetCONNECT, BAS-COM Global Internet Services, Inc.; OpenLinux eServer 2.3, Caldera Systems, Inc.; CommuniGate Pro, Super-Symmetric Dynamic Cluster Version, Stalker Software, Inc.; Embedix Linux 1.0, Lineo, Inc.; white dwarf linux 1.0, EMJ Embedded Systems; Linux-based Internal G.lite Modem, SAS; IntraLinux, STS International, Inc.; Linux End-to-End Security Solution, JAWS Technologies; JetForm Central for Linux, JetForm Corporation; JWAVE version 3.0, Visual Numerics, Inc.; RTEL, Lantronix; Japanese WordMage v.5.7, Lava Software Pty. Ltd.; TotalView 4.0 Parallel Debugger, Etnus; iNUX Consumer Linux Personal Computer systems, iNUX Inc. Linux Journal, 73:??, May 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:NPQ

[Ano00h]  Anonymous.  New prod-

ucts: Quick-Connect, Bascom Global Internet Services, Inc.; CoolKeyboards, CoolKeyboards Corporation; ESP Print Pro, Easy Software Products; EnlightenDSM version 3.4 Enlighten Software Solutions, Inc.; RTSP/RTP Server, Entera, Inc.; EZTerm 1.3 for Linux, CSV Technologies, Inc.; OpenMerchant, OpenSales, Inc.; High-Density T1, E1 WAN Cards and ImageStream Routers, ImageStream Internet Solutions; NetMax Thin Server Series, Cybernet Systems Corporation; OpenMail 6.0 for Linux, Hewlett–Packard; VerteX 400-2U Server, EIS Computers, Inc.; VSI-FAX 4.0, V-Systems, Inc.; Xpresso Linux 2000, Xpresso Ltd.; iMigo, Lineo, Inc.; Storm Linux Beta, Stromix Technologies Inc. Linux Journal, 69:??, January 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:SAH

[Ano00i] Anonymous. The school administrator’s handbook of essential Internet sites. Aspen Publishers, Gaithersburg, MD, USA, fourth edition, 2000. ISBN 0-8342-1809-7. ix + 224 pp. LCCN LB1044.87 .S34 2000.

Anonymous:2000:VWY

[Ano00j] Anonymous. Veterinary Web:

your guide to the Internet. Veterinary Learning Systems, Overland Park, KS, USA, 2000. 30 pp. LCCN SF610.6 V48 2000. Supplement to Compendium, Veterinary forum, Veterinary technician.

Anonymous:2001:CWR

[Ano01a]  Anonymous.  Child wel-

fare review: the Internet journal on child welfare. Child Welfare Research Institute, School of Public Policy and Social Research, UCLA, Los Angeles, CA, USA, 2001. LCCN HV741. URL

http://www.childwelfare.
com/kids/news.htm. Con-

tains links to articles related to child welfare on the Internet, and also original articles. This journal of edited by Prof. Duncan Lindsay, UCLA Social Welfare Dept.

Anonymous:2001:CRUa

[Ano01b] Anonymous. Conference report: The 3rd USENIX Symposium on Internet Technologies and Systems (USITS ’01). ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:DEI

[Ano01c] Anonymous. Derecho y estado en Internet: guia basica de direcciones. Civitas, White Plains, NY, USA, 2001. ISBN 84-205-3340-8. 151 pp. LCCN KE3649 .D47 2001.

Anonymous:2001:DCI

[Ano01d] Anonymous. Dictionary of computer and Internet words: [an A to Z guide to hardware, software, and cyberspace]. Houghton Mifflin, Boston, MA, USA, 2001. ISBN 0-618-10137-3 (paperback). vi + 298 pp. LCCN QA76.15 .D5255 2001.

Anonymous:2001:IBL

[Ano01e] Anonymous. Internet bigwigs: leading CEO’s and research analysts forecast the future of the Internet economy. Inside the minds. Aspatore Books, Bedford, MA, USA, 2001. ISBN 1-58762-010-3 (paperback). 206 pp.

LCCN HF5548.32 .I59 2001. US$27.95.

Anonymous:2001:L

[Ano01f] Anonymous. L’Internet. Loi au quotidien. Direction des journaux officiels. (French) [Direction of the Official Journals], Paris, France, 2001. ISBN 2-11-074971-7. viii + 137 pp. LCCN ???? SS 12806277.

Anonymous:2001:SIV

[Ano01g] Anonymous. Symposium: Internet voting and democracy. Loyola of Los Angeles Law Review, 34(3):979–1295, April 2001.

Anonymous:2002:IPC

[Ano02] Anonymous. Internet pornography: constitutionality of restricting access to adult Websites. Number 5(2) in Supreme Court debates. Congressional Digest Corp., Washington, DC, USA, February 2002. 33– 64 pp. LCCN R.R.

Anonymous:2003:CRUa

[Ano03a] Anonymous. Conference report: 4th USENIX Symposium on Internet Technologies and Systems (USITS ’03). ;login: the USENIX Association newsletter, 28 (3):??, June 2003. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/
login/2003-06/openpdfs/USITSconf.
pdf.

Anonymous:2003:NSI

[Ano03b] Anonymous. Networking and Sharing Information Super Series. Butterworth-Heinemann, Boston, MA, USA, fourth edition, 2003. ISBN 0-7506-5885-1. ???? pp. LCCN ???? US$34.99.

Anonymous:2003:PJU

[Ano03c] Anonymous. Products: Jet-Brains upgrades IntelliJ Java IDE; Catalyst’s USB analyzer supports device emulation; VMware releases Enterprise Server VM software; Motorola offers free soft modem reference design; Real-Networks releases source for Helix DNA Server; Packeteer accelerates intranet and Internet applications. Computer, 36(3):80–81, March 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/03/r3080.
htm;   http://csdl.computer.
org/dl/mags/co/2003/03/r3080.
pdf.

Anonymous:2005:CRSc

[Ano05a] Anonymous. Conference report: Steps to Reducing Unwanted Traffic on the Internet Workshop (SRUTI ’05). ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/

login/2005-10/openpdfs/sruti05.
pdf.

Anonymous:2005:FIN

[Ano05b] Anonymous, editor. Future Internet Network Design Meeting, December 5, 2005, Hilton Washington Dulles Airport, Herndon, VA, USA. ????, ????, 2005.

Asseraf-Olivier:2000:DDM

[AOB00] Frederique Asseraf-Olivier and Eric Barbry. Le droit du multimedia: du CD-Rom à l’internet. (French) [Multimedia law: from CD-ROM to the Internet]. Que saisje? Presses Universitaires de France, Paris, France, deuxième edition, 2000. ISBN 2-13-050630-5. 127 pp. LCCN KJV333.C65 A87 2000.

Alvoni:2001:ADD

[AR01] Giovanna Alvoni and Ulrich Rausch. Altertumswissenschaften digital: Datenbanken, Internet und e-Ressourcen in der altertumswissenschaftlichen Forschung. (German) [Antiquity sciences digitally: Data bases, Internet and e-resources in antiquity science research], volume 5 of Studienbucher Antike; ISSN: 1436-3526; Bd. 5. Olms, New York, NY, USA, 2001. ISBN 3-487-11248-5. ISSN 1436-3526. 191 pp. LCCN ????

Armitage:2000:QSI

[Arm00] Grenville Armitage. Quality of service in IP networks: foundations for a multi-service Internet. Technology series. MTR, Indianapolis, IN, USA, 2000. ISBN 1-57870-189-9. viii + 309 pp. LCCN TK5105.55 .A76 2000.

Arnaldo:2001:CAI

[Arn01] Carlos A. Arnaldo. Child abuse on the Internet: ending the silence. UNESCO Pub., Paris, France, 2001. ISBN 1-57181-245-8, 1-57181-246-6, 92-3-103728-5. xix + 220 pp. LCCN HV6570 .C545 2001.

Al-Salman:2002:NTD

[AS02] AbdulMalik S. Al-Salman. New trends and developments in computer and Internet technology for visually impaired people: a survey and usefulness in the Arab subregion. Lecture Notes in Computer Science, 2398:437– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2398/23980437.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2398/23980437.
pdf.

Ashbrook:2000:LML

[Ash00] Tom Ashbrook. The leap: a memoir of love and madness in the Internet gold rush. Houghton Mifflin, Boston, MA, USA, 2000. ISBN 0-395-83934-3. 295 pp. LCCN HD9696.8.U62 A84 2000.

Al-Shaer:2001:MMI

[ASP01a] Ehab S. Al-Shaer and Giovanni Pacifici, editors. Management of multimedia on the Internet: 4th IFIP/ IEEE International Conference on Management of Multimedia Networks and Services, MMNS 2001, Chicago, IL, USA, October 29–November 1, 2001: proceedings, volume 2216 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42786-4 (softcover). LCCN QA267.A1 L43 no.2216.

Al-Shaer:2001:MMN

[ASP01b] Ehab S. Al-Shaer and Giovanni Pacifici, editors. Management of multimedia on the Internet: 4th IFIP/ IEEE International Conference on Management of Multimedia Networks and Services, MMNS 2001, Chicago, IL, USA, October 29– November 1, 2001: proceedings, volume 2216 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42786-4 (softcover). LCCN QA267.A1 L43 no.2216; TK5103.75 .I58

2001. URL http://link.
springer-ny.com/link/service/
series/0558/tocs/t2216.htm.

Asakura:2002:XK

[ASY02] Hiroshi Asakura, Takeyuki Shimura, and Yohei Yamamoto. XML kyokasho. Sofuto Risachi Senta, Tokyo, 2002. ISBN 4-88373-167-7. 457 pp. LCCN ????

Afuah:2001:IBM

[AT01] Allan Afuah and Christopher L. Tucci. Internet business models and strategies: text and cases. Irwin/McGraw Hill, Boston, MA, USA, 2001. ISBN 0-07-239724-1. xxii + 358 pp. LCCN HD30.37 .A33 2001.

Antoniou:2004:SWP

[AV04] G. (Grigoris) Antoniou and Frank Van Harmelen. A Semantic Web primer. Cooperative information systems. MIT Press, Cambridge, MA, USA, 2004. ISBN 0-262-01210-3. xx + 238 pp. LCCN TK5105.88815 . A58 2004.

Akdeniz:2000:ILS

[AWW00] Yaman Akdeniz, David Wall, and Clive Walker. The Internet, law and society. Longman, New York, NY, USA, 2000. ISBN 0-582-35656-3. xx + 388 pp. LCCN K564.C6 A54 2000.

Brewer:2001:FFG

[B^{+}01] Ernest W. Brewer et al. Finding funding: grantwriting from start to finish, including project management and Internet use. Corwin Press, Thousand Oaks, CA, USA, fourth edition, 2001. ISBN 0-7619-7798-8 (paperback), 0-7619-7797-X. viii + 383 pp. LCCN LB2825 .F522 2001. Rev. ed. of: Finding funding / Ernest W. Brewer. 3rd ed. c1998. Contents: I. Exploring in the Grants World.

1. Unraveling the Mystique in Grant Applications. 2. How to Explore Grant Possibilities in Education. 3. Using the Federal Register (FR) and the Code of Federal Regulations (CFR). 4. Using the Catalog of Federal Domestic Assistance (CFDA). 5. Using the Internet to Access Funding Resources — II. Writing Grant Proposals. 6. What Are the Components of a Proposal? 7. Components of a Foundation Proposal. 8. Writing a Proposal. 9. Helpful Hints From Grantwriting Professionals. 10. Reviewing a Funded Proposal. 11. Understanding How Grants Are Awarded. 12. The Players and Their Roles in the Process — III. Implementing, Operating, and Terminating a Project. 13. Implementing a Funded Project. 14. Understanding EDGAR and GPRA.

15. Understanding Site Visits and Audits. 16. Closing Out a Project.

Bernstein:2002:VPT

[B^{+}02] Philip A. Bernstein et al., editors. VLDP 2002: proceedings of the Twenty-Eighth International Conference on Very Large Data Bases, Hong Kong SAR, China, 20–23 August 2002. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-869-9. LCCN ????

Bailey:2001:NTD

[Bai01] Andy Bailey. Network Technology for Digital Audio. Focal Press, Boston, MA, USA and Oxford, UK, 2001. ISBN 0-240-51588-9. xiii + 275 pp. LCCN TK5103.7 .B355 2001. US$49.99.

Baker:2000:IPO

[Bak00a] Mark Baker. Internet programming with OmniMark. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-7237-9. xviii + 391 pp. LCCN QA76.625 .I58 2000.

Bakos:2000:ECR

[Bak00b] Yannis Bakos. Electronic commerce in the retail brokerage industry: trading costs of Internet versus full service firms. Technical report, New York University Salomon Center, New York, NY, USA, 2000. 30 + 16 pp. November 1999.

Ballon:2001:CIL

[Bal01] Ian C. Ballon. E-commerce and Internet law: treatise with forms. Glasser Legal-Works, Little Falls, NJ, USA, 2001. ISBN 1-888075-94-5 (v.1), 1-888075-74-0 (set). various pp. LCCN KF390.5.C6 B35 2001. URL http:

//ballononecommerce.com.

Three volumes.

Baran:2000:NVf

[Bar00a] Nicholas Baran. News and views: Freenet: More anarchy for the Internet?; magnetic properties key to nanoengineering; nanoseconds not fast enough? here come femtoseconds; Caltech leads U.S. field in ACM programming contest; robotic surgeons may make fewer mistakes; free software for designing ICs. Dr. Dobbs Journal, 25(6):18, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVIa

[Bar00b] Nicholas Baran. News and views: Internet security in “infantile state”; Hedy Lamarr: Spread spectrum pioneer; new weather supercomputer not so super; researchers like armies of cheap robots; Kurzweil predicts human brain uploads. Dr. Dobbs Journal, 25(4): 18, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVNa

[Bar00c] Nicholas Baran. News and views: New modem standards should speed up Internet access; robocopter: AI lifts off; feet don’t fail me now; IBM claims world’s fastest supercomputer; new color displays based on light-emitting polymers; W3C moves forward with XLink. Dr. Dobbs Journal, 25(9):18, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Bark:2000:PBA

[Bar00d] Dennis L. Bark. Professor Bark’s amazing digital adventure: five faces of the Internet. Woodford Press, Emeryville, CA, USA, 2000. ISBN 0-942627-75-X. 280 pp. LCCN ZA4201 .B36 2000.

Baran:2001:NVA

[Bar01a] Nicholas Baran. News and views: Anonymity and the Internet; is industry hogging computer science talent?; relief from acronyms; OpenML spec released; C# not just a Java killer, says anders; and the winners are …. Dr. Dobbs Journal, 26(7):18, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/.

Baran:2001:NVM

[Bar01b] Nicholas Baran. News and views: Mail server benchmark released; new traveling salesman algorithm; ACLU challenges ICANN on freedom of speech; Microsoft opts for decaf; store your data on the Internet; and mind your electronic trash. Dr. Dobbs Journal, 26(4):18, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.cs.wustl.edu/
                                        ;   http://www.ddj.com/;

http://www.spec.org/osg/
mail2001/docs/FAQ.html;
http:/www.research.att.com/
~alb/ALENEX01.

Barnes:2001:OCI

[Bar01c] Susan B. Barnes. Online connections: Internet interpersonal relationships. The Hampton Press communication series. Hampton Press, Cresskill, NJ, USA, 2001. ISBN 1-57273-376-4 (paperback), 1-57273-375-6 (hardcover). xiv + 295 pp. LCCN HM1106.B37 2001.

Bartlett:2002:BPH

[Bar02] John G. Bartlett. Bioterrorism and public health: an Internet resource guide. eMedguides.com. Thomson Medical Economics, Montvale, NJ, USA, 2002. ISBN 1-56363-427-9. xi + 305 pp. LCCN QW 39 B616 2002.

Barken:2003:HSY

[Bar03] Lee Barken. How Secure is Your Wireless Network? Safeguarding Your Wi-Fi LAN. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-140206-4. xx + 199 pp. LCCN QA76.9.A25B3775 2003. US$34.99.

Bayne:2000:IMP

[Bay00] Kim M. Bayne. The Internet marketing plan: the complete guide to instant Web presence. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 2000. ISBN 0-471-35598-4 (paperback/CD-ROM). xix + 411 pp. LCCN HF5415.1265 .B39 2000.

Bear:2000:CDM

[BB00a] John Bear and Mariah P. Bear. College degrees by mail and Internet 2000: 100 accredited schools that offer Bachelor’s, Master’s, Doctorates, and law degrees by home study. Ten Speed Press, Berkeley, CA, USA, 2000. ISBN 1-58008-109-6. 205 pp. LCCN L901 .B36 2000.

Bunnell:2000:MCC

[BB00b] David Bunnell and Adam Brate. Making the Cisco connection: the story behind the real Internet superpower. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2000. ISBN 0-471-35711-1. xxi + 218 pp. LCCN HD9696.T444 C57 2000.

Briggs:2002:SHM

[BB02] Asa Briggs and Peter Burke. A social history of the media: from Gutenberg to the Internet. Blackwell, Malden, MA, USA, 2002. ISBN 0-7456-2374-3, 0-7456-2375-1 (paperback). viii + 374 pp. LCCN P90.B725 2002.

Ballon:2000:FAI

[BBL00a] Ian C. Ballon, James F. Brelsford, and Stuart D. Levi.

Fourth annual Internet Law Institute, volume G-610, G-611 of Intellectual property course handbook series; no. G-610, G-611. Practising Law Institute, New York, NY, USA, 2000. ISBN 0-87224-802-X (paperback). ???? pp. LCCN KF390.5.C6 I573 2000. Prepared for distribution at the Fourth annual Internet Law Institute Program, June–July 2000.

Barbier:2000:HMD

[BBL00b] Frederic Barbier and Catherine Bertho-Lavenir. Histoire des medias: de Diderot à Internet. (French) [History of the media: from Diderot to the Internet]. Collection U. Histoire. A. Colin, Paris, France, deuxième edition, 2000. ISBN 2-200-25226-9. 351 pp. LCCN P90 .B287 2000.

Bassi:2003:IBP

[BBM^{+}03] Alessandro Bassi, Micah Beck, Terry Moore, James S. Plank, Martin Swany, Rich Wolski, and Graham Fagg. The Internet Backplane Protocol: a study in resource sharing. Future Generation Computer Systems, 19(4):551–561, May 2003. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bickerton:2000:CEH

[BBP00] Pauline Bickerton, Matthew Bickerton, and Upkar Pardesi. Cybermarketing (2nd Edition): How to Use the Internet to Market Your Goods and Services. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-7506-4704-3. xvi + 355 pp. LCCN HF5415.1265 .B53 2000. US$44.95.

Biehl:2000:LGI

[BC00] Kathy Biehl and Tara Calishain. The lawyer’s guide to Internet research. Scarecrow Press, Metuchen, NJ, USA, 2000. ISBN 0-8108-3885-0 (paperback). x + 350 pp. LCCN KF242.A1 B46 2000. Includes “Glossary”–P. [307]-309.

Blumenthal:2001:RDI

[BC01] Marjory S. Blumenthal and David D. Clark. Rethinking the design of the Internet: the end-to-end arguments vs. the brave new world. ACM Transactions on Internet Technology (TOIT), 1(1):70–109, August 2001. CODEN ???? ISSN 1533-5399.

Bang:2002:MMC

[BC02] Y.-C. Bang and H. Choo. On multicasting with minimum costs for the Internet topology. Lecture Notes in Computer Science, 2400:736– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2400/24000736.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/

0558/papers/2400/24000736.
pdf.

Black:2006:SAI

[BCG06] J. Black, M. Cochran, and R. Gardner. A security analysis of the Internet Chess Club. IEEE Security & Privacy, 4 (1):46–52, January/February 2006. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33481/01588825.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33481&arnumber=1588825.

Blank:2000:BP

[BCK^{+}00] Jon Blank, Wankyu Choi, Allan Kent, Ganesh Prasad, and Chris Ullman. Beginning PHP4. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-373-0. ca. 800 pp. LCCN ???? US$39.99, UK£28.99, CDN$59.95. URL http:/

/www.wrox.com/Books/Book_
Details.asp?ISBN=1861003730.

Bellovin:2006:WPS

[BCK06] Steven M. Bellovin, Bill Cheswick, and Angelos D. Keromytis. Worm propagation strategies in an IPv6 Internet. ;login: the USENIX Association newsletter, 31(1): ??, February 2006. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/
login/2006-02/pdfs/bellovin.
pdf.

Braun:2005:WWI

[BCKT05] Torsten Braun, Georg Carle, Yevgeni Koucheryavy, and Vassilis Tsaoussidis, editors. Wired/wireless Internet communications: third international conference, WWIC 2005, Xanthi, Greece, May 11–13, 2005. Proceedings, volume 3510 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CO-DEN LNCSD9. ISBN 3-540-25899-X (soft cover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.
springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3510;   http:
//www.springerlink.com/openurl.
asp?genre=volume&id=doi:
10.1007/b136247.

Banks:2000:LI

[BD00] Cate Banks and Heather Douglas. Law on the Internet. Federation Press, Leichhardt, NSW, Australia, 2000. ISBN 1-86287-354-2. vii + 128 pp. LCCN KH49 .B36 2000.

Ballew:2005:DW

[BD05] Joli Ballew and Jeff Duntemann. Degunking Windows. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-933097-07-8. 376 (est) pp. LCCN ???? US$24.99, CDN$36.99.

Baxevanidis:2002:GRN

[BDFG02] Kyriakos Baxevanidis, Howard Davies, Ian Foster, and Fabrizio Gagliardi. Grids and research networks as drivers and enablers of future Internet architectures. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):5–17, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/97/27/28/
abstract.html.

Beaulieu:2002:WIA

[Bea02] Mark Beaulieu. Wireless Internet applications and architecture: building professional wireless applications worldwide. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-73354-4. 632 pp. LCCN TK5103.485.B43 2002.

Bejtlich:2005:TNS

[Bej05] Richard Bejtlich. The Tao of network security monitoring: beyond intrusion detection. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-24677-2 (paperback). xxxiv + 798 pp. LCCN TK5105.59 .B44 2005. URL http:/

/www.loc.gov/catdir/toc/
ecip0416/2004007857.html.

Bejtlich:2006:EDS

[Bej06] Richard Bejtlich. Extrusion detection: security monitoring for internal intrusions. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-34996-2 (paperback). xxviii + 385 pp. LCCN TK5105.59 .B43 2006.

BeMent:2000:AGI

[BeM00] T. J. BeMent. AILA’s guide to immigration and legal research on the Internet. American Immigration Lawyers Association, Washington, DC, USA, revised edition, 2000. ISBN 1-57370-065-7. iv + 119 pp. LCCN KF242.I42 B46 2000.

Bentham:2000:TIL

[Ben00] Jeremy Bentham. TCP/IP Lean: Web Servers for Embedded Systems. CMP Books, 6600 Silacci Way, Gilroy, CA 95020, USA, 2000. ISBN 1-929629-11-7. 416 (est.) pp. LCCN QA76.625 .B46 2000. US$49.95. URL http://www.

                                        cmpbooks.com/.

Benson:2001:NSC

[Ben01]  Allen C. Benson.  Neal-

Schuman complete Internet companion for librarians. Neal-Schuman Publishers, New York, NY, USA, second edition, 2001. ISBN 1-55570-414-X. xxxi + 566 pp. LCCN TK5105.875.I57 B46 2001.

Benantar:2002:IPK

[Ben02] Messaoud Benantar. Introduction to the Public Key Infrastructure for the Internet. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-060927-7. xii + 254 pp. LCCN QA76.9.A25 B45 2002. US$54.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130609277.html.

Berry:2000:NII

[Ber00] Nina M. Berry, editor. Network intelligence: Internetbased manufacturing: 8 November 2000, Boston, USA, volume 4208 of SPIE proceedings series; v. 4208 4007355562. SPIE Optical Engineering Press, Bellingham, WA, USA, 2000. ISBN 0-8194-3873-1. LCCN TS156.8 .N48 2000.

Berghel:2003:DVD

[Ber03]  Hal Berghel.  Digital vil-

lage: The discipline of Internet forensics. Communications of the ACM, 46(8):15–20, August 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Besing:2000:WBA

[Bes00] Ray G. Besing. Who broke up AT&T?: from Ma Bell to the Internet. 1stBooks, Bloomington, IN, USA, 2000. ISBN 1-58820-010-8. xiv + 413 pp. LCCN HE8846.A55 B47 2000.

Beyda:2000:DCB

[Bey00]  William J. Beyda.  Data

Communications: From Basics to Broadband. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-096139-6. 368 pp. LCCN TK5105.B485

2000.  URL http://www.

phptr.com/ptrbooks/esm_0130961396.
html.

Burgoyne:2001:IDP

[BF01] Patrick Burgoyne and Liz Faber. The Internet Design Project, user: the best of graphic art on the Web. Universe Publishers and St. Martin’s Press, New York, NY, USA, 2001. ISBN 0-7893-0636-0 (paperback). 144 pp. LCCN TK5105.888 .B87 2001.

Benameur:2002:QSF

[BFOBR02] N. Benameur, S. Ben Fredj, S. Oueslati-Boulahia, and J. W. Roberts. Quality of service and flow level admission control in the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):57–71, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/97/27/31/
abstract.html.

Bell:2002:FWA

[BFS^{+}02] Eric Bell, Hao Howard Feng, Edward L. W. Soong, David Zhang, and Shijia Sam Zhu. Fundamentals of Web Applications Using .NET and XML. The integrated .NET series from Object Innovations and Prentice Hall PTR. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-041790-4. xxvi + 564 pp. LCCN TK5105.888 .F86 2002.

US$44.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130417904.html.

Battersby:2000:DIA

[BG00a] Gregory J. Battersby and Charles W. Grimes. Drafting Internet agreements. Aspen Law and Business, Gaithersburg, MD, USA, 2000. ISBN 0-7355-1717-7. various pp. LCCN KF905.C6 B38. Looseleaf for updating.

Bressler:2000:CCB

[BG00b] Stacey E. Bressler and Charles E.

Grantham. Communities of commerce: building Internet business communities to accelerate growth, minimize risk, and increase customer loyalty. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-136115-4. x + 324 pp. LCCN HD30.37 .B74 2000.

Brown:2000:DIM

[BG00c] Jeffrey R. (Jeffrey Robert)

Brown and Austan Goolsbee. Does the Internet make markets more competitive?: evidence from the life insurance industry. NBER working paper series 7996, National Bureau of Economic Research, Cambridge, MA, USA, November 2000. 28 pp.

Bourke:2001:IQA

[BG01a] Jane Bourke and Judy Gabrovec.

Internet Quests: animals. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3406-7. 48 pp. LCCN QL50 .B68 2001.

Boyer:2001:DSU

[BG01b] Henri Boyer and Philippe Gardy. Dix siécles d’usages et d’images de l’occitan: des troubadours à l’Internet. (French) [Ten centuries of use and images of the Occitan: troubadors on the Internet]. Sociolinguistique. Harmattan, Paris, France, 2001. ISBN 2-7475-1324-6. 469 pp. LCCN PC3215 .D59 2001.

Burdett:2000:IOT

[BGE00] David Burdett, Marcus Goncalves, and Donald E. Eastlake. Internet Open Trading Protocol. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-135501-

4. xiii + 354 pp. LCCN

HF5548.32 .B87 2000.

Basta:2008:CSP

[BH08] Alfred Basta and Wolf Halton. Computer security and penetration testing. Thomson, Boston, MA, USA, 2008. ISBN 1-4180-4826-7. xvii + 414 pp. LCCN TK5105.59 .B366

                                        2008. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1302/2008270481-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1302/2008270481-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1302/2008270481-t.html.

Bhasin:2005:II

[Bha05] K. B. Bhasin. Interplanetary Internet. Computer Networks (Amsterdam, Netherlands: 1999), 47(5):599–601, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Basdogan:2000:ESR

[BhHSS00] Cagatay Basdogan, Chih hao Ho, Mandayam A. Srinivasan, and Mel Slater. An experimental study on the role of touch in shared virtual environments. ACM Transactions on Computer-Human Interaction, 7(4):443– 460, 2000. CODEN AT-CIF4. ISSN 1073-0516. URL

http://www.acm.org/pubs/
articles/journals/tochi/
2000-7-4/p443-basdogan/p443-
basdogan.pdf;   http://www.
acm.org/pubs/citations/journals/
tochi/2000-7-4/p443-basdogan/

.

Basha:2002:ANG

[BI02] S. Jeelani Basha and Romin Irani. AXIS: the next generation of Java SOAP. Wrox Press, Chicago, IL, USA, 2002. ISBN 1-86100-715-9. v + 275 pp. LCCN QA76.76.H94 B37 2002 Bar.

Biegel:2001:BOC

[Bie01] Stuart Biegel. Beyond our control?: confronting the limits of our legal system in the age of cyberspace. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-02504-3. xiv + 452 pp. LCCN KF390.5.C6 B495 2001. US$34.95, UK£23.95.

Birnbaum:2000:PEI

[Bir00] Michael H. Birnbaum. Psychological experiments on the Internet. Academic Press Inc., New York, NY, USA, 2000. ISBN 0-12-099980-3. xx + 317 pp. LCCN BF198.7 .P79 2000.

Birbeck:2001:PX

[Bir01] Mark Birbeck. Professional XML. Programmer to programmer. Wrox Press, Chicago, IL, USA, second edition, 2001. ISBN 1-86100-505-9 (paperback). xiii + 1269 pp. LCCN QA76.76.H94.P76 2001.

Birman:2003:IIC

[Bir03]  Ken Birman.  IEEE In-

ternet computing: Peer to peer: The League of SuperNets. IEEE Distributed Systems Online, 4(10), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0310/f/w5p2pp.htm.

Basedow:2000:LAG

[BK00a] Jurgen Basedow and Toshiyuki Kono. Legal aspects of globalization: conflict of laws, Internet, capital markets, and insolvency in a global economy. Kluwer Law International, Boston, MA, USA, 2000. ISBN 90-411-1332-0. xiii + 230 pp. LCCN K7340.A6 L44 2000.

Berger:2000:DBE

[BK00b] Arthur W. Berger and Yaakov Kogan. Dimensioning bandwidth for elastic traffic in high-speed data networks. IEEE/ACM Transactions on Networking, 8(5):643–654, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). URL

http://www.acm.org/pubs/
citations/journals/ton/2000-
8-5/p643-berger/.

Boni:2000:NGT

[BK00c] William Boni and Gerald L.

Kovacich. Netspionage: the global threat to information. Butterworth-Heinemann, Boston, MA, USA, 2000. ISBN 0-7506-7257-9. xx + 260 pp. LCCN HV6773 .B665 2000. US$34.95.

Breese:2000:GJI

[BK00d] Pierre Breese and Gautier Kaufman. Guide juridique de l’internet et du commerce électronique. (French) [Legal guide to the Internet and electronic commerce]. Entreprendre. Informatique. Vuibert, Paris, France, 2000. ISBN 2-7117-7995-5. 459 pp. LCCN KJV333.C65 B74 2000.

Bayles:2005:ICH

[BK05] Aaron W. Bayles and Christopher W. Klaus. InfoSec career hacking: sell your skillz, not your soul. Syngress Publishing, Inc., Rockland, MA, USA, 2005. ISBN 1-59749-011-3. xxix + 441 pp. LCCN QA76.25 .B39 2005.

Banerjee:2004:EPL

[BKB04] Suman Banerjee, Chris Kommareddy, and Bobby Bhattacharjee. Efficient peer location on the Internet. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):5– 17, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

BuSung:2003:DIJ

[BKY^{+}03] L. BuSung, Y. C. Kiat, S. I. Yann, L. K. Kee, and S. Wei. Design and implementation of a Java-based meeting space over Internet. Multimedia Tools and Applications, 20 (2):179–195, 2003. CODEN MTAPFB. ISSN 1380-7501.

Braunstein:2000:DBI

[BL00a] Marc Braunstein and Edward H. Levine. Deep branding on the Internet: applying heat and pressure online to ensure a lasting brand. Prima venture, Roseville, CA, USA, 2000. ISBN 0-7615-2532-7. xx + 380 pp. LCCN HF5415.1265 .B73 2000.

Bunnell:2000:BPB

[BL00b] David Bunnell and Richard A. Luecke. The e-Bay phenomenon: business secrets behind the world’s hottest Internet company. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2000. ISBN 0-471-38490-9 (hardcover). xi + 210 pp. LCCN HF5478 .B86 2000.

Bourke:2001:IQM

[BL01] Jane Bourke and Kellie Lloyd.

Internet quests: my community. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3404-0. 48 pp. LCCN HM131 .B666 2001.

Black:2000:IAI

[Bla00a]  Uyless D. Black.  Inter-

net Architecture: An Introduction to IP Protocols. Prentice Hall series in advanced communications technologies. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-019906-0. xviii + 380 pp. LCCN TK5105.585 .B536 2000. US$49.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130199060.html.

Black:2000:ISP

[Bla00b] Uyless D. Black. Internet Security Protocols: Protecting IP Traffic. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-014249-2. xviii + 286 pp. LCCN TK5105.59. B56 2000. US$49.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130142492.html.

Black:2001:ITC

[Bla01] Uyless D. Black. Internet telephony: call processing protocols. Prentice Hall series in advanced communications technologies. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-025565-3. xix + 373 pp. LCCN TK5105.8865.B52 2001.

Black:2002:TLI

[Bla02a] Sharon K. Black. Telecommunications law in the Internet age. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-546-0. xix + 516 pp. LCCN KF2761 .B54 2002.

Black:2002:ITR

[Bla02b] Uyless Black. IP Telephony Resource Kit. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-066062-0. ???? pp. LCCN ???? US$84.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130660620.html. Boxed

set containing Voice over IP, Second Edition, and Internet Telephony: Call Processing Protocols.

Black:2002:ONT

[Bla02c] Uyless Black. Optical Networks: Third Generation Transport Systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-060726-6. 352 (est.) pp. LCCN ???? US$55.00. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130607266.html.

Blaze:2002:ERS

[Bla02d] Matt Blaze. Efficient, DoSresistant, secure key exchange for Internet protocols (transcript of discussion). Lecture Notes in Computer Science, 2467:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2467/24670040.htm;
http://link.springer.de/
link/service/series/0558/
papers/2467/24670040.pdf.

Blanchet:2005:MIP

[Bla05] Marc Blanchet. Migrating to IPv6: a practical guide to implementing IPv6 in mobile and fixed networks. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2005. ISBN 0-471-49892-0 (paperback). xxxii + 418 pp. LCCN TK5105.585 .B5437 2005. URL http:

//www.loc.gov/catdir/toc/
ecip0518/2005025352.html.

Bradford:2002:LWI

[BM02] Ed Bradford and Lou Mauget.

Linux and Windows Interoperability Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-032477-9. 618 (est.) pp. LCCN ???? US$44.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130324779.html.

Boavida:2007:WWI

[BMM07] Fernando Boavida, Edmundo Monteiro, and Saverio Mascolo, editors. Wired/wireless Internet communications: 5th international conference, WWIC 2007, Coimbra, Portugal, May 23–25, 2007. proceedings, volume 4517 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-72694-2 (softcover). LCCN ????

Bauknecht:2000:ECW

[BMP00] Kurt Bauknecht, Sanjay Kumar Madria, and G. Pernul, editors. Electronic commerce and Web technologies: first international conference, EC-Web 2000, London, UK, September 4–6, 2000 proceedings, volume 1875 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67981-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN HF5548.32 .E35 2000.

Bauknecht:2001:ECW

[BMP01] Kurt Bauknecht, Sanjay Kumar Madria, and G. Pernul, editors. Electronic commerce and Web technologies: second international conference, EC-WEB 2001, Munich, Germany, September 4–6, 2001: proceedings, volume 2115 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42517-9 (softcover). LCCN QA267.A1 L43 no.2115; HF5548.32 .E35

2001. URL http://link.
springer-ny.com/link/service/
series/0558/tocs/t2115.htm.

The refereed proceedings of the Second International Conference on Electronic Commerce and Web Technologies.

Bhatnagar:2000:RCI

[BMR00] Amit Bhatnagar, Sanjog Misra, and H. Raghav Rao. On risk, convenience, and Internet shopping behavior. Communications of the ACM, 43(11):98–105, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-11/p98-bhatnagar/

.

Bissell:2001:CIW

[BMR01] Joan S. Bissell, Anna Manring, and Veronica Rowland. CyberEducator: the Internet and World Wide Web for K–12 and teacher education. McGraw-Hill College, Boston, MA, USA, second edition, 2001. ISBN 0-07-238056-X (book), 0-07-241668-8 (CD-ROM). vi + 122 pp. LCCN LB1044.87 .B57 2001. Accompanying CD-ROM includes: an overview of the book; a list of links featured in the book; Web tours; video mini-cases; Web activities; mailing lists; CyberLibrary links; and Web search tools.

Barcelo:2004:SIA

[BNHGV04] José M. Barceló, Juan I. Nieto-Hipólito, and Jorge García-Vidal. Study of Internet autonomous system interconnectivity from BGP routing tables. Computer Networks (Amsterdam, Netherlands: 1999), 45(3):333–344, June 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Bodow:2000:WFB

[Bod00] Steve Bodow. Websights: the future of business and design on the Internet. RC Publications, New York, NY, USA, 2000. ISBN 1-883915-07-4. 188 pp. LCCN TK5105.888 .W43 2000.

Bonnett:2000:IGD

[Bon00] Kendra Bonnett. An IBM guide to doing business on the Internet. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-031846-8. x + 261 pp. LCCN HF5548.32 .B66 2000.

Borner:2000:GIL

[Bor00] Fritjof Borner. The German Internet law adviser: legal and tax implications for e-commerce. Bundesanzeiger, Cologne, Germany, 2000. ISBN 3-88784-950-7 (paperback). 193 pp. LCCN KK164.C66 G47 2000.

Bourke:2001:IQP

[Bou01] Jane Bourke. Internet quests:

plants. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3401-6. 48 pp. LCCN QK49 .B68 2001.

Boukerche:2006:HAW

[Bou06] Azzedine Boukerche, editor. Handbook of algorithms for wireless networking and mobile computing. Chapman and Hall/CRC computer and information science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2006. ISBN 1-58488-465-7. various pp. LCCN TK5103.2 .H335

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0664/2005049375-d.html;
 http://www.loc.gov/catdir/
toc/fy0707/2005049375.html.

Bovik:2000:HIV

[Bov00] Alan Bovik, editor. Handbook of Image and Video Processing. Academic Press, New York, NY, USA, 2000. ISBN 0-12-119790-5. xv + 891 pp. LCCN TA1637 .H36 2000. US$105.00.

Bowman:2000:NXW

[Bow00] Charles F. Bowman. The new X Window system: a complete Internet architecture. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-18463-X. xx + 153 pp. LCCN QA76.76.W56 B6785 2000.

Brinson:2000:ILB

[BR00] J. Dianne Brinson and Mark F. Radcliffe. Internet law and business handbook. Ladera Press, Port Huron, MI, USA, 2000. ISBN 0-9639173-3-1 (paperback). xiii + 554 pp. LCCN KF390.5.C6 B76 2000.

Bradley:2000:BEI

[Bra00] Phil Bradley. The business and economy Internet resource handbook. Library Association Publishing, London, UK, 2000. ISBN 1-85604-351-7. xiv + 302 pp. LCCN ZA4201 .B87 2000.

Braun:2001:IIY

[Bra01] Linda W. Braun. Introducing the Internet to young learners: ready-to-go activities and lesson plans. Neal-Schuman Publishers, New York, NY, USA, 2001. ISBN 1-55570-404-2 (paperback). xvii + 147 pp. LCCN LB1044.87 .B72 2001.

Braithwaite:2003:UNP

[Bra03] Chris Braithwaite. UMTS Network Planning and Development: Design and Implementation of the 3G CDMA Infrastructure. Newnes, Amsterdam, The Netherlands and Boston, MA, USA, 2003. ISBN 0-7506-6082-1. ???? pp. LCCN ???? US$59.99.

Briere:2008:WHN

[Bri08]  Danny Briere.  Wireless

home networking for dummies. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, third edition, 2008. ISBN 0-470-25889-6 (paperback). ???? pp. LCCN

???? URL http://www.loc.
gov/catdir/enhancements/
fy0814/2008923598-b.html;
http://www.loc.gov/catdir/
enhancements/fy0814/2008923598-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0833/2008923598-t.html.

Brondmo:2000:ECN

[Bro00a] Hans Peter Brondmo. The eng@ged customer: the new rules of Internet direct marketing. HarperBusiness, New York, NY, USA, 2000. ISBN 0-06-662078-3. xxiii + 257 pp. LCCN HF5415.1265 .B76 2000.

Brown:2000:BBG

[Bro00b] Stuart M. Brown. Bioinformatics: a biologist’s guide to biocomputing and the Internet. Eaton, Natick, MA, USA, 2000. ISBN 1-881299-18-X. xi + 188 + 12 pp. LCCN QH506 .B767 2000.

Brown:2001:LRI

[Bro01] Valerie J. Atkinson Brown.

Legal research via the Internet. West Legal Studies series. West/Thomson Learning, Albany, NY, USA, 2001. ISBN 0-8273-7450-X. xv + 96 pp. LCCN KF242.A1 B76 2001; KF242.A1 B77 2001.

Browning:2002:EDU

[Bro02]  Graeme Browning.  Elec-

tronic democracy: using the Internet to transform American politics. CyberAge Books/Information Today, Inc., Medford, NJ, USA, second edition, 2002. ISBN 0-910965-49-8. xv + 190 pp. LCCN JK1764 .B78 2002.

Brush:2000:HTS

[Bru00] Kathleen E. Brush. Hightech strategies in the Internet era. Word Association Publishers, Tarentum, PA, USA, 2000. ISBN 1-891231-24-3. vi + 281 pp. LCCN HD62.37 .B78 2000.

Beretta:2000:IIJ

[BS00] Giordano B. Beretta and Raimondo Schettini, editors. Internet imaging: 26–28 January 2000, San Jose, California, volume 3964 of SPIE proceedings series; v. 3964 4006651037. SPIE Optical Engineering Press, Bellingham, WA, USA, 2000. ISBN 0-8194-3582-1. LCCN TS510.S63 v.3964.

Beretta:2001:III

[BS01a] Giordano B. Beretta and Raimondo Schettini, editors. Internet imaging II: 24–26 January, 2001, San Jose, [California] USA, volume 4311 of SPIE proceedings series; v. 4311 4007367116. SPIE Optical Engineering Press, Bellingham, WA, USA, 2001. ISBN 0-8194-3989-4. LCCN TS510.S63 v.4311.

Burgess:2001:IQWb

[BS01b] Betsy Burgess and Laurie Swiryn. Internet quests: Web activities across the curriculum. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3446-6. 80 pp. LCCN LB1044.87 .B87 2001.

Beretta:2002:III

[BS02a] Giordano B. Beretta and Raimondo Schettini, editors. Internet imaging III: 21–23 January 2002, San Jose, [California] USA, volume 4672 of SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 2002. ISBN 0-8194-4412-X. LCCN TS510.S63 v.4672.

Brassil:2002:SIM

[BS02b] Jack Brassil and Henning Schulzrinne. Structuring Internet media streams with cueing protocols. IEEE/ACM Transactions on Networking, 10(4): 466–476, August 2002.

Bray:2002:BCC

[BS02c] Jennifer Bray and Charles F.

Sturman. Bluetooth 1.1: Connect without Cables. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-066106-6. xxviii + 593 pp. LCCN TK5103.3 .B72 2002. US$48.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130661066.html.

Browner:2000:LIG

[BSP00] Stephanie Browner, Richard Sears, and Stephen Pulsford. Literature and the Internet: a guide for students, teachers, and scholars. Wellesley studies in critical theory, literary history, and culture; Garland reference library of the humanities; vol. 21. vol. 2167. Garland, New York, NY, USA, 2000. ISBN 0-8153-3453-2. ix + 191 pp. LCCN PN73 .B76 2000.

Besen:2002:ECE

[BSS02] Stanley M. Besen, Jeffrey S. Spigel, and Padmanabhan Srinagesh. Evaluating the competitive effects of mergers of Internet backbone providers. ACM Transactions on Internet Technology (TOIT), 2(3): 187–204, August 2002. CO-DEN ???? ISSN 1533-5399.

Becker:2000:SLI

[BSW00] Brandon Charles Becker, Stephen J. Schulte, and Michelle C. Wallach. Securities law and the Internet: doing business in a rapidly changing environment, 2000, volume B-1188, B-1189 of Corporate law and practice course handbook series; numbers B-1188, B-1189. Practising Law Institute, New York, NY, USA, 2000. ISBN 0-87224-792-9 (paperback). ???? pp. LCCN KF1440 .S44 2000. Prepared for distribution at the Securities Law and the Internet: Doing Business in a Rapidly Changing Environment Program, June-July 2000.

Bohme:2001:IIC

[BU01] Thomas Bohme and Herwig Unger, editors. Innovative Internet computing systems: International Workshop IICS 2001, Ilmenau, Germany, June 21–22, 2001: proceedings, volume 2060 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42275-7 (paperback). LCCN QA76.59 .I59 2001; QA267.A1 L43 no.2060. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t2060.htm.

Brandt:2003:IMM

[BU03] D. Scott Brandt and Lorna Uden. Insight into mental models of novice Internet searchers. Communications of the ACM, 46(7):133–136, July 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Buri:2000:VID

[Bur00]  Ueli Buri.  Die Verwech-

selbarkeit von Internet Domain Names: nach schweizerischen Firmen-, Marken- , Namens- und Lauterkeitsrecht. (German) [The confusion of Internet domain names: Swiss companies- , marks-, names- and integrity right], volume 633 of Abhandlungen zum schweizerischen Recht. Stampfli, Bern, Switzerland, 2000. ISBN 3-7272-0363-3. xliv + 252 pp. LCCN KE3879 .B86 2000.

Burr:2001:MCW

[Bur01] Alister Burr. Modulation and Coding for Wireless Communications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-201-39857-5. xx + 360 pp. LCCN TK5102.9 .B87 2001. US$83.00. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0201398575.html.

Burkhardt:2002:PCT

[Bur02] Jochen Burkhardt. Pervasive computing: technology and architecture of mobile Internet applications. Addison-Wes-ley, Reading, MA, USA, 2002. ISBN 0-201-72215-1. xxi + 410 pp. LCCN TK5105.875.I57 P472 2002.

Buttenfield:2002:TVG

[But02] Barbara P. Buttenfield. Transmitting vector geospatial data across the Internet. Lecture Notes in Computer Science, 2478:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/

service/series/0558/bibs/
2478/24780051.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2478/24780051.pdf.

Burgess:2001:IQWa

[BV01] Betsy Burgess and Judy Vandegrift. Internet Quests: world history. Teacher created materials. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3443-1. 80 pp. LCCN LB1584.7 .B87 2001.

Ballintijn:2000:CIP

[BvS00] Gerco Ballintijn and Maarten van Steen. Characterizing Internet performance to support wide-area application development. Operating Systems Review, 34(4):41–47, October 2000. CODEN OSRED8. ISSN 0163-5980.

Banes:2000:IA

[BW00a] David Banes and Richard Walter. Internet for all. David Fulton, London, UK, 2000. ISBN 1-85346-693-X. 108 pp. LCCN LB1044.87 .B36 2000.

Borriello:2000:EIE

[BW00b] Gaetano Borriello and Roy Want. Embedding the Internet: embedded computation meets the World Wide Web. Communications of the ACM, 43(5):59, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        2000-43-5/p59-borriello/.

Black:2002:STA

[BW02] Uyless Black and Sharleen Waters. SONET and T1: Architectures for Digital Transport Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2002. ISBN 0-13-065416-7. 384 (est.) pp. LCCN ???? US$55.00. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130654167.html.

Byrne:2003:MNI

[Byr03] John Byrne. From Molecules to Networks: An Introduction to Cellular and Molecular Neuroscience. Academic Press, New York, NY, USA, 2003. ISBN 0-12-148660-5. ???? pp. LCCN ???? US$79.95.

Cain:2001:IDS

[Cai01] Bruce E. Cain. The Internet in the (dis)service of democracy. Loyola of Los Angeles School of Law, Los Angeles, CA, USA, 2001. 1005–1021 pp. LCCN 2001 0885. Offprint: Loyola of Los Angeles law review. Vol. 34, no. 3 (Apr. 2001).

CCEB:2000:ECI

[Cal00] California Continuing Education of the Bar. Electronic commerce and Internet law and practice: 3rd annual recent developments program: program handbook, January 2000. Continuing Education of the Bar, Berkeley, CA, USA, 2000. v + 223 pp. LCCN KF390.5.C6 E44 2000.

CCEB:2001:ECI

[Cal01] California Continuing Education of the Bar. Electronic commerce and Internet law and practice: 4th annual recent developments program: program handbook. Continuing Education of the Bar, Oakland, CA, USA, January 2001. vii + 277 pp. LCCN KF390.5.C6 E44 2001.

CCEB:2002:ECI

[Cal02] California Continuing Education of the Bar. Electronic commerce and Internet law and practice: 5th annual recent developments program: program handbook, January 2002. Continuing Education of the Bar, Oakland, CA, USA, 2002. vi + 284 pp. LCCN KF390.5.C6 E44 2002.

Calvert:2003:TIS

[Cal03] Kenneth Calvert. TCP/IP Sockets in Java: Practical Guide for Programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-685-8. xii + 116 pp. LCCN QA76.625 .C35 2002. US$15.95.

Camp:2000:TRI

[Cam00] L. Jean Camp. Trust and risk in Internet commerce. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-03271-6. xiii + 279 pp. LCCN HF5548.32 .C35 2000.

Camp:2001:TRI

[Cam01] L. Jean Camp. Trust and Risk in Internet Commerce. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-53197-6. 296 (est.) pp. LCCN ???? US$19.95, UK£13.95.

Crawford:2004:FCL

[CAN^{+}04] Diane Crawford, Scott Ambler, Peter G. Neumann, Peter Turk, Hartmut Wedekind, John Stanley, Julian V. Noble, Hong-Lok Li, Michael Blaha, and Dean Wakerley. Forum: Consider long-term implications now and forever; the real philosophical roots of literacy; give me common carrier Internet service; for science and math, skip the online course; prepare CS graduates for industrial reality; elegant code produces maximal comprehension. Communications of the ACM, 47(10):11– 13, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Capron:2000:CPI

[Cap00a] H. L. Capron. Capron’s pocket Internet: 4001 sites. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-201-61195-3. 384 pp. LCCN ZA4226 .C39 2000.

Capron:2000:CTI

[Cap00b]  H. L. Capron.  Comput-

ers: tools for an Information Age. Prentice-Hall, Upper Saddle River, NJ 07458, USA, sixth edition, 2000. ISBN 0-201-47659-2, 0-201-61192-9 (instructor’s edition), 0-201-61211-9 (paperback/CD-ROM). xxxi + 576 pp. LCCN QA76.5 .C363 2000.

Carpenter:2000:EBI

[Car00a]  Phil Carpenter.  eBrands:

building an Internet business at breakneck speed. Harvard Business School Press, Boston, MA, USA, 2000. ISBN 0-585-16220-4 (electronic book), 0-87584-929-6. x + 301 pp. LCCN HD69.B7 C37 2000. US$25.95.

Carter:2000:L

[Car00b]  Sarah Carter.  Lawlinks.

Cavendish, London, UK, 2000. ISBN 1-85941-607-1 (paperback). 80 pp. LCCN KD394 .C37 2000. US$20.00.

Carlomagno:2001:FRT

[Car01a] Marcos Cantera Carlomagno.

El Foro Rodelu: la tertulia virtual del Uruguay en Internet. Torre del Vigia-Ediciones, Montevideo, Uruguay, 2001. ISBN 9974-7593-2-3. 223 pp. LCCN ZA4480 .F67 2001.

Carlson:2001:DYD

[Car01b] Gustav Carlson. Do you dotcom?: a field guide to understanding life at an Internet company. AMACOM, New York, NY, USA, 2001. ISBN 0-8144-7103-X (paperback). viii + 232 pp. LCCN HF5549 .C2935 2001.

Carr:2009:ICW

[Car09] Jeffrey Carr. Inside cyber warfare. Mapping the cyber underworld. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2009. ISBN 0-596-80215-3. xviii + 212 pp. LCCN ????

Casey:2000:DEC

[Cas00a] Eoghan Casey. Digital evidence and computer crime: forensic science, computers and the Internet. Academic, London, UK, 2000. ISBN 0-12-162885-X. 279 pp. LCCN K5475 .C37 2000.

Castro:2000:HWW

[Cas00b] Elizabeth Castro. HTML 4 for the World Wide Web. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, fourth edition, 2000. ISBN 0-201-35493-4. 384 pp. LCCN QA76.76.H94 C3698 2000.

Castells:2001:IGR

[Cas01] Manuel Castells. The Internet galaxy: reflections on the Internet, business, and society. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-924153-8. xi + 292 pp. LCCN HD79.I55 C38 2001.

Cheswick:2003:FIS

[CBR03] William R. Cheswick, Steven M.

Bellovin, and Aviel D. Rubin. Firewalls and Internet Security: Repelling the Wily Hacker. Addison-Wes-ley, Reading, MA, USA, second edition, 2003. ISBN 0-201-63466-X. xiv + 433 pp. LCCN TK5105.875.I57C44 2003. US$49.99, CDN$77.99.

Carlton:2001:FRS

[CC01] Dennis W. Carlton and Judith A. Chevalier. Free riding and sales strategies for the Internet. NBER working paper series; no. 8067 4007425215 8067, National Bureau of Economic Research, Cambridge, MA, USA, January 2001. 20 + [13] pp.

Costa:2005:VEE

[CCC^{+}05] Manuel Costa, Jon Crowcroft, Miguel Castro, Antony Rowstron, Lidong Zhou, Lintao Zhang, and Paul Barham. Vigilante: end-to-end containment of Internet worms. Operating Systems Review, 39(5): 133–147, December 2005. CO-DEN OSRED8. ISSN 0163-5980.

Chiueh:2000:CMD

[cCP00] Tzi cker Chiueh and Prashant Pradhan. Cache memory design for Internet processors. IEEE Micro, 20 (1):28–33, January/February 2000. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2000/pdf/m1028.
pdf.

Chiu:2001:IQP

[CdBvdM01] Angela L. Chiu, Frank Huebner-Szabo de Bucs, and Robert D. van der Mei, editors. Internet quality and performance and control of network systems: 6–7 November 2000, Boston, USA, volume 4211 of SPIE proceedings series; v. 4211 4007384133. SPIE Optical Engineering Press, Bellingham, WA, USA, 2001. ISBN 0-8194-3876-6. LCCN TK5105.5 .I626 2001. A merger of Internet: quality of service and future directions and the Performance and control of network systems conferences.

Chen:2002:SID

[CDT02] Y. Chen, A. K. Datta, and S. Tixeuil. Stabilizing inter-domain routing in the Internet (research note). Lecture Notes in Computer Science, 2400:749–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2400/24000749.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2400/24000749.
pdf.

CF:2000:EBI

[Cen00] Century Foundation. Enforcing a ban on Internet gambling. Idea brief 16, Century Foundation, New York, NY, USA, September 2000.

4 pp. URL http://www.
ideas2000.org/IdeaBriefs.
html.

Cerf:2001:BPI

[Cer01] Vinton Cerf. Beyond the post-PC Internet. Communications of the ACM, 44(9):34– 37, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/articles/journals/cacm/
2001-44-9/p34-cerf/p34-cerf.
pdf;   http://www.acm.org/
pubs/citations/journals/
cacm/2001-44-9/p34-cerf/.

Crowley:2002:NPDa

[CFHO02a] Patrick Crowley, Mark A.

Franklin, Haldun Hadimioğlu, and Peter Z. Onufryk, editors. Network Processor Design: Issues and Practices, Volume 1. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-875-3. 340 (est.) pp. LCCN ???? US$59.95.

Crowley:2002:NPDb

[CFHO02b] Patrick Crowley, Mark A.

Franklin, Haldun Hadimioğlu, and Peter Z. Onufryk, editors. Network Processor Design: Issues and Practices, Volume 2. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 0-232-98267-6. 340 (est.) pp. LCCN ???? US$64.95.

Feng:2003:OGE

[cFJH^{+}03] Wu chun Feng, Justin, Hurwitz, Harvey B. Newman, Sylvain Ravot, Roger Les Cottrell, Olivier Martin, Fabrizio Coccetti, Cheng Jin, David Wei, and Steven Low. Optimizing 10-gigabit Ethernet in networks of workstations, clusters, and grids: a case study. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.
sc-conference.org/sc2003/
inter_cal/inter_cal_detail.
php?eventid=10702#1;   http:
//www.sc-conference.org/
sc2003/paperpdfs/pap293.
pdf.

Ciciora:2004:MCT

[CFLA04] Walter Ciciora, James Farmer, David Large, and Michael Adams, editors. Modern cable television technology: video, voice and data communications. Elsevier/Morgan Kaufmann, San Francisco, CA, USA, second edition, 2004. ISBN 1-55860-828-1. xxxviii + 1053 pp. LCCN TK6675 .M64 2004. US$84.95. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els051/2003107478.
html;   http://www.loc.gov/
catdir/toc/els051/2003107478.
html.

Compaine:2001:CPT

[CG01] Benjamin M. Compaine and Shane M. Greenstein, editors. Communications policy in transition: the Internet and beyond. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-03292-9 (hardcover). LCCN P95.82.U6 C66 2001.

Chang:2004:TCR

[CGJ^{+}04] Hyunseok Chang, Ramesh Govindan, Sugih Jamin, Scott J. Shenker, and Walter Willinger. Towards capturing representative AS-level Internet topologies. Computer Networks (Amsterdam, Netherlands: 1999), 44(6):737–755, April 22, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Cowan:2000:RIR

[CH00] Douglas E. Cowan and Jeffrey K. Hadden. Religion on the Internet: research prospects and promises, volume 8 of Religion and the social order; vol.8. JAI, New York, NY, USA, 2000. ISBN 0-7623-0535-5. xi + 365 pp. LCCN BL37 .H32 2000. Contents: 1. Introduction — 2. Internet research: studying religion on the Web — 3. Internet faith: religions in Cyberspace — 4. Webs of deceit: religious propaganda on the Net — 5. Internet teaching: pedagogy and the World Wide Web.

Chaffey:2000:IMS

[Cha00] Dave Chaffey. Internet marketing: strategy, implementation, and practice. Financial Times Prentice Hall, New York, NY, USA, 2000. ISBN 0-273-64309-6. xix + 508 pp. LCCN HF5415.1265 .I57 2000.

Chang:2002:AIS

[Cha02] Whie Chang, editor. Advanced Internet services and applications: first international workshop, AISA 2002, Seoul, Korea, August 1–2, 2002: proceedings, volume 2402 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-43968-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 A34 2002.

Chandra:2005:BWS

[Cha05] Praphul Chandra. Bulletproof wireless security: GSM, UMTS, 802.11 and ad hoc security. Communications engineering series. Newnes, Amsterdam, The Netherlands, 2005. ISBN 0-7506-7746-5 (paperback). xxxiii + 237 pp. LCCN TK5103.2 .C445 2005. URL http:

                                        //books.elsevier.com/us/
                                        mk/us/subindex.asp?isbn=
                                        0750677465.

Cilia:2002:PIC

[CHB02] Mariano Cilia, P. Hasselmayer, and Alejandro P. Buchmann. Profiling and Internet connectivity in automotive environments. In Bernstein et al. [B^{+}02], pages 1071–1074. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.
org/conf/2002/S33P08.pdf.

Chang:2007:VMS

[CHCC07] Da-Wei Chang, Cheng-En Hsieh, Yan-Pai Chen, and Kwo-Cheng Chiu. Virtual machine support for zero-loss Internet service recovery and upgrade. Software—Practice and Experience, 37(13):1349– 1376, November 10, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chellen:2000:EGI

[Che00] Sydney S. Chellen. The essential guide to the Internet. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2000. ISBN 0-415-22747-X (paperback). xiv + 215 pp. LCCN W26.5 C455 2000.

Chen:2001:IOI

[Che01] Thomas M. Chen. Increasing the observability of Internet behavior. Communications of the ACM, 44(1): 93, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
cacm/2001-44-1/p93-chen/.

Chesta:2002:GVE

[Che02] C. Chesta. Globalization of voice-enabled Internet access solutions. Lecture Notes in Computer Science, 2411:398– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2411/24110398.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2411/24110398.pdf.

Cheah:2005:FS

[Che05] Chu Yeow Cheah. Firefox secrets. A need-to-know guide. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2005. ISBN 0-9752402-4-2. xvi + 297 pp. LCCN TK5105.883.F55 C44 2005.

Chien:2001:PPC

[Chi01] Andrew A. Chien. Parallel programming challenges for Internet-scale computing (entropia). ACM SIG-PLAN Notices, 36(7):82, July 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/proceedings/ppopp/
                                        379539/p82-chien/p82-chien.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/proceedings/
                                        ppopp/379539/p82-chien/.

Chon:2001:FID

[Cho01a] Kilnam Chon. The future of the Internet digital divide. Communications of the ACM, 44(3):116–117, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2001-44-3/p116-chon/.

Chorafas:2001:ISC

[Cho01b] Dimitris N. Chorafas. The Internet supply chain: impact on accounting and logistics. Palgrave, New York, NY, USA, 2001. ISBN 0-333-94963-3 (hardcover). xxii + 298 pp. LCCN HF5548.32 .C478 2001.

Churchhouse:2002:CCJ

[Chu02] Robert F. Churchhouse. Codes and ciphers: Julius Caesar, the Enigma, and the Internet. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-81054-X (hardcover), 0-521-00890-5 (paperback). x + 240 pp. LCCN Z103 .C48 2002. US$55.00 (hardcover), US$20.00 (paperback).

Caro:2003:RFS

[CIA^{+}03] Armando L. Caro Jr., Janardhan R. Iyengar, Paul D. Amer, Sourabh Ladha, Gerard J. Heinz II, and Keyur C. Shah. Research feature: SCTP: a proposed standard for robust Internet data transport. Computer, 36(11):56–63, November 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/co/2003/11/ry056.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/11/ry056.
                                        pdf;   http://www.computer.
                                        org/computer/co2003/ry056abs.
                                        htm.

Cisco:2000:ITO

[Cis00] Cisco Online. Internetworking technology overview: SNMP. World-Wide Web document.,

                                        2000.  URL http://www.
                                        cisco.com/univercd/cc/td/
                                        doc/cisintwk/ito_doc/snmp.
                                        htm.

Chappell:2002:JWS

[CJ02] David A. Chappell and Tyler Jewell. Java Web services. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00269-6. xii + 262 pp. Using Java in service-oriented architectures.

Choi:2002:CPE

[CJL02] Yoonsuk Choi, Howon Jung, and Kyungshik Lim. A comparative performance evaluation of the transport layer protocols for wireless transaction processing in mobile Internet environments. Lecture Notes in Computer Science, 2344:11– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2344/23440011.htm;
                                        http://link.springer.de/

link/service/series/0558/
papers/2344/23440011.pdf.

Clarke:2000:VSP

[CJM00] E. M. Clarke, S. Jha, and W. Marrero. Verifying security protocols with Brutus. ACM Transactions on Software Engineering and Methodology, 9(4):443–487, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tosem/
2000-9-4/p443-clarke/p443-
clarke.pdf;   http://www.
acm.org/pubs/citations/journals/
tosem/2000-9-4/p443-clarke/

.

Campbell:2000:MQV

[CK00] Jennifer (Jennifer Ruth) Campbell and Michael Keene. The Mayfield quick view guide to the Internet for students of English. Mayfield Pub. Co., Mountain View, CA, USA, version 2.0 edition, 2000. ISBN 0-7674-1146-3. iv + 54 pp. LCCN PE1066 .C36 2000. Version 2.0.

Chircu:2001:MVI

[CKK01] Alina M. Chircu, Robert J.

Kauffman, and Doug Keskey. Maximizing the value of Internet-based corporate travel reservation systems. Communications of the ACM, 44(11): 57–63, November 2001. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cheneau-Loquay:2000:ETC

[CL00] Annie Cheneau-Loquay. Enjeux des technologies de la communication en Afrique: du téléphone à Internet: . (French) [Stakes of communication technologies in Africa: from the telephone to the Internet]. Hommes et sociétés. REGARDS, Talence, France, 2000. ISBN 2-86537-989-2. 402 pp. LCCN HE8464 .E55 2000.

Comer:2002:HNI

[CL02] Douglas E. Comer and W. David Laverell. Hands-on Networking with Internet Technologies. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-048003-7. 192 (est.) pp. LCCN ???? US$30.00. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130480037.html.

Clark:2003:IIC

[Cla03a] David D. Clark. IEEE Internet computing: Peer to peer: Economics and the design of open systems. IEEE Distributed Systems Online, 4(4), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0304/f/p2p_print.htm.

Clark:2003:DNI

[Cla03b] Martin P. Clark. Data Networks, IP and the Internet: Protocols, Design and Operation. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2003. ISBN 0-470-84856-1. 864 (est.) pp. UK£85.00, EUR 140.30.

URL http://www.wiley.co.
uk/clarkdata/.

Clift:2000:VID

[Cli00] Steven Clift. Viewpoint: an Internet of democracy. Communications of the ACM, 43 (11):31–32, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://

www.acm.org/pubs/citations/
journals/cacm/2000-43-11/
p31-clift/.

Clifford:2001:CBG

[Cli01] Karen Clifford. The complete beginner’s guide to genealogy, the Internet, and your genealogy computer program. Genealogical Pub. Co., Baltimore, MD, USA, 2001. ISBN 0-8063-1636-5. various pp. LCCN CS14 .C56 2001.

Crawford:2000:FOS

[CLWR00] Diane Crawford, Robert Levine, John A. Wills, and James L. Rash. Forum: The other side of embedding the Internet; A cure for lost programming companionship;l licensing software developers. Communications of the ACM, 43(8):11–14, August 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/

                                        citations/journals/cacm/
                                        2000-43-8/p11-crawford/.

Cabri:2000:MAC

[CLZ00] Giacomo Cabri, Letizia Leonardi, and Franco Zambonelli. Mobileagent coordination models for Internet applications. Computer, 33(2):82–89, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://dlib.computer.org/
                                        co/books/co2000/pdf/r2082.
                                        pdf.

Cahoon:2000:EPD

[CML00a] Brendon Cahoon, Kathryn S. McKinley, and Zhihong Lu. Evaluating the performance of distributed architectures for information retrieval using a variety of workloads. ACM Transactions on Information Systems, 18(1):1– 43, January 2000. CODEN ATISET. ISSN 1046-8188.

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        tois/2000-18-1/p1-cahoon/

.

Choi:2000:DII

[CML00b] James J. Choi, Andrew Metrick, and David I. Laibson. Does the Internet increase trading?: evidence from investor behavior in 401(k) plans. NBER working paper series 7878, National Bureau of Economic Research, Cambridge, MA, USA, 2000. 46 pp.

Capling:2000:DDT

[CN00] Ann Capling and Kim Richard Nossal. Death of distance or tyranny of distance?: the Internet, deterritorialisation, and the anti-globalisation movement in Australia. Technical Report 2000/3, Research School of Pacific and Asian Studies, Australian National University, Canberra, ACT 0200, Australia, October 2000. ISBN 0-7315-3108-6. 28 pp.

Chandrasiri:2004:ICU

[CNI^{+}04] Naiwala P. Chandrasiri, Takeshi Naemura, Mitsuru Ishizuka, Hiroshi Harashima, and István Barakonyi. Internet communication using real-time facial expression analysis and synthesis. IEEE MultiMedia, 11(3):20–29, July/September 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL

http://csdl.computer.org/
dl/mags/mu/2004/03/u3020.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/03/u3020.
pdf;   http://www.hc.t.u-
tokyo.ac.jp/~cds/IEEEmultimedia/

.

Cochran:2003:NVd

[Coc03a] Shannon Cochran. News and views: Caltech “Turing Tournament”; science and engineering Ph.D.s getting scarcer; E. Coli for data storage; Swiss accept Internet ballots; extensible resource identifiers. Dr. Dobbs Journal, 28(4):10, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7827/ddj0304o/.

Cochran:2003:NVC

[Coc03b]  Shannon Cochran.  News

and views: Cryptographers [Ronald Rivest, Adi Shamir, and Leonard Adleman] receive Turing Award; computerscience pioneer [John G. “Jack” Herriot] passes away; programming quantum computers; the demography of the Internet. Dr. Dobbs Journal, 28(7):14, July 2003. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Cohen:2000:AGI

[Coh00a] Eric E. Cohen. Accountant’s guide to the Internet. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 2000. ISBN 0-471-35834-7 (paperback). xx + 364 pp. LCCN HF5625.7 .C64 2000.

Cohen:2000:CDI

[Coh00b] Jonathan Cohen. Communication and design with the Internet. W.W. Norton, New York, NY, USA, 2000. ISBN 0-393-73043-3. 288 pp. LCCN NA2728 .C62 2000.

Collora:2000:CCF

[Col00]  Michael A. Collora.  Cy-

bercrime and corporate fraud: rights and remedies: a cutting edge discussion of computer fraud, from stock swindles via the Internet to theft of trade secrets by employees. Massachusetts Continuing Legal Education, 10 Winter Place, Boston 02108-4751, MA, USA, 2000. xiv + 528 pp. LCCN KF9350.Z9 C93 2000.

Collins:2001:LDI

[Col01] Matthew Collins. The law of defamation and the Internet. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-924468-5. lvi + 430 pp. LCCN KF1266 .C65 2001.

Cole:2002:ITV

[Col02] Marion Cole. Introduction to telecommunications: voice, data, and the Internet. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-060890-4. 608 pp. LCCN TK5101 .C63496 2002.

Comer:2000:IBE

[Com00a] Douglas Comer. The Internet book: everything you need to know about computer networking and how the Internet works. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2000. ISBN 0-13-030852-8. xxvi + 351 pp. LCCN TK5105.875.I57 C65 2000.

Comer:2000:ITI

[Com00b] Douglas E. Comer. Internetworking with TCP/IP: Volume 1. Principles, Protocols, and Architectures. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2000. ISBN 0-13-018380-6 (hardcover). xxix + 750 pp. LCCN TK5105.585 .C66 2000 v.1, 3 (2000-c2001).

Compaine:2001:DDF

[Com01] Benjamin M. Compaine, editor. The digital divide: facing a crisis or creating a myth? MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-53193-3. xvi + 357 pp. LCCN HN90.I56 D54 200. US$29.95, UK£20.00.

Connolly:2002:LIS

[Con02] Kevin J. Connolly. Law of Internet security and privacy. Aspen Law and Business, New York, NY, USA, 2002. ISBN 0-7355-2211-1. xxiii + 394 pp. LCCN KF390.5.C6 C68 2002.

Cooklev:2004:WCS

[Coo04] Todor Cooklev. Wireless communication standards: a study of IEEE 802.11, 802.15, and 802.16. Standards Information Network, IEEE Press, New York, NY, USA, 2004. ISBN 0-7381-4066-X. 325 (est.) pp. LCCN TK5105.5668 .C66 2004. US$65.00. URL http:/

/shop.ieee.org/ieeestore/
product.aspx?product_no=

SP1134. IEEE Product Number SP1134.

Cothran:2002:IOV

[Cot02] Helen Cothran. The Internet: opposing viewpoints. Opposing viewpoints series. Greenhaven Press, San Diego, CA, USA, 2002. ISBN 0-7377-0779-8 (paperback), 0-7377-0780-1 (lib.). 218 pp. LCCN HM851 .I57 2002.

Crowcroft:2001:TIL

[CP01] Jon Crowcroft and Iain Phillips. TCP/IP and the Linux protocol implementation: systems code for the Linux Internet. Wiley Networking Council series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-40882-4 (hardcover). xlix + 925 pp. LCCN TK5105.585 .T34 2001.

Chen:2003:IIC

[CP03] Yih-Farn Robin Chen and Charles Petrie. IEEE Internet computing: Guest Editors’ introduction: Ubiquitous mobile computing. IEEE Distributed Systems Online, 4(4), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0304/f/w2gei.htm.

Chin:2001:EBE

[CPM01] Wendy Chin, Frederic Patricelli, and Veljko Milutinovic. Electronic business and education: recent advances in Internet infrastructures. Multimedia systems and applications series. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-7508-4. xii + 400 pp. LCCN TK5105.875.I57 E42 2002.

Calhoun:2001:CNT

[CR01a] Karen Calhoun and John J. Riemer. CORC: new tools and possibilities for cooperative electronic resource description. Haworth Information Press, New York, NY, USA, 2001. ISBN 0-7890-1304-5, 0-7890-1305-3 (paperback). xvi + 184 pp. LCCN Z695.24 .C67 2001. “Co-published simultaneously as Journal of Internet cataloging, volume 4, numbers 1/2, 2001.”.

Cheyne:2001:TAI

[CR01b] Tanya L. Cheyne and Frank E. Ritter. Targeting audiences on the Internet. Communications of the ACM, 44(4): 94–98, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        cacm/2001-44-4/p94-cheyne/

.

Crane:2000:TIS

[Cra00] Beverley E. Crane. Teaching with the Internet: strategies and models for K–12 curricula. Neal-Schuman Publishers, New York, NY, USA, 2000. ISBN 1-55570-375-5. xxi + 381 pp. LCCN LB1044.87 .C73 2000.

Cramer:2007:SPP

[Cra07]  Curt Cramer.  Structured

peer-to-peer overlays for wireless mobile ad-hoc networks. Berichte aus der Informatik. Shaker, Aachen, Germany, 2007. ISBN 3-8322-0672-8. xvii + 190 pp. LCCN ????

URL http://www.shaker.
eu/Catalogue/details.asp?
ID=1858968&CC=21&ISBN=3-
8322-6072-2&Reihe=-20&ReiheUR=

-2. Universität Karlsruhe dissertation.

Criddle:2000:PLG

[Cri00] Sally Criddle. The public librarian’s guide to the Internet. Library Association, London, UK, 2000. ISBN 1-85604-328-

2. xi + 249 pp. LCCN ZA4201 .P83 2000. Published in association with UKOLN.

Crisp:2003:ICC

[Cri03]  John Crisp.  Introduction

to Copper Cabling: Applications for Telecommunications, Data Communications and Networking. Newnes, Amsterdam, The Netherlands and Boston, MA, USA, 2003. ISBN 0-7506-5555-0. ???? pp. LCCN ???? US$29.99.

Crow:2000:LHN

[Cro00]  Preston F. Crow.  The

Linux home network. Linux Journal, 72:??, April 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue72/3575.
html.

Crowcroft:2000:QFI

[CRS00] Jon Crowcroft, James Roberts, and Mikhail I. Smirnov, editors. Quality of future Internet services: first COST 263 international workshop, QofIS 2000, Berlin, Germany, September 25–26, 2000: proceedings, volume 1922 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-41076-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 Q36 2000; QA267.A1 L43 no.1922.

Carzaniga:2001:DEW

[CRW01] Antonio Carzaniga, David S. Rosenblum, and Alexander L. Wolf. Design and evaluation of a wide-area event notification service. ACM Transactions on Computer Systems, 19(3): 332–383, 2001. CODEN AC-SYEC. ISSN 0734-2071. URL

                                        http://www.acm.org/pubs/
                                        citations/journals/tocs/
                                        2001-19-3/p332-carzaniga/

.

Crystal:2001:LI

[Cry01] David Crystal. Language and the Internet. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80212-1. ix + 272 pp. LCCN P107 .C78 2001.

Comer:2001:ITI

[CS01] Douglas E. Comer and David L.

Stevens. Internetworking with TCP/IP, Vol. 3: Client-Server Programming and Applications, Linux/Posix Sockets Version. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-032071-4. ???? pp. LCCN TK5105.585 .C66 2001. URL http:/

/www.phptr.com/ptrbooks/
esm_0130320714.html.

Casaca:2002:GET

[CS02] Augusto Casaca and Amardeo C.

Sarma. Guest editorial: Towards a new Internet architecture. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):1–4, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/97/27/27/
abstract.html.

Chung:2000:ECM

[CSK00] C.-M. Chung, P.-S. Shin, and S.-D. Kim. An effective configuration method for Java– Internet computing environment. Parallel Processing Letters, 10(1):73–??, March 2000. CODEN PPLTEE. ISSN 0129-6264. URL http:

//ejournals.wspc.com.sg/
ppl/10/1001/S0129626400000093.
html.

Cao:2002:IPG

[CSMW02] Xi-Ren Cao, Hong-Xia Shen, Rodolfo Milito, and Patrica Wirth. Internet pricing with a game theoretical approach: concepts and examples. IEEE/ ACM Transactions on Networking, 10(2):208–216, April 2002.

Cristofori:2000:RAC

[CSS00] Alessandro Cristofori, Carla Salvaterra, and Ulrich Schmitzer. La rete di Arachne: contributi su nuove tecnologie, didattica ed antichita classica nell’ambito del progetto “Telemaco” = Arachnes Netz: Beitrage zu Antike, EDV und Internet im Rahmen des Projekts “Telemachos”, volume 71 of Palingenesia; ISSN: 0552-9638; Bd. 71. F. Steiner, Stuttgart, Germany, 2000. ISBN 3-515-07821-5. ISSN 0552-9638. 281 pp. LCCN LC1021 .R48 2000. Italian, German, English, Spanish or French.

Crawford:2001:FPV

[CTBA^{+}01] Diane Crawford, Thomas Tiahrt, Moti Ben-Ari, Matt West, Hans A. von Spakovsky, and Deborah Phillips. Forum: Participatory vs. representative democracy; why store everything?; Emulex hoax; correction. Communications of the ACM, 44(4):13–15, April 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2001-44-4/p13-crawford/.

See [Sch00a, PvS01].

Chang:2007:KSZ

[CTLL07] Da-Wei Chang, Chuan-Ming Tsai, Wei-Kou Li, and Tzu-Rung Lee. Kernel support for zero-loss Internet service restart. Software—Practice and Experience, 37(8):833– 855, July 10, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chu:2003:ODA

[CTYS03] Lingkun Chu, Hong Tang, Tao Yang, and Kai Shen. Optimizing data aggregation for cluster-based Internet services. ACM SIGPLAN Notices, pages 119–130, 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Curtis:2000:UAS

[Cur00] Todd Curtis. Understanding aviation safety data: using the Internet and other sources to analyze air travel risk. Society of Automotive Engineers, Warrendale, PA, USA, 2000. ISBN 0-7680-0602-3. x + 209 pp. LCCN TL553.5 .C87 2000.

Cockcroft:2000:CPI

[CW00] Adrian Cockcroft and Bill Walker. Capacity Planning for Internet Services. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-089402-8. xx + 222 pp. LCCN TK5105.5.C557 2001.

URL books/caphi.pdf;
http://www.sun.com/books/
catalog/capplan/. Part No.

806-3684-10 May 2000, Revision 01.

Chen:2005:SBP

[CWZ^{+}05] Songqing Chen, Haining Wang, Xiaodong Zhang, Bo Shen, and Susie Wee. Segmentbased proxy caching for internet streaming media delivery. IEEE MultiMedia, 12(3): 59–67, July/September 2005. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic).

Chen:2004:IBN

[CZ04] Jyh-Cheng Chen and Tao Zhang. IP-Based Next-Generation Wireless Networks: Systems, Architectures and Protocols. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2004. ISBN 0-471-23526-1. 438 (est.) pp. LCCN TK5105.78.C49

2004. URL http://www.loc.
gov/catdir/toc/wiley032/
2003012945.html.

Crawford:2001:FHC

[CZB^{+}01] Diane Crawford, Mick Zraly, Hal Berghel, Ken Pugh, Mat H. West, Conrad Weisert, Terry Steyaert, and Richard Johnson. Forum: How can the Web advance Western democracies? who needs digital signatures; misinformation and the Emulex hoax; OOSCD not really so unified; go back to non-OOSD. Communications of the ACM, 44(2):11–13, February 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2001-44-2/p11-crawford/.

See [Sch00a].

Alvarez:2000:DII

[dAA00] Enrique de Alarcon Alvarez.

Diccionario de informatica e Internet. Guias practicas. Anaya Multimedia, Madrid, Spain, 2000. ISBN 84-415-1048-2. 351 pp. LCCN QA76.15 .A398 2000.

Dara-Abrams:2001:SWS

[DADAPW01] Benay Dara-Abrams, Drew Dara-Abrams, Trevor Peace, and Bebo White. Supporting Web Servers Interactive Workbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-085899-4. xxiii + 388 pp. LCCN TK5105.888 .S87 2001. US$44.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130858994.html.

Dangerfield:2002:CPF

[Dan02] Steven Dangerfield. Check Point Firewall-1 Administration and CCSA Study Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-093863-7. 370 pp. LCCN ???? US$54.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130938637.html.

Dasgupta:2001:MII

[Das01] Subhasish Dasgupta. Managing Internet and intranet technologies in organizations: challenges and opportunities. Idea Group Pub., Hershey, PA, USA, 2001. ISBN 1-878289-95-0 (paperback). iii + 224 pp. LCCN HD30.37 .M356 2001.

Davis:2000:ACG

[Dav00] James B. Davis. The Annual Consumer’s Guide to Health and Medicine on the Internet 2000. Health Information Press, Los Angeles, CA, USA, updated [ed.] edition, 2000. ISBN 1-885987-18-8 (paperback). x + 958 pp. LCCN W 26.5 A615 2000.

Davis:2001:HMI

[Dav01a] James B. Davis. Health and medicine on the Internet: a comprehensive guide to medical information on the World Wide Web. PMIC, Los Angeles, CA, USA, third edition, 2001. ISBN 1-57066-207-X. x + 610 pp. LCCN W 26.5 H432 2001.

Davis:2001:HPO

[Dav01b] Kate Davis. Health policy online: Internet 101. National Conference of State Legislatures, Denver, CO, USA, January 2001. ISBN 1-58024-136-0. v + 27 pp. LCCN Z674.75.I58 H43 2001 CRB. “January 2001.”.

Davenport:2002:VAI

[Dav02a] David Davenport. Viewpoint:

Anonymity on the Internet: why the price may be too high. Communications of the ACM, 45(4):33–35, April 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Davison:2002:LPL

[Dav02b] Andrew Davison. Logic programming languages for the Internet. Lecture Notes in Computer Science, 2407:66– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2407/24070066.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2407/24070066.
pdf.

Day:2008:PNA

[Day08] John D. Day. Patterns in Network Architecture: a Return to Fundamentals. Pearson Education, Upper Saddle River, NJ, USA, 2008. ISBN 0-13-225242-2 (hardcover). xxxi + 429 pp. LCCN TK5105.52 .D39 2008.

deBroglie:2001:DAI

[dB01] Gabriel de Broglie. Le droit d’auteur et l’internet: rapport du groupe de travail de l’Académie des sciences morales et politiques. (French) [Author rights and the Internet: report of the working group of the Academy of Moral and Political Sciences]. Cahier des sciences morales et politiques. Presses universitaires de France, Paris, France, 2001. ISBN 2-13-051468-5. viii + 105 pp. LCCN KJV3270 .D76 2001.

Divitini:2000:IBO

[DBS00] Monica Divitini, Terje Brasethvik, and David G. Schwartz, editors. Internet-based organizational memory and knowledge management. Idea Group Pub., Hershey, PA, USA, 2000. ISBN 1-878289-82-9 (paperback). LCCN HD30.2 .S39 2000.

Donahoo:2001:TIS

[DC01] Michael J. Donahoo and Kenneth L. Calvert. TCP/ IP sockets in C: practical guide for programmers. The Morgan Kaufmann practical guides series. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-826-5 (paperback). xv + 130 pp. LCCN QA76.76.A63 D66 2001b. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els031/2001095844.
html;   http://www.loc.gov/
catdir/toc/els031/2001095844.
html.

Downing:2000:DCI

[DCC00] Douglas Downing, Michael A. Covington, and Melody Mauldin Covington. Dictionary of computer and Internet terms. Barron’s, Hauppauge, NY, USA, seventh edition, 2000. ISBN 0-7641-1265-1. iv + 554 pp. LCCN QA76.15 .D667 2000. First-4th eds. published under title: Dictionary of computer terms.

Deitel:2000:IWW

[DDN00] Harvey M. Deitel, Paul J. Deitel, and T. R. Nieto. Internet and World Wide Web: how to program. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-016143-8. xliv + 1157 pp. LCCN QA76.625 .D45 2000.

Deitel:2002:IWW

[DDN02] Harvey M. Deitel, Paul J. Deitel, and T. R. Nieto. Internet and World Wide Web: how to program. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2002. ISBN 0-13-030897-8. lvii + 1428 pp. LCCN QA76.625 .D47 2002. CD-ROM contains Microsoft’s Internet Explorer 5.5 and Agent 2.0, Adobe Acrobat Reader 5.0, MySQL 3.23, (90day evaluation version), Active State Languages and more.

Decasper:2000:RPS

[DDPP00] Dan Decasper, Zubin Dittia, Guru Parulkar, and Bernhard Plattner. Router plugins: a software architecture for next-generation routers. IEEE/ACM Transactions on Networking, 8(1):2– 15, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        ton/2000-8-1/p2-decasper/

.

DeMuth:2002:DEI

[DE02] Brian DeMuth and Dan Eisenreich. Designing Embedded Internet Devices. Butterworth-Heinemann, Boston, MA, USA, 2002. ISBN 1-878707-98-1. xii + 582 pp. LCCN ???? US$59.95.

Dean:2000:BYO

[Dea00] Tony Dean. Building your own Internet site. Linux Journal, 72:??, April 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

deCarmo:2000:MGC

[deC00] Linden deCarmo. The Media Gateway Control Protocol. Dr. Dobbs Journal, 25(5):23– 24, 26–28, 32, May 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Deitel:2001:WIM

[Dei01a] Harvey M. Deitel. Wireless Internet and mobile business: how to program. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-062226-5. lv + 1328 pp. LCCN QA76.625 .W58 2001.

Deitel:2001:XHP

[Dei01b] Harvey M. Deitel. XML: how to program. How to program series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-028417-3. xxxviii + 934 pp. LCCN QA76.76.H94 X45 2001. Includes CD-ROM.

Denning:2002:IT

[Den02] Peter Denning. Internet time out. Communications of the ACM, 45(3):15–18, March 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dutt:2000:GBG

[DF00] Christoph Dutt and Joachim Freiburg. GiMP: Bilder gestalten, Fotos retuschieren; [Grundlagen der professionellen Bildbearbeitung, der Umgang mit Fotos, Grafiken und Text, Bilder f?r das Internet richtig vorbereiten; auf der CD: GIMP f?r Windows, SCO Unix, Debian GNU Linux, Solaris, OS/2 und BSD, Quelltext aller GIMP- und GTK-Versionen, ?ber 300 Plugins in C, Perl, tcl, Phyton und Scheme, XFree86/23.3.6, GIMP User Manual als PDF-Dateien]. C und L, Böblingen, Germany, 2000. ISBN 3-932311-64-7. 522 + 98 pp. LCCN ???? Includes CD-ROM.

Diamond:2001:WUP

[DF01] Wayne J. Diamond and Richard B. Freeman. Will unionism prosper in cyberspace?: the promise of the Internet for employee organization. NBER working paper series 8483, National Bureau of Economic Research, Cambridge, MA, USA, September 2001. 41 pp.

Douglis:2003:IIC

[DF03] Fred Douglis and Ian Foster. IEEE Internet computing: Guest Editors’ introduction: The Grid grows up. IEEE Distributed Systems Online, 4(7), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0307/0307/f/w4gei.htm.

Djilali:2004:RVT

[DHL^{+}04] Samir Djilali, Thomas Herault, Oleg Lodygensky, Tangui Morlier, Gilles Fedak, and Franck Cappello. RPC-V: Toward fault-tolerant RPC for Internet connected desktop grids with volatile nodes. In ACM [ACM04], page 39. ISBN 0-7695-2153-3. LCCN ????

Diffley:2002:IPE

[Dif02]  Mary Diffley.  Internet

prophets: enlightened Ebusiness strategies for every budget. CyberAge Books, Medford, NJ, USA, 2002. ISBN 0-910965-55-2. xxiii + 342 pp. LCCN HF5548.32 .D534 2002.

Dillman:2000:MIS

[Dil00] Don A. Dillman. Mail and Internet surveys: the tailored design method. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 2000. ISBN 0-471-32354-3. xvi + 464 pp. LCCN HM538 .D55 2000. Rev. ed. of: Mail and telephone surveys. c1978.

Dang:2002:PVU

[DJK02] Luan Dang, Cullen Jennings, and David Kelly. Practical VoIP Using VOCAL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00078-2. xvii + 502 pp. LCCN TK5105.8865 .D35 2002. US$44.95. URL http:/

/www.oreilly.com/catalog/
voip.

Laubier:2000:PI

[dL00] Charles de Laubier. La presse sur Internet. (French) [The Press on the Internet], volume 3582 of Que sais-je?; 3582 4007463854. Presses universitaires de France, Paris, France, 2000. ISBN 2-13-051339-5. 127 pp. LCCN PN4784.E5 L28 2000.

DiGeorgio:2001:CTG

[DL01] Marvin DiGeorgio and Sylvia Lesage. The 21st century teachers’ guide to recommended Internet sites. Neal-Schuman netguide series. Neal-Schuman Publishers, New York, NY, USA, 2001. ISBN 1-55570-401-8 (paperback). xix + 283 pp. LCCN ZA4201.D54 2001.

Durand:2002:SRU

[DL02] Nicolas Durand and Luigi Lancieri. Study of the regularity of the users’ Internet accesses. Lecture Notes in Computer Science, 2412:173– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2412/24120173.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2412/24120173.
pdf.

Demaine:2002:FEI

[DLOM02] Erik D. Demaine, Alejandro López-Ortiz, and J. Ian Munro. Frequency estimation of Internet packet streams with limited space. Lecture Notes in Computer Science, 2461:348–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer-ny.com/link/service/
series/0558/bibs/2461/24610348.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2461/24610348.
pdf.

Delta:2002:LI

[DM02] George B. Delta and Jeffrey H. Matsuura. Law of the Internet. Aspen Law and Business, New York, NY, USA, second edition, 2002. ISBN 0-7355-2219-7 (loose-leaf). various pp. LCCN KF390.5.C6 D45 2002.

Davis:2000:DEU

[DMP00] Kelly Davis, Tom Miller, and Charles Price. Distance education using Linux and the MBone. Linux Journal, 78: ??, October 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Dochartaigh:2002:IRH

[Doc02] Niall O. Dochartaigh. The Internet research handbook: a practical guide for students and researchers in the social sciences. SAGE, London, UK, 2002. ISBN 0-7619-6439-8, 0-7619-6440-1 (paperback). xx + 274 pp. LCCN H61.95 .O23 2002. UK£16.99.

Donahoo:2003:TIS

[Don03] Michael Donahoo. TCP/IP Sockets in C: Practical Guide for Programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-826-5. xv + 130 pp. LCCN QA76.76.A63 D66 2001b. US$15.95.

DCC:2001:DCB

[Dot01] Dot-Com Consulting. Dotcom and beyond: breakthrough Internet-based architectures and methodologies. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-062297-4. xvii + 314 pp.

LCCN TK5105.5.D68 2001. “A Prentice Hall title.”.

Dutta:2003:AID

[DR03] Amitava Dutta and Rahul Roy. Anticipating Internet diffusion. Communications of the ACM, 46(2):66–71, February 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Draetta:2001:ICE

[Dra01] Ugo Draetta. Internet e commercio elettronico: nel diritto internazionale e dei privati. Contratti and commercio internazionale. Giuffre Editore, Milano, Italy, 2001. ISBN 88-14-08998-1. xii + 224 pp. LCCN K1005 D73 2001.

Drews:2000:MNI

[Dre00] Jorg Drews. Mitwelt, Nachwelt, Internet: Vortrage und Materialien zur Rezeption Goethes zwischen 1800 und 2000. (German) [Fellow man, future generations, Internet: Lecture and materials for the reception of Goethe between 1800 and 2000], volume 14 of Bielefelder Schriften zu Linguistik und Literaturwissenschaft; Bd. 14. Aisthesis, Bielefeld, Germany, 2000. ISBN 3-89528-299-5 (paperback). 172 pp. LCCN PT2165 .M58 2000.

Dreyfus:2001:I

[Dre01a] Hubert L. Dreyfus. On the Internet. Thinking in action.

Routledge & Kegan Paul, London, UK and New York, NY, USA, 2001. ISBN 0-415-22806-9, 0-415-22807-7 (paperback). ix + 127 pp. LCCN HM851 .D74 2001. Contents: Introduction — The hype about hyperlinks — How far is distance learning from education? — Disembodied telepresence and the remoteness of the real — Nihilism on the Information Highway: anonymity vs. commitment in the present age — Conclusion.

Dreyfus:2001:MND

[Dre01b] Nathalie Dreyfus. Marques et noms de domaine de l’internet. (French) [Trademarks and domain names of the Internet]. Hermes Science Publications, Paris, France, 2001. ISBN 2-7462-0287-5. 254 pp. LCCN ????

Dhanjani:2009:HNG

[DRH09] Nitesh Dhanjani, Billy Rios, and Brett Hardin. Hacking: the next generation. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2009. ISBN 0-596-80630-2. xiii + 279 pp. LCCN ????

DMH:2000:DW

[DRI00]  DRI/McGraw-Hill.  DRI

Webstract. DRI/McGraw-Hill, Lexington, MA, USA, 2000. LCCN HC59.15. URL

http://sun3.lib.uci.edu/~
dtsang/ssda/driweb.html.

Dominick:2000:BCI

[DSM00] Joseph R. Dominick, Barry L. Sherman, and Fritz Messere. Broadcasting, cable, the Internet, and beyond: an introduction to modern electronic media. McGraw-Hill, New York, NY, USA, fourth edition, 2000. ISBN 0-07-290441-0 (softcover). xvii + 330 pp. LCCN HE8689.8 .D66 2000. Rev. ed. of: Broadcasting/cable and beyond. 3rd ed. c1996.

Dally:2003:PPI

[DT03] William Dally and Brian Towles. Principles and Practices of Interconnection Networks. Academic Press, New York, NY, USA, 2003. ISBN 0-12-200751-4. ???? pp. LCCN ???? US$59.95.

Dudeney:2000:ILC

[Dud00] Gavin Dudeney. The Internet and the language classroom. Cambridge handbooks for language teachers. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-78373-9 (paperback). x + 181 pp. LCCN P53.285 .D83 2000.

Dudney:2008:CAM

[Dud08] Bill Dudney. Core animation for Mac OS X and the iPhone: Creating compelling dynamic user interface. Pragmatic Bookshelf, Raleigh, NC, USA, 2008. ISBN 1-934356-10-7. ix + 182 pp. LCCN ????

Durrant:2002:GIC

[Dur02] Fay Durrant. e-government and the Internet in the Caribbean: An initial assessment. Lecture Notes in Computer Science, 2456:101– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2456/24560101.htm;
http://link.springer.de/
link/service/series/0558/
papers/2456/24560101.pdf.

Dietrich:2002:MIA

[DV02] Brenda Dietrich and Rakesh V.

Vohra, editors. Mathematics of the Internet: e-auction and markets, volume 127 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95359-0. LCCN HF5478 .M38 2002. Based on the proceedings of a workshop held in December 2000.

Dwivedi:2002:BRO

[Dwi02] Neelam Dwivedi. Book review: One-stop reference for networking fundamentals: a review of Computer Networks and Internets with Internet Applications, 3rd ed. by Douglas E. Comer. IEEE Distributed Systems Online, 3(10), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0210/d/bks_b.htm.

Dwivedi:2009:HVP

[Dwi09] Himanshu Dwivedi. Hacking VoIP: protocols, attacks, and countermeasures. No Starch Press, San Francisco, CA, USA, 2009. ISBN 1-59327-163-8. xii + 211 pp. LCCN TK5105.8865 .P37

                                        2009. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0913/2008038559-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0915/2008038559-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/ecip0827/
                                        2008038559.html.

Dai:2002:IBC

[DWX02] Kaiyu Dai, Yinglin Wang, and Xiqing Xu. An Internetbased conference system for real-time distributed design evaluation. Lecture Notes in Computer Science, 2480:499– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2480/24800499.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2480/24800499.
                                        pdf.

Duato:2003:INE

[DYN03] José Duato, Sudhakar Yalamanchili, and Lionel Ni. Interconnection Networks: An Engineering Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-852-4. xxiii + 600 pp. LCCN TK5105.5 .D78 2003. US$59.95.

Eady:2003:NIM

[Ead03] Fred Eady. Networking and Internetworking with Microcontrollers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-7506-7698-1. ???? pp. LCCN ???? US$59.95.

Earle:2006:WSH

[Ear06] Aaron E. Earle. Wireless security handbook. Auerbach, Boca Raton, FL, USA, 2006. ISBN 0-8493-3378-4. xv + 366 pp. LCCN TK5105.78 .E23 2006eb; TK5105.78 .E23 2006.

Ein-Dor:2000:IPM

[EDGW00] Phillip Ein-Dor, Seymour E.

Goodman, and Peter Wolcott. International perspectives: From Via Maris to electronic highway: the Internet in Canaan. Communications of the ACM, 43(7):19–23, July 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-7/p19-ein-dor/.

Edwards:2002:INA

[Edw02] Margaret J. A. Edwards. The Internet for nurses and allied health professionals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2002. ISBN 0-387-95236-5 (softcover). ix + 138 pp. LCCN WY26.5 E26i 2002.

Estrin:2000:EII

[EGH00] Deborah Estrin, Ramesh Govindan, and John Heidemann. Embedding the Internet: introduction. Communications of the ACM, 43(5):38, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-5/p38-estrin/.

Eisenmann:2002:IBM

[EHT02] Thomas R. Eisenmann, Roger Hallowell, and Mary Tripsas. Internet business models: texts and cases. Mc-Graw-Hill/Irwin, Boston, MA, USA, 2002. ISBN 0-07-249868-4, 0-07-112363-6 (international ed.), 0-07-250672-5 (disc). xviii + 650 pp. LCCN HD9696.8.A2 I57 2002. Title on disc: Teledesic case to accompany Internet business models.

Enzmann:2002:NIU

[EKS02] Matthias Enzmann, Thomas Kunz, and Markus Schneider. A new infrastructure for user tracking prevention and privacy protection in Internet shopping. Lecture Notes in Computer Science, 2437:199– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2437/24370199.htm;
http://link.springer.de/
link/service/series/0558/
papers/2437/24370199.pdf.

Elfenbein:2001:LHE

[EL01] Dan Elfenbein and Joshua Lerner. Links and hyperlinks: an empirical analysis of Internet portal alliances, 1995– 1999. NBER working paper series; no. 8251 4007554648 8251, National Bureau of Economic Research, Cambridge, MA, USA, April 2001. 34 + [27] pp.

Elliott:2000:GSI

[Ell00]  Alan C. Elliott.  Getting

started in Internet auctions. The getting started in series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2000. ISBN 0-471-38087-3 (paper). xv + 234 pp. LCCN HF5478 .E45 2000.

Emshwiller:2000:SDM

[Ems00] John R. Emshwiller. Scam dogs and mo-mo Mamas: inside the wild and woolly world of Internet stock trading. HarperBusiness, New York, NY, USA, 2000. ISBN 0-06-019620-3 (hardcover). xxiv + 295 pp. LCCN HG4515.95 .E47 2000.

Engh:2000:SCY

[Eng00] Rohn Engh. Sellphotos.com: your guide to establishing a successful stock photography business on the Internet. Writer’s Digest Books, Cincinnati, OH, USA, 2000. ISBN 0-89879-944-9. 276 pp. LCCN TR690 .E534 2000.

Englander:2002:JS

[Eng02]  Robert Englander.  Java

and SOAP. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00175-4. xv + 258 pp. LCCN QA76.76.H94 E54 2002. US$39.95. URL http://

                                        safari.oreilly.com/0596001754;
                                         http://www.oreilly.com/
                                        catalog/javasoap.

Engels:2006:BFW

[EP06] Marc Engels and Frederik Petré. Broadband fixed wireless access: a system perspective. Signals and communication technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-33956-6. xviii + 211 pp. LCCN TK5103.2 .E57 2006. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        fy0708/2006926213.html.

Ellis:2003:VVD

[EPR03] Juanita Ellis, Charles Pursell, and Joy Rahman. Voice, Video, and Data Network Convergence: Architecture and Design, From VoIP to Wireless. Academic Press, New York, NY, USA, 2003. ISBN 0-12-236542-9. ???? pp. LCCN TK5101 .E48 2003. US$59.99.

Elling:2002:DES

[ER02] Richard Elling and Tim Read.

Designing Enterprise Solutions with Sun Cluster 3.0. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-008458-1. 302 (est.) pp. LCCN ???? URL http:

//www.phptr.com/ptrbooks/
ptr_0130084581.html.

Erbschloe:2003:SRI

[Erb03] Michael Erbschloe. Socially responsible IT management. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2003. ISBN 1-55558-290-7. xviii + 362 pp. LCCN HD30.2.E73 2003. US$39.99.

Ellis:2001:ISG

[ES01] Juanita Ellis and Tim Speed.

The Internet security guidebook: from planning to deployment. The Korper and Ellis E-Commerce Books Series. Academic Press Inc., New York, NY, USA, 2001. ISBN 0-12-237471-1. xxiv + 320 pp. LCCN TK5105.875.I57 E46 2001.

Ellis:2003:ISJ

[ES03] Juanita Ellis and Tim Speed.

Internet Security: a Jumpstart for Systems Administrators and IT Managers. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2003. ISBN 1-55558-298-2. 207 (est.) pp. LCCN ???? US$44.99.

Ebner:2000:FMF

[EST00] Ezra Ebner, Weiguang Shao, and Wei-Tek Tsai. The fivemodule framework for Internet application development. ACM Computing Surveys, 32 (1es), March 2000. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
surveys/2000-32-1/p40-ebner/

. Article No. 40.

Estevez-Tapiador:2003:BRC

[ET03] Juan M. Estevez-Tapiador. Book review: Concepts and attitudes for Internet security: a review of Firewalls and Internet Security: Repelling the Wily Hacker, second edition by William R. Cheswick, Steven M. Bellovin, and Aviel D. Rubin. IEEE Distributed Systems Online, 4(10), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0310/d/bks_b.htm.

Euroconfidential:2000:GEI

[Eur00] Euroconfidential (Firm). The guide to EU information sources on the Internet. Euroconfidentiel, Genval, Belgium, 2000. ISBN 2-930066-54-7. 424 pp. LCCN JN27 .G85 2000.

Everard:2000:VSI

[Eve00] Jerry Everard. Virtual states:

the Internet and the boundaries of the nation state. Technology and the global political economy. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2000. ISBN 0-415-17213-6, 0-415-17214-4 (paperback). xviii + 174 pp. LCCN HM851 .E94 2000.

Edwards:2000:LIF

[EW00] Lilian Edwards and Charlotte Waelde. Law and the Internet: a framework for electronic commerce. Hart, Portland, OR, USA, second edition, 2000. ISBN 1-84113-141-5 (paperback). xxxvii + 396 pp. LCCN K564.C6 L37 2000. Earlier edition published under: Law and the Internet: regulating cyberspace (Hart, 1997).

Fallon:2000:IT

[Fal00] Thomas J. Fallon. The Internet today. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-010139-7. x + 277 pp. LCCN TK5105.875.I57 F354 2000.

Farace:2001:PIH

[Far01] Joe Farace. The photographer’s Internet handbook. Allworth Press, New York, NY, USA, rev. ed edition, 2001. ISBN 1-58115-084-9 (paperback). xv + 215 pp. LCCN TR11 .F37 2001.

Farrel:2003:IPC

[Far03] Adrian Farrel. The Internet and Its Protocols: a Comparative Approach. Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-913-X. xxix + 809 pp. LCCN TK5105.55 .F37

2004.  US$49.95.  URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/els051/2003027812.
                                        html;   http://www.loc.gov/
                                        catdir/toc/ecip0412/2003027812.
                                        html.

Finke:2004:RAS

[FB04] Matthias Finke and Dirk Balfanz. A reference architecture supporting hypervideo content for ITV and the Internet domain. Computers and Graphics, 28(2):179–191, April 2004. CODEN COGRD2. ISSN 0097-8493.

Farrel:2005:GAA

[FB05] Adrian Farrel and Igor Bryskin. GMPLS: Architecture and Applications. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2005. ISBN 0-12-088422-4. 400 (est.) pp. LCCN ???? US$49.95.

Feitsma:2001:EIT

[Fei01] E. F. Feitsma. Effectenverkeer via Internet: een tussenstand vanuit effectenrechtelijk perspectief, volume 66 of Vennootschaps- en rechtspersonenrecht. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 90-268-3830-1. ix + 89 pp. LCCN KE2721 .V45 v.66.

Ferdinand:2000:IDD

[Fer00] Peter Ferdinand. The Internet, democracy and democratization. Democratization studies. ISSN: 1465-4601. Frank Cass, Portland, OR, USA, 2000. ISBN 0-7146-5065-X (hardcover), 0-7146-8114-8 (paperback). 193 pp. LCCN JC421 .I57 2000. This group of studies first appeared in a special issue … Democratization. Contents: The Internet, Democracy and Democratization / Peter Ferdinand — Paradoxical Partners: Electronic Communication and Electronic Democracy / Beth Simone Noveck — Democratizing Democracy: Strong Democracy, US Political Campaigns and the Internet / Jennifer Stromer-Galley — Revitalizing the Party System or Zeitgeist-on-Line? Virtual Party Headquarters and Virtual Party Branches in Germany / Christoph Bieber — The Politics of African America On-Line / Rohit Lekhi — Neo-Nazis and Taliban On-Line: Anti-Modern Political Movements and Modern Media / Peter Chroust — The Internet in Indonesia’s New Democracy / David T. Hill and Krishna Sen — The Electronic Republic? The Role of the Internet in Promoting Democracy in Africa / Dana Ott and Melissa Rosser — Tibet, Democracy and the Internet Bazaar / John Bray — Conclusion / Peter Ferdinand.

Ferdinandi:2002:RPS

[Fer02] Patricia L. Ferdinandi. A requirements pattern: succeeding in the Internet economy. Addison-Wesley information technology series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-73826-0 (paperback). xviii + 506 pp. LCCN QA76.76.D47 F49 2002.

Fries:2000:MIA

[FF00] Bruce Fries and Marty Fries. The MP3 and Internet audio handbook: your guide to the digital music revolution. TeamCom Books, Silver Spring, MD, USA, 2000. ISBN 1-928791-10-7. xvi + 302 pp. LCCN ML74.4.M6 .F75 2000.

Fensel:2003:SSW

[FHLW03] Dieter Fensel, James Hendler, Henry Lieberman, and Wolfgang Wahlster, editors. Spinning the Semantic Web: bringing the World Wide Web to its full potential. MIT Press, Cambridge, MA, USA, 2003. ISBN 0-262-06232-1, 0-262-56212-X. xxiii + 479 pp. LCCN ???? Foreword by Tim Berners-Lee.

Fiander:2000:CIF

[FHM00] P. Michelle Fiander, Joseph C.

Harmon, and Jonathan David Makepeace, editors. From Carnegie to Internet2: forging the serials future: proceedings of the North American Serials Interest Group, Inc.: 14th annual conference, June 10–13, 1999, Carnegie Mellon University, Pittsburgh, Pennsylvania. Haworth Information Press, New York, NY, USA, 2000. ISBN 0-7890-1007-0, 0-7890-1035-6. LCCN Z692.S5 N67 1999. Co-published simultaneously as The serials librarian, v. 38, no. 1/2 and 3/4, 2000.

FCIA:2001:FCS

[Fib01] Fibre Channel Industry Association. Fibre Channel Storage Area Networks. LLH Technology Publishing, Eagle Rock, VA, USA, 2001. ISBN 1-878707-97-3. ???? pp. LCCN ???? US$19.99.

Fielden:2001:IRT

[Fie01] Ned L. Fielden. Internet research: theory and practice. McFarland, Jefferson, NC, USA, second edition, 2001. ISBN 0-7864-0538-4 (softcover), 0-7864-1099-X (paperback). x + 205 pp. LCCN ZA4201 .F54 2001. URL

http://userwww.sfsu.edu/~
fielden/internetresearch.

htm. Updates and errata available on the World Wide Web.

Fitzgerald:2000:GDL

[Fit00] Anne Fitzgerald. Going digital 2000: legal issues for ecommerce, software and the Internet. Prospect Media, St. Leonards, NSW, Australia, second edition, 2000. ISBN 1-86316-150-3. xviii + 420 pp. LCCN K564.C6 G64 2000.

Francis:2001:IGI

[FJJ^{+}01] Paul Francis, Sugih Jamin, Cheng Jin, Yixin Jin, Danny Raz, Yuval Shavitt, and Lixia Zhang. IDMaps: a global Internet host distance estimation service. IEEE/ACM Transactions on Networking, 9(5): 525–540, October 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fromm:2001:RII

[FK01] Rainer Fromm and Barbara Kernbach. Rechtsextremismus im Internet: die neue Gefahr. (German) [Right-wing extremism in the Internet: the new danger]. Olzog, München, Germany, 2001. ISBN 3-7892-8055-0 (paperback). 288 pp. LCCN JC481 .F76 2001.

Foster:2003:GBN

[FK03] Ian Foster and Carl Kesselman. The Grid: Blueprint for a new computing infrastructure. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-933-4. 800 (est.) pp. LCCN ???? US$59.95.

Foster:2004:GBN

[FK04] Ian Foster and Carl Kesselman. The Grid 2: Blueprint for a new computing infrastructure. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2004. ISBN 1-55860-933-4. xxvii + 748 pp. LCCN QA76.9.C58 G75 2004. US$59.95.

Feldman:2000:NSD

[FKC00] Alan Feldman, Cliff Konold, and Bob Coulter. Network science, a decade later: the Internet and classroom learning. Lawrence Erlbaum Associates, Mahwah, NJ, USA, 2000. ISBN 0-8058-3425-7, 0-8058-3426-5 (paperback). xxiv + 186 pp. LCCN LB1583.3 .N48 2000. This project was undertaken by TERC, Inc., Cambridge, Massachusetts.

Fabrikant:2002:HOT

[FKP02] Alex Fabrikant, Elias Koutsoupias, and Christos H. Papadimitriou. Heuristically optimized trade-offs: a new paradigm for power laws in the Internet. Lecture Notes in Computer Science, 2380:110– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2380/23800110.
htm;   http://link.springer-
ny.com/link/service/series/

                                        0558/papers/2380/23800110.
                                        pdf.

Freytag:2003:VPI

[FLA^{+}03] Johann Christoph Freytag, Peter C. Lockemann, Serge Abiteboul, Michael J. Carey, Patricia G. Selinger, and Andreas Heuer, editors. VLDB 2003: Proceedings of 29th International Conference on Very Large Data Bases, September 9–12, 2003, Berlin, Germany. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-12-722442-4. LCCN ????

Flinn:2000:GBO

[Fli00] Cherri Melton Flinn. Genealogy basics online: a stepby-step introduction to finding your ancestors through the Internet. Muska and Lipman, Cincinnati, OH, USA, 2000. ISBN 1-929685-00-9. x + 212 pp. LCCN CS21 .F55 2000.

Flor:2001:WBE

[Flo01] Nick V. Flor. Web business engineering: using offline activities to drive Internet strategies. Addison-Wesley information technology series. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-60468-X (paperback). xxiii + 219 pp. LCCN HF5548.32 F58 2000.

Flynn:2001:PHD

[Fly01] Nancy Flynn. The e-policy handbook: designing and implementing effective e-mail, Internet, and software policies. AMACOM, New York, NY, USA, 2001. ISBN 0-8144-7091-2 (paperback). xviii + 285 pp. LCCN HE7551 .F58 2001.

Fitzpatrick:2000:IRN

[FM00] Joyce J. Fitzpatrick and Kristen S. Montgomery. Internet resources for nurses. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-8261-1371-0 (softcover). xvii + 250 pp. LCCN WY 26.5 I61-2 2000.

Fairhurst:2005:IMT

[FM05] Gorry Fairhurst and Marie-José Montpetit. Internet over MPEG-2 transmission networks. Computer Networks (Amsterdam, Netherlands: 1999), 48(1):1–3, May 16, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Ferris:2000:NCS

[FMM00] Michael C. Ferris, Michael P.

Mesnier, and Jorge J. Moré. NEOS and Condor: solving optimization problems over the Internet. ACM Transactions on Mathematical Software, 26(1):1–18, March 2000. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://doi.acm.org/
10.1145/347837.347842;
http://www.acm.org/pubs/
citations/journals/toms/
2000-26-1/p1-ferris/;   http:

                                        //www.acm.org/pubs/citations/
                                        journals/toms/2000-26-1/
                                        p1-ferris/p1-ferris.pdf.

Farhoomand:2003:CSC

[FN03] Ali F. Farhoomand and Pauline S. P. Ng. Creating sustainable competitive advantage through internetworked communities. Communications of the ACM, 46(9):83– 88, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fair:2005:CSH

[FNRC05] Ted Fair, Michael Nordfelt, Sandra Ring, and Eric Cole. Cyber spying: tracking your family’s (sometimes) secret online lives. Syngress Publishing, Inc., Rockland, MA, USA, 2005. xxiii + 439 pp. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://site.ebrary.com/
                                        lib/ucsc/Doc?id=10074965.

Focazio:2001:FBC

[Foc01]  Martin T. Focazio.  The

e-factor: building a 24/7 customer-centric, electronic business for the Internet age. AMACOM/American Management Association, New York, NY, USA, 2001. ISBN 0-8144-0489-8. viii + 242 pp. LCCN HF5548.32 .F63 2001.

Floyd:2001:DSI

[FP01] Sally Floyd and Vern Paxson. Difficulties in simulating the Internet. IEEE/ACM Transactions on Networking, 9(4):

392–403, August 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Foster:2005:SSP

[FPM05] James C. Foster, Mike Price, and Stuart McClure. Sockets, shellcode, porting and coding: reverse engineering exploits and tool coding for security professionals. Syngress Publishing, Inc., Rockland, MA, USA, 2005. ISBN 1-59749-005-9 (paperback). xxviii + 667 pp. LCCN QA76.9.A25 F65 2004.

FDAP:2000:BPL

[Fra00] France. Délégation des arts plastiques. Bibliographie pour l’art contemporain: [250 titres, ouvrages généraux, philosophie, esthétique et critique, domaines d’expression plastiques, collections, expositions, jeunes lecteurs, site Internet, périodiques, éditeurs]. Guides pour l’art contemporain; 2. Editions 00h00, Paris, France, 2000. ISBN 2-7454-0304-4. 98 pp. LCCN Z5935.5 .B52 2000. At head of title: Ministère de la culture et de la communication, Délégation aux arts plastiques.

Franda:2001:GIE

[Fra01a] Marcus F. Franda. Governing the Internet: the emergence of an international regime. Ipolitics. L. Rienner Publishers, Boulder, CO, USA, 2001.

ISBN 1-55587-999-3. xi + 255 pp. LCCN K564.C6 F73 2001.

Frank:2001:NSI

[Fra01b]  Stephen E. Frank.  Net-

worth: successful investing in the companies that will prevail through Internet booms and busts (they’re not always the ones you expect). Simon and Schuster, 1230 Ave. of the Americas, New York, NY 10020, USA, 2001. ISBN 0-7432-1093-X. 236 pp. LCCN HD9696.8.A2 F73 2001. “A Wall Street Journal Book.”.

Frankel:2001:FCI

[Fra01c] Yair Frankel, editor. Financial cryptography: 4th international conference, FC 2000, Anguilla, British West Indies, February 20–24, 2000: proceedings, volume 1962 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42700-7 (softcover). LCCN QA267.A1 L43 no.1962; HG1710 .F35

                                        2000. URL http://link.
                                        springer-ny.com/link/service/
                                        series/0558/tocs/t1962.htm.

Franda:2002:LCI

[Fra02] Marcus F. Franda. Launching into cyberspace: Internet development and politics in five world regions. Ipolitics. Lynne Rienner Publishers, Boulder, CO, USA, 2002. ISBN 1-58826-012-7, 1-58826-037-2 (paperback). xiv + 297 pp. LCCN JZ1254 .F73 2002.

Francis:2003:IIC

[Fra03a]  Paul Francis.  IEEE In-

ternet computing: Peer to peer: Is the Internet going NUTSS? IEEE Distributed Systems Online, 4(11), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0311/f/w6p2pp.htm.

Franklin:2003:NPDa

[Fra03b] Mark Franklin. Network Processor Design: Issues and Practices. Academic Press, New York, NY, USA, 2003. ISBN 0-12-198157-6. ???? pp. LCCN ???? US$64.95.

FF:2001:RNW

[Fre01] Freedom Forum. Report of the National Workshop on Internet Voting: issues and research agenda. Technical report, Internet Policy Institute, Washington, DC, USA, March 2001. iii + 52 pp. URL http://www.

netvoting.org/resources/
internetvotingreport.pdf.

Frieden:2001:MID

[Fri01]  Rob Frieden.  Managing

Internet-driven change in international telecommunications. The Artech House telecommunications library. Artech House Inc., Norwood, MA, USA, 2001. ISBN 1-58053-019-2. xxii + 480 pp. LCCN HE7645 .F744 2001.

Feral-Schuhl:2000:CDE

[FS00] Christiane Feral-Schuhl. Cyberdroit: le droit à l’épreuve de l’Internet. (French) [Cyberlaw: Law under the test of the Internet]. Dunod, Paris, France, deuxième edition, 2000. ISBN 2-10-005285-3. xii + 323 pp. LCCN KJV333.C65 F47 2000.

Fan:2000:IFF

[FSW00] Ming Fan, Jan Stallaert, and Andrew B. Whinston. The Internet and the future of financial markets. Communications of the ACM, 43(11):82– 88, November 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
cacm/2000-43-11/p82-fan/.

Fenoll-Trousseau:2000:IPD

[FTH00] Marie-Pierre Fenoll-Trousseau and Gerard Haas. Internet et protection des données personnelles. (French) [Internet and Protection of Personal Data]. Decouvrir. Litec, Paris, France, 2000. ISBN 2-7111-3179-3. 206 pp. LCCN KJV2094 .F46 2000.

Furht:2000:HIC

[Fur00] Borivoje Furht. Handbook of Internet computing. Internet and communications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, November 2000. ISBN 0-8493-0086-X. 532 pp. LCCN TK5105.875.I57 F88 2000. US$99.95. Also available via the World Wide Web in the database, ENGnet-BASE engineering handbooks online.

Freund:2000:EII

[FW00] Caroline L. Freund and Diana Marie Weinhold. On the effect of the Internet on international trade. International finance discussion papers 693, Board of Governors of the Federal Reserve System, Washington, DC, USA, December 2000. 44 pp.

Flickenger:2006:WH

[FW06] Rob Flickenger and Roger Weeks. Wireless hacks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2006. ISBN 0-596-10144-9. xxii + 440 pp. LCCN TK5103.2 .F55 2006.

URL http://www.loc.gov/
catdir/toc/fy0606/2006272661.
html.

Gabrovec:2001:IQMb

[Gab01a]  Judy Gabrovec.  Internet

quests: mammals. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3411-3. 48 pp. LCCN QL50 .G33 2001.

Gabrovec:2001:IQMa

[Gab01b]  Judy Gabrovec.  Internet

quests: my body. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3403-2. 48 pp. LCCN QP434 .G33 2001.

Gabrovec:2001:IQO

[Gab01c]  Judy Gabrovec.  Internet

Quests: oceans. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3402-4. 48 pp. LCCN QH91 .G32 2001. Primary grades.

Gabrovec:2001:IQW

[Gab01d]  Judy Gabrovec.  Internet

quests: weather. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3407-5. 48 pp. LCCN QC981.3 .G33 2001.

Gagne:2000:LES

[Gag00] Marcel Gagné. A Linux email server. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):53–58, February 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Nicholas:2002:CID

[Gal02a] Nicholas Galbreath. Cryptography for Internet and database applications: developing secret and public key techniques with Java. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-21029-3. 400 pp. LCCN QA76.9.A25 G35 2002.

Gallo:2002:NE

[Gal02b] Michael Gallo. Networking Explained. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 2002. ISBN 1-55558-252-4. xvi + 672 pp. LCCN TK5105.5 .G35 2002. US$54.95.

Gao:2001:IAS

[Gao01] Lixin Gao. On inferring autonomous system relationships in the Internet. IEEE/ACM Transactions on Networking, 9 (6):733–745, December 2001. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Garber:2000:TND

[Gar00a]  Lee Garber.  Technology

news: Denial-of-service attacks rip the Internet. Computer, 33(4):12–17, April 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co2000/pdf/r4012.
pdf.

LagaresGarcia:2000:IDT

[Gar00b] Diego Lagares Garcia. Internet y el derecho: tecnologia y jurisprudencia, dos conceptos obligados a entenderse. (Spanish) [Internet and the law: technology and jurisprudence, two concepts to be understood]. Ediciones Carena, Barcelona, Spain, 2000. ISBN 84-88944-57-8. 145 pp. LCCN KE3656 .L34 2000.

Gartz:2000:ALO

[Gar00c]  Joachim Gartz.  America

Latina online: teoria y practica de eficientes pesquisas de Internet en El Dorado virtual. IHILA, Cologne, Germany, 2000. ISBN 980-07-6535-2. 144 pp. LCCN F1408 .G37 2000. Translated from the German edition published by A. Monch, Cologne, Germany, 1999.

Garzaniti:2000:TBI

[Gar00d] L. J. H. F. (Laurent J. H. F.) Garzaniti. Telecommunications, broadcasting, and the Internet: E.U. competition law and regulation. Sweet and Maxwell, London, UK, 2000. ISBN 0-421-65150-4. xxxiv + 365 pp. LCCN KJE6964 .G37 2000.

Gattiker:2001:IDC

[Gat01] Urs E. Gattiker. The Internet as a diverse community: cultural, organizational, and political issues. Telecommunications. Lawrence Erlbaum Associates, Hillsdale, NJ, 2001. ISBN 0-8058-2489-8 (paperback), 0-8058-2488-X. xxiv + 272 pp. LCCN K564.C6 G38 2001.

Goolsbee:2000:DIM

[GB00] Austan Goolsbee and Jeffrey R. Brown. Does the Internet make markets more competitive?: evidence from the life insurance industry. ABF working paper 2018, American Bar Foundation, Chicago, IL, USA, 2000. 28 pp.

Gast:2004:RIS

[GB04] Jim Gast and Paul Barford.

Representing the Internet as a succinct forest. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):35– 44, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gribble:2000:PDD

[GBCH00] Steven D. Gribble, Eric A.

Brewer, David Culler, and Joseph M. Hellerstein. Persistent distributed data structures to simplify cluster-based Internet services. Operating Systems Review, 34(2):37–38, April 2000. CODEN OSRED8. ISSN 0163-5980.

Guillon:2002:SIS

[GBM02] Benoit Guillon, Dominique Burger, and Bruno Marmol. A secure Internet service for delivering documents for the blind. Lecture Notes in Computer Science, 2398:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2398/23980360.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2398/23980360.
pdf.

Gutmans:2005:PPP

[GBR05] Andi Gutmans, Stig Sæther Bakken, and Derick Rethans. PHP 5 power programming. Bruce Perens’ Open Source series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-147149-X (paperback). xxviii + 689 pp. LCCN QA76.73.P224 G88 2005.

Gutierrez:2004:LRW

[GCB04] José A. Gutiérrez, Edgar H. Callaway, and Raymond L. Barrett. Low-rate wireless personal area networks: enabling wireless sensors with IEEE 802.15.4. IEEE standards wireless networks series. Standards Information Network, IEEE Press, New York, NY, USA, 2004. ISBN 0-7381-3557-7. xviii + 155 pp. LCCN TK5105.78 .G88 2004. URL http:/

                                        /shop.ieee.org/ieeestore/
                                        product.aspx?product_no=
                                        SP1131.

Gopalan:2002:IRL

[GcC02] Kartik Gopalan and Tzi cker Chiueh. Improving route lookup performance using network processor cache. In IEEE [IEE02c], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.
                                        org/paperpdfs/pap.pap272.
                                        pdf.

Gips:2002:AIC

[GDC^{+}02] James Gips, Philip A. Di-Mattia, Marialice Curran, Debra Lees, and Maureen Gates. Accessing Internet courses by eye movement and head movement. Lecture Notes in Computer Science, 2398:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2398/23980236.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2398/23980236.
pdf.

Geist:2000:ILC

[Gei00] Michael A. (Michael Allen)

Geist. Internet law in Canada. Captus Press, North York, ON, Canada, 2000. ISBN 1-896691-79-X. xx + 747 pp. LCCN KE452.C6 G44 2000.

Germann:2000:GSI

[Ger00] Michael Germann. Gefahrenabwehr und Strafverfolgung im Internet. (German) [Danger warning and prosecution in the InterNet], volume 812 of Schriften zum offentlichen Recht; ISSN: 0582-0200; Band 812. Duncker and Humblot, Berlin, Germany, 2000. ISBN 3-428-09985-0. ISSN 0582-0200. 757 pp. LCCN KK8825 .G47 2000. Originally presented as the author’s Inaugural-Dissertation (doctoral) — Friedrich-Alexander-Universitat Erlangen-Nürnberg,

Goss:2003:WIT

[GG03] Ernest Goss and Uma Gupta. Women and the Internet: is there an economic payoff? Communications of the ACM, 46(9):160–166, September 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Golomb:2005:SDG

[GG05] Solomon W. (Solomon Wolf) Golomb and Guang Gong. Signal design for good correlation for wireless communication, cryptography, and radar. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-82104-5 (hardcover). xviii + 438 pp. LCCN TK5102.92 .G65 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/ecip057/2005002719.html.

Gibson:2001:MCD

[Gib01] Jerry Gibson, editor. Multimedia Communications: Directions and Innovations. Academic Press, New York, NY, USA, 2001. ISBN 0-12-282160-2. xviii + 318 pp. LCCN TK5105.2 .M85 2000. US$95.00.

Gilbey:2000:MID

[Gil00] Chris Gilbey. MP3 and the infinite digital jukebox: a stepby-step guide to accessing and downloading CD-quality music from the Internet. Seven Stories Press, New York, NY, USA, 2000. ISBN 1-58322-034-8 (paperback). xix + 169 pp. LCCN ML74.4.M6 G55 2000.

Gilb:2004:WMG

[Gil04] James P. K. Gilb. Wireless multimedia: a guide to the IEEE 802.15.3 standard. IEEE standards wireless networks series. Standards Information Network, IEEE Press, New York, NY, USA, 2004. ISBN 0-7381-3668-9. xix + 314 pp. LCCN TK5105.5668 .G55 2004. URL http:/

/shop.ieee.org/ieeestore/
product.aspx?product_no=
SP1132.

Gilb:2005:IWD

[Gil05]  James P. K. Gilb.  The

IEEE wireless dictionary. IEEE standards wireless networks series. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7381-4766-4. ???? pp. LCCN TK5103.2 .G54 2005. URL http:/

/shop.ieee.org/ieeestore/
product.aspx?product_no=
SP1142.

Giladi:2007:NPA

[Gil07] Ran Giladi. Network Processors: Architecture, Programming, and Implementation. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2007. ISBN 0-12-370891-5. ???? pp. LCCN ????

Goni:2002:AMS

[GIM^{+}02] A. Goñi, A. Illarramendi, E. Mena, Y. Villate, and J. Rodríguez. ANTARC-TICA: a multiagent system for Internet data services in a wireless computing framework. Lecture Notes in Computer Science, 2538:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2538/25380119.htm;
http://link.springer.de/
link/service/series/0558/
papers/2538/25380119.pdf.

Giova:2000:CDC

[Gio00] Stefania Giova. La conclusione del contratto via Internet: evoluzione tra normativa e prassi. Quaderni della Rassegna di diritto civile. Edizioni scientifiche italiane, Napoli, Italy, 2000. ISBN 88-495-0195-1. 204 pp. LCCN KJC2045 .G56 2000.

Glazier:2000:PSS

[Gla00]  Stephen C. Glazier.  E-

patent strategies for software, e-commerce, the Internet, telecom services, financial services, and business methods (with case studies and forecasts). LBI Institute, Washington, DC, USA, 2000. ISBN 0-9661437-8-7. xxix + 193 pp. LCCN KF3114 .G549 2001. “Companion volume to Patent strategies for business, third edition.”.

Glazer:2001:UIP

[Gla01] Evan Glazer. Using Internet primary sources to teach critical thinking skills in mathematics. Greenwood professional guides in school librarianship. ISSN: 1074-150X. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2001. ISBN 0-313-31327-X. xi + 222 pp. LCCN QA11.5 .G54 2001.

Glisic:2006:AWN

[Gli06] Savo G. Glisic. Advanced wireless networks: 4G technologies. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2006. ISBN 0-470-01593-4. xx + 863 pp. LCCN TK5103.2 .G553

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0657/2006298111-d.html.

Gong:2005:SSL

[GLMT05] Wei-Bo Gong, Yong Liu, Vishal Misra, and Don Towsley. Self-similarity and long range dependence on the Internet: a second look at the evidence, origins and implications. Computer Networks (Amsterdam, Netherlands: 1999), 48(3): 377–399, June 21, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Grant:2000:CHG

[GM00a] David Grant and Peter McBride. Caterer and Hotelkeeper Guide to the Internet: Getting your business online. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-7506-4896-1. vi + 168 pp. LCCN TX911.3.M3 G73 2000. US$31.95.

Groth:2000:CCG

[GM00b] David Groth and Jim McBee. Cabling: the complete guide to network wiring. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2645-6. xxiv + 821 pp. LCCN TK5103.12 .G3 2000.

Grossman:2001:DGN

[GM01] Lawrence K. Grossman and Newton N. Minow. A digital gift to the nation: fulfilling the promise of the digital and Internet age. Century Foundation report. Century Foundation Press, New York, NY, USA, 2001. ISBN 0-87078-466-8 (pamphlet). xiii + 280 pp. LCCN T58.5 .G78 2001. “A publication of the Digital Promise Project.” Sponsored by the Carnegie Corporation of New York and others. Partial Contents: Foreword by Richard C. Leone — Executive Summary — The recommendation — The need — The new landscape for lifelong learning — A new role for public broadcasting — Conclusion — Notes — Digital perspectives background papers. The nascent age of broadband by Les Brown — Justin Smith Morrill and the politics and legacy of the Land-Grant College Acts by Craig L. LaMay — Governance models for digital opportunity investment trust by Marion R. Fremont-Smith — Teaching the promise: the research university in the Information Age by Cathy N. Davidson — Fulfilling the digital promise in nonprofit education on a national level by Thomas R. Martin — The promise of the digital age for the renewal of regional civic culture by Mary L. Walshok — Fathom: a case study of a public/private partnership in higher education by Ann Kirschner — Bridging the digital divide: public service telecommunication in the Internet and Digital Age by Martin Gomez — How to put the Web to work for nonprofits by Susan A. Rogers Partial Contents: Its 2001. Do you know what your children are watching?: the critical role of children’s television in the Age of New Media by Rita Weisskoff — Older Americans and the digital revolution by Everette E. Dennis — The digital promise of Lincoln Center by John Goberman — (Re)imagining BAM in the age of broadband technologies by Wayne Ashley — You will find no profit in winning, but if you lose, you lose it all by Richard Kimball — Whose voices count: a proposal to strengthen the communications capability of community groups by Mark Lloyd — The public radio station of the future by Laura Walker — Public television in the digital age by Richard Somerset-Ward.

Graham:2003:CTS

[GM03] Leah Graham and Panagiotis Takis Metaxas. “of course it’s true; I saw it on the Internet!”: critical thinking in the Internet era. Communications of the ACM, 46(5):70–75, May 2003. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ghosh:2004:DCI

[GM04] R. K. Ghosh and Hrushikesha Mohanty, editors. Distributed computing and Internet technology: first international conference, ICDCIT 2004, Bhubaneswar, India, December 22–24, 2004: Proceedings, volume 3347 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-24075-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I24 2004. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/
                                        ;   http://springerlink.
                                        metapress.com/openurl.asp?

genre=issue&issn=0302-9743&
volume=3347;   http://www.
springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3347;   http:
//www.springerlink.com/openurl.
asp?genre=volume&id=doi:
10.1007/b104418.

Graham:2000:UWD

[GMP00] David T. Graham, Jane Mc-Neil, and Lloyd Pettiford. Untangled Web: developing teaching on the Internet. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-582-41854-2. xii + 308 pp. LCCN LB1044.87 .G73 2000.

Gstalder:2001:GII

[GN01] Steven Gstalder and Greg R.

Notess. Government information on the Internet. Bernan Press, Lanham, MD, USA, fourth edition, 2001. xxi + 911 pp. LCCN ZA5075 .G68 2001.

Goertzel:2002:CII

[Goe02] Ben Goertzel. Creating Internet intelligence: wild computing, distributed digital consciousness, and the emerging global brain, volume 18 of IFSR international series on systems science and engineering. Kluwer Academic / Plenum Publishers, New York, NY, USA, 2002. ISBN 0-306-46735-6. xvi + 330 pp. LCCN TK5105.875.I57 G64 2002.

Goldberg:2000:RGT

[Gol00a] Ken Goldberg. The robot in the garden: telerobotics and telepistemology in the age of the Internet. Leonardo. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-07203-3. xix + 366 pp. LCCN TJ211 .R537 2000.

Goldstein:2000:HHP

[Gol00b] Douglas E. Goldstein. Ehealthcare: harness the power of Internet e-commerce and e-care. Aspen Publishers, Gaithersburg, MD, USA, 2000. ISBN 0-8342-1365-6. xxxi + 491 pp. LCCN R859.7.E43 G65 2000.

Goldstein:2000:APS

[Gol00c] Norm Goldstein. The Associated Press stylebook and briefing on media law: fully revised and updated with a new Internet guide and glossary. Perseus, Cambridge, MA, USA, revised edition, 2000. ISBN 0-7382-0308-4. 413 pp. LCCN PN4783.A793 2000 H/SS. Previous eds. have title: “The Associated Press stylebook and libel manual”.

LlanezaGonzalez:2000:ICD

[Gon00] Paloma Llaneza Gonzalez. Internet y comunicaciones digitales: regimen legal de las tecnologias de la informacion y la comunicacion. (Spanish) [Internet and digital communications: legal regime of technologies of information and communication]. Bosch, Barcelona, Spain, 2000. ISBN 84-7676-587-8. 450 pp. LCCN KE3671 .L53 2000.

Goolsbee:2000:EI

[Goo00] Austan Goolsbee. Education and the Internet. ABF working paper 2021, American Bar Foundation, Chicago, IL, USA, 2000. 23 pp.

Goodrich:2002:IA

[Goo02a] Michael T. Goodrich, editor.

Internet algorithmics, volume 33(3) of Algorithmica. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CO-DEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic). 271–409 pp.

Goodwill:2002:AJT

[Goo02b]  James Goodwill.  Apache

Jakarta-Tomcat. Apress, Berkeley, CA, USA, 2002. ISBN 1-893115-36-4. 264 (est.) pp. LCCN ???? US$34.95.

Gordon:2001:TIL

[Gor01] Rachel Singer Gordon. Teaching the Internet in libraries. American Library Association, Chicago, IL, USA, 2001. ISBN 0-8389-0799-7. vii + 143 pp. LCCN ZA4201 .G64 2001.

Goth:2004:WSM

[Got04] G. Goth. When is a “Little in the Middle” ok? The Internet’s end-to-end principle faces more debate. IEEE Distributed Systems Online, 5(3):??, ???? 2004. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28687/01285878.pdf?
isnumber=28687&prod=JNL&
arnumber=1285878&arSt=&ared=
&arAuthor=Goth%2C+G.;   http:
//ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28687&
arnumber=1285878&count=7&
index=1.

Gough:2000:IIS

[Gou00] Leo Gough. Investing in Internet stocks: the global gold rush of the new economy. John Wiley and Sons (Asia), New York, NY, USA, 2000. ISBN 0-471-83971-X. xii + 254 pp. LCCN HD9696.8.A2 G68 2000.

Gowan:2001:EIG

[Gow01] Mary Gowan. eHRM: an Internet guide for human resource management. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-091283-2. 45 pp. LCCN HF5549.A27 G69 2001. Includes exercises for students.

Ghini:2005:HGD

[GPR^{+}05] V. Ghini, G. Pau, M. Roccetti, P. Salomoni, and M. Gerla. For here or to go? Downloading music on the move with an ultra reliable wireless Internet application. Computer Networks (Amsterdam, Netherlands: 1999), 49(1):4– 26, September 15, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Gao:2001:SIR

[GR01] Lixin Gao and Jennifer Rexford. Stable Internet routing without global coordination. IEEE/ACM Transactions on Networking, 9(6):681– 692, December 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Gardner:2002:XXG

[GR02] John Robert Gardner and Zarella L. Rendon. XSLT and XPath: a guide to XML transformations. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-040446-2. xxii + 558 pp. LCCN QA76.76.H94 R46 2002.

Graber:2001:PPL

[Gra01a] Doris A. (Doris Appel) Graber.

Processing politics: learning from television in the Internet age. Studies in communication, media, and public opinion. University of Chicago Press, Chicago, IL, USA, 2001. ISBN 0-226-30575-9 (hardcover), 0-226-30576-7 (paper). xiii + 231 pp. LCCN PN4888.T4 G73 2001.

Graham:2001:TIA

[Gra01b] Buck Graham. TCP/IP Addressing: Designing and Optimizing your IP Addressing Scheme. Academic Press, New York, NY, USA, second edition, 2001. ISBN 0-12-295021-6. xv + 455 pp.

LCCN TK5105.585 .G73 2001.

US$34.95.

Gralla:2002:HIW

[Gra02] Preston Gralla. How the Internet works. How it works. Que Corporation, Indianapolis, IN, USA, sixth edition, 2002. ISBN 0-7897-2582-

7. xi + 340 pp. LCCN TK5105.875.I57 G72423 2002. Previous ed.: 1999.

Gratton:2005:DPW

[Gra05] Dean A. Gratton. Developing Practical Wireless Applications. Elsevier/Digital Press, Amsterdam, The Netherlands; Boston, MA, USA, 2005. ISBN 1-55558-310-5. 320 (est.) pp. LCCN ????

Graves:2007:COC

[Gra07] Kimberly Graves. CEH: official certified ethical hacker review guide. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2007. ISBN 0-7821-4437-3 (paperback/cdrom). xxii + 238 pp. LCCN QA76.3 .G69

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0653/2006101131-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0741/2006101131-
                                        b.html;   http://www.loc.
                                        gov/catdir/toc/ecip0718/
                                        2006101131.html.

Greenberg:2000:MOE

[Gre00a] Jane Greenberg. Metadata and organizing educational resources on the Internet. Haworth Information Press, New York, NY, USA, 2000. ISBN 0-7890-1178-6, 0-7890-1179-4. 302 pp. LCCN QA76.9.W43 M48 2000. “Co-published simultaneously as Journal of Internet cataloging, volume 3, numbers 1 and 2/3 2000.”.

Greenblatt:2000:IDH

[Gre00b]  Bruce Greenblatt.  Inter-

net directories: how to build and manage applications for LDAP, DNS, and other directories. Prentice Hall series in computer networking and distributed systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-974452-5. xix + 290 pp. LCCN TK5105.595 .G74 2000. US$44.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0139744525.html.

Greenstein:2000:BDV

[Gre00c] Shane M. Greenstein. Building and delivering the virtual world: commercializing services for Internet access. NBER working paper series 7690, National Bureau of Economic Research, Cambridge, MA, USA, May 2000. 38 + [1] pp.

Greenberg:2001:CSL

[Gre01a] Paul Greenberg. CRM at the speed of light: capturing and keeping customers in Internet real time. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001.

ISBN 0-07-212782-1. xxiii + 360 pp. LCCN HF5548.32 .G74 2001.

Grey:2001:IS

[Gre01b] Duncan Grey. The Internet in school. Continuum, New York, NY, USA, second edition, 2001. ISBN 0-8264-5365-1. 220 pp. LCCN LB1044.87 .G749 2001.

Greenblatt:2002:BLE

[Gre02a]  Bruce Greenblatt.  Build-

ing LDAP-enabled Applications with Microsoft’s Active Directory and Novell’s NDS. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-062145-5. 196 pp. LCCN ???? US$39.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130621455.html.

Greenfield:2002:EGO

[Gre02b] David Greenfield. The Essential Guide to Optical Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-042956-2. xvii + 229 pp. LCCN TK5103.59 .G744 2002. US$39.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130429562.html.

Greenstein:2002:MEW

[Gre02c]  Shane Greenstein.  Micro

economics: Which industries use the Internet? IEEE Micro, 22(6):70–72, November/December 2002. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://dlib.
computer.org/mi/books/mi2002/
pdf/m6070.pdf.

Greenstein:2003:MEW

[Gre03] Shane Greenstein. Micro economics: Where did the Internet go? IEEE Micro, 23 (1):84–86, January/February 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2003/pdf/m1084.
pdf;   http://www.computer.
org/micro/mi2003/m1084abs.
htm.

Griffiths:2000:MYI

[Gri00] Peter Griffiths. Managing your Internet and intranet services: the information and library professional’s guide to strategy. Library Association, London, UK, 2000. ISBN 1-85604-340-1. xi + 223 pp. LCCN TK5105.888 .G75 2000.

Ghazaryan:2000:IIA

[GS00] E. A. (Eduard Arthashesi) Ghazaryan and A. A. Step’anyan. Informatika, Internet: angleren-ruseren-hayeren bats’atrakan bararan. Makmilan Armenia, Erevan, Armenia, 2000. ISBN 99930-72-09-5. 319 pp. LCCN QA76.15 .G49 2000.

Garfinkel:2003:PUI

[GSS03] Simson Garfinkel, Gene Spafford, and Alan Schwartz. Practical Unix & Internet Security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2003. ISBN 0-596-00323-4. 1000 (est.) pp. LCCN ???? US$54.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        puis3.

Gabrovec:2001:IQF

[GT01] Judy Gabrovec and Ken Tunell. Internet Quests: fairy tales. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3405-9. 48 pp. LCCN PZ8 .G33 2001. Primary grades.

Goodrich:2002:ADF

[GT02] Michael T. Goodrich and Roberto Tamassia. Algorithm design: foundations, analysis, and Internet examples. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-38365-1. xii + 708 pp. LCCN QA76.9.A43 G661 2002.

Gurak:2001:CNI

[Gur01] Laura J. Gurak. Cyberliteracy: navigating the Internet with awareness. Yale University Press, New Haven, CT, USA, 2001. ISBN 0-300-08979-1. viii + 194 pp. LCCN TK5105.875.I57 G88 2001.

Gibson:2000:RDB

[GW00] Rachel Kay Gibson and Stephen Ward. Reinvigorating democracy?: British politics and the Internet. Ashgate, Burlington, VT, USA, 2000. ISBN 1-84014-793-8. xii + 239 pp. LCCN JN329.E4 R45 2000.

Gunasekaran:2001:IBE

[GY01] A. Gunasekaran and Yahaya Y. Yusuf, editors. Internet-based enterprise integration and management: 31 October–1 November 2001, Newton, USA, volume 4566 of SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 2001. ISBN 0-8194-4294-1. LCCN TS510.S63 v.4566.

Hakanson:2000:MRL

[Hak00] Lars Hakanson. Modelling radiocesium in lakes and coastal areas: new approaches for ecosystem modellers: a textbook with Internet support. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-6245-4. x + 215 pp. LCCN TD427.R3 H36 2000.

Halabi:2000:IRA

[Hal00a] Bassam Halabi. Internet routing architectures. Cisco Press, Indianapolis, IN, USA, second edition, 2000. ISBN 1-57870-233-X. xvii + 477 pp. LCCN ???? Edited by John Kane.

Hall:2000:ICP

[Hal00b] Eric A. Hall. Internet Core Protocols: The Definitive Guide: Help for Network Administrators. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, February 2000. ISBN 1-56592-572-6. 469 pp. LCCN TK5105.875.I57 H355 2000. US$39.95. URL http://

www.oreilly.com/catalog/

coreprot. Includes CD-ROM.

Hanson:2000:PIM

[Han00] Ward A. Hanson. Principles of Internet marketing. South-Western College Pub., Cincinnati, OH, USA, 2000. ISBN 0-538-87573-9 (softcover). xix + 467 pp. LCCN HF5415.1265 .H36 2000.

Hargittai:2000:RLI

[Har00a]  Eszter Hargittai.  Radio’s

lessons for the Internet. Communications of the ACM, 43 (1):50, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/articles/journals/cacm/
2000-43-1/p50-hargittai/
p50-hargittai.pdf;   http:
//www.acm.org/pubs/citations/
journals/cacm/2000-43-1/
p50-hargittai/.

Harmon:2000:UIO

[Har00b]  Charles Harmon.  Using

the Internet, online services, and CD-ROMs for writing research and term papers. Neal-Schuman netguide series. Neal-Schuman Publishers, New York, NY, USA, second edition, 2000. ISBN 1-55570-374-7. x + 126 pp. LCCN Ordered for Health Sci.

Harrison:2000:DWP

[Har00c] Graham (Graham Paul) Harrison. Dynamic Web programming: using Java, JavaScript, and Informix. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-086184-7. xviii + 995 pp. LCCN QA76.625.H37 2000.

Hart:2000:AIA

[Har00d] Michael Hart. The American Internet advantage: global themes and implications of the modern world. University Press of America, Lanham, MD, USA, 2000. ISBN 0-7618-1672-0 (paperback). xix + 131 pp. LCCN HE7583.U6 H37 2000.

Hardy:2001:SLI

[Har01] Stephen T. Hardy. Studying law on the Internet: how to use the Internet for learning and study, exams and career development. Internet Handbooks, Plymouth, UK, 2001. ISBN 1-84025-370-3. 128 pp. LCCN KD394 .H37 2001.

Harnedy:2002:MPI

[Har02] Sean Harnedy. The MPLS Primer: An Introduction to Multiprotocol Label Switching. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-032980-0. xxv + 484 pp. LCCN TK5105.573 .H37 2002. US$44.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130329800.html.

Harrington:2005:NSP

[Har05]  Jan L. Harrington.  Net-

work security: a practical approach. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-12-311633-3. xv + 365 pp. LCCN TK5105.59 .H357 2005.

Hawke:2001:CIU

[Haw01] Constance S. Hawke. Computer and Internet use on campus: a legal guide to issues of intellectual property, free speech, and privacy. The Jossey-Bass higher and adult education series. Jossey-Bass, San Francisco, CA, USA, 2001. ISBN 0-7879-5516-7. xix + 172 pp. LCCN KF390.5.C6 H39 2001. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        onix06/00009572.html.

Hayward:2000:PUI

[Hay00] Allison Hayward. E-politics: using the Internet to participate in politics and interact with your government. Sams teach yourself today. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31981-0 (paperback). 292 pp. LCCN JK1764 .H39 2000. “Web site directory”: p. [233]- 259.

Hardy:2001:LLI

[HB01] Stephen T. Hardy and Michael Base. Law and lawyers on the Internet: an essential guide and resource for legal practitioners. Internet handbooks. Internet Handbooks, Plymouth, UK, USA, 2001. ISBN 1-84025-345-2 (paperback). 176 pp. LCCN K88.I58 H37 2001. UK£12.99.

Haines:2003:LSO

[HB03] Michael Haines and Tom Bialaski. LDAP in the Solaris Operating Environment: Deploying Secure Directory ServiceS. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-145693-8. 704 (est.) pp. LCCN ???? US$54.00.

Hofmann:2005:CNA

[HB05] Markus Hofmann and Leland R. Beaumont. Content networking: architecture, protocols, and practice. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2005. ISBN 1-55860-834-6. xiv + 358 pp. LCCN TK5105.5 .H63 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip056/2005001732.html.

Hoglund:2006:RSW

[HB06] Greg Hoglund and James Butler. Rootkits: subverting the Windows kernel. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-29431-9 (paperback). xxiii + 324 pp. LCCN QA76.9.A25 H637

                                        2006. URL http://www.loc.
                                        gov/catdir/toc/ecip0512/
                                        2005013061.html.

Hoffman:2001:IVP

[HC01] Lance J. Hoffman and Lorrie Cranor. Internet voting for public officials: introduction. Communications of the ACM, 44(1):69, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        2001-44-1/p69-hoffman/.

Hiller:2002:ILP

[HC02] Janine Hiller and Ronnie Cohen. Internet law and policy. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-033428-6. xv + 377 pp. LCCN KF390.5.C6 H55 2002.

He:2004:TUP

[HC04] Jingyi He and S.-H. Gary Chan. TCP and UDP performance for Internet over optical packet-switched networks. Computer Networks (Amsterdam, Netherlands: 1999), 45 (4):505–521, July 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hogie:2005:USI

[HCP05] Keith Hogie, Ed Criscuolo, and Ron Parise. Using standard Internet Protocols and applications in space. Computer Networks (Amsterdam, Netherlands: 1999), 47(5): 603–650, April 5, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hart-Davis:2001:IPS

[HD01] Guy Hart-Davis. Internet pricay [sic.] exposed. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2920-X. xxvi + 485 pp. LCCN TK5105.875.I57 H375 2001.

Abdelsalam:2002:WTE

[Hed02] Abdelsalam “Solom” Heddaya. Web technologies: An economically scalable Internet. Computer, 35(9):93– 95, September 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/09/r9093.htm;   http:
//csdl.computer.org/dl/mags/
co/2002/09/r9093.pdf.

Held:2000:NDP

[Hel00] Gilbert Held, editor. Network Design: Principles and Applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-0859-3. xvi + 852 pp. LCCN TK5105.5 .N4654 2000. US$79.95.

Held:2001:TIP

[Hel01] Gilbert Held. TCP/IP Professional Reference Guide. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-0824-0. xv + 239 pp. LCCN TK5105.585 .H448 2001. US$49.95.

Held:2002:QSC

[Hel02] Gilbert Held. Quality of Service in a Cisco Networking Environment. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-470-84425-6. 343 (est.) pp. LCCN ???? US$55.00.

Herlein:2000:OPI

[Her00] Greg Herlein. The Open-Phone project — Internet telephony for everyone! Linux Journal, 69:??, January 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue69/3512.
html.

Herler:2001:USA

[Her01] Brita Kristina Herler. Upphovsrättsligt skydd av digitala musikaliska verk vid marknadsföring i Internet. Acta Wasaensia (ISSN 0355-2667); Rättsvetenskap (ISSN 1457-7992); Handelsrätt no. 89. no. 2. Universitas Wasaensis, Vasa, Sweden, 2001. 385 pp. LCCN Z653 .H47 2001.

Heim:2000:FSG

[HH00] Judy Heim and Gloria Hansen. Free stuff for gardeners on the Internet. C and T Pub., Lafayette, CA, USA, 2000. ISBN 1-57120-097-5 (paper). 176 pp. LCCN SB450.955 .H45 2000.

Heflin:2003:IIC

[HH03] Jeff Heflin and Michael N.

Huhns. IEEE Internet computing: Guest Editors’ introduction: The zen of the Web. IEEE Distributed Systems Online, 4(9), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0309/f/w5geip.htm.

Hick:2000:HRI

[HHH00] Steven Hick, Eric Hoskins, and Edward F. Halpin. Human rights and the Internet. St. Martin’s Press, New York, NY, USA, 2000. ISBN 0-333-77733-6. xviii + 257 pp. LCCN JC571 .H68854 2000.

Huebsch:2003:QIP

[HHL^{+}03] Ryan Huebsch, Joseph M.

Hellerstein, Nick Lanham, Boon Thau Loo, Scott Shenker, and Ion Stoica. Querying the Internet with PIER. In Freytag et al. [FLA^{+}03], pages 321–332. ISBN 0-12-722442-4. LCCN ????

URL http://www.vldb.org/
dblp/db/indices/a-tree/h/
Huebsch:Ryan.html.

Hillier:2000:MCF

[Hil00] Susan Hillier. Making connections: family and relationship studies on the Internet. Mc-Graw-Hill Higher Education, Boston, MA, USA, 2000. ISBN 0-07-235947-1. v + 42 pp. LCCN HQ518 .H55 2000.

Hill:2001:CSS

[Hil01] Sonya D. Hill. Consumer sourcebook: a subject guide to over 15,000 federal, state, and local government agencies and offices, national, regional, and grassroots associations and organizations, information centers, clearinghouses, publications, Internet databases, multimedia resources, media contacts, and related consumer resources. Gale Research, Detroit, MI, USA, fourteenth edition, 2001. ISBN 0-7876-3346-1. lxxxiii + 1245 pp. LCCN HC110.C63 H54 2001.

Hird:2000:LCS

[Hir00] Anne Hird. Learning from cyber-savvy students: how Internet-age kids impact classroom teaching. Stylus, Sterling, VA, USA, 2000. ISBN 1-57922-030-4, 1-57922-031-2 (paper). xi + 176 pp. LCCN LB1028.43.H57 2000.

Howard:2004:SOI

[HJ04] Philip N. Howard and Steve Jones, editors. Society online: the Internet in context. Sage Publications, Thousand Oaks, CA, USA, 2004. ISBN 0-7619-2707-7 (Cloth), 0-7619-2708-5 (Paper). xxxiv + 350 pp. LCCN HM851 .S655

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0658/2003008611-d.html;

http://www.loc.gov/catdir/
toc/ecip042/2003008611.html.

Hofmann:2005:FTG

[HKH05] Chris Hofmann, Marcia Knous, and John Hedtke. Firefox and Thunderbird Garage. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-187004-1. 384 (est.) pp. LCCN ????

Huang:2003:TID

[HKLS03] H. Huang, C. Keser, J. Leland, and J. Shachat. Trust, the Internet, and the digital divide. IBM Systems Journal, 42(3):507–518, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/423/huang.pdf.

Hofmann:2004:GI

[HKT04] Markus Hofmann, Robin Kravets, and Joseph D. Touch. The Global Internet. Computer Networks (Amsterdam, Netherlands: 1999), 45(1):1– 3, May 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Hui:2000:IAI

[HL00] Lucas Chi-Kwong Hui and Dik Lun Lee, editors. Internet applications: 5th International Computer Science Conference, ICSC ’99, Hong Kong, China, December 13– 15, 1999: proceedings, volume 1749 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66903-5. LCCN QA267.A1 L43 no.1749.

Hiltunen:2002:API

[HL02a] Yrjö Hiltunen and Mika Lappalainen. Automated personalisation of Internet users using self-organising maps. Lecture Notes in Computer Science, 2412:31– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2412/24120031.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2412/24120031.
pdf.

Huang:2002:JCA

[HL02b] B. Huang and H. Lin. A Java/ CGI approach to developing a geographic virtual reality toolkit on the Internet. Computers and Geosciences, 28(1): 13–19, February 2002. CO-DEN CGEODT, CGOSDN. ISSN 0098-3004.

Howard:2003:WSC

[HL03] Michael Howard and David LeBlanc. Writing secure code. Microsoft Press, Redmond, WA, USA, second edition, 2003. ISBN 0-7356-1722-8. xxviii + 768 pp. LCCN QA76.9.A25 H698 2003.

Howard:2005:DSS

[HLV05] Michael Howard, David LeBlanc, and John Viega. 19 deadly sins of software security: programming flaws and how to fix them. McGraw-Hill, New York, NY, USA, 2005. ISBN 0-07-226085-8. LCCN QA76.9.A25 H69 2005. URL

http://library.books24x7.
com/library.asp?B&bookid=
11971.

Huet:2000:CC

[HM00] Jerome Huet and Herbert Maisl. Code de la communication. (French) [Communication policy]. Dalloz, Paris, France, 2000. ISBN 2-247-02956-6. xiii + 1345 + 26 pp. LCCN KJV6220 .A24 2000.

Hick:2002:AAI

[HM02a] Steven Hick and John G. Mc-Nutt. Advocacy, activism, and the Internet: community organization and social policy. Lycecum Books, Chicago, IL, USA, 2002. ISBN 0-925065-60-9. xiii + 241 pp. LCCN HV29.2 .A38 2002.

Honda:2002:EGI

[HM02b] Gail Honda and Kipp Martin.

The essential guide to Internet business technology. Essential guide series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-042820-5. ???? pp. LCCN HD30.2 .H659 2002.

Hoglund:2004:ESH

[HM04] Greg Hoglund and Gary Mc-Graw. Exploiting Software: How to Break Code. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-201-78695-8 (paperback). xxxv + 471 pp. LCCN QA76.9.A25 H635

2004. URL http://www.loc.
gov/catdir/toc/ecip0411/
2003025556.html.

Hoffman:2004:IBI

[HNV04] Donna L. Hoffman, Thomas P. Novak, and Alladi Venkatesh. Has the Internet become indispensable? Communications of the ACM, 47(7):37–42, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hofacker:2001:IM

[Hof01] Charles F. Hofacker. Internet marketing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, third edition, 2001. ISBN 0-471-39051-8 (paperback). x + 165 pp. LCCN HF5415.1265 .H638 2001. URL http:

//www.wiley.com/college/

hofacker. Supplemented by a companion Web site and complimentary 10 week subscription to the Wall Street journal.

Hohloch:2001:RID

[Hoh01] Gerhard Hohloch, editor. Recht und Internet: 6. “Deutsch-Schwedisches Juristentreffen” vom 31. Marz bis 2. April 2000 in Lund. (German) [Law and the Internet: Sixth “German-Swedish lawyer meeting” of 31. March

— 2 April 2000 in Lund], volume 197 of Arbeiten zur Rechtsvergleichung. Nomos, Baden-Baden, Germany, 2001. ISBN 3-7890-7125-0. LCCN K564.C6 D49 2000.

Hooke:2001:II

[Hoo01] Adrian Hooke. The interplanetary Internet. Communications of the ACM, 44(9):38– 40, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/articles/journals/cacm/
2001-44-9/p38-hooke/p38-
hooke.pdf;   http://www.
acm.org/pubs/citations/journals/
cacm/2001-44-9/p38-hooke/

.

Houlding:2001:VVD

[Hou01] David Houlding. VoiceXML and the voice-driven Internet. Dr. Dobbs Journal, 26(4):88, 90–92, 94–95, April 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/ftp/2001/2001_04/
voicexml.zip;   http://www.
ddj.com/ftp/2001/2001_04/
voicexmltxt.zip.

Howells:2001:CLC

[How01] Cyndi Howells. Cyndi’s list:

a comprehensive list of 70,000 genealogy sites on the Internet. Genealogical Pub. Co., Baltimore, MD, USA, second edition, 2001. ISBN 0-8063-1681-0 (vol. 1), 0-8063-1680-2 (vol. 2), 0-8063-1678-0 (set). ???? pp. LCCN CS21 .H747 2001 v.1-2. URL http://www.

                                        cyndislist.com.

Heinle:2002:DJC

[HP02] Nick Heinle and Bill Pena. Designing with JavaScript: creating dynamic Web pages. O’Reilly Web studio. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 1-56592-360-X. xi + 216 pp. LCCN QA76.73.J39 H45 2002.

Huske:2001:ICE

[HPS01] Lynne Huske, Nancy Protheroe, and Elizabeth Shellard. The Internet connection: exploring the power of the Web as a teaching tool. Educational Research Service, Arlington, VA, USA, 2001. ISBN 1-931762-01-5. vi + 136 pp. LCCN LB1044.87 .H87 2001.

Hanzo:2007:VAC

[HSW07] Lajos Hanzo, F. Clare A. Somerville, and Jason P. Woodard. Voice and audio compression for wireless communications. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 2007. ISBN 0-470-51581-3 (cloth). xxxv + 843 pp. LCCN TK7882.S65 H35

                                        2007. URL http://www.loc.

gov/catdir/toc/ecip0713/
2007011025.html.

Huber:2002:CTM

[Hub02]  Josef F. Huber.  Com-

munications: Toward the mobile Internet. Computer, 35(10):100–102, October 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/10/rx100.
htm;   http://csdl.computer.
org/dl/mags/co/2002/10/rx100.
pdf.

Hudson:2006:PN

[Hud06] Paul Hudson. PHP in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10067-1 (paperback). xv + 352 pp. LCCN QA76.73.P224 H83 2006.

Huhns:2003:IIC

[Huh03] Michael N. Huhns. IEEE Internet computing: Agents on the Web: The Sentient Web. IEEE Distributed Systems Online, 4(11), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0311/d/w6age.htm.

Huitema:2000:RI

[Hui00] Christian Huitema. Routing in the Internet. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2000. ISBN 0-13-022647-5. xiii + 384 pp. LCCN TK5105.875.I57 H85 2000. US$42.00.

Hunter:2002:PGI

[Hun02] Tracy S. Hunter. e-Pharmacy: a guide to the Internet Care-Zone. American Pharmaceutical Association, Washington, DC, USA, 2002. ISBN 1-58212-029-3. xii + 321 pp. LCCN QV 26.5 E11 2002.

Huston:2000:IPS

[Hus00] Geoff Huston. Internet performance survival guide: QoS strategies for multi-service networks. Wiley Networking Council series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2000. ISBN 0-471-37808-9 (paper). xxiv + 551 pp. LCCN TK5105.875.I57 H86 2000.

Huseby:2004:ICS

[Hus04]  Sverre H. Huseby.  Inno-

cent code: a security wakeup call for Web programmers. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2004. ISBN 0-470-85744-7 (paperback). xix + 226 pp. LCCN QA76.9.A25 H88

2004. URL http://www.loc.
gov/catdir/bios/wiley047/
2003015774.html;   http:/
/www.loc.gov/catdir/description/
wiley0310/2003015774.html;
 http://www.loc.gov/catdir/
toc/wiley032/2003015774.
html.

Hutchby:2001:CTT

[Hut01] Ian Hutchby. Conversation and technology: from the telephone to the Internet. Blackwell, Malden, MA, USA, 2001. ISBN 0-7456-2110-4, 0-7456-2111-2 (paperback). 221 pp. LCCN P96.T42 H88 2001. Contents: 1. Introduction: Technologies for Communication — 2. The Communicative Affordances of Technological Artifacts — 3. Communication as Computation?

— 4. Talk-in-interaction — 5. The Telephone: Technology of Sociability — 6. Telephone Interaction and Social Identity — 7. Technological Mediation and Asymmetrical Interaction — 8. Computers, Humans and Conversation — 9. Virtual Conversation — 10. Conclusion: A Reversion to the Real?

Hilton:2000:BDA

[HW00] Craig Hilton and Jeff Willis.

Building Database Applications on the Web Using PHP3. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-65771-6. xix + 597 pp. LCCN QA76.9.D26 H56 2000. US$39.95. Includes CD-ROM.

Ishibashi:2002:CTS

[IAI02] Keisuke Ishibashi, Masaki Aida, and Makoto Imase. Characteristics of temporal and spatial locality of Internet access patterns. Lecture Notes in Computer Science, 2376:15– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2376/23760015.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2376/23760015.pdf.

Iannotta:2001:NSR

[Ian01] Joah G. Iannotta. Nontechnical strategies to reduce children’s exposure to inappropriate material on the Internet: summary of a workshop. The compass series. National Academy Press, Washington, DC, USA, 2001. ISBN 0-309-07591-2. xi + 90 pp. LCCN HQ784.I58 N6 2001.

IEEE:2001:IETa

[IEE01a] IEEE, editor. 2000 IEEE Emerging Technologies Symposium: Broadband, wireless Internet access: April 10 and 11, 2000, Richardson, TX, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7803-6364-7.

IEEE:2001:IETb

[IEE01b] IEEE, editor. 2001 IEEE Emerging Technologies Symposium on Broadband Communications for the Internet Era: symposium digest, September 10th and 11th, 2001, Radisson Hotel, Richardson, TX. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7803-7161-5. LCCN 49395.

IEEE:2001:JII

[IEE01c]  IEEE, editor.  Joint 4th

IEE[sic] International Conference on ATM (ICATM’01) and High Speed Intelligent Internet Symposium: April 22– 25, 2001, Seoul, Korea. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7803-7093-7. LCCN 48698.

IEEE:2001:PIC

[IEE01d] IEEE, editor. Proceedings, 39th International Conference and Exhibition on Technology of Object-Oriented Languages and Systems: TOOLS 39: 29 July-3 August 2001, Santa Barbara, California, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1251-8 (paperback), 0-7695-1253-4 (microfiche). LCCN QA76.64.T66 2001a.

IEEE:2001:PSA

[IEE01e] IEEE, editor. Proceedings:

Symposium on Applications and the Internet (1st: 2001: San Diego, Calif.). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-0942-8, 0-7695-0943-6 (case), 0-7695-0944-4 (microfiche). LCCN TK5105.875.I57 S95 2001.

IEEE:2001:WSI

[IEE01f] IEEE, editor. WIAPP 2001: the Second IEEE Workshop on Internet Applications: proceedings: July 23–24, 2001, San Jose, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1137-6, 0-7695-1138-4 (case), 0-7695-1139-2 (microfiche).

IEEE:2002:SAIb

[IEE02a] IEEE, editor. 2002 Symposium on Applications and the Internet (SAINT 2002): proceedings: January 28–February 1, 2002, Nara City, Nara, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1447-2, 0-7695-1448-0 (case), 0-7695-1449-9 (microfiche). LCCN 49384.

IEEE:2002:SAIa

[IEE02b] IEEE, editor. 2002 Symposium on Applications and the Internet (SAINT) Workshops: proceedings: January 28–February 1, 2002, Nara City, Nara, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1450-2, 0-7695-1451-0 (case), 0-7695-1452-9 (microfiche). LCCN 49383.

IEEE:2002:STI

[IEE02c] IEEE, editor. SC2002: From Terabytes to Insight. Proceedings of the IEEE ACM SC 2002 Conference, November 16–22, 2002, Baltimore, MD, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1524-X. LCCN ????

Immenga:2000:IWI

[ILS00] Ulrich Immenga, Natalie Lubben, and Hans-Peter Schwintowski. Das internationale Wirtschaftsrecht des Internet. (German) [The international commercial law of the Internet], volume 3 of Internationale Berliner Wirtschaftsrechtsgesprache. Nomos, Baden-Baden, Germany, 2000. ISBN 3-7890-6701-6 (paperback). 243 pp. LCCN K1005 .I67 2000.

Imparato:2000:PPI

[Imp00] Nicholas Imparato. Public policy and the Internet: privacy, taxes, and contract, volume 481 of Hoover Institution Press publication; no. 481. Hoover Institution Press, Stanford, CA, USA, 2000. ISBN 0-8179-9892-6. xvi + 159 pp. LCCN KF390.5.C6 P83 2000.

Iniewski:2003:ON

[Ini03] Krzysztof Iniewski. Optical Networking. Academic Press, New York, NY, USA, 2003.

ISBN 0-12-371051-0. ???? pp. LCCN ???? US$59.95.

Iniewski:2007:EWT

[Ini07] Krzysztof Iniewski. Emerging wireless technologies: circuits, systems, and devices. Taylor and Francis, Boca Raton, FL, USA, 2007. ISBN 0-8493-7996-2. ???? pp. LCCN TK5103.2 .I515

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0710/2007011165-d.html;
                                         http://www.loc.gov/catdir/
                                        toc/ecip0713/2007011165.
                                        html.

Insam:2003:TIE

[Ins03] Edward Insam. TCP/IP Embedded Internet Applications. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-7506-5735-9. 384 (est.) pp. LCCN ???? US$49.99.

ISIRS:2000:EAO

[ISI00a] ISI ResearchSoft. EndNote: the all-in-one solution; search bibliographic databases on the Internet, organize your references in a snap, create bibliographies in one simple step. ISI ResearchSoft, Berkeley, CA, USA, 2000. xvi + 445 pp. LCCN QA76.8.E6 E56 2000.

ISIRS:2000:EAS

[ISI00b] ISI ResearchSoft. EndNote: the all-in-one solution; search bibliographic databases on the Internet, organize your references in a snap, create bibliographies in one simple step. ISI ResearchSoft, Berkeley, CA, USA, 2000. xvi + 445 pp. LCCN QA76.8.E6 E56 2000. Enables users to do online searching of remote bibliographic databases (including MEDLINE, Library of Congress and university library catalogs), allowing users to create a database of references and create bibliographies.

ISIRS:2001:EAO

[ISI01] ISI ResearchSoft. EndNote:

the all-in-one solution; search bibliographic databases on the Internet, organize your references in a snap, create bibliographies in one simple step. ISI ResearchSoft, Berkeley, CA, USA, Windows version 5.0 edition, 2001. xvi + 493 pp. LCCN QA76.9.D3 .E53 2001.

Ives:2000:AIT

[Ive00] Gary Ives. Ariel: Internet transmission software for document delivery. Haworth Information Press, London, UK, 2000. ISBN 0-7890-1041-0 (hardcover), 0-7890-1042-9 (paperback). 134 pp. LCCN TK5105.875.I57 A74 2000. Copublished simultaneously as Journal of interlibrary loan, document delivery and information supply, volume 10, number 4, 2000.

Jacobson:2000:CTW

[Jac00] Trudi Jacobson. Critical thinking and the Web: teaching users to evaluate Internet resources, volume 4 of Active learning series. Library Instruction Publications, Pittsburg, PA, USA, 2000. ISBN 0-9652711-5-3. 259 pp. LCCN ZA4375 .C75 2000.

Jakubowski:2000:NEL

[Jak00] Maxim Jakubowski. The New English Library book of Internet stories. New English Library, London, UK, 2000. ISBN 0-340-76973-4. 384 pp. LCCN PR1309.I57 N49 2000.

Janczewski:2000:IIS

[Jan00] Lech Janczewski. Internet and intranet security management: risks and solutions. Idea Group Pub., Hershey, PA, USA, 2000. ISBN 1-878289-71-3. 302 pp. LCCN TK5105.875.I57 I684 2000. Contents: Pt. I. State of the Art. Ch. 1. Security Risk Assessment and Electronic Commerce: A Cross-Industry Analysis / Jonathan W. Palmer, Jamie Kliewer and Mark Sweat. Ch. 2. Securing the Internet in New Zealand: Threats and Solutions / Jairo A. Gutierrez — Pt. II. Managing Intranet and Internet Security. Ch. 3. Developing Trust for Electronic Commerce / Dieter Fink. Ch. 4. Managing Security Functions Using Security Standards / Lech Janczewski. Ch. 5. Managing Security in the World Wide Web: Architecture, Services and Techniques / Fredj Dridi and Gustaf Neumann — Pt. III. Cryptography and Technical Security Standards. Ch.

6. Cryptography: Protecting Confidentiality, Integrity and Availability of Data / Henry B. Wolfe. Ch. 7. Foundations for Cryptography / Dieter Gollmann. Ch. 8. Developments in Security Mechanism Standards / Chris Mitchell — Pt. IV. Security and the Law. Ch. 9. Electronic Mail, Employee Privacy and the Workplace / Charles Prysby and Nicole Prysby. Ch. 10. Protecting Personal Privacy in Cyberspace: The Limitations of Third Generation Data Protection Laws Such as the New Zealand Privacy Act 1993 / Gehan Gunasekara.

Janulaitis:2001:CHU

[Jan01] Michael Janulaitis. Creating HTML user interfaces for server programs. Dr. Dobbs Journal, 26(6):101–102, 104, 106, 108–109, June 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/ftp/2001/2001_06/
htmlui.txt;   http://www.
ddj.com/ftp/2001/2001_06/
htmlui.zip.

Jones:2005:RDF

[JBR05] Keith J. (Keith John) Jones, Richard Bejtlich, and Curtis W. Rose. Real digital forensics: computer security and incident response. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-24069-3 (paperback). xxx + 650 pp. LCCN HV8079.C65 J66 2005.

Jayasumana:2001:TPS

[JC01] Anura P. Jayasumana and V. Chandrasekar, editors. Technologies, protocols, and services for next-generation Internet: 21 and 23 August 2001, Denver, USA, volume 4527 of SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 2001. ISBN 0-8194-4251-8. LCCN TK5105.55 .T43 2001.

Jeffery:2000:CWY

[Jef00] Lawrence Jeffery. China’s wired!: your guide to the Internet in China. Asia Law and Practice, Euromoney Publications (Jersey), Hong Kong, 2000. ISBN 962-936-086-1. various pp. LCCN HF5548.325.C5 J43 2000.

Jenkins:2001:BTC

[Jen01] Philip Jenkins. Beyond tolerance: child pornography on the Internet. New York University Press, New York, NY, USA, 2001. ISBN 0-8147-4262-9 (hardcover). xii + 260 pp. LCCN HQ471 .J45 2001; K5293 .J46 2001.

Jepsen:2003:DSN

[Jep03] Thomas C. Jepsen. Distributed Storage Networks: Architecture, Protocols and Management. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2003. ISBN 0-470-85020-5. xxiv + 314 pp. LCCN ???? US$59.95, UK£34.95.

Jennings:2000:HWP

[JF00] Charles Jennings and Lori Fena. The hundredth window: protecting your privacy and security in the age of the Internet. Free Press, New York, NY, USA, 2000. ISBN 0-684-83944-X. xxv + 278 pp. LCCN K3264.C65 J46 2000.

Jacobi:2006:PJA

[JF06] Jonas Jacobi and John R. Fallows. Pro JSF and Ajax: building rich Internet components. The Expert’s voice in Java. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-580-7. xxiii + 435 pp. LCCN ????

Jarrett:2008:FWD

[JG08] Caroline Jarrett and Gerry Gaffney. Forms that work: designing Web forms for usability. Interactive technologies. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2008. ISBN 1-55860-710-2 (paperback). ???? pp. LCCN ???? US$22.99.

Jul:2000:HPC

[JKH^{+}00] Eric Jul, Povl Koch, Jørgen S.

Hansen, Michael Svendsen, Kim Henriksen, Kenn Nielsen, and Mads Dydensborg. Highperformance cluster-based Internet servers. Operating Systems Review, 34(2):38, April 2000. CODEN OSRED8. ISSN 0163-5980.

Jiao:2001:IRS

[JL01] Qun G. Jiao and Lewis-Guodo Liu. Internet resources and services for international finance and investment. Global guides to Internet business resources. Oryx Press, Westport, CT, USA, 2001. ISBN 1-57356-346-3. xiii + 540 pp. LCCN HG64 .J5 2001. es.

Jia:2001:FM

[JLHD01] Xiaohua Jia, Deying Li, Xiaodong Hu, and DingZhu Du. Optimal placement of Web proxies for replicated Web servers in the Internet. The Computer Journal, 44(5):329– 339, ???? 2001. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_44/Issue_05/440329.
sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_44/Issue_05/pdf/
440329.pdf.

Joch:2001:HFM

[Joc01] Alan Joch. How to find money online: an Internet-based capital guide for entrepreneurs. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-136060-3 (paperback). x + 274 pp. LCCN HG4027.7 .J63 2001.

Johnson:2000:IEP

[Joh00]  Kevin Johnson.  Internet

email protocols: a developer’s guide. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-43288-9. xxvi + 478 pp. LCCN TK5105.73.J64 2000.

Johnson:2003:WBC

[Joh03] Jeff Johnson. Web bloopers: 60 common Web design mistakes, and how to avoid them. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-840-0. 329 pp. LCCN ???? US$49.95.

Johnson:2006:FCC

[Joh06] Thomas A. Johnson, editor.

Forensic Computer Crime Investigation. Auerbach Publications, Boca Raton, FL, USA, 2006. ISBN 0-8247-2435-6. 338 (est.) pp. LCCN ???? US$99.95.

Jolitz:2003:WSD

[Jol03] William F. Jolitz. Web services and datacenter environments. Dr. Dobbs Journal, 28 (4):60, 62–65, April 2003. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/documents/s=7827/
ddj0304g/.

Jones:2000:JTF

[Jon00a] Bill Jones. Jones task force outlines the parameters for secure Internet voting. News release BJ00:06, California Secretary of State, Sacramento, CA, USA, January 18, 2000. 2 pp.

Jones:2000:IE

[Jon00b] J. Christopher (John Christopher) Jones. The Internet and everyone. Ellipsis, London, UK, 2000. ISBN 1-899858-20-2. 592 pp. LCCN TK5105.875.I57 J734 2000.

Jones:2005:PRT

[Jon05] Hollin Jones. Propellerhead Reason tips and tricks. PC, Merton, UK, 2005. ISBN 1-870775-97-X (paperback). 160 pp.

Jefferson:2004:AIV

[JRSW04] David Jefferson, Aviel D. Rubin, Barbara Simons, and David Wagner. Analyzing Internet voting security. Communications of the ACM, 47 (10):59–64, October 2004. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jagannathan:2002:ICM

[JSK02] Sridhar Jagannathan, Jay Srinivasan, and Jerry Kalman. Internet commerce metrics and models in the new era of accountability. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-028186-7. xiii + 390 pp. LCCN HF5548.32 .J33 2002.

Jackson:2000:IHC

[JT00] M. Drew Jackson and Timothy L. Taylor. The Internet handbook for Canadian lawyers. Carswell, Scarborough, ON, Canada, third edition, 2000. ISBN 0-459-23908-2. xvi + 252 pp. LCCN TK5105.875.I57 J33 2000. US$63.00.

Johari:2001:EEC

[JT01] Ramesh Johari and David Kim Hong Tan. End-to-end congestion control for the Internet: delays and stability. IEEE/ACM Transactions on Networking, 9(6):818– 832, December 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Jukan:2002:TOI

[Juk02] Admela Jukan, editor. Towards an optical Internet: new visions in optical network design and modelling: IFIP TC6 Fifth Working Conference on Optical Network Design and Modelling (ONDM 2001), February 5–7, 2001, Vienna, Austria, volume 76 of International Federation for Information Processing; 76. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 0-7923-7541-6. LCCN TK5103.59 .I35 2001.

Jackson:2004:IIU

[JvEB^{+}04] Linda A. Jackson, Alexander von Eye, Gretchen Barbatsis, Frank Biocca, Hiram E. Fitzgerald, and Yong Zhao. The impact of Internet use on the other side of the digital divide. Communications of the ACM, 47(7):43–47, July 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnson:2008:EAS

[JWC08] Dave Johnson, Alexei White, and Andre Charland. Enterprise AJAX: strategies for building high performance Web applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-224206-0 (paperback). xix + 468 pp. LCCN TK5105.8885.A52 J64

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/ecip0715/
                                        2007015974.html.

Jeong:2002:IAO

[JY02] Sangjin Jeong and Chan-Hyun Youn. Instability analysis for OBGP routing convergence in optical Internet. Lecture Notes in Computer Science, 2343:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2343/23430346.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2343/23430346.pdf.

Jeffay:2002:RMC

[JZ02] Kevin Jeffay and HongJiang Zhang. Readings in Multimedia Computing and Networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-651-3. xx + 863 pp. LCCN QA76.585 .R43 2002. US$69.95.

Kador:2000:IJC

[Kad00] John Kador. Internet jobs!:

the complete guide to finding the hottest Internet jobs. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-135261-9. xxii + 184 pp. LCCN HD9696.8.U62 K336 2000. Contents: pt. 1. Netenabled careers — ch. 1. New worlds of work — ch. 2. Net economy rules — ch. 3. Contract work — pt. 2. Internet jobs — ch. 4. Content jobs — ch. 5. Infrastructure jobs — ch. 6. Business jobs — ch. 7. Emerging net jobs and skills — pt. 3. Internet job resources — ch. 8. One hundred hottest companies on the net.

Kadrich:2007:ES

[Kad07] Mark Kadrich. Endpoint security. Addison Wesley Professional, Indianapolis, IN, USA, 2007. ISBN 0-321-43695-4 (paperback). ???? pp. LCCN QA76.9.A25 K325 2007.

Kahney:2005:CI

[Kah05] Leander Kahney. The cult of iPod. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-066-6. 256 (est.) pp. LCCN ML74.4.I48 K34 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;

                                         http://www.loc.gov/catdir/
                                        toc/ecip0517/2005023404.
                                        html;   http://www.oreilly.
                                        com/catalog/1593270666/.

Kaiser:2006:BBP

[Kai06a] Shirley Kaiser. Building Best Practice Websites: 101 Must Have Checklists. Build Your Own. Pagina förlags AB, 2006. ISBN 0-9758419-0-4. 400 (est.) pp. LCCN ????

Kaiser:2006:EWC

[Kai06b] Shirley Kaiser. Essential Website Checklists. SitePoint Pty Ltd, Collingwood, Victoria, Australia, 2006. ISBN 0-9758419-0-4. 400 (est.) pp. LCCN ????

Kamath:2000:LRC

[Kam00] Nandan Kamath. Law relating to computers, Internet, and e-commerce: a guide to cyberlaws and the Information Technology Act, 2000. W. Gaunt and Co., ???, FL, USA, second edition, 2000. ISBN 81-7534-194-7. xl + 752 pp. LCCN KNS80.C65 L39 2000.

Kamoroff:2001:OOB

[Kam01] Bernard Kamoroff. Online operator: business, legal, and tax guide to the Internet. Bell Springs Publishing, Willits and Laytonville, CA, USA, 2001. ISBN 0-917510-20-8. various pp. LCCN KF889 .K36 2001. Partial Contents: The Internet — Starting an Internet business — Web sites — E-commerce — Law of the Internet — Intellectual property — International commerce — Home business.

Kanalakis:2002:WSJ

[Kan02] John M. Kanalakis, Jr. Web services and Java server pages. Dr. Dobbs Journal, 27(1):28, 30–34, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/ftp/2002/2002_01/
wsjsp.txt;   http://www.
ddj.com/ftp/2002/2002_01/
wsjsp.zip.

Kaplan:2000:IVB

[Kap00] Michael S. Kaplan. Internationalization with Visual Basic. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31977-2. 672 (est.) pp. LCCN ???? US$49.99. URL http://www.

i18nWithVB.com/;   http:/
/www.mcp.com/sams/detail_
sams.cfm?item=0672319772.

Katz:2000:GHT

[Kat00] Jon Katz. Geeks: how two lost boys rode the Internet out of Idaho. Villard Books, New York, NY, USA, 2000. ISBN 0-375-50298-X. xlii + 207 pp. LCCN TK7885.54 .K38 2000.

Katz:2001:GHT

[Kat01] Jon Katz. Geeks: how two lost boys rode the Internet out of Idaho. Broadway Books, New York, NY, USA, 2001. ISBN 0-7679-0699-3 (reprint), 0-375-50298-X (original). xlii + 208 pp. LCCN TK7885.54 .K38 2001. Republication of [Kat00].

Kaye:2000:CIU

[Kay00] H. Stephen Kaye. Computer and Internet use among people with disabilities. Disability statistics report 13, U.S. Dept. of Education, National Institute on Disability and Rehabilitation Research, Washington, DC, USA, March 2000. iv + 13 pp.

Kovacich:2000:HTC

[KB00] Gerald L. Kovacich and William C. Boni. Hightechnology-crime investigator’s handbook: working in the global information environment. Butterworth-Heinemann, Boston, MA, USA, 2000. ISBN 0-7506-7086-X. xix + 298 pp. LCCN HV8079.C65 K68 2000. US$39.95.

Kovacs:2000:HFM

[KC00] Diane Kovacs and Ann L. Carlson. How to find medical information on the Internet: a print and online tutorial for the healthcare professional and consumer, volume 10 of Internet workshop series; no. 10. Library Solutions Press, 2137 Oregon St., Berkeley, CA, USA, 2000. ISBN 1-882208-27-7. xii + 128 pp. LCCN R859.7.I58 K68 2000.

Kim:2002:NPM

[KCC02] Byung-Chul Kim, Young-Soo Choi, and You-Ze Cho. A new packet marking strategy for explicit congestion notification in the Internet. Lecture Notes in Computer Science, 2343:431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2343/23430431.htm;
http://link.springer.de/
link/service/series/0558/
papers/2343/23430431.pdf.

Kikta:2002:WIC

[KCF02] Roman Kikta, Michael P.

Courtney, and Al Fisher. Wireless Internet crash course. McGraw-Hill telecom portable consultant. McGraw-Hill, New York, NY, USA, 2002. ISBN 0-07-138212-7. xxi + 231 pp. LCCN TK5105.55 .K57 2002.

Kusmierek:2005:SVD

[KD05] Ewa Kusmierek and David H. C. Du. Streaming video delivery over Internet with adaptive end-to-end QoS. The Journal of Systems and Software, 75(3):237–252, March 2005. CODEN JSSODM. ISSN 0164-1212.

Korper:2001:CBB

[KE01] Steffano Korper and Juanita Ellis. The E-Commerce Book: Building the E-Empire. Academic Press, New York, NY, USA, second edition, 2001. ISBN 0-12-421161-5. xiv + 248 pp. LCCN HF5548.32 .K667 2001. US$41.95.

Kenyon:2002:DNR

[Ken02a] Tony Kenyon. Data Networks: Routing, Security, and Performance Optimization. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2002. ISBN 1-55558-271-0. xvi + 807 pp. LCCN TK5105.543 .K46 2002. US$59.99.

Kenyon:2002:HPD

[Ken02b] Tony Kenyon. High Performance Data Network Design: Design Techniques and Tools. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2002. ISBN 1-55558-207-9. xiii + 623 pp. LCCN TK5105.5 .K45 2002. US$54.99.

Kereliuk:2003:IVI

[Ker03] Ryan Kereliuk. Introducing Voice Over Internet Protocol technology. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/
                                        login/2003-10/pdfs/kereliuk.
                                        pdf.

Kettell:2002:VIT

[Ket02]  Brian Kettell.  Valuation

of Internet and Technology Stocks: Implications for Investment Analysis. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002.

ISBN 0-7506-5383-3. xii + 209 pp. LCCN ???? US$74.99.

Keyes:2000:IM

[Key00] Jessica Keyes. Internet management. Best practices series. Auerbach, Princeton, NJ, USA, 2000. ISBN 0-8493-9987-4. xv + 758 pp. LCCN HD30.37 .I56 2000.

Klewitz-Hommelsen:2002:IPV

[KH02] Sayeed Klewitz-Hommelsen.

Indicators for privacy violation of Internet sites. Lecture Notes in Computer Science, 2456:456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2456/24560456.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2456/24560456.pdf.

Khan:2002:PLP

[Kha02] Ashfaq A. Khan. Practical Linux programming: device drivers, embedded systems, and the Internet. Charles River Media, Inc., Hingham, MA, USA, 2002. ISBN 1-58450-096-4 (paperback). xv + 420 pp. LCCN QA76.76.O63 K497 2002.

Kiam:2000:SRB

[KI00] Alex Kiam and Tim Ireland.

Starting and running a business on the Internet. Net-Works, Harrogate, UK, 2000.

ISBN 1-873668-78-3. 109 pp. LCCN HF5548.32 .K52 2000.

Kim:2001:HSI

[Kim01] Won Kim, editor. The Human Society and the Internet: Internet-related socioeconomic issues, First International Conference Human.Society@Internet 2001, Seoul, Korea, July 4–6, 2001: proceedings, volume 2105 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42313-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 I5345 2001; QA267.A1 L43 no.2105. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t2105.htm.

Kim:2002:NRA

[Kim02]  Youngyong Kim.  A no-

ble routing algorithm for the Internet computational GRID. Lecture Notes in Computer Science, 2402:33– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2402/24020033.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2402/24020033.
pdf.

Kipper:2007:WCF

[Kip07] Gregory Kipper. Wireless crime and forensic investigation. Auerbach Publications, Boca Raton, FL, USA, 2007. ISBN 0-8493-3188-9. xxviii + 251 pp. LCCN TK5103.2 .K565

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0647/2006040493-d.html.

Kistler:2000:ARS

[Kis00] John M. Kistler. Animal rights: a subject guide, bibliography, and Internet companion. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2000. ISBN 0-313-31231-1. xviii + 227 pp. LCCN HV4708 .K57 2000.

Kalyanaraman:2000:ESA

[KJF^{+}00] Shivkumar Kalyanaraman, Raj Jain, Sonia Fahmy, Rohit Goyal, and Bobby Vandalore. The ERICA switch algorithm for ABR traffic management in ATM networks. IEEE/ACM Transactions on Networking, 8(1):87– 98, February 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
ton/2000-8-1/p87-kalyanaraman/

.

Kim:2002:CAR

[KJL02] Hagyoung Kim, Sungin Jung, and Cheolhoon Lee. Contentadaptive request distribution strategy for Internet server cluster. Lecture Notes in Computer Science, 2402:85– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2402/24020085.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2402/24020085.
pdf.

Klusch:2000:CIA

[KK00] Matthias Klusch and Larry Kerschberg, editors. Cooperative information agents IV: the future of information agents in cyberspace: 4th International Workshop, CIA 2000, Boston, MA, USA, July 7–9, 2000: proceedings, volume 1860 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67703-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1860.

Kumar:2003:PBD

[KKT04] C. Bala Kumar, Paul J. Kline, and Timothy J. Thompson. Bluetooth application programming with the Java APIs. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2004. ISBN 1-55860-934-2. xxii + 498 pp. LCCN TK5103.3 .K86 2004. US$44.95. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els051/2003107476.
html;   http://www.loc.gov/
catdir/toc/els051/2003107476.
html.

Kim:2004:IRI

[KKYK04] Won Kim, Il-Ju Ko, Jin-Sung Yoon, and Gye-Young Kim. Inference of recommendation information on the Internet using improved FAM. Future Generation Computer Systems, 20(2):265–273, February 16, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Klein:2002:SEO

[Kle02]  Michel Klein.  Supporting

evolving ontologies on the Internet. Lecture Notes in Computer Science, 2490:597– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2490/24900597.htm;
http://link.springer.de/
link/service/series/0558/
papers/2490/24900597.pdf.

Klevinsky:2002:HST

[KLG02] T. J. Klevinsky, Scott Laliberte, and Ajay Gupta. Hack I.T.: security through penetration testing. Addison-Wes-ley, Reading, MA, USA, 2002.

ISBN 0-201-71956-8 (paperback). xxiv + 512 pp. LCCN QA76.9.A25 K56 2002.

Keen:2000:EEC

[KM00] Peter G. W. Keen and Mark McDonald. The eProcess edge: creating customer value and business wealth in the Internet era. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212626-4. xx + 300 pp. LCCN HD30.28 .K44 2000.

Kaye:2001:JCA

[KM01a] Barbara K. Kaye and Norman J. Medoff. Just a click away: advertising on the Internet. Allyn and Bacon, Inc., Needham Heights, MA, USA, 2001. ISBN 0-205-31875-4. x + 146 pp. LCCN HF6146.I58 K39 2001.

Keen:2001:FEG

[KM01b] Peter G. W. Keen and Ron Mackintosh. The freedom economy: gaining the Mcommerce edge in the era of the wireless Internet. Osborne/ McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213367-8. xii + 259 pp. LCCN HF5548.32 .K428 2001.

Kevan:2001:MIL

[KM01c] Tim Kevan and Paul Mc-Grath. E-mail, the Internet and the law: essential knowledge for safer surfing. EMIS Professional Publishing, Welwyn Garden City, Hertfordshire, UK, 2001. ISBN 1-85822-268-0, 1-85811-268-

0. xvi + 223 pp. LCCN KD667.C65 K48 2001.

Kumar:2004:CNA

[KMK04] Anurag Kumar, D. Manjunath, and Joy Kuri. Communication networking: an analytical approach. The Morgan Kaufmann series in networking. Elsevier/Morgan Kaufmann, San Francisco, CA, USA, 2004. ISBN 0-12-428751-

4. xxvii + 929 pp. LCCN TK5101 .K865 2004. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els051/2004274429.
html;   http://www.loc.gov/
catdir/toc/els051/2004274429.
html.

Knuckles:2001:IIP

[Knu01] Craig D. Knuckles. Introduction to interactive programming on the Internet: using HTML and JavaScript. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-38366-X (paperback). xviii + 423 pp. LCCN QA76.625 .K58 2001.

Kohavi:2002:WML

[Koh02] Ron Kohavi, editor. WE-BKDD 2001 — mining log data across all customer touch points: third international workshop, San Francisco, CA, USA, August 26, 2001: revised papers, volume 2356 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-43969-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ZA4235 .W43 2001 Bar.

Kitaguchi:2002:DHA

[KOS^{+}02] Yoshiaki Kitaguchi, Haruo Okazawa, Syunsuke Shinomiya, Yutaka Kidawara, Kaisuya Hakozaki, and Shin ichi Nakagawa. Development of a high-accurate time server for measurements of the Internet. Lecture Notes in Computer Science, 2344:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2344/23440351.htm;
http://link.springer.de/
link/service/series/0558/
papers/2344/23440351.pdf.

Kozlowski:2000:AGI

[Koz00] Rob Kozlowski. The actor’s guide to the Internet. Heinemann, Portsmouth, NH, USA, 2000. ISBN 0-325-00162-6. vii + 136 pp. LCCN PN2071.C58 K69 2000.

Kozlowski:2001:IGL

[Koz01] Ken Kozlowski. The Internet guide for the legal researcher: the complete resource guide to finding legal information on the Internet. Infosources Pub., Teaneck, NJ, USA, third edition, 2001. ISBN 0-939486-63-6. iv + 360 pp. LCCN KF242.A1 K69 2001 Ref. URL http:

//www.infosourcespub.com/
legresearchers.cfm.

Kozierok:2005:TIG

[Koz05] Charles Kozierok. The TCP/ IP guide: a comprehensive, illustrated Internet protocols reference. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-047-X. 1648 (est.) pp. LCCN TK5105.585 .K69 2005. US$79.95, CDN$107.99. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip0417/2004008863.
html;   http://www.oreilly.
com/catalog/159327047X/.

Kuglin:2001:SCN

[KR01a] Fred A. Kuglin and Barbara A. Rosenbaum. The supply chain network @ Internet speed: preparing your company for the E-commerce revolution. AMACOM, New York, NY, USA, 2001. ISBN 0-8144-0595-9. xxiv + 246 pp. LCCN HD39.5 .K83 2001.

Kurose:2001:CNT

[KR01b] James F. Kurose and Keith W.

Ross. Computer networking: a top-down approach featuring the Internet. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-47711-4. xxiv + 712 pp. LCCN TK5105.875.I57 K876 2001.

Kurose:2003:CNT

[KR03] James F. Kurose and Keith W. Ross. Computer Networking: a Top-Down Approach Featuring the Internet. Addison-Wes-ley, Reading, MA, USA, second edition, 2003. ISBN 0-201-97699-4. xxvi + 752 pp. LCCN TK5105.875.I57 K88 2002.

Kurose:2005:CNT

[KR05] James F. Kurose and Keith W. Ross. Computer Networking: a Top-Down Approach Featuring the Internet. Addison-Wes-ley, Reading, MA, USA, third edition, 2005. ISBN 0-321-22735-2. ???? pp. LCCN TK5105.875.I57 K88 2005.

Kurose:2008:CNT

[KR08] James F. Kurose and Keith W. Ross. Computer networking: a top-down approach. Pearson/Addison Wesley, Boston, MA, USA, fourth edition, 2008. ISBN 0-321-49770-

8. xxiv + 852 pp. LCCN TK5105.875.I57 K88 2008.

                                        URL http://www.loc.gov/
                                        catdir/toc/ecip078/2007002094.
                                        html.

Kretchmar:2003:OSN

[Kre03]  James Kretchmar.  Open

Source Network Administration. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-046210-1. 256 (est.) pp. LCCN TK5105.5 .K74 2003. US$44.99.

Kroeker:2000:PCL

[Kro00]  Kirk L. Kroeker.  Prod-

ucts: Controlling the look and feel of Windows; sophisticated project estimation for $99; Web applications for the Internet-challenged; network-based digital asset manager; Linux middleware framework; XML-based information portal; Java development for small computing devices; build installations for Windows 2000. Computer, 33(3):105–107, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co2000/pdf/r3105.
pdf.

Katchaounov:2002:OOM

[KRZ02] Timour Katchaounov, Tore Risch, and Simon Zürcher. Object-oriented mediator queries to Internet search engines. Lecture Notes in Computer Science, 2426:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2426/24260176.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2426/24260176.pdf.

Kuhner:2000:MLF

[KS00] Anja Kuhner and Thilo Sturm.

Das Medien-Lexikon: Fachbegriffe von A–Z aus Print, Radio, TV und Internet. (German) [The media encyclopedia: Technical terms of A–Z from print, radio, TV and Internet]. Verlag Moderne Industrie, Landsberg, Germany, 2000. ISBN 3-478-37680-7. 288 pp. LCCN P87.5 .K84 2000.

Kare-Silver:2001:SNR

[KS01] Michael De Kare-Silver. Eshock: the new rules: Internet strategies for retailers and manufacturers. AMACOM, New York, NY, USA, second edition, 2001. ISBN 0-8144-7116-1. xiv + 335 + 12 pp. LCCN HF5548.325.U6 D45x 2001.

Kotsis:2005:WSM

[KS05] Gabriele Kotsis and Otto Spaniol, editors. Wireless systems and mobility in next generation Internet: first international workshop of EURO-NGI network of excellence, Dagstuhl Castle, Germany, June 7–9, 2004. Revised selected papers, volume 3427 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-25329-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5103.2 .I59 2004. URL http://www.

springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3427;   http:
//www.springerlink.com/openurl.

asp?genre=volume&id=doi:
10.1007/b107131.

Keidl:2002:SDS

[KSSK02] Markus Keidl, Stefan Seltzsam, Konrad Stocker, and Alfons Kemper. ServiceGlobe: Distributing E-services across the Internet. In Bernstein et al. [B^{+}02], pages 1047–1050. ISBN 1-55860-869-9. LCCN

???? URL http://www.vldb.
org/conf/2002/S33P02.pdf.

Kobayashi:2000:IRW

[KT00] Mei Kobayashi and Koichi Takeda. Information retrieval on the web. ACM Computing Surveys, 32(2):144–173, 2000. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

http://www.acm.org/pubs/
articles/journals/surveys/
2000-32-2/p144-kobayashi/
p144-kobayashi.pdf;   http:
//www.acm.org/pubs/citations/
journals/surveys/2000-32-
2/p144-kobayashi/.

Kuo:2001:DBM

[Kuo01] J. David Kuo. Dot.bomb: my days and nights at an Internet Goliath. Little, Brown, Boston, MA, USA, 2001. ISBN 0-316-50749-0. 311 pp. LCCN HF5548.325.U6 K86 2001. BUY-A-BOOK gift of Penny Gillam.

Kuribayashi:2000:AUY

[Kur00a] Akira Kuribayashi. Anguraintanetto ura yogo jiten = A dictionary for Internet’s underground world. Foresuto Shuppan, Tokyo, Japan, shohan edition, 2000. ISBN 4-89451-096-0. 318 pp. LCCN TK5105.875.I57 K87 2000. In Japanese.

Kurniawan:2000:IPV

[Kur00b] Budi Kurniawan. Internet programming with Visual Basic. Books for professionals by professionals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-893115-75-5. xiii + 521 pp. LCCN QA76.625 .K87 2000.

Kurbel:2002:VTL

[Kur02] Karl Kurbel. Virtual teaching and learning based on multimedia techniques: Experiences from a master program delivered over the Internet. Lecture Notes in Computer Science, 2436:252– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2436/24360252.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2436/24360252.
pdf.

Kutais:2001:IPI

[Kut01] B. G. Kutais. Internet policies and issues, volume II. Nova Science Publishers, Huntington, NY, USA, 2001. ISBN 1-56072-927-9. vi + 182 pp. LCCN TK5105.875.I57 I5692 2001.

Kahin:2000:IPB

[KV00] Brian Kahin and Hal R. Varian. Internet publishing and beyond: the economics of digital information and intellectual property. Harvard Information Infrastructure Project. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-61159-7 (paperback). vi + 243 pp. LCCN Z286.I57 I58 2000. “A publication of the Harvard Information Infrastructure Project in collaboration with the School of Information Management and Systems at the University of California at Berkeley”.

Krutz:2008:CPG

[KV08] Ronald L. Krutz and Russell Dean Vines. The CEH prep guide: the comprehensive guide to certified ethical hacking. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2008. ISBN 0-470-13592-1 (hardcover. xxviii + 738 pp. LCCN QA76.9.A25 K785

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0741/2007033354-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0741/2007033354-
t.html;   http://www.loc.
gov/catdir/enhancements/
fy0743/2007033354-b.html.

Contains CD-ROM.

Kotok:2002:ENG

[KW02] Alan Kotok and David R. R. Webber. ebXML: the new global standard for doing business over the Internet. New Riders Publishing, Carmel, IN, USA, 2002. ISBN 0-7357-1117-8. xxxix + 339 pp. LCCN QA76.76.H94 K686 2002.

Kahn:2006:FDD

[KW06] Robert Kahn and Robert Wilensky. A framework for distributed digital object services. International Journal on Digital Libraries, 6(2): 115–123, April 2006. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s00799-005-0128-x.

Klusch:2001:CIA

[KZ01] Matthias Klusch and Franco Zambonelli, editors. Cooperative information agents V: 5th international workshop, CIA 2001, Modena, Italy, September 6–8, 2001: proceedings, volume 2182 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42545-4 (paperback). LCCN Q337 .C66 2001; QA267.A1 L43 no.2182.

La:2000:CST

[LA00a] Richard J. La and Venkatachalam Anantharam. Chargesensitive TCP and rate control in the Internet. Memorandum UCB/ERL M00/13, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, Berkeley, CA, March 24, 2000. i + 31 pp.

La:2000:CTR

[LA00b] Richard J. La and Venkatachalam Anantharam. Chargesensitive TCP and rate control in the Internet. Technical report, Electronics Research Laboratory, College of Engineering, University of California, March 24, 2000. i + 31 pp.

Lindstrom:2000:BBI

[LA00c] Martin Lindstrom and Tim Frank Andersen. Brand building on the Internet. Kogan Page, Dover, NH, USA, 2000. ISBN 0-7494-3313-2. 319 pp. LCCN HF5415.1265 .L56 2000.

La:2002:UBR

[LA02] Richard J. La and Venkat Anantharam. Utility-based rate control in the Internet for elastic traffic. IEEE/ACM Transactions on Networking, 10(2):272–286, April 2002.

Labiod:2007:WAH

[Lab07] Houda Labiod, editor. Wireless ad hoc and sensor networks. ISTE, London, UK, 2007. ISBN 1-905209-86-X. 352 (est) pp.

Labovitz:2001:DIR

[LABJ01] Craig Labovitz, Abha Ahuja, Abhijit Bose, and Farnam Jahanian. Delayed Internet routing convergence. IEEE/ACM Transactions on Networking, 9 (3):293–306, June 2001. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Lai:2001:LTP

[Lai01] Kwok-Wing Lai. E-learning: teaching and professional development with the Internet. University of Otago Press, Dunedin, New Zealand, 2001. ISBN 1-877276-03-0. 206 pp. LCCN LB1044.87 .E43 2001.

Langford:2000:IEa

[Lan00a] Duncan Langford. Internet ethics. St. Martin’s Press, New York, NY, USA, 2000. ISBN 0-312-23279-9. xiv + 257 pp. LCCN TK5105.875.I57 I547 2000.

Langford:2000:IEb

[Lan00b] Duncan Langford. Internet ethics. Macmillan Publishing Company, New York, NY, USA, 2000. ISBN 0-333-77626-7. xiv + 281 pp. LCCN TK5105.875.I57 I547 2000b.

Landau:2005:SWI

[Lan05a] S. Landau. Security, wiretapping, and the Internet. IEEE Security & Privacy, 3(6): 26–33, November/December 2005. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8013/33104/01556533.pdf;
http://ieeexplore.ieee.
org/xpls/abs_all.jsp?isnumber=
33104&arnumber=1556533.

Langridge:2005:DUM

[Lan05b] Stuart Langridge. DHTML utopia: modern Web design using JavaScript and DOM. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2005. ISBN 0-9579218-9-6. xiv + 318 pp. LCCN TK5105.888 .L36 2005.

Landoll:2006:SRA

[Lan06] Douglas J. Landoll. The security risk assessment handbook: a complete guide for performing security risk assessments. Auerbach Publications, Boca Raton, FL, USA, 2006. ISBN 0-8493-2998-1. xxi + 473 pp. LCCN HF5548.37 .L358

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0647/2005050717-d.html.

Lavin:2006:OOP

[Lav06] Peter Lavin. Object-oriented PHP: concepts, techniques, and code. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-077-1. xx + 190 pp. LCCN QA76.73.P224 L38 2006. URL http:/

/www.loc.gov/catdir/toc/
ecip0613/2006015309.html;
http://www.nostarch.com/
frameset.php?startat=oophp.

Lawrence:2000:LIP

[Law00] Penelope Lawrence. Law on the Internet: a practical guide. Sweet and Maxwell, London, UK, 2000. ISBN 0-421-73780-8. ix + 227 pp. LCCN KD394 .L39 2000.

Lawson:2004:VSW

[Law04] K. Lawson. Visualizing the Semantic Web: XML-based Internet and Information Visualization [book review]. IEEE Distributed Systems Online, 5(2):6/1–6/3, February 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL http://

ieeexplore.ieee.org/iel5/
8968/28454/01270786.pdf?
isnumber=28454&prod=JNL&
arnumber=1270786&arSt=+6%
2F1&ared=+6%2F3&arAuthor=
+Lawson%2C+K.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=28454&
arnumber=1270786&count=6&
index=4.

Layton:2006:ISD

[Lay06] Timothy P. Layton. Information security: design, implementation, measurement, and compliance. Auerbach Publications, Boca Raton, FL, USA, 2006. ISBN 0-8493-7087-6. 264 (est.) pp. LCCN HF5548.37 .L393

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0661/2006047620-d.html.

Lee:2002:ACC

[LBJ^{+}02] B. P. Lee, R. K. Balan, L. Jacob, W. K. G. Seah, and A. L. Ananda. Avoiding congestion collapse on the Internet using TCP tunnels. Computer Networks (Amsterdam, Netherlands: 1999), 39(2):207–219, June 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/49/33/
abstract.html.

Leiner:2009:BHI

[LCC^{+}09] Barry M. Leiner, Vinton G.

Cerf, David D. Clark, Robert E. Kahn, Leonard Kleinrock, Daniel C. Lynch, Jon Postel, Larry G. Roberts, and Stephen Wolff. A brief history of the Internet. Computer Communication Review, 39(5):22–31, October 2009. CODEN CCRED2. ISSN 0146-4833 (print), 1943-5819 (electronic). URL https:

//dl.acm.org/doi/10.1145/
1629607.1629613.

Liu:2008:HPP

[LCH^{+}08] Duo Liu, Zheng Chen, Bei Hua, Nenghai Yu, and Xinan Tang. High-performance packet classification algorithm for multithreaded IXP network processor. ACM Transactions on Embedded Computing Systems, 7(2):16:1–16:??, February 2008. CODEN ???? ISSN 1539-9087.

Lee:2004:DEC

[LCL04] Choongseok Lee, Byounggu Choi, and Heeseok Lee. A development environment for customer-oriented Internet business: eBizBench. The Journal of Systems and Software, 72(2):163–178, July 2004. CODEN JSSODM. ISSN 0164-1212.

Champion:2000:TPT

[LD00] Remy Le Champion and Benoit Danard. Télévision de pénurie, télévision d’abondance: des origines à Internet. (French) [Television of poverty, television of plenty: the origins of the Internet], volume 5110-11 of Notes et études documentaires. Documentation Française, Paris, France, April 2000. ISBN 2-11-004458-6. 224 pp. LCCN HE8700.9.F7 T35 2000.

Lu:2003:SRC

[LD03] Dong Lu and Peter August Dinda. Synthesizing realistic computational grids. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

                                        ????  URL http://www.
                                        sc-conference.org/sc2003/
                                        inter_cal/inter_cal_detail.
                                        php?eventid=10690#0;   http:
                                        //www.sc-conference.org/
                                        sc2003/paperpdfs/pap180.
                                        pdf.

Lear:2000:NBP

[Lea00] Anne C. Lear. News briefs: PalmOS gains upper hand in handheld market; Internet viruses get more dangerous; standard offers interim net address solution; optical router promises faster net; translation technologies advance. Computer, 33(1):24–26, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co2000/pdf/r1024.
pdf.

Lee:2001:IMR

[Lee01] Ook Lee. Internet marketing research: theory and practice. Idea Group Pub., Hershey, PA, USA, 2001. ISBN 1-878289-97-7 (paperback). ii + 261 pp. LCCN HF5415.1265 .I569 2001.

Lehtisalo:2005:HNT

[Leh05] Kaarina Lehtisalo. The history of NORDUnet: Twenty-five years of networking cooperation in the Nordic countries. Web document, June 7, 2005.

URL https://nordu.net/
wp-content/uploads/2023/
11/TheHistoryOfNordunet_
simple.pdf.

Leiviska:2000:ISH

[Lei00] Karen Leiviska. Internet scavenger hunts for the topics you teach. Scholastic Professional Books, New York, NY, USA, 2000. ISBN 0-439-17034-6. 64 pp. LCCN LB1044.87 .L45

Lejeune:2000:CEJ

[Lej00] Philippe Lejeune. “Cher écran —”: journal personnel, ordinateur, Internet. (French) [“Dear screen —”: Personal diary, computer, Internet]. La couleur de la vie. Editions du Seuil, Paris, France, 2000. ISBN 2-02-041251-9. 443 pp. LCCN CT25 .L364 2000.

Lemley:2000:SIL

[Lem00] Mark A. Lemley. Software and Internet law. Aspen Law and Business, Gaithersburg, MD, USA, 2000. ISBN 0-7355-1312-0. xxviii + 1115 pp. LCCN KF3024.C6 S64 2000.

Leondes:2002:DDC

[Leo02] Cornelius T. Leondes. Database and Data Communication Network Systems, Three-Volume Set: Techniques and Applications. Academic Press, New York, NY, USA, 2002. ISBN 0-12-443895-4. ???? pp. LCCN ???? US$995.00. Three volumes.

Lerdorf:2000:PPR

[Ler00a] Rasmus Lerdorf. PHP pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-769-9. vi + 114 pp. LCCN QA76.73.P224 L47 2000.

Lerner:2000:MNC

[Ler00b] Michah Lerner. Middleware networks: concept, design, and deployment of Internet infrastructure. Kluwer international series on advances in database systems. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-7840-7. xxiv + 375 pp. LCCN TK5105.875.I57 M537 2000.

Levy:2000:GSI

[Lev00] Peter Levy. Getting started with the Internet: an easy and practical guide for teachers. Scholastic Professional Books, New York, NY, USA, 2000. ISBN 0-439-14114-1. 64 pp. LCCN LB1044.87 .L48 2000.

Leventhal:2001:IAW

[Lev01a] Lance A. Leventhal, editor.

Internet Appliance Workshop 2001: proceedings: February 20–21, 2001, Wyndham Hotel, San Jose, California. Internet Appliance Workshop, San Diego, CA, USA, 2001. LCCN TK5105.875.I57 I536 2001.

Levy:2001:VCT

[Lev01b] Michael Levy. E-Volve-or-Die.com: thriving in the Internet age through e-commerce management. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1028-7. xvi + 318 pp. LCCN HF5548.32 .L48 2000.

Lewin:2001:UIS

[Lew01] Larry Lewin. Using the Internet to strengthen curriculum. Association for Supervision and Curriculum Development, Alexandria, VA, USA, 2001. ISBN 0-87120-511-4. vii + 137 pp. LCCN LB1044.87 .L498 2001.

Lathrop:2000:SCP

[LF00] Ann Lathrop and Kathleen Foss. Student cheating and plagiarism in the Internet era: a wake-up call. Libraries Unlimited, Englewood, CO, USA, 2000. ISBN 1-56308-841-X (paperback). xiv + 255 pp. LCCN LB3609 .L28 2000.

Laubach:2001:DIC

[LFD01] Mark E. Laubach, David J. Farber, and Stephen D. Dukes. Delivering Internet connections over cable: breaking the access barrier. Wiley Networking Council series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-38950-1 (paperback). xxii + 312 pp. LCCN TK5103.4 .L38 2001.

Lazard:2000:WCW

[LH00] Dorothy Lazard and Mana Hayakawa. Women of color on the Web: an Internet supplement to the National Directory of Women of Color Organizations and Projects. Women of Color Resource Center, Berkeley, CA, USA, 2000. 14 pp. LCCN HQ1904 .L392 2000; HQ1904 .W65 1998.

Li:2004:PAC

[LIB04] Tianshu Li, Youssef Iraqi, and Raouf Boutaba. Pricing and admission control for QoS-enabled Internet. Computer Networks (Amsterdam, Netherlands: 1999), 46(1):87– 110, September 16, 2004. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Light:2003:BIT

[Lig03] Jennifer S. Light. Before the Internet, there was cable. IEEE Annals of the History of Computing, 25(2):96, 94–95, April 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL

http://csdl.computer.org/
dl/mags/an/2003/02/a2096.
htm;   http://csdl.computer.
org/dl/mags/an/2003/02/a2096.
pdf.

LBST:2006:WND

[Lim06] Limehouse Book Sprint Team.

Wireless Networking in the Developing World: a practical guide to planning and building low-cost telecommunications infrastructure. ????, ????, January 2006. ISBN ???? vi + 244 pp. LCCN ???? URL

http://wndw.net/.

Linten:2000:IBI

[Lin00] Allison J. Linten. The Internet: a bibliography with indexes. Nova Science Publishers, Huntington, NY, USA, 2000. ISBN 1-56072-813-2. 368 pp. LCCN ZA4195.5 .L56 2000.

Lipschultz:2000:FEA

[Lip00] Jeremy Harris Lipschultz. Free expression in the age of the Internet: social and legal boundaries. Westview Press, Boulder, CO, USA, 2000. ISBN 0-8133-9108-3 (hardcover), 0-8133-9113-X (paperback). xv + 331 pp. LCCN KF4772 .L57 2000.

Litman:2001:DCP

[Lit01] Jessica Litman. Digital copyright: protecting intellectual property on the Internet. Prometheus Books, Amherst, NY, USA, 2001. ISBN 1-57392-889-5 (hardcover). 208 pp. LCCN KF3030.1 .L58 2001.

Liu:2001:RII

[Liu01] Lewis-Guodo Liu. The role and impact of the Internet on library and information services, volume 96 of Contributions in librarianship and information science. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2001. ISBN 0-313-30920-5. ISSN 0084-9243. xvi + 334 pp. LCCN Z674.75.I58 R65 2001.

Luengo-Jones:2001:AOW

[LJ01] Steve Luengo-Jones. ‘All-to-one’: the winning model for marketing in the post-Internet economy. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-709799-8. xvii + 278 + 1 pp. LCCN HF5415.1265 .L84 2001.

Lee:2002:DIH

[LKK02] Jonghwan Lee, Yongho Kim, and Kyongsok Kim. Development of Internet home server to control information appliances remotely. Lecture Notes in Computer Science, 2343:581–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2343/23430581.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2343/23430581.pdf.

Sung:2005:BOI

[LKKS05] Ju Yong Lee, Sunggon Kim, Deokseong Kim, and Dan Keun Sung. Bandwidth optimization for Internet traffic in generalized processor sharing servers. IEEE Transactions on Parallel and Distributed Systems, 16 (4):324–334, April 2005. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Lee:2002:PIL

[LKMK02] Sangmoon Lee, Sinjun Kang, Byungseok Min, and Hagbae Kim. A personal Internet live-broadcasting system. Lecture Notes in Computer Science, 2402:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2402/24020263.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2402/24020263.
                                        pdf.

Lee:2002:AAP

[LKP02] ByungGil Lee, HyunGon Kim, and KilHoum Park. An AAA application protocol design and service for secure wireless Internet gateway roaming. Lecture Notes in Computer Science, 2344:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2344/23440123.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2344/23440123.pdf.

Larson:2001:DW

[LL01] Matt Larson and Cricket Liu. DNS on Windows 2000. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2001. ISBN 0-596-00230-0. xvi + 334 pp. LCCN TK5105.875.I57 A44 2001. US$39.95. URL http://

                                        safari.oreilly.com/0596002300;
                                         http://www.oreilly.com/
                                        catalog/dnswin2.

Llana:2000:CWT

[Lla00] Andres Llana. Convergence of wireless technology and enterprise networks: integrating the Internet. Computer Technology, Charleston, SC, USA, 2000. ISBN 1-56607-079-1 (paperback). iv + 127 pp. LCCN TK5105.875.I57 L55 2000.

Lee:2003:CIC

[LLL03] Younghwa Lee, Zoonky Lee, and Kai R. T. Larsen. Coping with Internet channel conflict. Communications of the ACM, 46(7):137–142, July 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lloyd:2000:CTL

[Llo00] Frank Lloyd. Cable television law, 2000: competition in video, Internet and telephony, volume G-593 of Intellectual property course handbook series; no. G-593. Practising Law Institute, New York, NY, USA, 2000. ISBN 0-87224-743-0. ???? pp. LCCN KF2844.Z9 C315 2000. Prepared for distribution at the Cable Television Law 2000: Competition in Video, Internet and Telephony Program, New York City, March 2-3, 2000, San Francisco, April 17-18, 2000.

Lloyd:2001:IQCa

[Llo01a] Kellie Lloyd. Internet Quests:

creepy crawlies. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3408-3. 48 pp. LCCN QL467.2 .L46 2001.

Lloyd:2001:IQCb

[Llo01b] Kellie Lloyd. Internet Quests: cultures of the world. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3410-5. 48 pp. LCCN CB429 .L46 2001.

Lloyd:2001:IQF

[Llo01c] Kellie Lloyd. Internet Quests: folk tales and tall tales. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3409-1. 48 pp. LCCN PZ8.1.L56 2001.

Lloyd:2001:IQN

[Llo01d] Kellie Lloyd. Internet quests: Native Americans. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3412-1. 48 pp. LCCN E77.4 .L56 2001.

Liu:2003:AVM

[LLZ03] Jiangchuan Liu, Bo Li, and Ya-Qin Zhang. Adaptive video multicast over the Internet. IEEE MultiMedia, 10(1):22–33, January–March 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

                                        /csdl.computer.org/comp/
                                        mags/mu/2003/01/u1022abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/01/u1022.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mu/2003/01/u1022.
                                        pdf.

Laurence:2000:ARI

[LM00] Helen Laurence and William Miller. Academic research on the Internet: options for scholars and libraries. Haworth Information Press, New York, NY, USA, 2000. ISBN 0-7890-1176-X, 0-7890-1177-8 (paperback). ISSN 0193-0826. 386 (or 378??) pp. LCCN ZA4201 .A216 2000. Published also as v.30, no. 1/2 and 3/4, 2000 of the Journal of Library Administration.

Lehr:2001:IT

[LMC01] William Lehr, Lee W. McKnight, and David D. Clark. Internet telephony. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-13385-7. xii + 395 pp. LCCN TK5105.8865 .I57 2001. US$39.95, UK£27.50.

Laahs:2002:MSP

[LMV02] Kevin Laahs, Emer McKenna, and Don Vickers. Microsoft SharePoint portal server: building knowledgesharing applications. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2002. ISBN 1-55558-244-3. ???? pp. LCCN TK5105.875.I6L33 2002. US$54.99.

Lockwood:2003:GEI

[Loc03] John W. Lockwood. Guest Editor’s introduction: Hot Interconnects 10—thinking beyond the Internet. IEEE Micro, 23(1):8–9, January/February 2003. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        http://dlib.computer.org/
                                        mi/books/mi2003/pdf/m1008.
                                        pdf;   http://www.computer.
                                        org/micro/mi2003/m1008abs.
                                        htm.

Logan:2000:SLL

[Log00] Robert K. Logan. The sixth language: learning a living in the Internet age. Stoddart, Toronto, ON, Canada, 2000. ISBN 0-7737-6104-7 (paperback), 0-7737-3251-9 (bound). x + 318 pp. LCCN P96.T42 L64 2000. US$32.95.

Long:2000:LRU

[Lon00] Judy A. Long. Legal research using the Internet. West Legal Studies/Thomson Learning, Albany, NY, USA, 2000. ISBN 0-7668-1335-5. xiv + 119 pp. LCCN KF242.A1 L66 2000.

Loshin:2000:BBIc

[Los00a] Peter Loshin. Big book of Internet file transfer RFCs. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-12-455845-3. various pp. LCCN TK5105.55 .L67 2000.

Loshin:2000:BBIb

[Los00b] Peter Loshin. Big book of Internet host standards. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-12-455844-5. various pp. LCCN TK5105.875.I57 B49 2000.

Loshin:2000:BBIa

[Los00c] Peter Loshin. Big book of IPsec RFCs. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-12-455839-9. various pp. LCCN TK5105.567.L68 2000.

Loshin:2003:ITP

[Los03a] Pete Loshin. IPv6: theory, protocol, and practice. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-810-9. xxvi + 709 pp. LCCN TK5105.585.L66

2003.  US$49.95.  URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els031/2002113772.
html;   http://www.loc.gov/
catdir/toc/els031/2002113772.
html.

Loshin:2003:TIC

[Los03b] Pete Loshin. TCP/IP Clearly Explained. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, fourth edition, 2003. ISBN 1-55860-782-X. xxvi + 709 pp. LCCN TK5105.585 .L668

2003.  US$49.95.  URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els031/2002113772.
html;   http://www.loc.gov/
catdir/toc/els031/2002113772.
html.

Loshin:2004:ITP

[Los04] Peter Loshin. IPv6: theory, protocol, and practice. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2004. ISBN 1-55860-810-9 (paperback). xxiv + 536 pp. LCCN TK5105.585 .L66 2004. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/els041/2003058303.
                                        html;   http://www.loc.gov/
                                        catdir/toc/els041/2003058303.
                                        html.

Louis:2001:MCC

[Lou01] P. J. Louis. M-commerce crash course: the technology and business of next generation Internet services. McGraw-Hill telecommunications. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-136994-5. xii + 345 pp. LCCN HF5548.32 .L68 2001.

Lowe:2005:HNM

[Low05] Scott Lowe. Home networking: the missing manual. The missing manual. Pogue Press, Sebastopol, CA, USA; Beijing, China, 2005. ISBN 0-596-00558-X (paperback). xvi + 245 pp. LCCN TK5105.75 .L69 2005.

Lee:2002:SBF

[LPK02] Changho Lee, Jaehyun Park, and Yoosung Kim. A SOAPbased framework for the Internetworked distributed control systems. Lecture Notes in Computer Science, 2402:195– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2402/24020195.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2402/24020195.
pdf.

LeVitus:2001:IID

[LPM01] Bob LeVitus and Natanya Pitts-Moultis. The Internet for iMacs for dummies. – For dummies. I D G Books Worldwide, Indianapolis, IN, USA, 2001. ISBN 0-7645-0796-

6. xxii + 371 pp. LCCN TK5105.875.I57 L397 2001.

Lacoste:2004:SSE

[LPSW04] Gérard Lacoste, Birgit Pfitzmann, Michael Steiner, and Michael Waidner. SEMPER— Secure Electronic Marketplace for Europe, volume 1854 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-67825-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xviii + 350 pp. LCCN HF5548.32 .S46 2000. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1854.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-

                                        9743&volume=1854;   http:
                                        //www.springerlink.com/openurl.
                                        asp?genre=volume&id=doi:
                                        10.1007/b75215.

Litan:2001:BDC

[LR01a] Robert E. Litan and Alice M. Rivlin. Beyond the dot.coms: the economic promise of the Internet. Brookings Institution Press, Washington, DC, USA, 2001. ISBN 0-8157-0002-4 (hardcover). ix + 130 pp. LCCN HE7583.U6 L58 2001. Contents: Is the Internet a big deal? — Why productivity is so important — Productivity and the Internet: what lies ahead? — Benefits of the Internet that may not show up in the GDP — Realizing the Internet’s potential — Concluding thoughts.

Litan:2001:EPI

[LR01b] Robert E. Litan and Alice M. Rivlin. The economic payoff from the Internet revolution. Internet Policy Institute/Brookings Institution Press, Washington, DC, USA, 2001. ISBN 0-8157-0065-2 (paperback). viii + 292 pp. LCCN HC 79.T4 E25 2001; HC79.T4 E25283 2001. “Brookings Task Force on the Internet.” Contents: The economy and the Internet: what lies ahead? / Robert E. Litan and Alice M. Rivlin — Manufacturing: lowering boundaries, improving productivity / Andrew McAfee — Automotive industry: Internet-driven innovation and economic performance / Charles H. Fine and Daniel M. G. Raff — Financial services: transparency, differential pricing, and disintermediation / Eric K. Clemons and Lorin M. Hitt — Trucking industry: challenges to keep pace / Anuradha Nagarajan, Enrique Canessa, Will Mitchell, and C. C. White III — Retail services: continuing the Internet success / Joseph P. Bailey — Health care: competition and productivity / Patricia M. Danzon and Michael Furukawa — Public sector: early stage of a deep transformation / Jane Fountain, with Carlos A. Osorio-Urzua — Higher education: promises for future delivery — Austan Goolsbee.

Lightman:2002:BNU

[LR02] Alex Lightman and William Rojas. Brave new unwired world: the digital big bang and the infinite Internet. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-44110-4 (hardcover). xx + 314 pp. LCCN HM851 .L54 2002.

Luo:2003:CFB

[LRJ^{+}03] Yue Luo, Juan Rubio, Lizy Kurian John, Pattabi Seshadri, and Alex Mericas. Cover feature: Benchmarking Internet servers on superscalar machines. Computer, 36(2):34–40, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/co/2003/02/r2034.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/02/r2034.
                                        pdf;   http://www.computer.
                                        org/computer/co2003/r2034abs.
                                        htm.

Low:2000:CID

[LS00] Su-Lin Low and Ron Schneider. CDMA Internetworking: Deploying the Open A-Interface. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-088922-9. xxxiii + 616 pp. LCCN TK5103.452.L69 2000. US$75.00. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130889229.html.

Lach-Smith:2001:PBD

[LS01a] Barbara Lach-Smith. Programmer’s bookshelf: Developing for wireless systems. Dr. Dobbs Journal, 26(10): 130–131, October 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/.

Leamer:2001:EGI

[LS01b] Edward E. Leamer and Michael Storper. The economic geography of the Internet age. NBER working paper series 8450, National Bureau of Economic Research, Cambridge, MA, USA, August 2001. 34 pp.

Lin:2004:MAU

[LSC^{+}04] Tsung-Ching Lin, Yeali S.

Sun, Shi-Chung Chang, Shao-I Chu, Yi-Ting Chou, and Mei-Wen Li. Management of abusive and unfair Internet access by quota-based priority control. Computer Networks (Amsterdam, Netherlands: 1999), 44(4):441–462, March 15, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Lemberg:2001:CPC

[LSS01] Ingrid Lemberg, Bernhard Schroder, and Angelika Storrer. Chancen und Perspektiven computergestutzter Lexikographie: Hypertext, Internet und SGML/XML für die Produktion und Publikation digitaler Worterbucher. (German) [Opportunities and perspectives of computer-aided lexicography: Hypertext, Internet and SGML/XML for the production and publication of digital dictionaries], volume 107 of Lexicographica. Niemeyer, Tübingen, Germany, 2001. ISBN 3-484-39107-3 (paperback). ISSN 0175-9264. vi + 270 pp. LCCN P327.5.T43 C43 2001; P327.5.T43 C436 2001. In German; abstracts in English and French.

Lam:2001:ICM

[LT01a] Calvin K. M. Lam and Bernard C. Y. Tan. The Internet is changing the music industry. Communications of the ACM, 44(8):62, August 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
articles/journals/cacm/2001-
44-8/p62-lam/p62-lam.html;
 http://www.acm.org/pubs/
articles/journals/cacm/2001-
44-8/p62-lam/p62-lam.pdf;
 http://www.acm.org/pubs/
citations/journals/cacm/
2001-44-8/p62-lam/.

LeBoudec:2001:NCT

[LT01b] Jean-Yves Le Boudec and Patrick Thiran. Network calculus: a theory of deterministic queuing systems for the Internet, volume 2050 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42184-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). xix + 274 pp. LCCN QA267.A1 L43 no.2050; TK5105.875.I57 L455 2001.

Loomis:2002:FIU

[LT02] David G. Loomis and Lester D. Taylor. Forecasting the Internet: understanding the explosive growth of data communications, volume 39 of Topics in regulatory economics and policy series; 39. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 0-7923-7546-7. xiv + 250 pp. LCCN TK5105.875.I57 F67 2002.

Lucas:2002:SEC

[Luc02] Henry C. Lucas. Strategies for electronic commerce and the Internet. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-12242-1 (hardcover). xiv + 265 pp. LCCN HF5548.32 .L83 2002.

Lugaresi:2000:IPP

[Lug00] Nicola Lugaresi. Internet, privacy e pubblici poteri negli Stati Uniti, volume 200 of Seminario giuridico della Universita di Bologna; 200. A. Giuffre, Milano, Italy, 2000. ISBN 88-14-08202-2. ix + 262 pp. LCCN KF1262.A73 L84 2000.

Lusa:2000:NMH

[Lus00] John M. Lusa, editor. The Network Manager’s Handbook. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2000. ISBN 0-8493-9841-X. xv + 572 pp. LCCN TK5105.5 .N48 2000. US$79.95.

Lea:2003:IIC

[LV03] Doug Lea and Steve Vinoski. IEEE Internet computing: Guest Editors’ introduction: Middleware for Web services. IEEE Distributed Systems Online, 4(1), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0301/f/w1gei.htm.

Litan:2000:GGC

[LW00] Robert E. Litan and Peter J. Wallison. The GAAP gap: corporate disclosure in the Internet age. AEI Press, La Vergne, TN, USA, 2000. ISBN 0-8447-4147-7 (paperback). ix + 96 pp. LCCN HF5616.U5 L58 2000.

Lee:2007:BDC

[LWD07] Wenke Lee, Cliff Wang, and David Dagon, editors. Botnet detection: countering the largest security threat, volume 36 of Advances in information security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-68766-1 (hardcover), 0-387-68768-8 (ebook). xii + 167 pp. LCCN TK5105.59 .B68 2008; QA76.9.A25 B64

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0825/2007936179-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0825/2007936179-
t.html.

Long:2008:NTH

[LWPM08] Johnny Long, Jack Wiles, Scott Pinzon, and Kevin D. Mitnick. No Tech Hacking: a Guide to Social Engineering, Dumpster Diving, and Shoulder Surfing. Syngress Publishing, Inc., Rockland, MA, USA, 2008. ISBN 1-59749-215-9. xxiv + 285 pp. LCCN QA76.9.A25 L65 2008. URL http:/

/www.loc.gov/catdir/toc/
fy0804/2008273519.html.

Menasce:2002:CPW

[MA02] Daniel A. Menasce and Virgilio A. F. Almeida. Capacity Planning for Web Services: Metrics, Models, and Methods. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-065903-7. xxiii + 572 pp. LCCN QA76.9.C63 M463 2002. US$54.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130659037.html.

Makofske:2003:MSP

[MA03] David Makofske and Kevin Almeroth. Multicast sockets: practical guide for programmers. The Morgan Kaufmann practical guides series. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-846-X. xii + 180 pp. LCCN TK5105.887 .M35 2003. US$19.95. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els051/2004555479.
html;   http://www.loc.gov/
catdir/toc/els051/2004555479.
html.

Mackay:2003:PID

[Mac03] Steve Mackay. Practical Industrial Data Networks: Design, Installation and Troubleshooting. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-7506-5807-X. ???? pp. LCCN ???? US$49.99.

Maitra:2000:ISP

[Mai00] Amit K. Maitra. Internet solutions for project managers. Wiley operations management series for professionals. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2000. ISBN 0-471-33027-2 (hardcover). xxi + 217 pp. LCCN HD69.P75 M35 2000.

Maier:2006:ATL

[Mai06] Phillip Q. Maier. Audit and trace log management: consolidation and analysis. Auerbach Publications, Boca Raton, FL, USA, 2006. ISBN 0-8493-2725-3. 192 (est.) pp. LCCN TK5105.59 .M34

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0647/2005057044-d.html.

Mandel:2000:CID

[Man00a] Michael J. Mandel. The coming Internet depression: why the high-tech boom will go bust, why the crash will be worse that you think, and how to prosper afterwards. Basic Books, New York, NY, USA, 2000. ISBN 0-465-04358-5. xi + 167 pp. LCCN HD9696.8.U62 M36 2000.

Mandel:2000:IDB

[Man00b] Michael J. Mandel. The Internet depression: the boom, the bust, and beyond. Basic Books, New York, NY, USA, 2000. ISBN 0-465-04359-3. xxi + 167 pp. LCCN HD9696.8.U62 M36 2000b.

Mandel:2001:CID

[Man01] Michael J. Mandel. The coming Internet depression: why the high-tech boom will go bust, why the crash will be worse that you think, and how to prosper afterwards. Financial Times Prentice Hall, London, UK, 2001. ISBN 0-273-65322-9. xi + 167 pp. LCCN ????

Manley:2002:MWA

[Man02] Kevin T. Manley. Monitoring Web application performance and ISAPI. Dr. Dobbs Journal, 27(7):65–68, July 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_07/
isapi.txt;   http://www.
ddj.com/ftp/2002/2002_07/
isapi.zip.

Martin:2000:AIA

[Mar00] James Martin. After the Internet: alien intelligence. Capital Press, Washington, DC, USA, 2000. ISBN 0-89526-280-0. 480 pp. LCCN QA76.9.C66 M363 2000.

Marco:2001:EJJ

[Mar01] Lou Marco. EJB and JSP:

Java on the edge, unlimited edition. Professional mindware. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-4802-6. xxiv + 452 pp. LCCN QA76.73.J38 M3482 2001.

Mason:2006:PVC

[Mas06] Mike Mason. Pragmatic Version Control using Subversion. Pragmatic, Lewisville, TX, USA, second edition, 2006. ISBN 0-9776166-5-7 (paperback). 250 (est.) pp. LCCN ???? US$20.99.

Matthews:2002:IOU

[Mat02] Joseph R. Matthews. Internet outsourcing using an application service provider: a how-to-do-it manual for librarians, volume 110 of How-to-do-it manuals for libraries. Neal-Schuman Publishers, New York, NY, USA, 2002. ISBN 1-55570-422-0. ix + 127 pp. LCCN Z678.93.C67 M38 2002.

Mauriello:2001:NWU

[Mau01] Carrie Mauriello. Net Worth: Using the Internet for Personal Financial Planning. Butterworth-Heinemann, Boston, MA, USA, 2001. ISBN 1-884133-83-5. 491 pp. LCCN HG179 .M355 2001. US$39.99.

Maufer:2003:FGW

[Mau03] Thomas Maufer. A Field Guide to Wireless LANs for Administrators and Power Users. Prentice Hall PTR series in computer networking and distributed systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003.

ISBN 0-13-101406-4. 368 (est.) pp. LCCN TK5105.78 .M38 2003. US$39.99.

Maxwell:2000:HTP

[Max00] Bruce Maxwell. How to track politics on the Internet. CQ Press, Washington, DC, USA, 2000. ISBN 1-58602-472-X (paperback) (??invalid ISBN checksum??). xx + 325 pp. LCCN JK1764 .M39 2000.

McKnight:2000:IWPa

[MB00a] Lee W. McKnight and Jahangir Boroumand. Internet watch: Pricing Internet services: Approaches and challenges. Computer, 33(2):128– 129, February 2000. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.
computer.org/co/books/co2000/
pdf/r2128.pdf.

McKnight:2000:IWPb

[MB00b] Lee W. McKnight and Jahangir Boroumand. Internet watch: Pricing Internet services: Proposed improvements. Computer, 33(3): 108–109, March 2000. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.
computer.org/co/books/co2000/
pdf/r3108.pdf.

Merkow:2002:PIP

[MB02] Mark S. Merkow and Jim Breithaupt. The e-privacy imperative: protect your customers’ Internet privacy and ensure your company’s survival in the electronic age. AMACOM, New York, NY, USA, 2002. ISBN 0-8144-0628-9. xi + 267 pp. LCCN HF5548.32 .M47 2002.

Miller:2000:LTE

[MC00] Philip Miller and Michael Cummins. LAN technologies explained. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2000. ISBN 1-55558-234-6. xxv + 745 pp. LCCN TK5105.7.M555 2000.

McBride:2002:SIP

[McB02] Patricia McBride. Secure Internet practices: best practices for securing systems in the Internet and e-Business age. METASeS, Cornelius, NC, USA, 2002. ISBN 0-8493-1239-6 (paperback). xv + 209 pp. LCCN TK5105.59 .S435 2002.

McCarty:2000:IGB

[McC00a] Ron McCarty. ICMP: The good, the bad, and the ugly. Sys Admin: The Journal for UNIX Systems Administrators, 9(4):69–71, April 2000. CODEN SYADE7. ISSN 1061-

2688.  URL http://www.
samag.com/.

McCormack:2000:SSS

[McC00b] Mark H. McCormack. Staying street smart in the Internet age; what hasn’t changed about the way we do business. Viking, New York, NY, USA, 2000. ISBN 0-670-89306-4. xvii + 286 pp. LCCN HD30.37 .M388 2000.

McCabe:2003:NAA

[McC03]  James McCabe.  Network

Analysis, Architecture and Design. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2003. ISBN 1-55860-887-7. ???? pp. LCCN ???? US$64.95.

McFarlane:2003:RAD

[McF03] Nigel McFarlane. Rapid Application Development with Mozilla. Bruce Perens’ Open source series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-142343-6. 800 (est.) pp. LCCN QA76.625 .M38 2003. US$44.99.

McFarlane:2005:FHT

[McF05]  Nigel McFarlane.  Firefox

Hacks: Tips & Tools for Next-Generation Web Browsing. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00928-3. 400 (est.) pp. LCCN ???? US$24.95; CDN$34.95; UK£17.50.

McGuire:2000:IHW

[McG00]  Mary McGuire.  The In-

ternet handbook for writers, researchers, and journalists. Guilford Press, New York, NY, USA, 2000/2001 ed edition, 2000. ISBN 1-57230-550-9 (paperback). xii + 276 pp. LCCN ZA4201 .I566 2000.

                                        URL http://www.guilford.
                                        com/writernet.

McGraw:2006:SSB

[McG06] Gary McGraw. Software Security: Building Security In. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-35670-5 (paperback). ???? pp. LCCN QA76.9.A25 M4286 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip062/2005031598.html.

Mascara:2001:IRG

[MCH01] Cynthia Mascara, Patricia Czar, and Toni Hebda. Internet resource guide for nurses and health care professionals. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2001. ISBN 0-13-031103-0. vi + 106 pp. LCCN RT50.5 .M37 2001.

Myhre:2002:CBC

[MD02] Robert N. Myhre and Steven Dangerfield. CCNP 2.0: Building Cisco Remote Access Networks (BCRAN) Exam 640-505. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-093672-3. 462 (est.) pp. LCCN ???? US$49.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130936723.html.

Mead:2000:DBS

[Mea00]  Donald C. Mead.  Direct

Broadcast Satellite Communications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-201-69582-0. xv + 300 pp. LCCN TK6677 .M43 2000. US$65.00. URL http:

//www.phptr.com/ptrbooks/
ptr_0201695820.html.

Mehra:2002:ICG

[Meh02] Pankaj Mehra. Internet computing: Guest Editor’s introduction: Global deployment of data centers. IEEE Distributed Systems Online, 3(10), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0209/f/wp5gei.htm.

Mertin:2000:IIR

[Mer00]  Andreas Mertin.  Internet

im Religionsunterricht. (German) [Internet in religious education]. Vandenhoeck und Ruprecht, Göttingen, Germany, 2000. ISBN 3-525-61380-6 (paperback). 186 pp.

Merriden:2001:IFB

[Mer01] Trevor Merriden. Irresistible forces: the business legacy of Napster and the growth of the underground Internet. Capstone, Oxford, UK, 2001. ISBN 1-84112-170-3. xiii + 178 pp. LCCN ML3790 .M44 2001.

Messerschmitt:2000:UNA

[Mes00] David Messerschmitt. Understanding Networked Applications: a First Course. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-537-1. xxx + 623 pp. LCCN TK5105.5 .M46 2000. US$74.95.

Metcalfe:2000:ICO

[Met00] Robert M. Metcalfe. Internet collapses and other InfoWorld punditry. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-3503-X (paperback). xxiii + 324 pp. LCCN ZA4201 .M48 2000.

Metz:2003:IIC

[Met03] Chris Metz. IEEE Internet computing: Guest Editor’s introduction: Moving toward an IPv6 future. IEEE Distributed Systems Online, 4(5), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0305/f/w3gei.htm.

Mewton:2001:AYN

[Mew01] Conrad Mewton. All you need to know about music and the Internet revolution. Sanctuary Pub., London, UK, 2001. ISBN 1-86074-325-0. 215 pp. LCCN ML74.7 .M49 2001.

MoreiraDeSouza:2001:TEI

[MFdSeS01] Jorge Moreira De Souza, Nelson L. S. Da Fonseca, and E. de Souza e. Silva, editors. Teletraffic engineering in the Internet era: proceedings of the International Teletraffic Congress — ITC-17, Salvador da Bahia, Brazil, 24–28 September 2001, volume 4 of Teletraffic science and engineering. Elsevier Science Publishers B.V., Amsterdam, The Netherlands, 2001. ISBN 0-444-50911-9 (hardcover). LCCN TK5101.A1 I55 2001.

Male:2000:QGI

[MG00] Mary Male and Doug Gotthoffer. Quick guide to the Internet for special education. Allyn and Bacon, Inc., Needham Heights, MA, USA, 2000. ISBN 0-205-31054-0. v + 122 pp. LCCN LC3965 M25 2000.

Mohen:2001:CIV

[MG01] Joe Mohen and Julia Glidden. The case for Internet voting. Communications of the ACM, 44(1): 72, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
cacm/2001-44-1/p72-mohen/

.

McDaniel:2002:MRI

[MG02] Carl D. McDaniel. Jr. and Roger H. Gates. Marketing research: the impact of the Internet. South-Western, Cincinnati, Ohio, fifth edition, 2002. ISBN 0-324-06797-6, 0-324-12581-X (disc). xxxii + 682 + 36 + 9 pp. LCCN HF5415.2 .M382 2002.

Michaels:2000:PC

[Mic00] Laura Michaels. POP3 for C++. Dr. Dobbs Journal, 25 (12):109–110, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2000/2000_12/
                                        cpppop3.zip.

Millan:2001:IE

[Mil01a] Jose Antonio Millan. Internet y el español. (Spanish) [Internet and Spanish]. Biblioteca Fundacion Retevision. Fundacion Retevision, Madrid, Spain, 2001. ISBN 84-931542-7-X. 289 pp. LCCN ZA4201 .M55 2001.

Miller:2001:LII

[Mil01b] Robert R. Miller. Leapfrogging?: India’s information technology industry and the Internet, volume 42 of IFC discussion paper; no. 42. World Bank, Washington, DC, USA, 2001. ISBN 0-8213-4950-3. vii + 31 pp. LCCN HC440.H53 M55 2001.

Mills:2006:CNT

[Mil06]  David L. Mills.  Com-

puter network time synchronization: the Network Time Protocol. CRC/Taylor and Francis, Boca Raton, FL, USA, 2006. ISBN 0-8493-5805-1. xvii + 286 pp. LCCN TK5105.5 .M564

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0664/2005056889-d.html.

Mirkovic:2005:IDS

[Mir05] Jelena Mirkovic, editor. Internet denial of service: attack and defense mechanisms. The Radia Perlman series in computer networking and security. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-147573-8 (paperback). xxii + 372 pp. LCCN TK5105.59 .I5455

2004. URL http://www.loc.
gov/catdir/toc/ecip0422/
2004020335.html.

Mitra:2002:IMS

[Mit02a] Sugata Mitra. Introduction to Multimedia Systems. Academic Press, New York, NY, USA, 2002. ISBN 0-12-500452-4. xi + 207 pp. LCCN QA76.575 .I686 2002. US$59.95.

Mitzenmacher:2002:GHT

[Mit02b] Michael Mitzenmacher. Good hash tables and multiple hash functions. Dr. Dobbs Journal, 27(5):28, 30–32, May 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_05/
hash.txt.

Mandyam:2002:TGC

[ML02] Giridhar Mandyam and Jersey Lai. Third Generation CDMA Systems for Enhanced Data Services. Academic Press, New York, NY, USA, 2002. ISBN 0-12-468041-0. viii + 275 pp. LCCN TK5103.452 .M35 2002. US$79.95.

McLaren:2000:CBI

[MM00a] Constance H. McLaren and Bruce J. McLaren. Ecommerce: business on the Internet. South-Western Educational Pub., Cincinnati, OH, USA, 2000. ISBN 0-538-68918-8. xii + 291 pp. LCCN HF5548.32 .M378 2000.

Moon:2000:FBB

[MM00b] Michael Moon and Doug Millison. Firebrands: building brand loyalty in the Internet age. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212449-0. xviii + 318 pp. LCCN HF6161.B4 M66 2000.

Murthy:2001:RMR

[MM01] C. Siva Ram Murthy and G. Manimaran. Resource management in real-time systems and network. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-13376-8. xiv + 450 pp. LCCN QA76.54 .M87 2001. US$55.00, UK£37.95.

Maynor:2007:MTP

[MM07] David Maynor and K. K. Mookhey. Metasploit toolkit for penetration testing, exploit development, and vulnerability research. Syngress Publishing, Inc., Rockland, MA, USA, 2007. ISBN 1-59749-074-1 (paperback). xvi + 272 pp. LCCN QA76.9.A25 M396

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0805/2007299879-d.html.

Modahl:2000:NNH

[Mod00] Mary Modahl. Now or never: how companies must change today to win the battle for Internet consumers. HarperBusiness, New York, NY, USA, 2000. ISBN 0-06-662012-0. xxviii + 237 pp. LCCN HF5415.1265 .M63 2000.

Mohammed:2002:IMB

[Moh02] Rafi Mohammed. Internet marketing: building advantage in the networked economy. McGraw-Hill/Irwin marketspaceU, Boston, MA, USA, 2002. ISBN 0-07-251022-6. xli + 726 pp. LCCN HF5415.1265 .I565 2002.

Molenaar:2002:FMP

[Mol02] Cor Molenaar. The future of marketing: practical strategies for marketers in the post-Internet age. Financial Times Prentice Hall, London, UK, 2002. ISBN 0-273-65497-7. xxviii + 235 pp. LCCN ????

Moore:2000:LFL

[Moo00] Geoffrey A. Moore. Living on the fault line: managing for shareholder value in the age of the Internet. HarperBusiness, New York, NY, USA, 2000. ISBN 0-88730-888-0. xx + 288 pp. LCCN HD31 .M625 2000.

Wendell:2002:IWB

[Moo02] James W. (James Wendell)

Moore. The Internet weather: balancing continuous change and constant truths. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-06408-4. ix + 246 pp. LCCN HM851 .M66 2002.

Morath:2000:SBP

[Mor00] Peter Morath. Success @ ebusiness: profitable Internet business and commerce. Mc-Graw-Hill, New York, NY, USA, 2000. ISBN 0-07-709625-8. xix + 342 pp. LCCN HF5548.32 .M67 2000.

Morrison:2008:ACK

[Mor08] Michael Morrison. Ajax construction kit: building plugand-play Ajax applications. Negus live Linux series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-235008-4 (paperback). xvii + 291 pp. LCCN TK5105.8885.A52 M67 2008.

Mosher:2005:CMO

[Mos05] Sue Mosher. Configuring Microsoft Outlook 2003. Elsevier/ Digital Press, Amsterdam, The Netherlands; Boston, MA, USA, 2005. ISBN 1-55558-326-1 (paperback). xv + 495 pp. LCCN HF5548.4.M5255 M668 2005.

Mouchtaris:2001:VIV

[Mou01] Petros Mouchtaris, editor. Voice over IP (VoIP) technology: 21 August 2001, Denver, USA, volume 4522 of SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 2001. ISBN 0-8194-4246-1. LCCN TK5105.8865 .V65 2001.

Martinez:2001:SWG

[MP01] Rey C. Martinez and Carol Clark Powell. The social worker’s guide to the Internet. Allyn and Bacon, Inc., Needham Heights, MA, USA, 2001. ISBN 0-205-29735-8. xi + 180 pp. LCCN HV29.2 .M37 2001.

McNamara:2000:IPL

[MR00] Elizabeth A. McNamara and Steven Eric Rayman. Internet publishing: the legal and business issues as traditional publishing moves to electronic media, volume G-601 of Intellectual property course handbook series; no. G-601. Practising Law Institute, New York, NY, USA, 2000. ISBN 0-87224-768-6. 456 pp. LCCN KF2750.Z9 I58 2000.

Medhi:2007:NRA

[MR07] Deepankar Medhi and Karthikeyan Ramasamy. Network routing: algorithms, protocols, and architectures. Elsevier/ Morgan Kaufmann, San Francisco, CA, USA, 2007. ISBN 0-12-088588-3 (hardcover). ???? pp. LCCN TK5105.5 .M425

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0668/2006028700-d.html;
http://www.loc.gov/catdir/
toc/ecip0620/2006028700.
html.

Milenkovic:2003:CFT

[MRK^{+}03] Milan Milenkovic, Scott H.

Robinson, Rob C. Knauerhase, David Barkai, Sharad Garg, Vijay Tewari, Todd A. Anderson, and Mic Bowman. Cover feature: Toward Internet distributed computing.

Computer, 36(5):38–46, May 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/05/r5038.
htm;   http://csdl.computer.
org/dl/mags/co/2003/05/r5038.
pdf;   http://www.computer.
org/computer/co2003/r5038abs.
htm.

ContrerasMontellano:2000:ICE

[MRM00] Oscar F. Contreras Montellano, Hiram Felix Rosas, and Ulises Ponce Mendoza. Internet y comercio electronico: una oportunidad para el desarrollo regional? (Spanish) [Internet and electronic commerce: an opportunity for regional development?]. Technical report, El Colegio de Sonora, Hermosillo, Sonora, Mexico, October 2000. 24 pp.

Mann:2000:ICQ

[MS00a] Chris Mann and Fiona Stewart. Internet communication and qualitative research: a handbook for researching online. New technologies for social research. Sage Publications, Thousand Oaks, CA, USA, 2000. ISBN 0-7619-6626-9, 0-7619-6627-7 (paperback). x + 258 pp. LCCN H61.95.M36 2000.

Masand:2000:WUA

[MS00b] Brij Masand and Myra Spiliopoulou, editors. Web usage analysis and user profiling: international WEBKDD’99 workshop, San Diego, CA, USA, August 15, 1999; revised papers, volume 1836 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67818-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN TK5105.875.I57 I479 1999.

Miller:2000:IEA

[MS00c] Daniel Miller and Don Slater.

The Internet: an ethnographic approach. Berg, New York, NY, USA, 2000. ISBN 1-85973-384-0 (hardcover), 1-85973-389-1 (paper). ix + 217 pp. LCCN HM851 .M65 2000. URL http://

ethnonet.gold.ac.uk.

Maxemchuk:2001:IMS

[MS01a] N. F. Maxemchuk and D. H.

Shur. An Internet multicast system for the stock market. ACM Transactions on Computer Systems, 19(3):384– 412, 2001. CODEN AC-SYEC. ISSN 0734-2071. URL

http://www.acm.org/pubs/
citations/journals/tocs/
2001-19-3/p384-maxemchuk/

.

Mitra:2001:NIP

[MS01b] Sanjit K. Mitra and Giovanni Sicuranza, editors. Nonlinear Image Processing. Academic Press, New York, NY, USA, 2001. ISBN 0-12-500451-6. xv + 455 pp. LCCN TA1632 .N633 2001. US$95.00.

McClellan:2000:LNG

[MSC00] Stan McClellan, Michael Stricklen, and Bob Cummings. Linux and the next generation Internet. Linux Journal, 72:??, April 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Maheu:2001:IIV

[MSCG01] Marlene M. Maheu, Rona Subotnik, Al Cooper, and Shirley Glass. Infidelity on the Internet: virtual relationships and real betrayal. Sourcebooks, Naperville, IL, USA, 2001. ISBN 1-57071-722-2. xviii + 254 pp. LCCN HQ806 .M34 2001. Contents: Prelude to infidelity. Cyber-affairs: lighthearted flirting, hot chatting, or infidelity? — Love at the speed of electricity: the power of virtual attraction — Cybersex — Cyber-sexual compulsivity: impulses gone haywire — The shock of infidelity on the Internet. Infidelity: virtual or real? — Discovery: real betrayal from the virtual world — Decisions: to stay or to go? — Path to recovery. Beginning steps to recovery — Concluding steps to recovery — Appendix A: Understanding the Internet — Appendix B: Emoticons and acronyms.

McClure:2003:HEN

[MSK03] Stuart McClure, Joel Scambray, and George Kurtz. Hacking exposed: network security secrets and solutions. Mc-Graw-Hill/Osborne, Berkeley, CA, USA, fourth edition, 2003. ISBN 0-07-222742-7. xxiv + 737 pp. LCCN TK5105359 .M48 2003. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
bios/mh051/2004351501.html;
 http://www.loc.gov/catdir/
description/mh051/2004351501.
html;   http://www.loc.gov/
catdir/toc/mh051/2004351501.
html.

Monnoyer-Smith:2002:EVI

[MSM02]  Laurence  Monnoyer-Smith

and Eric Maigret. Electronic vote and Internet campaigning: State of the art in Europe and remaining questions. Lecture Notes in Computer Science, 2456:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2456/24560280.htm;
http://link.springer.de/
link/service/series/0558/
papers/2456/24560280.pdf.

Maggs:2001:ICL

[MSS01] Peter B. Maggs, John T.

Soma, and James A. Sprowl. Internet and computer law: cases — comments — questions. American casebook series. West Publishing Company, St. Paul, MN, USA, 2001. ISBN 0-314-25295-9. xxv + 829 pp. LCCN KF390.5.C6 M345 2001.

Markopoulou:2003:AQV

[MTK03] Athina P. Markopoulou, Fouad A. Tobagi, and Mansour J. Karam. Assessing the quality of voice communications over Internet backbones. TON, 11 (5):747–760, October 2003.

Muchow:2002:CJT

[Muc02]  John W. Muchow.  Core

J2ME Technology and MIDP. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-066911-3. 737 pp. LCCN ???? US$49.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130669113.html.

Muffke:2001:CPE

[Müf01] Friedger Müffke. The Curl programming environment: Making the Internet interactive. Dr. Dobbs Journal, 26(9):66, 68, 70–71, September 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2001/2001_09/
                                        curllist.zip.

Murdock:2000:JYV

[Mur00a] Kelly Murdock. JavaScript: your visual blueprint for building dynamic Web pages. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-4730-5. xvii + 285 pp. LCCN QA76.73.J39 M87 2000.

Murphy:2000:WRH

[Mur00b] Tom Murphy. Web rules: how the Internet is changing the way consumers make choices. Dearborn, Dearborn, IL, USA, 2000. ISBN 0-7931-3613-X (hb). x + 303 pp. LCCN HF5415.1265 .M87 2000.

Morrow:2003:DIB

[MV03] Monique Morrow and Kateel Vijayananda. Developing IPbased services: solutions for service providers and vendors. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-779-X. xxi + 313 pp. LCCN TK5105.5 .M72 2003. US$49.95.

McKnight:2001:CDB

[MVK01] Lee W. McKnight, Paul M.

Vaaler, and Raul Luciano Katz, editors. Creative destruction: business survival strategies in the global Internet economy. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-13389-X (hardcover). LCCN HC79.T4 C74 2001. Papers from a symposium held at the Fletcher School of Law and Diplomacy, Tufts University.

Mates:2000:ATI

[MWD00] Barbara T. Mates, Doug Wakefield, and Judith M. Dixon. Adaptive technology for the Internet: making electronic resources accessible to all. American Library Association, Chicago, IL, USA, 2000.

ISBN 0-8389-0752-0. ix + 192 pp. LCCN Z675.B6 M38 2000.

Morton:2000:ICR

[MZSR00] Fiona Scott Morton, Florian Zettelmeyer, and Jorge M. Silva-Risso. Internet car retailing. NBER working paper series 7961, National Bureau of Economic Research, Cambridge, MA, USA, October 2000. 33 pp.

Nadeau:2003:MNM

[Nad03] Thomas D. Nadeau. MPLS Network Management: MIBs, Tools, and Techniques. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-751-X. xxxvi + 529 pp. LCCN TK5105.573 .N34 2003. US$59.95. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/els031/2002108920.
                                        html.

Nairn:2002:EMM

[Nai02] Alasdair Nairn. Engines that move markets: technology investing from railroads to the Internet and beyond. Wiley-Academy, New York, NY, USA, 2002. ISBN 0-471-20595-

8. xvii + 494 pp. LCCN

HC79.T4 N35 2002.

Naldi:2002:IAT

[Nal02] M. Naldi. Internet access traffic sharing in a multi-operator environment. Computer Networks (Amsterdam, Netherlands: 1999), 38(6):809–824, April 22, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/76/37/34/
abstract.html.

NRC:2000:NHP

[Nat00] National Research Council (U.S.). Committee on Enhancing the Internet for Health and Biomedical Applications: Technical Requirements and Implementation Strategies. Networking health: prescriptions for the Internet. National Academy Press, Washington, DC, USA, 2000. ISBN 0-309-06843-6 (casebound). xviii + 358 pp. LCCN R859.7.I58

N48 2000. URL http://www.
nap.edu/books/0309068436/
html.

NRC:2001:ICA

[Nat01] National Research Council (U.S.). Committee on the Internet in the Evolving Information Infrastructure. The Internet’s coming of age. National Academy Press, Washington, DC, USA, 2001. ISBN 0-309-06992-0. xvii + 236 pp. LCCN TK5105.875.I57 I5435 2001. URL http://www.

nap.edu/books/0309069920/
html.

Naughton:2000:BHF

[Nau00] John (John J.) Naughton. A brief history of the future: from radio days to Internet years in a lifetime. Overlook Press, Woodstock, NY, USA, 2000. ISBN 1-58567-032-4. xv + 327 pp. LCCN TK5102.2 .N38 2000.

Nixon:2000:ILP

[NB00] Carol Nixon and Jennifer Burmood, editors. Internet librarian 2000: proceedings-2000, Monterey, California, November 6–8, 2000. Information Today, Inc., Medford, NJ, USA, 2000. ISBN 1-57387-107-9. LCCN Z674.82.I59 I585 2000. Proceedings of the Fourth Internet Librarian Conference, Monterey, California, November 6-8, 2000.

Nassif:2002:ISE

[NCC^{+}02] Lilian N. Nassif, Luiz Henrique A. Correia, Carlos Frederico M. C. Cavalcanti, José Marcos Nogueira, Antonio A. F. Loureiro, and Geraldo Robson Mateus. InterQoS — strategy enterprise game for price and QoS negotiation on the Internet. Lecture Notes in Computer Science, 2511:293– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2511/25110293.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2511/25110293.pdf.

Nath:2003:IAI

[NDK^{+}03] Suman Nath, Amol Deshpande, Yan Ke, Phillip B. Gibbons, Brad Karp, and Srinivasan Seshan. IrisNet: An architecture for Internet-scale sensing services. In Freytag et al. [FLA^{+}03], pages 1137–1140. ISBN 0-12-722442-4. LCCN ???? URL

http://www.vldb.org/dblp/
db/indices/a-tree/n/Nath:
Suman.html.

Neuschotz:2000:IDL

[Neu00] Nilson Neuschotz. Introduction to Director and Lingo: multimedia and Internet application. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-690322-3. xvii + 617 pp. LCCN QA76.575 .N46 2000.

Newell:2000:LCC

[New00a] Frederick Newell. Loyalty.com:

customer relationship management in the new era of Internet marketing. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-135775-0. xx + 325 pp. LCCN HF5415.1265 .N495 2000.

Newton:2000:NTD

[New00b] Harry Newton. Newton’s telecom dictionary: the official dictionary of telecommunications and the Internet. Telecom, New York, NY, USA, sixteenth edition, 2000. ISBN 1-57820-053-9. xii + 974 + [17] pp. LCCN TK5102 .N485

Newburger:2001:HCI

[New01a] Eric C. Newburger. Home computers and Internet use in the United States: August 2000. Technical Report P23-207, U.S. G.P.O, Washington, DC, USA, September 2001. 9 pp.

Newton:2001:NTD

[New01b] Harry Newton. Newton’s telecom dictionary: the official dictionary of telecommunications networking and Internet. CMP Books, New York, NY, USA, seventeenth edition, 2001. ISBN 1-57820-069-5. viii + 787 pp. LCCN TK5102 .N486 2001.

Nguyen:2001:TAW

[Ngu01] Hung Quoc Nguyen. Testing applications on the Web: test planning for Internetbased systems. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-39470-X (paperback). xiv + 402 pp. LCCN QA76.76.A65 N48 2000. “Wiley Computer Publishing.”.

Nicoll:2001:CNN

[Nic01] Leslie H. Nicoll. Computers in nursing’s nurses’ guide to the Internet. Lippincott, Philadelphia, PA, USA, third edition, 2001. ISBN 0-7817-2459-7. ix + 242 pp. LCCN WY26.5 N53 2001. Disk includes: links to each of the Web sites listed in the book.

NiederstRobbins:2006:HXP

[Nie06] Jennifer Niederst Robbins.

HTML and XHTML Pocket Reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2006. ISBN 0-596-52727-6. 112 (est.) pp. LCCN ???? EUR 9.90, US$13.09.

Nisley:2001:ESD

[Nis01] Ed Nisley. Embedded space:

Dynamic attention. Dr. Dobbs Journal, 26(12):107– 109, December 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Nagaraja:2003:QIA

[NKB^{+}03] Kiran Nagaraja, Neeraj Krishnan, Ricardo Bianchini, Richard P. Martin, and Thu D. Nguyen. Quantifying and improving the availability of high-performance cluster-based Internet services. In ACM [ACM03], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.
sc-conference.org/sc2003/
inter_cal/inter_cal_detail.
php?eventid=10686#2;   http:
//www.sc-conference.org/
sc2003/paperpdfs/pap213.
pdf.

Nakamura:2004:DFI

[NLiMK04] Masahide Nakamura, Pattara Leelaprute, Ken ichi Matsumoto, and Tohru Kikuno. On detecting feature interactions in the programmable service environment of Internet telephony. Computer Networks (Amsterdam, Netherlands: 1999), 45(5):605–624, August 5, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Northcutt:2002:NID

[NN02] Stephen Northcutt and Judy Novak. Network intrusion detection. New Riders Publishing, Carmel, IN, USA, third edition, 2002. ISBN 0-7357-1265-4. xvii + 490 pp. LCCN TK5105.59 .N475 2003.

Northcutt:2001:NID

[NNM01] Stephen Northcutt, Judy Novak, and Donald McLachlan. Network intrusion detection: an analyst’s handbook. New Riders Publishing, Carmel, IN, USA, second edition, 2001. ISBN 0-7357-1008-2 (paperback). xxxii + 430 pp. LCCN TK5105.59 .N475 2001.

Norberg:2000:SWN

[Nor00a]  Stefan Norberg.  Securing

Windows NT/2000 servers for the Internet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-768-0. xiv + 199 pp. LCCN QA76.76.O63 N65 2001.

Nordquist:2000:ISP

[Nor00b] Joan Nordquist. The Internet: social, political, economic concerns: a bibliography. Contemporary social issues no. 60.

Reference and Research Services, Santa Cruz, CA, USA, 2000. ISBN 1-892068-18-4. 68 pp. LCCN ZA4201 .N67 2000.

Nordquist:2001:IIP

[Nor01a] Joan Nordquist. The Internet (II): psychological, security, race, class and gender issues: a bibliography, volume 62 of Contemporary social issues ISSN: 0887-3569; no. 62. Reference and Research Services, Santa Cruz, CA, USA, 2001. ISBN 1-892068-22-2. ISSN 0887-3569. 68 pp. LCCN H1 .C657 no.62.

Norris:2001:DDC

[Nor01b] Pippa Norris. Digital divide: civic engagement, information poverty, and the Internet worldwide. Communication, society, and politics. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-80751-4 (hardcover), 0-521-00223-0 (paperback). xv + 303 pp. LCCN HN49.I56 N67 2001.

Northcutt:2001:ISA

[Nor01c] Stephen Northcutt, editor. Intrusion signatures and analysis. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1063-5. xxi + 408 pp. LCCN QA76.9.A25 I59 2001; QA76.9.A25 I63 2001.

Notess:2000:GII

[Not00] Greg R. Notess. Government information on the Internet. Bernan Press, Lanham, MD, USA, third edition, 2000. ISBN 0-89059-247-0. xviii + 833 pp. LCCN K87 .N68 2000 Ref.

Negrino:2001:JWW

[NS01] Tom Negrino and Dori Smith. JavaScript for the World Wide Web. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, fourth edition, 2001. ISBN 0-201-73517-2. xiv + 430 pp. LCCN QA76.73.J39 N44 2001.

Nuaymai:2007:WTB

[Nua07] Loutfi Nuaymai. WiMAX: technology for broadband wireless access. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2007. ISBN 0-470-02808-4 (hardcover). xxv + 283 pp. LCCN TK5105.78 .N83 2007.

Orebaugh:2005:SC

[OBB05] Angela Orebaugh, Simon Biles, and Jacob Babbin. Snort cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00791-4 (paperback). xiii + 270 pp. LCCN TK5105.59 .O74 2005.

Oestges:2007:MWC

[OC07] Claude Oestges and Bruno Clerckx. MIMO wireless communication: from real-world propagation to space-time code design. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-12-372535-6. xxvi + 448 pp. LCCN ????

ODriscoll:2000:EGH

[O’D00] Gerard O’Driscoll. The Essential Guide to Home Networking Technologies. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-019846-3. 378 (est.) pp. LCCN ???? US$34.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130198463.html.

Odlyzko:2000:HCI

[Odl00] Andrew Odlyzko. The history of communications and its implications for the Internet. Report, AT&T Labs — Research, Murray Hill, NJ, USA, June 16, 2000. 160 pp.

URL https://www-users.
cse.umn.edu/~odlyzko/doc/
history.communications0.
pdf.

Osterle:2000:BNS

[OFA00] Hubert Osterle, Elgar Fleisch, and Rainer Alt. Business networking: shaping enterprise relationship on the Internet. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66612-5. xv + 376 pp. LCCN HD30.37 .O85 2000.

Olpin:2000:ABQ

[OG00] Michael Olpin and Doug Gotthoffer. Allyn and Bacon quick guide to the Internet for health. Allyn and Bacon, Inc., Needham Heights, MA, USA, 2000.

ISBN 0-205-30963-1. iv + 155 pp. LCCN R859.7.D36 O48 2000.

Ouyang:2002:FES

[OKB02] Chun Ouyang, Lars Michael Kristensen, and Jonathan Billington. A formal and executable specification of the Internet Open Trading Protocol. Lecture Notes in Computer Science, 2455:377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2455/24550377.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2455/24550377.
pdf.

OLooney:2000:AMY

[O’L00] John O’Looney. Access: making your community Internetready. IQ service report 32(5), International City/ County Management Association, Washington, DC, USA, May 2000. 19 pp.

Olexa:2003:IWN

[Ole03] Ron Olexa. Implementing 802.11, 802.16, and 802.20 Wireless Networks: Planning, Troubleshooting, and Maintenance. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-7506-7808-9. ???? pp. LCCN ???? US$59.99.

Ong:2002:UVC

[OLLY02] T. M. Ong, T. M. Lim, B. S.

Lee, and C. K. Yeo. Unicorn: voluntary computing over Internet. Operating Systems Review, 36(2):36–51, April 2002. CODEN OSRED8. ISSN 0163-5980.

Omicini:2001:CIA

[Omi01] Andrea Omicini. Coordination of Internet agents: models, technologies, and applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-41613-7. xxv + 523 pp. LCCN QA76.76.I58 C68 2001.

Ohsaki:2002:MRT

[OMM02] Hiroyuki Ohsaki, Mitsushige Morita, and Masayuki Murata. On modeling roundtrip time dynamics of the Internet using system identification. Lecture Notes in Computer Science, 2343:359– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2343/23430359.htm;
http://link.springer.de/
link/service/series/0558/
papers/2343/23430359.pdf.

Ojanpera:2001:WTI

[OP01] Tero Ojanpera and Ramjee Prasad. WCDMA: towards IP mobility and mobile Internet. Artech House universal personal communications series. Artech House Inc., Norwood, MA, USA, revised edition, 2001. ISBN 1-58053-180-6. xxii + 477 pp. LCCN TK5103.452 .W395 2001. Rev. ed.

OHara:2004:IHD

[OP04] Bob O’Hara and Al Petrick. The IEEE 802.11 handbook: a designer’s companion. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, second edition, 2004. ISBN 0-7381-4449-5. ???? pp. LCCN TK5105.78 .O43 2004. URL http:/

/shop.ieee.org/ieeestore/
product.aspx?product_no=
SP1136.

Oppliger:2000:STW

[Opp00] Rolf Oppliger. Security technologies for the World Wide Web. Artech House computer security series. Artech House Inc., Norwood, MA, USA, 2000. ISBN 1-58053-045-1. xxi + 419 pp. LCCN QA76.9.A25

O68 2000. URL http://www.
artechhouse.com/Detail.aspx?
strIsbn=978-1-58053-045-
3.

Oppliger:2001:IIS

[Opp01a] Rolf Oppliger. Internet and Intranet security. Artech House computer security series. Artech House Inc., Norwood, MA, USA, second edition, 2001. ISBN 1-58053-166-0. xxiii + 403 pp.

LCCN TK5105.875.I57 O67

2001.  URL http://www.
artechhouse.com/Detail.aspx?
strIsbn=978-1-58053-166-
5.

Oppliger:2001:SMP

[Opp01b] Rolf Oppliger. Secure messaging with PGP and S/MIME. Artech House computer security series. Artech House Inc., Norwood, MA, USA, 2001. ISBN 1-58053-161-X. xxiii + 305 pp. LCCN TK5102.85 .O67 2001. URL http://www.

artechhouse.com/Detail.aspx?
strIsbn=978-1-58053-161-
0.

Oppliger:2002:IIS

[Opp02] Rolf Oppliger. Internet and Intranet security. Artech House computer security series. Artech House Inc., Norwood, MA, USA, second edition, 2002. ISBN 1-58053-166-0. xxiii + 403 pp. LCCN TK5105.875.I57 O67 2002.

Oppliger:2003:STW

[Opp03] Rolf Oppliger. Security technologies for the World Wide Web. Artech House computer security series. Artech House Inc., Norwood, MA, USA, second edition, 2003. ISBN 1-58053-348-5. xxiii + 416 pp. LCCN QA76.9.A25

O67 2003. URL http://www.
artechhouse.com/Detail.aspx?
strIsbn=978-1-58053-348-
5;   http://www.loc.gov/
catdir/toc/fy035/2002032665.
html.

Oppliger:2007:PCM

[Opp07] Rolf Oppliger. Providing certified mail services on the Internet. IEEE Security & Privacy, 5(1):16–22, January/February 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Oppliger:2009:STT

[Opp09]  Rolf Oppliger.  SSL and

TLS: theory and practice. Artech House information security and privacy series. Artech House Inc., Norwood, MA, USA, 2009. ISBN 1-59693-447-6. xxi + 257 pp. LCCN TK5105.59 .O67

                                        2009.  URL http://www.
                                        artechhouse.com/Detail.aspx?
                                        strIsbn=978-1-59693-447-
                                        4.

OReilly:2000:VIP

[O’R00]  Tim O’Reilly.  Viewpoint:

The Internet patent land grab. Communications of the ACM, 43(6):29–31, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        2000-43-6/p29-o_reilly/.

Oravec:2002:CAI

[Ora02] Jo Ann Oravec. Constructive approaches to Internet recreation in the workplace. Communications of the ACM, 45 (1):60–63, January 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ohura:2002:EQE

[OTPK02] Yusuke Ohura, Katsumi Takahasii, Iko Pramudiono, and Masaru Kitsuregawa. Experiments on query expansion for Internet Yellow Page services using Web log mining. In Bernstein et al. [B^{+}02], pages 1008–1018. ISBN 1-55860-869-9. LCCN ????

URL http://www.vldb.org/
conf/2002/S32P01.pdf.

Owad:2005:ARC

[Owa05] Tom Owad. Apple I replica creation: back to the garage. Syngress Publishing, Inc., Rockland, MA, USA, 2005. xxv + 333 pp. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://site.ebrary.com/
lib/ucsc/Doc?id=10074968.

Owen:2002:HIC

[Owe02a] David J. Owen. The herbal Internet companion: herbs and herbal medicine online. Haworth Information Press, New York, NY, USA, 2002. ISBN 0-7890-1051-8, 0-7890-1052-6. xv + 193 pp. LCCN RM666.H33 O975 2002.

Owezarski:2002:DII

[Owe02b] Philippe Owezarski. Does IPv6 improve the scalability of the Internet? Lecture Notes in Computer Science, 2515:130– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2515/25150130.htm;
http://link.springer.de/
link/service/series/0558/
papers/2515/25150130.pdf.

Palazzo:2001:ETI

[Pal01] Sergio Palazzo, editor. Evolutionary trends of the Internet: 2001 Thyrrhenian International Workshop on Digital Communications, IWDC 2001, Taormina, Italy, September 2001: proceedings, volume 2170 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42592-6 (paperback). LCCN QA267.A1 L43 no.2170; TK5103.7 .T48

2001. URL http://link.
springer-ny.com/link/service/
series/0558/tocs/t2170.htm.

Also available via the World Wide Web.

Palacios:2002:SRA

[Pal02] Rafael Palacios. On site: Remote automatic doorman via the Internet. Communications of the ACM, 45(10):23– 25, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Panzarino:2002:TTS

[Pan02] Frank Panzarino. Telecommunications Technologies for Small Business. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-093742-8. ???? pp. LCCN ???? US$34.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130937428.html.

Panwar:2004:TIE

[Pan04] Shivendra S. Panwar, editor.

TCP/IP essentials: a labbased approach. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-84144-5 (hardcover), 0-521-60124-X (paperback). xxi + 262 pp. LCCN TK5105.585 .T355 2004. US$90.00 (hardcover), US$39.99 (paperback). URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam041/2004045664.
html;   http://www.loc.gov/
catdir/toc/cam041/2004045664.
html.

Papadimitriou:2002:LI

[Pap02] Christos Papadimitriou. Learning the Internet. Lecture Notes in Computer Science, 2375:396–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer-ny.com/link/service/
series/0558/bibs/2375/23750396.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2375/23750396.
pdf.

Parker:2005:SYE

[Par05] Jim Parker. Start your engines: developing driving and racing games. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-933097-01-9. xxii + 434 pp. LCCN GV1469.17.A98 P37 2005.

Pashtan:2005:MWS

[Pas05] Ariel Pashtan. Mobile Web services. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83049-4 (hardcover). xviii + 266 pp. LCCN TK5103.2 .P38 2005. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/description/
cam051/2004054623.html;
http://www.loc.gov/catdir/
toc/cam051/2004054623.html.

Paulson:2003:NBS

[Pau03] Linda Dailey Paulson. News briefs: Spammers begin sending viruses with their junk mail; “ballistic” technique boosts storage capacity; FAST approach could accelerate Internet downloads; software cuts the “bull” out of documents. Computer, 36(8):26–28, August 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/08/r8026.
htm;   http://csdl.computer.
org/dl/mags/co/2003/08/r8026.
pdf.

Poirier:2000:SCU

[PB00] Charles C. Poirier and Michael J.

Bauer. E-supply chain: using the Internet to revolutionize your business: how market leaders focus their entire organization on driving value to customers. Berrett-Koehler, San Francisco, CA, USA, 2000. ISBN 1-57675-117-1. xv + 236 pp. LCCN HF5548.32 .P65 2000.

Podlipnig:2003:SWC

[PB03] Stefan Podlipnig and Laszlo Böszörmenyi. A survey of Web cache replacement strategies. ACM Computing Surveys, 35(4):374–398, December 2003. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Phillips:2001:TIL

[PC01] Iain Phillips and Jon Crowcroft.

TCP/IP and the Linux protocol implementation: systems code for the Linux Internet. Wiley Networking Council series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-40882-4 (hardcover). 925 pp. LCCN TK5105.585 .T34 2002.

Pradhan:2002:IEQ

[PcC02] Prashant Pradhan and Tzi cker Chiueh. Implementation and evaluation of a QoS-capable cluster-based IP router. In IEEE [IEE02c], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http://

www.sc-2002.org/paperpdfs/
pap.pap320.pdf.

Peterson:2003:CNS

[PD03] Larry Peterson and Bruce S. Davie. Computer Networks: a Systems Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, third edition, 2003. ISBN 1-55860-832-X. viii + 176 pp. LCCN TK5105.5 .P479 2003. US$89.95.

Peterson:2007:CNS

[PD07] Larry L. Peterson and Bruce S. Davie. Computer networks: a systems approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, fourth edition, 2007. ISBN 0-12-370548-7 (hardcover), 0-12-374013-4 (paperback). xxvii + 806 pp. LCCN TK5105.5 .P479

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0704/2006102454-d.html.

Perlman:2000:ISE

[Per00] Radia Perlman. Interconnections, Second Edition: Bridges, Routers, Switches, and Internetworking Protocols. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-63448-1. xiii + 537 pp. LCCN TK5105.5 .P474 2000. US$59.95.

Peterson:2001:GIC

[Pet01] Susan Lee Peterson. The grantwriter’s Internet companion: a resource for educators and others seeking grants and funding. Corwin Press, Thousand Oaks, CA, USA, 2001. ISBN 0-7619-7746-5 (paper), 0-7619-7808-9 (hardcover). vi + 122 pp. LCCN LC243.A1 P48 2001.

PRCPP:2000:ISB

[Pew00] Pew Research Center for the People and the Press. Internet sapping broadcast news audience: investors now go online for quotes, advice. News release. Pew Research Center for the People and the Press, Washington, DC, USA, June 11, 2000. 107 pp. LCCN A8292 June 11, 2000.

PCS:2001:SRI

[Pew01] Pew Center on the States.

Statehouse reporters: Internet resource guide. The Center, Washington, DC, USA, 2001. 32 pp. LCCN JK2408 .P446 2001.

Pfaffenberger:2001:LNC

[Pfa01] Bryan Pfaffenberger. Linux networking clearly explained. Academic Press, New York, NY, USA, 2001. ISBN 0-12-533171-1. ix + 390 pp. LCCN QA76.76.O63 P533 2001.

Pourmohammadi-Fallah:2003:IDM

[PFAHA03] Yasser Pourmohammadi-Fallah, Kambiz Asrar-Haghighi, and Hussein Alnuweiri. Internet delivery of MPEG-4 object-based multimedia. IEEE MultiMedia, 10(3):68– 78, July–September 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/comp/
mags/mu/2003/03/u3068abs.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/u3068.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/u3068.
pdf.

Prins:2000:PAV

[PG00] C. (Corien) Prins and S. J. H. Gijrath. Privaatrechtelijke aspecten van elektronische handel: juridische aandachtspunten voor Internet service providers, volume 61 of Studiepockets privaatrecht. W. E. J. Tjeenk Willink, Deventer, The Netherlands, 2000. ISBN 90-271-5220-9 (paperback). xvii + 310 pp. LCCN KE2717 .P75 2000.

Pang:2001:MEC

[PG01] Valerie Ooka Pang and Paul Gorski. Multicultural education: a caring-centered, reflective approach. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-242329-3 (set), 0-07-236953-1 (Pang), 0-07-242102-9 (Gorski). ???? pp. LCCN LC1099.3 .P35 2001.

Park:2005:LRD

[PHCMS05] Cheolwoo Park, Félix Hernández-Campos, J. S. Marron, and F. Donelson Smith. Longrange dependence in a changing Internet traffic mix. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):

401–422, June 21, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Phillips:2001:VVW

[Phi01] Lee Anne Phillips. VoiceXML and the voice/Web environment. Dr. Dobbs Journal, 26(10):91–92, 94, 96, October 2001. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/ftp/2001/2001_10/
voicexml.zip;   http://www.
ddj.com/ftp/2001/2001_10/
voicexml_txt.zip.

Polk:2003:IIC

[PHM03] William T. Polk, Nelson E.

Hastings, and Ambarish Malpani. IEEE Internet computing: Security track: Public key infrastructures that satisfy security goals. IEEE Distributed Systems Online, 4(7), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0307/f/wp4sec.htm.

Phoha:2001:SDI

[Pho01] Vir V. Phoha. The Springer Dictionary of Internet Security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95261-6. 329 (est.) pp. LCCN ???? US$39.95.

Poulter:2000:LIP

[PHT00] Alan Poulter, Debra Hiom, and Gwyneth Tseng. The library and information professional’s guide to the Internet. Library Association, London, UK, third edition, 2000. ISBN 1-85604-376-2. xiii + 146 pp. LCCN TK5105.875.I57 T73 2000. Previous ed.: 1997.

Pickens:2003:SOM

[Pic03] Dave Pickens. Sun ONE Messaging Server: Practices and Techniques for Enterprise Customers. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-145496-X. 288 (est.) pp. LCCN ???? US$48.00.

Pansier:2000:IIJ

[PJ00a] Frederic-Jerome Pansier and Emmanuel Jez. Initiation à l’Internet juridique: Connexion, utilisation, protection. (French) [Initiation to the legal Internet: Connection, use, protection]. Droit@Litec. Litec, Paris, France, deuxième edition, 2000. ISBN 2-7111-3149-1. x + 127 pp. LCCN KJV143.I58 P36 2000.

Pansier:2000:CI

[PJ00b] Frederic-Jerome Pansier and Emmanuel Jez. La criminalité sur l’Internet. (French) [Criminality on the Internet], volume 3546 of Que saisje? Presses universitaires de France, Paris, France, 2000. ISBN 2-13-050510-4. 127 pp. LCCN KJV143.I58 P36 2000.

Poyser:2005:FGP

[PJ05] Debbie Poyser and Derek Johnson. Fast guide to Propellerhead Reason. PC, Merton, UK, second edition, 2005. ISBN 1-870775-93-7 (paperback). 441 pp. LCCN ML74.4.R43 P69 2005.

Park:2002:RDE

[PJN^{+}02] Kyung-Joon Park, Won Jong Jeon, Klara Nahrstedt, Tamer Basar, and Chong-Ho Choi. Robust delay estimation for Internet multimedia applications. Lecture Notes in Computer Science, 2515:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2515/25150255.htm;
http://link.springer.de/
link/service/series/0558/
papers/2515/25150255.pdf.

Pottie:2000:EIW

[PK00] G. J. Pottie and W. J.

Kaiser. Embedding the Internet: wireless integrated network sensors. Communications of the ACM, 43(5):51, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-5/p51-pottie/.

Pahlavan:2002:PWN

[PK02] Kaveh Pahlavan and Prashant Krishnamurthy. Principles of Wireless Networks: a Unified Approach. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-093003-2. 600 (est.) pp. LCCN ???? US$84.00. URL http:

//www.phptr.com/ptrbooks/
ptr_0130930032.html.

Papadopoulos:2004:RSI

[PKN^{+}04] Christos Papadopoulos, Chris Kyriakakis, Ulrich Neumann, Alexander A. Sawchuk, Cyrus Shahabi, and Roger Zimmermann. RMI system: Internet meets the future home theater. IEEE MultiMedia, 11 (2):48–57, April/June 2004. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/comp/
mags/mu/2004/02/u2048abs.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/02/u2048.
htm;   http://csdl.computer.
org/dl/mags/mu/2004/02/u2048.
pdf.

Pashtan:2003:IIC

[PKP03] Ariel Pashtan, Shriram Kollipara, and Michael Pearce. IEEE Internet computing: Industry report: Adapting content for wireless Web services. IEEE Distributed Systems Online, 4(9), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0309/d/w5indp.htm.

Pahlavan:2005:WIN

[PL05] Kaveh Pahlavan and Allen H.

Levesque. Wireless information networks. Wiley series in telecommunications and signal processing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 2005. ISBN 0-471-72542-0 (hardcover). xv + 722 pp. LCCN TK5103.2 .P34

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0623/2005041792-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0623/2005041792-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0623/2005041792-t.html.

Pfleeger:2007:IBC

[PLW07] Shari Lawrence Pfleeger, Martin Libicki, and Michael Webber. I’ll buy that! cybersecurity in the Internet marketplace. IEEE Security & Privacy, 5(3):25–31, May/June 2007. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Pioro:2004:RFC

[PM04] Michał Pióro and Deepankar Medhi. Routing, flow, and capacity design in communication and computer networks. The Morgan Kaufmann series in networking. Elsevier/Morgan Kaufmann, San Francisco, CA, USA, 2004. ISBN 0-12-557189-5. xxviii + 765 pp. LCCN TK5105.5 .P56 2004. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/els051/2004301909.
                                        html;   http://www.loc.gov/
                                        catdir/toc/els051/2004301909.
                                        html.

Pasch:2001:UIP

[PN01] Grete Pasch and Kent Norsworthy. Using Internet primary sources to teach critical thinking skills in world languages. Greenwood professional guides in school librarianship. ISSN: 1074-150X. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2001. ISBN 0-313-31259-1. vii + 135 pp. LCCN P53.285 .P37 2001.

Patil:2003:IIC

[PN03] Sanjay Patil and Eric Newcomer. IEEE Internet computing: Spotlight ebXML and Web services. IEEE Distributed Systems Online, 4(5), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0305/f/wp3spot.htm.

Politycki:2000:MID

[Pol00] Matthias Politycki. Marietta — die Idee, der Datensatz und der Strohhut: Schreiben und Schreiben-Lassen im Internet. (German) [Marietta — the idea, the data record and the straw hut: Writing and writing loss in the Internet], volume 1 of Abhandlungen der Klasse der Literatur ISSN: 0002-2985; Jahrg. 2000, Nr.

1. F. Steiner, Stuttgart, Germany, 2000. ISBN 3-515-07794-4. ISSN 0002-2985. 20 pp. LCCN AS182 .M281 2001 no.1.

Poole:2003:NSP

[Poo03] Owen Poole. Network Security: a Practical Guide. Butterworth-Heinemann, Boston, MA, USA, 2003. ISBN 0-7506-5033-8. xi + 212 pp. LCCN TK5105.59 .P66 2003. US$47.95.

Powell:2001:HGI

[Pow01] Carol Clark Powell. The Harcourt guide to the Internet. Harcourt College Publishers, Fort Worth, TX, USA, 2001. ISBN 0-15-506701-X. 96 pp. LCCN TK5105.875.I57 P688 2001.

Powers:2002:ILG

[Pow02] Dennis M. Powers. The Internet legal guide: everything you need to know when doing business online. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-16423-2 (paperback). xiv + 272 pp. LCCN KF390.5.C6 P69 2002.

Pottruck:2000:CMP

[PP00] David S. Pottruck and Terry Pearce. Clicks and mortar: passion-driven growth in an Internet-driven world. The Jossey-Bass business and management series. Jossey-Bass Publishers, San Francisco, CA, USA, 2000. ISBN 0-7879-5273-7. xxxiii + 314 pp. LCCN HD30.37 .P68 2000.

Perkins:2001:IBI

[PP01a] Anthony B. Perkins and Michael C. Perkins. The Internet bubble: the inside story on why it burst — and what you can do to profit now. HarperBusiness, New York, NY, USA, rev. ed edition, 2001. ISBN 0-06-664001-6. viii + 327 pp. LCCN HD9696.8.A2 P45 2001.

Pottruck:2001:CMP

[PP01b] David S. Pottruck and Terry Pearce. Clicks and mortar: passion-driven growth in an Internet-driven world. The Jossey-Bass business and management series. Jossey-Bass, San Francisco, CA, USA, second edition, 2001. ISBN 0-7879-5688-0. xxxv + 314 pp. LCCN HD30.37 .P68 2001.

Palaia:2005:DMO

[PP05] Wayne Palaia and Christina Palaia. Degunking Microsoft Office. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. xxvi + 385 pp. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://site.ebrary.com/
                                        lib/ucsc/Doc?id=10080002.

Price:2000:IEB

[Pri00] Christopher Price. The Internet entrepreneurs: business rules are good: break them. FT.com, London, UK, 2000. ISBN 0-273-64921-3 (paperback). xxiv + 196 pp. LCCN HD30.37 .P74 2000.

Priestman:2002:WRR

[Pri02] Chris Priestman. Web Radio: Radio Production for Internet Streaming. Focal Press, Boston, MA, USA and Oxford, UK, 2002. ISBN 0-240-51635-4. xx + 275 pp. LCCN ???? US$29.99.

Proctor:2000:PID

[Pro00] Paul E. Proctor. The Practical Intrusion Detection Handbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-025960-8. xxiv + 359 pp. LCCN QA76.9.A25 P763 2001. US$49.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130259608.html.

Provenzo:2002:IWW

[Pro02] Eugene F. Provenzo. The Internet and the World Wide Web for teachers. Allyn and Bacon, Inc., Needham Heights, MA, USA, second edition, 2002. ISBN 0-205-34349-X. vi + 154 pp. LCCN LB1044.87 .P763 2002. First edition published under title: “The Internet and the World Wide Web for preservice teachers”.

Paschalidis:2000:CDP

[PT00] Ioannis Ch. Paschalidis and John N. Tsitsiklis. Congestiondependent pricing of network services. IEEE/ACM Transactions on Networking, 8(2): 171–184, 2000. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        ton/2000-8-2/p171-paschalidis/

.

Purba:2002:NDI

[Pur02] Sanjiv Purba. New directions in Internet management. The Auerbach best practices series. Auerbach, Princeton, NJ, USA, second edition, 2002. ISBN 0-8493-1160-8. xviii + 804 pp. LCCN HD30.37 .I56 2002. Rev. ed. of: Internet management. 2000 ed. c2000.

Puzmanova:2003:BRD

[Puz03a] Rita Puzmanova. Book review: Delivering voice over IP and multimedia with session initiation protocol: a review of Internet Communications Using SIP by Henry Sinnreich and Alan B. Johnston. IEEE Distributed Systems Online, 4(5), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

                                        http://dsonline.computer.
                                        org/0305/d/bks_a.htm.

Puzmanova:2003:BRI

[Puz03b] Rita Puzmanova. Book review: Internet for the mobile world: a review of The Wireless Mobile Internet: Architectures, Protocols and Services by Abbas Jamalipour. IEEE Distributed Systems Online, 4(11), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0311/d/bks_b.htm.

Phillips:2001:GRI

[PvS01] Deborah M. Phillips and Hans A. von Spakovsky. Gauging the risks of Internet elections. Communications of the ACM, 44(1):73, January 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2001-44-1/p73-phillips/.

See correction [CTBA^{+}01].

Qian:2000:SFI

[Qia00] Zhenyu Qian. Standard fixpoint iteration for Java bytecode verification. ACM Transactions on Programming Languages and Systems, 22(4): 638–672, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
toplas/2000-22-4/p638-qian/

.

Raisanen:2001:QSV

[Räi01] Vilho Räisänen. Quality of service and voice-over-IP. Dr. Dobbs Journal, 26(5):50, 54, 56, 58, May 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Rajagopal:2000:MOS

[Raj00] Raj Rajagopal, editor. Multi-Operating System Networking: Living with UNIX, NetWare, and NT. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-9831-2. 1360 (est.) pp. LCCN ???? US$79.95.

Rao:2002:NEE

[Rao02] Nageswara S. V. Rao. Netlets for end-to-end delay minimization in distributed computing over the Internet using two-paths. The International Journal of High Performance Computing Applications, 16 (3):285–292, Fall 2002. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Rappaport:2002:WCP

[Rap02] Theodore S. Rappaport. Wireless Communications: Principles and Practice. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2002. ISBN 0-13-042232-0. xxiii + 707 pp. LCCN TK5103.2.R366 2002. US$92.00. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130422320.html.

Rash:2005:IPA

[Ras05a] Michael Brandon Rash, editor. Intrusion prevention and active response: deploying network and host IPS. Syngress Publishing, Inc., Rockland, MA, USA, 2005. ISBN 1-932266-47-X (paperback). xx + 402 pp. LCCN TK5105.59 .I58 2005.

Rasmussen:2005:MRI

[Ras05b] Soren Rasmussen. Managing remote installation services. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2005. ISBN 1-55558-337-7. ???? pp. LCCN QA76.76.O63 R375 2005.

Rahman:2001:ICS

[RB01] Syed Mahbubur Rahman and Robert J. Bignall. Internet commerce and software agents: cases, technologies, and opportunities. Idea Group Pub, Hershey, PA, USA, 2001. ISBN 1-930708-01-7 (paper). v + 408 pp. LCCN HF5548.32 .I588 2001.

Rodriguez:2002:DPA

[RB02] Pablo Rodriguez and Ernst W.

Biersack. Dynamic parallel access to replicated content in the Internet. IEEE/ACM Transactions on Networking, 10(4):455–465, August 2002.

Reiter:2001:ILB

[RBB01] David Reiter, Elizabeth Blumenfeld, and Mark Boulding. Internet law for the business lawyer. Section of Business Law, American Bar Association, Chicago, IL, USA, 2001. ISBN 1-57073-989-7 (softcover). xxiv + 555 pp. LCCN KF390.5.C6 I57 2001.

Raman:2002:IIT

[RBS02] Suchitra Raman, Hari Balakrishnan, and Murari Srinivasan. ITP: an Image Transport Protocol for the Internet. IEEE/ACM Transactions on Networking, 10(3):297–307, June 2002.

Rogers:2005:NSE

[RC05] Russ Rogers and Bryan Cunningham, editors. Network security evaluation using the NSA IEM. Syngress Publishing, Inc., Rockland, MA, USA, 2005. ISBN 1-59749-035-0. xxvi + 437 pp. LCCN TK5105.59 .N33 2005.

Rothman:2005:BCD

[RD05] Johanna Rothman and Esther Derby. Behind Closed Doors: Secrets of Great Management. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9766940-2-6. 176 (est.) pp. LCCN ???? US$24.95, CDN$34.95, UK£17.50. URL http://

www.oreilly.com/catalog/
0976694026/.

Redeker:2000:EPZ

[Red00] Helmut Redeker. Der EDV-Prozess: zivilrechtliche Probleme von Software und Internet. (German) [The EDP process: civil problems of software and Internet]. NJW-Schriftenreihe. Beck, München, Germany, second edition, 2000. ISBN 3-406-46714-8. xxvii + 417 pp. LCCN KK164.C66 .R43 2000.

Reed:2000:ILT

[Ree00] Chris Reed. Internet law: text and materials. Law in context. Butterworths, London, UK, 2000. ISBN 0-406-98141-8. xxviii + 280 pp. LCCN KD667.C65 R44 2000.

Reynders:2003:PTI

[Rey03] Deon Reynders. Practical TCP/IP and Ethernet Networking for Industry. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-7506-5806-1. ???? pp. LCCN ???? US$49.99.

Rolia:2002:SCC

[RFP02] Jerry Rolia, Rich Friedrich, and Chandrakant Patel. Service centric computing — next generation Internet computing. Lecture Notes in Computer Science, 2459:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2459/24590463.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2459/24590463.
pdf.

Rangel:2002:GDI

[RG02a] Paul Rangel and Jeremy Giovannetti. Genomes and databases on the Internet: a practical guide to functions and applications. Horizon Scientific, Wymondham, UK, 2002. ISBN 1-898486-31-X. xi + 223 pp. LCCN QH447 R26 2002.

Rodriguez:2002:EIU

[RG02b] Carlos Gregorio Rodríguez and Pedro Palao Gostanza. Ergotracer: An Internet user behaviour tracer. Lecture Notes in Computer Science, 2346:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2346/23460121.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2346/23460121.pdf.

Richardson:2005:SIP

[RG05] Jared R. Richardson and William Atwill Gwaltney. Ship it!: a practical guide to successful software projects. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9745140-4-7. xiii + 198 pp. LCCN Qa76.76.D47 R53 2005.

Rose:2000:ICL

[RGE00] Gene Rose, Pam Greenberg, and Brenda Erickson. Internet coverage of legislative proceedings. NCSL legisbrief; v. 8, no. 48 8(48), National Conference of State Legislatures, Denver, CO, USA, November/ December 2000. 2 pp. Title from caption. “Nov./Dec. 2000.”.

Rash:2005:ITF

[RHC05] James Rash, Keith Hogie, and Ralph Casasanta. Internet technology for future space missions. Computer Networks (Amsterdam, Netherlands: 1999), 47(5):651–659, April 5, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Renganarayanan:2002:IAE

[RHN02] Vidya Renganarayanan, Abdelsalam (Sumi) Helal, and Amar Nalla. Internet agents for effective collaboration. Lecture Notes in Computer Science, 2521:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2521/25210033.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2521/25210033.pdf.

Rhoton:2000:PGI

[Rho00] John Rhoton. Programmer’s guide to Internet mail: SMTP, POP, IMAP, and LDAP. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2000. ISBN 1-55558-212-5 (paperback). xix + 291 pp. LCCN TK5105.73 .R48 2000.

Rhoton:2002:WIE

[Rho02] John Rhoton. The wireless Internet explained. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2002. ISBN 1-55558-257-5 (paperback). xv + 248 pp. LCCN TK5103.485 .R52 2002.

Richardson:2000:I

[Ric00] Marjorie Richardson. The Internet. Linux Journal, 72:??, April 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

                                        //noframes.linuxjournal.
                                        com/lj-issues/issue72/3932.
                                        html.

Riccardi:2001:PDS

[Ric01a] Greg Riccardi. Principles of database systems with Internet and Java applications. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-61247-X. xxi + 518 pp. LCCN QA76.9.D3 R52 2001.

Richardson:2001:IMR

[Ric01b] Paul Richardson. Internet marketing: readings and online resources. The McGraw-Hill/Irwin series in marketing. McGraw-Hill/Irwin, New York, NY, USA, 2001. ISBN 0-07-242793-0 (paperback). xii + 350 pp. LCCN HF5415.1265 .I568 2001.

Riekert:2002:ARI

[Rie02] W.-F. Riekert. Automated retrieval of information in the Internet by using thesauri and gazetteers as knowledge sources. J.UCS: Journal of Universal Computer Science, 8(6):581–??, June 28, 2002. CODEN ???? ISSN 0948-6968. URL

http://www.jucs.org/jucs_
8_6/automated_retrieval_
of_information.

Rischpater:2001:IA

[Ris01] Ray Rischpater. Internet appliances. A Wiley tech brief. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-44111-2 (paperback). xiv + 210 pp. LCCN TK510.875.I57 R49 2001.

Ristic:2005:AS

[Ris05] Ivan Ristic. Apache security.

O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00724-8. xxii + 396 pp. LCCN TK5105.8885.A63 R57 2005.

Ritke:2000:IT

[Rit00] Ronn Ritke. An Internet toolbox. Dr. Dobbs Journal, 25(6): 84, 86–87, 89, June 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Reddick:2001:OJU

[RK01a] Randy Reddick and Elliot King. The online journ@list: using the Internet and other electronic resources. Harcourt College Publishers, Fort Worth, TX, USA, third edition, 2001. ISBN 0-15-506752-4. x + 277 pp. LCCN PN4784.E5 R43 2001.

Rice:2001:IHC

[RK01b] Ronald E. Rice and James E. Katz. The Internet and health communication: experience and expectations. Sage Publications, Thousand Oaks, CA, USA, 2001. ISBN 0-7619-2232-6 (hardcover). xvi + 459 pp. LCCN R859.7.E43 I56 2001.

Ramy:2000:NIL

[RM00] Herbert N. Ramy and Samantha A. Moppett. Navigating the Internet: legal research on the World Wide Web. Fred B. Rothman and Co., Littleton, CO, USA, 2000. ISBN 0-8377-1077-4 (paper). xi + 112 pp. LCCN KF242.A1 R36 2000.

Roze:2002:TCA

[RM02] Maris Roze and Simon Maxwell. Technical communication in the age of the Internet. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2002. ISBN 0-13-020574-5. x + 305 pp. LCCN T10.5 .R65 2002.

Roth:2000:LMB

[RO00] Alvin E. Roth and Axel Ockenfels. Last minute bidding and the rules for ending second-price auctions: theory and evidence from a natural experiment on the Internet. NBER working paper series 7729, National Bureau of Economic Research, Cambridge, MA, USA, 2000. 47 pp.

Roberts:2000:IWB

[Rob00] Lawrence G. Roberts. Internet watch: Beyond Moore’s law: Internet growth trends. Computer, 33(1):117–119, January 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co2000/pdf/r1117.
pdf.

Roederer:2001:RSI

[Roe01] Charlotte Roederer. Regulation of securities on the Internet. Aspen Law and Business, Gaithersburg, MD, USA, 2001. ISBN 0-7355-1652-9 (set), 0-7355-2434-3 (vol. 1), 0-7355-2435-1 (vol. 2). various pp. LCCN K1114 .R64 2001. Two volumes. Loose-leaf for updating.

Rosen:2000:TIJ

[Ros00] Jonathan Rosen. The Talmud and the Internet: a journey between worlds. Farrar, Straus and Giroux, New York, NY, USA, 2000. ISBN 0-374-27238-7. 132 pp. LCCN BM504 .R657 2000.

Rose:2001:IIE

[Ros01] Richard Rose. The Internet’s impact on East Asian governance: openness, impersonal rules and accountability, volume 354 of Studies in public policy. Centre for the Study of Public Policy, University of Strathclyde, Scotland, 2001.

ISSN 0140-8240. 56 pp. LCCN DS504.5 .R67 2001.

Rose:2002:BDG

[Ros02] Marshall T. Rose. BEEP: The Definitive Guide: Developing New Applications for the Internet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00244-0. xv + 204 pp. LCCN TK5105.55 .R66 2002. US$34.95. URL http://www.

oreilly.com/catalog/beep.

Ross:2003:BWF

[Ros03] John Ross. The book of Wi-Fi: install, configure, and use 802.11b wireless networking. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-45-X (paperback). xvii + 260 pp. LCCN TK5105.78 .R67

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0716/2003000711-d.html.

Ross:2008:BWP

[Ros08] John Ross. The book of wireless: a painless guide to wi-fi and broadband wireless. No Starch Press, San Francisco, CA, USA, second edition, 2008. ISBN 1-59327-169-7. xxi + 326 pp. LCCN TK5105.78 .R67

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0913/2007048461-b.html.

Ries:2000:ILI

[RR00] Al Ries and Laura Ries. The 11 immutable laws of Internet branding. HarperBusiness, New York, NY, USA, 2000. ISBN 0-06-019621-1. viii + 175 pp. LCCN HF5415.1265 .R53 2000.

Reilly:2002:JNP

[RR02] David Reilly and Michael Reilly. Java network programming and distributed computing. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-71037-4. xvii + 464 pp. LCCN QA76.73.J38 R45 2002.

Ratner:2003:NGI

[RR03a] Daniel Ratner and Mark Ratner. Nanotechnology: a gentle introduction to the next big idea. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-101400-5. xiv + 188 pp. LCCN T174.7 .R38 2003.

Ratner:2003:NHS

[RR03b] Daniel Ratner and Mark Ratner. Nanotechnology and Homeland Security: New Weapons for New Wars. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-145307-6. 176 (est.) pp. LCCN UA927.R38 2004. US$24.95.

Ransome:2005:VS

[RR05a] James F. Ransome and John W. Rittinghouse. VoIP security. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 1-55558-332-

6. xxvii + 402 pp. LCCN TK5105.8865 .R36 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els051/2004059935.
html.

Rittinghouse:2005:IIM

[RR05b] John Rittinghouse and James Ransome. IM Instant Messaging Security. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 1-55558-338-5. 432 (est.) pp. LCCN ???? URL http:

//books.elsevier.com/us/
mk/us/subindex.asp?isbn=
1555583385.

Rao:2000:NGN

[RS00a] S. Rao and Kaare Ingar Sletta, editors. Next generation networks: networks and services for the information society: 5th IFIP TC6 international symposium, INTERWORK-ING 2000, Bergen, Norway, October 3–6, 2000: proceedings, volume 1938 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-41140-2 (softcover). LCCN QA267.A1 L43 no.1938.

Runkehl:2000:ZII

[RS00b] Jens Runkehl and Torsten Siever. Das Zitat im Internet: ein electronic style guide zum Publizieren, Bibliografieren, Zitieren. (German) [The quotation in the Internet: an electronic style guide for publishing, bibliography work, citation]. Revonnah, Hannover, Germany, 2000. ISBN 3-927715-83-2 (paperback). 128 pp. LCCN PN171.F56 R86 2000. DM25.00.

Ramaswami:2002:ONP

[RS02a] Rajiv Ramaswami and Kumar N. Sivarajan. Optical Networks: a Practical Perspective. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2002. ISBN 1-55860-655-6. xxix + 831 pp. LCCN TK5103.59 .R36 2002. US$69.95.

Rinat:2002:MIP

[RS02b] Ran Rinat and Scott Smith.

Modular Internet programming with cells. Lecture Notes in Computer Science, 2374:257–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer-ny.com/link/service/
series/0558/bibs/2374/23740257.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2374/23740257.
pdf.

Rubin:2001:SCR

[Rub01] Avi Rubin. Security considerations for remote electronic voting over the Internet. ;login: the USENIX Association newsletter, 26(1):

??, February 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL http://www.
                                        usenix.org/publications/
                                        login/2001-02/pdfs/rubin.
                                        pdf.

Retana:2000:EIB

[RWS00] Alvaro Retana, Russ White, and Don Slice. EIGRP for IP: basic operation and configuration. The Addison-Wesley networking basics series. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-65773-2. xi + 123 pp. LCCN TK5105.543.R48 2000.

Ryan:2000:VUI

[Rya00]  Steve Ryan.  The virtual

university: the Internet and resource-based learning. Open and distance learning series. Kogan Page, London, UK, 2000. ISBN 0-7494-2508-3. xiii + 204 pp. LCCN LB1044.87 U66 2000.

Ribeiro:2005:STS

[RZMD05] Vinay J. Ribeiro, Zhi-Li Zhang, Sue Moon, and Christophe Diot. Smalltime scaling behavior of Internet backbone traffic. Computer Networks (Amsterdam, Netherlands: 1999), 48(3): 315–334, June 21, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Shorey:2006:MWS

[S^{+}06] Rajeev Shorey et al., editors. Mobile, wireless, and sensor networks: technology, applications, and future directions. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-471-71816-5 (hardcover). xvii + 430 pp. LCCN TK5103.2 .M634

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0622/2005010792-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0654/2005010792-
b.html;   http://www.loc.
gov/catdir/toc/ecip0511/
2005010792.html.

Soref:2000:ESD

[SA00] Jeremy Soref and Sheridan Adjei. Electronic software distribution. Dr. Dobbs Journal, 25(9):100, 102–106, September 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_09/license.
txt.

Samoriski:2002:ICC

[Sam02] Jan H. Samoriski. Issues in cyberspace: communication, technology, law, and society on the Internet frontier. Allyn and Bacon, Inc., Needham Heights, MA, USA, 2002. ISBN 0-205-32262-X. xv + 384 pp. LCCN KF390.5.C6 S25 2002.

Santini:2001:EID

[San01] Simone Santini. Exploratory Image Databases: Content-Based Retrieval. Academic Press, New York, NY, USA, 2001. ISBN 0-12-619261-8. xviii + 613 pp. LCCN ???? US$79.95.

Sandhu:2002:IIC

[San02] Ravi Sandhu. IEEE Internet computing: Guest Editor’s introduction: The technology of trust. IEEE Distributed Systems Online, 3(11), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0211/f/w6gei.htm.

Sandhu:2003:IIC

[San03]  Ravi Sandhu.  IEEE In-

ternet computing: Security track: Good enough security. IEEE Distributed Systems Online, 4(1), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0301/f/wp1trak.htm.

Stoica:2002:III

[SAR^{+}02] Ion Stoica, Daniel Adkins, Sylvia Ratnasamy, Scott Shenker, Sonesh Surana, and Shelley Zhuang. Internet indirection infrastructure. Lecture Notes in Computer Science, 2429:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2429/24290191.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2429/24290191.pdf.

Sauers:2001:UIR

[Sau01a] Michael Sauers. Using the Internet as a reference tool: a how to-do-it manual for librarians, volume 109 of How-to-do-it manuals for librarians; no. 109. Neal-Schuman Publishers, New York, NY, USA, 2001. ISBN 1-55570-417-4. xi + 143 pp. LCCN Z711.45 .S28 2001.

Saunders:2001:PIL

[Sau01b] Kurt M. Saunders. Practical Internet law for business. The Artech House telecommunications library. Artech House Inc., Norwood, MA, USA, 2001. ISBN 1-58053-003-6. xiii + 162 pp. LCCN KF390.5.C6 S26 2001.

Sayood:2003:LCH

[Say03] Khalid Sayood, editor. Lossless Compression Handbook. Academic Press, New York, NY, USA, 2003. ISBN 0-12-620861-1. xx + 455 pp. LCCN TK5102.5 .L69 2003. US$89.95.

Stamps:2001:TIH

[SB01] Robert F. Stamps and Peter Barach. The therapist’s Internet handbook. Norton, New York, NY, USA, 2001. ISBN 0-393-70342-8 (paperback/CD-ROM). xvi + 368 pp. LCCN RC489.I54 S73 2001.

Smith:2002:ILR

[SC02] Graham J. H. Smith and Simon Chalton. Internet law and regulation. Sweet and Maxwell, London, UK, third edition, 2002. ISBN 0-421-70590-6. xlvii + 737 pp. LCCN KD667.C65 I57 2002.

Sotomayor:2005:GTP

[SC05] Borja Sotomayor and Lisa Childers. Globus Toolkit 4: Programming Java Services. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2005. ISBN 0-12-369404-3. 320 (est.) pp. LCCN ???? US$39.95.

Smith:2007:WN

[SC07] Clint Smith and Daniel Collins. 3G wireless networks. McGraw-Hill communications. McGraw-Hill, New York, NY, USA, second edition, 2007. ISBN 0-07-226344-X. xxiv + 695 pp. LCCN TK6570.M6 S59

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0713/2007272081-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0713/2007272081-
d.html;   http://www.loc.
gov/catdir/toc/fy0706/2007272081.
html.

Schneier:2000:IRS

[Sch00a]  Bruce Schneier.  Inside

risks: semantic network attacks. Communications of the ACM, 43(12):168, December 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/

2000-43-12/p168-schneier/

. See letters [CZB^{+}01, CTBA^{+}01].

Schwartau:2000:CSH

[Sch00b] Winn Schwartau. CyberShock:

surviving hackers, phreakers, identity thieves, Internet terrorists, and weapons of mass disruption. Thunder’s Mouth Press, New York, NY, USA, 2000. ISBN 1-56025-246-4. xxvi + 470 pp. LCCN QA76.9.A25 S3537 2000.

Schachter:2001:LIS

[Sch01] Madeleine Schachter. Law of Internet speech. Carolina Academic Press, Durham, NC, USA, 2001. ISBN 0-89089-944-4. xvi + 544 pp. LCCN KF4772.A7 S33 2001.

Schiesser:2002:ISM

[Sch02a]  Rich Schiesser.  IT Sys-

tems Management: Designing, Implementing, and Managing World-Class Infrastructures. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-087678-X. 528 pp. LCCN ???? US$39.99. URL http:

//www.phptr.com/ptrbooks/
ptr_013087678X.html.

Schulzrinne:2002:IMS

[Sch02b] Henning Schulzrinne. Internet multimedia — streaming media (tutorial). Lecture Notes in Computer Science, 2515:365– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2515/25150365.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2515/25150365.pdf.

Schwartz:2005:MWC

[Sch05] Mischa Schwartz. Mobile wireless communications. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-84347-2. xi + 457 pp. LCCN TK5103.2 .S37 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/cam041/2004049268.
                                        html;   http://www.loc.gov/
                                        catdir/toc/cam041/2004049268.
                                        html.

Schlossnagle:2006:SIA

[Sch06]  Theo Schlossnagle.  Scal-

able Internet Architectures. Howard W. Sams, Indianapolis, IN 46268, USA, 2006. ISBN 0-672-32699-X (print), 0-7686-6618-X (electronic). 288 (est.) pp. LCCN ???? US$49.99. URL http:/

                                        /www.samspublishing.com/
                                        bookstore/product.asp?isbn=
                                        067232699X.

Score:2000:LI

[Sco00] Score. Learning and the Internet. SCORE!, Chicago, IL, USA, 2000. 168 pp. LCCN LB1044.87 .L42 2000 CRB.

Scott:2001:ITL

[Sco01] Michael D. (Michael Dennis) Scott. Internet and technology law desk reference. Aspen Law and Business, Gaithersburg, MD, USA, 2001 ed edition, 2001. ISBN 0-7355-1526-3. xvii + 663 pp. LCCN KF390.5.C6 S383 2001.

Schwartz:2000:IBO

[SDB00] David G. Schwartz, Monica Divitini, and Terje Brasethvik, editors. Internet-based organizational memory and knowledge management. Idea Group Pub., Hershey, PA, USA, 2000. ISBN 0-585-29241-8 (electronic book). LCCN ????

Sharma:2000:UIP

[SE00] Martha B. Sharma and Gary S. Elbow. Using Internet primary sources to teach critical thinking skills in geography. Greenwood professional guides in school librarianship. ISSN: 1074-150X. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2000. ISBN 0-313-30899-3. xv + 165 pp. LCCN G73.S393 2000.

Shirmohammadi:2003:JJT

[SEGS03] S. Shirmohammadi, A. ElSaddik, N. D. Georganas, and R. Steinmetz. JASMINE: a Java tool for multimedia collaboration on the Internet. Multimedia Tools and Applications, 19(1):5–28, 2003. CO-DEN MTAPFB. ISSN 1380-7501.

Sheth:2001:IM

[SEK01a] Jagdish N. Sheth, Abdolreza Eshghi, and Balaji C. Krishnan. Internet marketing. Harcourt College Publishers, Fort Worth, TX, USA, 2001. ISBN 0-03-032133-6. 419 pp. LCCN HF5415.1265 .S533 2001.

Speed:2001:PIS

[SEK01b] Timothy Speed, Juanita Ellis, and Steffano Korper. The Personal Internet Security Guidebook: Keeping Hackers and Crackers out of Your Home. Academic Press, New York, NY, USA, 2001. ISBN 0-12-656561-9. xxiv + 202 pp. LCCN ???? US$44.95.

Spindler:2000:VIP

[SF00] Gerald Spindler and Andreas Fuchs. Vertragsrecht der Internet-Provider. (German) [Treaty rights of the Internet provider]. O. Schmidt, Köln, Germany, 2000. ISBN 3-504-56034-7. xlvi + 713 pp. LCCN KK164.C66 V47 2000.

Strauss:2001:E

[SF01] Judy Strauss and Raymond Frost. Emarketing. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2001. ISBN 0-13-032264-4. xxiii + 519 pp. LCCN HF5415.1265 .S774 2001. Updated ed. of: “Marketing on the Internet” 1999.

Siganos:2003:PLL

[SFFF03] Georgos Siganos, Michalis Faloutsos, Petros Faloutsos, and Christos Faloutsos. Power laws and the AS-level Internet topology. TON, 11(4):514– 524, August 2003.

Street:2001:LI

[SG01] F. Lawrence Street and Mark P. Grant. Law of the Internet. LexisNexis, Charlottesville, VA, USA, 2001. ISBN 0-327-15935-9. various pp. LCCN KF390.5.C6 S76 2001.

Saroiu:2002:AIC

[SGD^{+}02] Stefan Saroiu, Krishna P.

Gummadi, Richard J. Dunn, Steven D. Gribble, and Henry M. Levy. An analysis of Internet content delivery systems. Operating Systems Review, 36(5S):315–327, December 2002. CODEN OSRED8. ISSN 0163-5980.

Sharpe:2000:PTC

[Sha00] Charles C. Sharpe. Patent, trademark, and copyright searching on the Internet. Mc-Farland, Jefferson, NC, USA, 2000. ISBN 0-7864-0757-3 (softcover). ix + 229 pp. LCCN T210 .S53 2000.

Sharma:2001:WIE

[Sha01] Chetan Sharma. Wireless Internet enterprise applications: a Wiley tech brief. Wiley tech brief series. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-39382-7 (paperback). xx + 236 pp. LCCN TK5103.2 .S52

Seok:2002:AFC

[SHK02] Seung-Joon Seok, Sung-Min Hong, and Chul-Hee Kang. An adaptive flow control scheme for improving TCP performance in wireless Internet. Lecture Notes in Computer Science, 2510:458–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2510/25100458.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2510/25100458.pdf.

Shuman:2001:ILI

[Shu01] Bruce A. Shuman. Issues for libraries and information science in the Internet age. Libraries Unlimited, Inc., Englewood, CO, USA, 2001. ISBN 1-56308-805-3. xviii + 228 pp. LCCN Z665 .S486 2001; Z674.75.I58 S58 2001.

Sadiku:2001:SLA

[SI01] Matthew N. O. Sadiku and Mohammad Illyas. Simulation of Local Area Networks. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-2473-4. xi + 200 pp. LCCN TK5105.7 .S22 1995. US$104.95.

Silverstein:2000:BBI

[Sil00] Barry Silverstein. Business-to-business Internet marketing: seven proven strategies for increasing profits through direct marketing. Maximum Press, Gulf Breeze, FL, USA, second edition, 2000. ISBN 1-885068-38-7 (paperback). xxxiii + 490 pp. LCCN HF5415.1265 .S535 2000.

Simmers:2002:AIU

[Sim02a] Claire A. Simmers. Aligning Internet usage with business priorities. Communications of the ACM, 45(1):71– 74, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Simpson:2002:XXL

[Sim02b] John E. Simpson. XPath and XPointer: Locating Content in XML Documents. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00291-2. x + 196 pp. LCCN QA76.76.H94 S578 2002 Stacks. US$24.95. URL

http://safari.oreilly.com/
0596002912;   http://www.
oreilly.com/catalog/xpathpointer.

Sindell:2001:LMI

[Sin01a] Kathleen Sindell. Loyalty marketing for the Internet Age: how to identify, attract, serve, and retain customers in an ecommerce environment. Dearborn Financial Pub., Chicago, IL, USA, 2001. ISBN 0-7931-4033-1. xx + 345 pp. LCCN HF5415.525 .S53 2000.

Singhal:2001:WWA

[Sin01b] Sandeep Singhal. WAP — The Wireless Application Protocol: writing applications for the mobile Internet. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70311-4. xx + 678 pp. LCCN TK5105.55 .W37 2001.

Six:2000:PNV

[Six00] Jann Six. Der privatrechtliche Namensschutz von und vor Domanennamen im Internet. (German) [Name protection under private law of and before domain names in the Internet], volume 158 of Zurcher Studien zum Privatrecht. Schulthess, Zürich, Switzerland, 2000. ISBN 3-7255-4099-3. xx + 203 pp. LCCN ????

Sinnreich:2001:ICU

[SJ01] Henry Sinnreich and Alan Johnston. Internet communications using SIP: delivering VolP and multimedia services with Session Initiation Protocol. Networking Council series. Wiley Computer Publishers, New York, NY, USA, 2001. ISBN 0-471-41399-2 (hardcover). xx + 298 pp. LCCN TK5105.55 .S56 2001.

Soares:2000:IGH

[SK00] John Soares and Daniel J. Kurland. Internet guide for history. Wadsworth Publishing Company, Pacific Grove, CA, USA, second edition, 2000. ISBN 0-534-56906-4. 63 pp. LCCN D16.255.C65 S62 2000. First ed., c1997, by D. J. Kurland and J. Soares.

Shinomiya:2002:NNA

[SKS^{+}02] Shunsuke Shinomiya, Yutaka Kidawara, Takeshi Sakurada, Seiji Tsuchiike, and Shin ichi Nakagawa. NA-DIA: Network accessible device on the Internet architecture. Lecture Notes in Computer Science, 2343:146– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2343/23430146.htm;
http://link.springer.de/
link/service/series/0558/
papers/2343/23430146.pdf.

Scharwaechter:2007:AAE

[SKW^{+}07] Hanno Scharwaechter, David Kammler, Andreas Wieferink, Manuel Hohenauer, Kingshuk Karuri, Jianjiang Ceng, Rainer Leupers, Gerd Ascheid, and Heinrich Meyr. ASIP architecture exploration for efficient IPSec encryption: a case study. ACM Transactions on Embedded Computing Systems, 6(2):12:1–12:??, May 2007. CODEN ???? ISSN 1539-9087.

Sloane:2001:MIB

[SL01] Andy Sloane and Dave Lawrence.

Multimedia Internet broadcasting: quality, technology and interface. Computer communications and networks. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 1-85233-283-2. xviii + 233 pp. LCCN QA76.575 .M83329 2001.

Shim:2002:WTI

[SL02] Simon S. Y. Shim and Yen-Jen Lee. Web technologies: Interactive TV: VoD meets the Internet. Computer, 35(7):108–109, July 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/07/r7108.
htm;   http://csdl.computer.
org/dl/mags/co/2002/07/r7108.
pdf;   http://www.computer.
org/computer/co2002/r7108abs.
htm.

Skoudis:2006:CHR

[SL06] Ed Skoudis and Tom Liston. Counter hack reloaded: a stepby-step guide to computer attacks and effective defenses. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2006. ISBN 0-13-148104-5 (paperback). ???? pp. LCCN TK5105.59 .S57

2006. URL http://www.loc.
gov/catdir/toc/ecip0519/
2005027164.html.

Slevin:2000:IS

[Sle00] James Slevin. The Internet and society. Blackwell, Malden, MA, USA, 2000.

ISBN 0-7456-2086-8, 0-7456-2087-6. xi + 266 pp. LCCN HM851 .S58 2000.

Slone:2000:LAN

[Slo00] John P. Slone, editor. Local Area Network Handbook. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, sixth edition, 2000. ISBN 0-8493-9838-X. xiii + 676 pp. LCCN TK5105.7 .L5816 2000. US$95.00.

Simpson:2000:ISP

[SM00a] Carol Mann Simpson and Sharron L. McElmeel. Internet for schools: a practical guide. Professional growth series. Linworth Pub., Worthington, OH, USA, third edition, 2000. ISBN 0-938865-98-6 (paperback). viii + 317 pp. LCCN Z675.S3 S5945 2000.

Stanley-Marbell:2000:IAD

[SM00b] Phillip Stanley-Marbell. Inferno application development with Limbo. Dr. Dobbs Journal, 25(12):113–114, 116, 118, 120, December 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_12/inferno.
txt.

Sukhatme:2000:EIE

[SM00c] Gaurav S. Sukhatme and Maja J. Mataric. Embedding the Internet: embedding robots into the Internet. Communications of the ACM, 43(5):67, May 2000. CODEN CACMA2.

ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-5/p67-sukhatme/.

Shearer:2002:DPI

[SM02a] J. Shearer and H. Maurer. Is democracy possible in the Internet? J.UCS: Journal of Universal Computer Science, 8(3):396–??, March 28, 2002. CODEN ???? ISSN 0948-6968. URL http://

www.jucs.org/jucs_8_3/is_
democracy_possible_in.

Sundaresan:2002:APM

[SM02b] Neel Sundaresan and Reshad Moussa. Algorithms and programming models for efficient representation of XML for Internet applications. Computer Networks (Amsterdam, Netherlands: 1999), 39(5):681–697, August 5, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/96/52/40/
abstract.html.

Smith:2000:KGI

[Smi00a] David E. Smith. Knowledge, groupware, and the Internet. Resources for the knowledgebased economy. Butterworth-Heinemann, Boston, MA, USA, 2000. ISBN 0-7506-7111-4 (paperback). xvi + 353 pp. LCCN HD30.2 .K6364 2000.

Smith:2000:IMM

[Smi00b] John R. Smith, editor. Internet multimedia management systems: 6–7 November 2000, Boston, USA, volume 4210 of SPIE proceedings series; v. 4210 4007195700. SPIE Optical Engineering Press, Bellingham, WA, USA, 2000. ISBN 0-8194-3875-8. LCCN QA76.575 .I684 2000.

Smith:2000:ERO

[Smi00c] Rebecca Smith. Electronic resumes and online networking: how to use the Internet to do a better job search, including a complete, up-to-date resource guide. Career Press, Franklin Lakes, NJ, USA, second edition, 2000. ISBN 1-56414-511-5 (paperback). 223 pp. LCCN HF5383 .S634 2000.

Smith:2000:BFW

[Smi00d] Robert Ellis Smith. Ben Franklin’s Web site: privacy and curiosity from Plymouth Rock to the Internet. Privacy Journal, Providence, RI, USA, 2000. ISBN 0-930072-14-6. 407 pp. LCCN JC596.2.U5 S644 2000. Contents: Watchfulness, 1620-1720 — Serenity, 1760-1800 — Mistrust, 1790-1880 — Space, 1800-1870 — Curiosity, 1820-1890 — Brandeis, 1890-1930 — Wiretaps, 1910-1990 — Sex, 1920-1990 — Torts, 1930-1995 — The Constitution, 1965-1995 — Numbers, 1935-2000 — Databanks, 1965-2000 — Cyberspace, 1990-2000 — Ben Franklin’s Web site: an epilogue.

Smirnov:2001:QFI

[Smi01a] Mikhail I. Smirnov, editor. Quality of future Internet services: Second COST 263 International Workshop, QofIS 2001, Coimbra, Portugal, September 24–26, 2001: proceedings, volume 2156 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42602-7 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.2156; TK5105.875.I57 C67 2001; TK5105.875.I57 Q36

                                        2001. URL http://link.
                                        springer-ny.com/link/service/
                                        series/0558/tocs/t2156.htm.

Smith:2001:FII

[Smi01b] Drew Smith. Food industry and the Internet. Blackwell Science, Malden, MA, USA, 2001. ISBN 0-632-05753-X. xxiv + 208 pp. LCCN HD9000.5 .S595 2001.

Smith:2001:IMM

[Smi01c] John R. Smith, editor. Internet multimedia management systems II: 22–23 August 2001, Denver, USA, volume 4519 of SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 2001. ISBN 0-8194-4243-7. LCCN QA76.575 .I684 2001.

Smith:2001:MIC

[Smi01d] Mark Smith. Managing the Internet controversy. Neal-Schuman netguide series. Neal-Schuman Publishers, New York, NY, USA, 2001. ISBN 1-55570-395-X (paperback). xii + 226 pp. LCCN Z692.C65 M365 2001.

Smirnow:2002:IMT

[Smi02a]  Mikhail Smirnow.  Inter-

net middleware (tutorial). Lecture Notes in Computer Science, 2515:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2515/25150366.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2515/25150366.pdf.

Perry:2002:IP

[Smi02b] Roger P. (Roger Perry) Smith.

The Internet for physicians. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2002. ISBN 0-387-95312-4 (softcover). xi + 288 pp. LCCN W 26.5 S658i 2002.

Stiller:2004:IEP

[SMKR04] Burkhard Stiller, Lee McKnight, Martin Karsten, and Peter Reichl. Internet economics: pricing and policies. Computer Networks (Amsterdam, Netherlands: 1999), 46 (1):1–2, September 16, 2004.

CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Stoica:2003:CSP

[SMLN^{+}03] Ion Stoica, Robert Morris, David Liben-Nowell, David R. Karger, M. Frans Kaashoek, Frank Dabek, and Hari Balakrishnan. Chord: a scalable peer-to-peer lookup protocol for Internet applications. IEEE/ACM Transactions on Networking, 11(1):17– 32, February 2003.

Soloway:2000:LEK

[SNB^{+}00] Elliot Soloway, Cathleen Norris, Phyllis Blumenfeld, Barry Fishman, Joseph Krajcik, and Ronald Marx. Log on education: K–12 and the Internet. Communications of the ACM, 43(1):19, January 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        2000-43-1/p19-krajcik/.

Snow:2001:IAL

[Sno01] Ray Snow. An information assembly line in Perl. Dr. Dobbs Journal, 26(6):72, 76, 78, 80, 82, June 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.
                                        ddj.com/ftp/2001/2001_06/
                                        infoline.txt.

Siau:2002:AIU

[SNT02] Keng Siau, Fiona Fui-Hoon Nah, and Limei Teng. Acceptable Internet use policy. Communications of the ACM, 45 (1):75–79, January 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sobrinho:2002:AAQ

[Sob02] João Luís Sobrinho. Algebra and algorithms for QoS path computation and hopby-hop routing in the Internet. IEEE/ACM Transactions on Networking, 10(4):541–550, August 2002.

SIBBCB:2000:BCS

[Soc00] Sociedad Israelita de Beneficencia Beit Chabad do Brasil. Beit Chabad: a sua referencia judaica na Internet. Sociedade Israelita de Beneficencia Beit Chabad do Brasil, Sao Paulo, Brasil, 2000. LCCN F2659.J5.

SIBC:2001:APA

[Soc01] Société internationale de bibliographie classique. L’année philologique: APh. (French) [The philological year: APh]. Société internationale de bibliographie classique, Paris, France, 2001.

Sookman:2001:CIE

[Soo01] Barry B. Sookman. Computer, Internet and electronic commerce terms: judicial, legislative and technical definitions. Carswell, Toronto, ON, Canada, 2001. ISBN 0-459-26196-7. 324 pp. LCCN K564.C6 S66 2001.

Sorfa:2001:OSI

[Sor01] Petr Sorfa. Open source IDEs for Linux/Unix. C/C++ Users Journal, 19(3):8–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Sikes:2000:FFA

[SP00] Alfred C. Sikes and Ellen Pearlman. Fast forward: America’s leading experts reveal how the Internet is changing your life. William Morrow, New York, NY, USA, 2000. ISBN 0-380-97828-8. xii + 298 pp. LCCN HM851 .S545 2000.

Samaras:2002:FPA

[SP02] George Samaras and Christoforos Panayiotou. A flexible personalization architecture for wireless Internet based on mobile agents. Lecture Notes in Computer Science, 2435:120–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer-ny.com/link/service/
series/0558/bibs/2435/24350120.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2435/24350120.
pdf.

Spar:2001:RWC

[Spa01] Debora L. Spar. Ruling the waves: cycles of discovery, chaos, and wealth from compass to the Internet. Harcourt, New York, NY, USA, 2001. ISBN 0-15-100509-5. 403 pp. LCCN HM846 .S63 2001.

Sporre:2001:AWW

[Spo01] Dennis J. Sporre. Artsguide:

World and Web: a guide to the arts on the Internet and in the world. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-089305-6 (paperback). xii + 226 + 8 pp. LCCN NX620 .S678 2001.

Shacham:2007:EAS

[SPP^{+}07] Hovav Shacham, Matthew Page, Ben Pfaff, Eu-Jin Goh, Nagendra Modadugu, and Dan Boneh. On the effectiveness of address-space randomization. Report, Department of Computer Science, Stanford University, Stanford, CA, USA, September 15, 2007. 18 pp.

URL https://cseweb.ucsd.
edu/~hovav/dist/asrandom.
pdf.

Stull:2000:HIP

[SR00] Andrew T. Stull and John P.

Rossi. History on the Internet, 1999–2000: a Prentice Hall guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-013821-5. xi + 115 pp. LCCN D16.255.C65 S78 2000.

Scribner:2002:ASI

[SS02] Kenn Scribner and Mark C.

Stiver. Applied SOAP: implementing .NET XML Web services. Howard W. Sams, Indianapolis, IN 46268, USA, 2002. ISBN 0-672-32111-4. ix + 415 pp. LCCN QA76.76.H94 S2575 2002.

Scambray:2006:HEW

[SSS06] Joel Scambray, Mike Shema, and Caleb Sima. Hacking exposed: Web applications. Mc-Graw-Hill, New York, NY, USA, second edition, 2006. ISBN 0-07-226299-0 (paperback). xxix + 520 pp. LCCN TK5105.59 .S32 2006.

Santry:2000:AI

[ST00] Patrick Santry and Mitch Tulloch. Administering IIS 5.0. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-212328-1. xxvi + 546 pp. LCCN TK5105.875.I57 S237 2000.

Stallings:2000:DCC

[Sta00a] William Stallings. Data and Computer Communications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, sixth edition, 2000. ISBN 0-13-084370-9. xx + 810 pp. LCCN TK5105 .S73 2000. US$57.00. URL http:

                                        //www.phptr.com/ptrbooks/
                                        esm_0130843709.html.

Stallings:2000:PB

[Sta00b] William Stallings. Programmer’s bookshelf: Frame relay wrap-up. Dr. Dobbs Journal, 25(7):147, July 2000. CODEN DDJOEB. ISSN 1044-789X.

Stallings:2000:SSC

[Sta00c] William Stallings. The SET Standard and E-commerce. Dr. Dobbs Journal, 25(11):30, 32, 34, 36, November 2000.

CODEN DDJOEB. ISSN 1044-789X.

SPC:2000:XX

[Sta00d] Standard and Poor’s Corporation. DRI Webstract. DRI/ McGraw-Hill, Lexington, MA, USA, 2000. LCCN HC59.15.

Standing:2000:ICD

[Sta00e] Craig Standing. Internet commerce development. Artech House computing library. Artech House Inc., Norwood, MA, USA, 2000. ISBN 1-58053-051-6. xiv + 236 pp. LCCN HF5548.32 .S73 2000.

RBCRPLS:2001:RPP

[Sta01a] State Bar of California. Real Property Law Section. Real property practice and the Internet: program handbook, March 2001. Continuing Education of the Bar, Berkeley, CA, USA, 2001. v + 291 pp. LCCN KFC74 .R43 2001.

SBCRPLS:2001:RPPa

[Sta01b] State Bar of California. Real Property Law Section. Real property practice and the Internet: program handbook, March 2001. Continuing Education of the Bar, Berkeley, CA, USA, 2001. v + 291 pp. LCCN KFC74 .R43 2001. “Prepared for use in conjunction with the March 2001 program, Real property practice and the Internet, presented jointly by CEB and the State Bar Real Property Law Section.”–Pref. “RE-45542 (program RE-05542)”– T.p. verso. Contents: Making the most of Internet e-mail / Albert Barsocchini — Making the most of your Web browser / Jean Murrell Adams — California legal research on the Internet / Adrienne Adan, Laura A. Cadra — Legal ethics, malpractice, and the Internet — Electronic recording: step into the future / Elizabeth Corbett — Selected Internet resources.

SBCRPLS:2001:RPPb

[Sta01c] State Bar of California. Real Property Law Section. Real property practice and the Internet: program handbook, March 2001. Continuing Education of the Bar, Berkeley, CA, USA, 2001. v + 116 pp. LCCN KFC140.A75 R48 2001.

Stallings:2002:HSN

[Sta02a]  William Stallings.  High-

Speed Networks and Internets: Performance and Quality of Service. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-032221-0. xvii + 715 pp. LCCN TK5105.585.S73 2002. US$74.00. URL http:

//www.phptr.com/ptrbooks/
esm_0130322210.html.

Stanton:2002:CPF

[Sta02b] Jeffrey M. Stanton. Company profile of the frequent Internet user. Communications of the ACM, 45(1):55– 59, January 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Stallings:2003:CNI

[Sta03a] William Stallings. Computer Networking with Internet Protocols. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-141098-9. 656 (est) pp. LCCN ???? US$96.67.

Stanley:2003:MNJ

[Sta03b] Nigel Stanley. Microsoft.NET:

jumpstart for systems administrators and developers. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2003. ISBN 1-55558-285-0. xv + 206 pp. LCCN QA76.76.M52N54 2003. US$49.99.

Steinbock:2000:BIM

[Ste00a] Dan Steinbock. The birth of Internet marketing communications. Quorum, Westport, CT, USA, 2000. ISBN 1-56720-303-5. xvi + 308 pp. LCCN HF5415.1265 .S735 2000.

Sterne:2000:CSI

[Ste00b] Jim Sterne. Customer service on the Internet: building relationships, increasing loyalty, and staying competitive. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, second edition, 2000. ISBN 0-471-38258-2 (paperback). xvi + 351 pp. LCCN HF5415.5 .S737 2000.

Stephens:2001:LIT

[Ste01] Michael T. Stephens. The library Internet trainer’s toolkit. Neal-Schuman netguide series. Neal-Schuman Publishers, New York, NY, USA, 2001. ISBN 1-55570-415-8 (paperback). xvi + 223 pp. LCCN ZA4201 .S74 2001.

Sterling:2003:FEE

[Ste03] Christopher Sterling. Focal Encyclopedia of Electronic Media (CD ROM Network Version). Focal Press, Boston, MA, USA and Oxford, UK, 2003. ISBN 0-240-80310-8. ???? pp. LCCN ???? US$320.95.

Schilit:2002:CWW

[STHK02a] Bill N. Schilit, Jonathan Trevor, David M. Hilbert, and Tzu Khiau Koh. Computer: Wireless: Web interaction using very small Internet devices. IEEE Distributed Systems Online, 3(10), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0210/f/sch_print.htm.

Schilit:2002:CFW

[STHK02b] Bill N. Schilit, Jonathan Trevor, David M. Hilbert, and Tzu Khiau Koh. Cover feature: Web interaction using very small Internet devices. Computer, 35(10): 37–45, October 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/10/rx037.htm;   http:
//csdl.computer.org/dl/mags/
co/2002/10/rx037.pdf;   http:
//www.computer.org/computer/
co2002/rx037abs.htm.

Stiller:2002:QPQ

[Sti02] Burkhard Stiller, editor. From QoS provisioning to QoS charging: third COST 263 International Workshop on Quality of Future Internet Services, QofIS 2002 and second International Workshop on Internet Charging and QoS Technologies, ICQT 2002, Zurich, Switzerland, October 16–18, 2002: proceedings, volume 2511 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-44356-8 (softcover). LCCN TK5105.875.I57 C67 2002.

Stirling:2003:BRT

[Sti03] Dale Stirling. Book review:

a thorough look at Internet robots: A review of Beyond Webcams: An Introduction to Online Robots, Ken Goldberg and Roland Siegwart, eds. IEEE Distributed Systems Online, 4(3), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0303/d/bks_a.htm.

Snell:2002:PWS

[STK02] James Snell, Doug Tidwell, and Pavel Kulchenko. Programming Web Services with SOAP. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00095-2. xiii + 244 pp. LCCN QA76.625 .S64 2002 Bar. US$34.95. URL http://

                                        safari.oreilly.com/0596000952;
                                         http://www.oreilly.com/
                                        catalog/progwebsoap.

Stojmenovic:2002:HWN

[Sto02] Ivan Stojmenovic. Handbook of wireless networks and mobile computing. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-41902-8. xxix + 630 pp. LCCN TK5103.2 .H36 2002. US$110.00.

Stone:2003:NII

[Sto03] Adam Stone. News: Internet invasion—is the Web secure? IEEE Distributed Systems Online, 4(5), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0305/f/News_print.htm.

Stoev:2005:WSD

[STPM05] Stilian Stoev, Murad S. Taqqu, Cheolwoo Park, and J. S. Marron. On the wavelet spectrum diagnostic for Hurst parameter estimation in the analysis of Internet traffic. Computer Networks (Amsterdam, Netherlands: 1999), 48(3):

423–445, June 21, 2005. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Strauss:2003:PEN

[Str03]  Cobus Strauss.  Practical

Electrical Network Automation and Communication Systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-7506-5801-

0. ???? pp. LCCN ???? US$49.99.

Strassner:2004:PBN

[Str04] John Strassner. Policy-Based Network Management: Solutions for the Next Generation. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2004. ISBN 1-55860-859-1. xvii + 516 pp. LCCN TK5105.5 .S777 2004. US$59.95. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els031/2003107477.
html.

Silber:2000:GFP

[STT00] William L. Silber, Robert M.

Tumarkin, and Scott T. Taylor. Glucksman Fellowship Program student research reports. Working paper S-00-26, New York University Salomon Center, Leonard N. Stern School of Business, New York, NY, USA, 2000. 44 + 17 pp.

Shen:2002:IRM

[STYC02] Kai Shen, Hong Tang, Tao Yang, and Lingkun Chu. Integrated resource management for cluster-based Internet services. Operating Systems Review, 36(5S):225–238, December 2002. CODEN OSRED8. ISSN 0163-5980.

SPS:2001:DCB

[Sun01] Sun Professional Services. Dot-com & beyond: breakthrough Internet-based architectures and methodologies. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-062297-4. xvii + 314 pp. LCCN TK5105.5.D68 2001.

Surfless:2000:WLL

[Sur00] SurfLess LLC. World’s leading law Internet sites, quickly and easily! SurfLess LLC, Rockville, MD, USA, 2000. ISBN 1-893957-10-1. 34 + vi pp. LCCN KF242.A1 W67 2000. Title from cover. Rev. ed. of: Finding law quickly and easily on the Internet. c1999. Contents: General law — Federal law resources — State law resources — Law by practice area — General Internet resources.

Surratt:2001:ISC

[Sur01] Carla G. Surratt. The Internet and social change. Mc-Farland, Jefferson, NC, USA, 2001. ISBN 0-7864-1019-1 (softcover). v + 233 pp. LCCN HM851 .S89 2001.

Sutherland:2000:SDE

[Sut00a]  Keith Sutherland.  Sub-

net Design for Efficient Networks. Butterworth-Heinemann, Boston, MA, USA, 2000. ISBN 0-7506-4465-6. vi + 196 pp. LCCN : TK5105.5 .S89 2000. US$59.95.

Sutherland:2000:UIC

[Sut00b] Keith Sutherland. Understanding the Internet: a clear guide to Internet technologies. Computer weekly professional series. Butterworth-Heinemann, Boston, MA, USA, 2000. ISBN 0-7506-4555-5. x + 153 pp. LCCN TK5105.875.I57 S88 2000.

Shiveley:2001:UIP

[SV01] James M. Shiveley and Phillip J.

VanFossen. Using Internet primary sources to teach critical thinking skills in government, economics, and contemporary world issues. Greenwood professional guides in school librarianship. ISSN: 1074-150X. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2001. ISBN 0-313-31283-4. xxv + 244 pp. LCCN H61.95 .S49 2001. This is a compilation of over 150 primary-source Web sites covering American government from pre-colonial times to the present day, as well as economics and world issues. Questions and activities for the student are included.

Sculley:2001:BEK

[SW01] Arthur B. Sculley and W. William A. Woods. B2B exchanges: the killer application in the business-to-business Internet revolution. HarperBusiness, New York, NY, USA, 2001. ISBN 0-06-662108-9 (paperback). xv + 288 pp. LCCN HF5415.1263 .S38 2001. Originally published: Hamilton, Bermuda: ISI Publications, 1999.

Swaine:2000:PPg

[Swa00a] Michael Swaine. Programming paradigms: The cluetrain antipasto. Dr. Dobbs Journal, 25 (7):107–110, July 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:2000:RMS

[Swa00b] Michael Swaine. Rebol and E-mail services. Dr. Dobbs Journal, 25(1):28, 30, 32– 34, January 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.
                                        ddj.com/ftp/2000/2000_01/
                                        rebol.txt.

Swaine:2001:PPA

[Swa01] Michael Swaine. Programming paradigms: Alternatives to Java. Dr. Dobbs Journal, 26(2):135–137, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/.

Sweeney:2001:WPY

[Swe01] Susan Sweeney. 101 ways to promote your Web site: filled with proven Internet marketing tips, tools, techniques, and resources to increase your Web site traffic. Maximum Press, Gulf Breeze, FL, USA, third edition, 2001. ISBN 1-885068-57-3. xxii + 464 pp. LCCN HF5415.1265 .S93 2001.

Sung:2003:ITB

[SX03] Minho Sung and Jun Xu.

IP traceback-based intelligent packet filtering: a novel technique for defending against internet DDoS attacks. IEEE Transactions on Parallel and Distributed Systems, 14(9): 861–872, September 2003. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
td/2003/09/l0861abs.htm;
 http://csdl.computer.org/
dl/trans/td/2003/09/l0861.
pdf.

SMI:2000:HPY

[Syn00] Syngress Media, Inc. Hack proofing your network: Internet tradecraft. Syngress Publishing, Inc., Rockland, MA, USA, 2000. ISBN 1-928994-15-6. xxix + 450 pp. LCCN QA76.9.A25 H33 2000.

Schneier:2001:BIP

[SZ01] Bruce Schneier and Elizabeth Zwicky. Banners and Internet protocols. Dr. Dobbs Journal, 26(11):38, 41, November 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/.

Sun:2004:ASC

[SZGbC04] Chengzhung Sun, Wanlei Zhou, Andrzej Goscinski, and Xue bin Chi. Advanced services for Clusters and Internet computing. Future Generation Computer Systems, 20 (4):501–503, May 3, 2004. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

ScottMorton:2000:ICR

[SZSR00] Fiona Scott Morton, Florian Zettelmeyer, and Jorge M. Silva-Risso. Internet car retailing. NBER working paper series; working paper 7961 4007312768. National Bureau of Economic Research, Cambridge, MA, USA, October 2000. 33 pp. LCCN H62.5.U5 N351 no.7961.

Tarallo:2007:WNT

[T^{+}07] Joseph A. Tarallo et al., editors. Wireless network technology, volume 11(4) of Bell Labs Technical Journal. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2007. ISBN ???? 348 pp. LCCN ????

Townsend:2000:LSL

[TAG00] Anthony M. Townsend, Robert J. Aalberts, and Steven A. Gibson. Legally speaking: libel and slander on the Internet. Communications of the ACM, 43(6):15–17, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-6/p15-townsend/.

Tanenbaum:2003:CN

[Tan03] Andrew S. Tanenbaum. Computer networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 2003. ISBN 0-13-066102-3. xx + 891 pp. LCCN TK5105.5 .T36 2003.

Taylor:2000:PBS

[Tay00] Jeffrey L. Taylor. Programmer’s bookshelf: Second time around. Dr. Dobbs Journal, 25 (4):127–128, 130, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Theobald:2000:IRL

[TD00] William Theobald and H. E.

Dunsmore. Internet Resources for Leisure and Tourism. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 0-7506-4644-6. xvi + 236 pp. LCCN G155.A1 T475 2000. US$62.95.

TeleGeography:2000:HST

[Tel00] TeleGeography, Inc. Hubs and spokes: a TeleGeography Internet reader. TeleGeography, Inc., Washington, DC, USA, 2000. ISBN 1-886142-23-8. 106 pp. LCCN TK5105.875.I57

                                        H8 2000. URL http://www.
                                        telegeography.com. “This is

copy HS00-104904.”.

Tennenhouse:2000:EIP

[Ten00] David Tennenhouse. Embedding the Internet: proactive computing. Communications of the ACM, 43(5):43, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        2000-43-5/p43-tennenhouse/

.

Tennison:2005:BXN

[Ten05]  Jeni Tennison.  Beginning

XSLT 2.0: from novice to professional. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-324-3 (paperback). xxvi + 797 pp. LCCN QA76.73.X58 T458

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0625/2006297782-d.html.

Terplan:2000:IPM

[Ter00] Kornel Terplan. Intranet Performance Management. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-9200-4. 316 pp. LCCN TK5105.875.I6 T46 2000. US$69.95.

Teeler:2000:HUI

[TG00] Dede Teeler and Peta Gray. How to use the Internet in ELT. Longman, Harlow, UK, 2000. ISBN 0-582-33931-6. viii + 120 pp. LCCN LB1044.87 .T44 2000.

Thomas:2005:AWD

[THB05] David Thomas, David Heinemeier Hansson, and Leon Breedt. Agile Web development with rails: a Pragmatic guide. The facets of Ruby series. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9766940-0-X. x + 558 pp. LCCN TK5105.888 .T46 2005.

Thiran:2000:SMI

[Thi00] Yves Thiran. Sexe, mensonges et Internet: réseau et transparence. (French) [Sex, lies, and the Internet: network and transparency]. Collection Quartier libre. Castells, Paris, France, 2000. ISBN 2-8040-1493-2, 2-912587-44-1. 91 pp. LCCN HM851 .T55 2001; TK5105.875.I57 T465 2000.

Tiller:2001:TGI

[Til01] James S. Tiller. A Technical Guide to IPSec Virtual Private Networks. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-0876-3. xxv + 345 pp. LCCN TK5105.875.E87 T55 2001. US$54.95.

Tiller:2005:EHF

[Til05] James S. Tiller. The ethical hack: a framework for business value penetration testing. Auerbach Publications, Boca Raton, FL, USA, 2005.

ISBN 0-8493-1609-X. 322 pp. LCCN TK5105.59 .T55

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0646/2003052467-d.html.

Timmons:2000:IAS

[Tim00] Mary E. Timmons. The Internet and acquisitions: sources and resources for development. Haworth Information Press, New York, NY, USA, 2000. ISBN 0-7890-0687-1. 137 pp. LCCN Z689.5.U6 I58 2000. Co-published simultaneously as The Acquisitions Librarian, number 23, 2000.

Tirri:2003:WTS

[Tir03] Henry Tirri. Web technologies: Search in vain: Challenges for Internet search. Computer, 36(1):115–116, January 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/01/r1115.
htm;   http://csdl.computer.
org/dl/mags/co/2003/01/r1115.
pdf.

Tipton:2006:ISM

[TK06] Harold F. Tipton and Micki Krause, editors. Information security management handbook. Auerbach Publications, Boca Raton, FL, USA, fifth edition, 2006. ISBN 0-8493-1997-8, 0-8493-9561-5. various pp. LCCN QA76.9.A25 I54165

2003. URL http://www.loc.

gov/catdir/enhancements/
fy0659/2003061151-d.html.

Tonoike:2002:DIW

[TKF^{+}02] Masatsugu Tonoike, Takaaki Komura, Kenji Fujikawa, Yasuo Okabe, and Katsuo Ikeda. Design and implementation of a Web-based Internet broadcasting system on QoS multicast routers. Lecture Notes in Computer Science, 2550:195– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2550/25500195.htm;
http://link.springer.de/
link/service/series/0558/
papers/2550/25500195.pdf.

Traunmuller:2002:EGF

[TL02] R. Traunmuller and Klaus Lenk, editors. Electronic government: first international conference, EGOV 2002, Aixen-Provence, France, September 2–6, 2002: proceedings, volume 2456 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-44121-2 (softcover). LCCN JF1525.A8 E584 2002.

Trinkle:2000:HHG

[TM00] Dennis A. Trinkle and Scott A.

Merriman. The history highway 2000: a guide to Internet resources. M. E. Sharpe, Armonk, NY, USA, second edition, 2000. ISBN 0-7656-0477-9. xiv + 600 pp. LCCN D16.255.C65 H58 2000. Rev. ed. of: The history highway.

Toran-Marti:2002:SNI

[TMVTdM02] Felix Toran-Marti, Javier Ventura-Traveset, and Juan Carlos de Mateo. Satellite navigation and the Internet. Dr. Dobbs Journal, 27(3):17–18, 20, 22, 24, 26, March 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2002/2002_03/
                                        satelite.txt;   http://www.
                                        ddj.com/ftp/2002/2002_03/
                                        satelite.zip.

Tobin:2001:GPE

[Tob01] James Tobin. Great projects: the epic story of the building of America: from the taming of the Mississippi to the invention of the Internet. Free Press, New York, NY, USA, 2001. ISBN 0-7432-1064-6. viii + 322 pp. LCCN TA23 .T63 2001.

Toh:2002:AHM

[Toh02] C.-K. Toh. Ad Hoc Mobile Wireless Networks: Protocols and Systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-007817-4. xx + 302 pp. LCCN TK5103.2 .T64 2002. US$89.00. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130078174.html.

Toigo:2003:HGN

[Toi03] Jon William Toigo. The Holy Grail of Network Storage Management. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-028416-5. 304 (est.) pp. LCCN QA76.9.D3 T62 2004. US$49.99. With illustrations By Margaret Romao Toigo.

Tosi:2000:CDD

[Tos00] Emilio Tosi. Il codice del diritto dell’informatica e di Internet: normativa nazionale e comunitaria. I codici vigenti e le leggi speciali. La Tribuna, Piacenza, Italy, 2000. ISBN 88-8294-092-6. 1406 pp. LCCN KE1807 .T67 2000.

Tvede:2001:DBM

[TPB01] Lars Tvede, Peter Pircher, and Jens Bodenkamp. Data broadcasting: merging digital broadcasting with the Internet. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, rev. ed edition, 2001. ISBN 0-471-48560-8. xvi + 274 pp. LCCN HE8689.8 .T93 2001.

Travis:2000:IAL

[Tra00] Jeffrey Travis. Internet applications in LabVIEW. National instruments virtual instrumentation series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-014144-5 (paperback/CD-ROM). xix + 601 pp. LCCN Q185 .T83 2000.

Trent:2001:IBG

[Tre01] Rod Trent. IIS 5.0: a beginner’s guide, Windows 2000. Network professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213372-4. xviii + 486 pp. LCCN TK5105.875.I57 T72 2001.

Trulove:2000:BN

[Tru00] James Trulove, editor. Broadband Networking. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-9821-5. xvii + 544 pp. LCCN TK5103.4 .B76525 2000. US$79.95.

Thomas:2000:ISI

[TS00] Alan R. Thomas and James R. Shearer. Internet searching and indexing: the subject approach. Haworth Information Press, New York, NY, USA, 2000. ISBN 0-7890-1030-5, 0-7890-1031-3 (paperback). 217 pp. LCCN Z695.24.I58 2000. Published also as v. 2, no. 3/4 2000 of the Journal of Internet Cataloging.

Thiruvathukal:2002:WPT

[TSC02] George K. (George Kuriakose) Thiruvathukal, John P. Shafaee, and Thomas W. Christopher. Web programming: techniques for integrating Python, Linux, Apache, and MySQL. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-041065-9. xviii + 745 pp. LCCN QA76.625 .T48 2002. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130410659.html.

Tsou:2002:OMF

[Tso02] Ming-Hsiang Tsou. An operational metadata framework for searching, indexing, and retrieving distributed geographic information services on the Internet. Lecture Notes in Computer Science, 2478:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2478/24780313.htm;
http://link.springer.de/
link/service/series/0558/
papers/2478/24780313.pdf.

Tedesco:2000:OMW

[TT00] Anthony Tedesco and Paul Tedesco. Online markets for writers: how to make money by selling your writing on the Internet. Holt, Rinehart and Winston, New York, NY, USA, 2000. ISBN 0-8050-6226-2 (paperback). viii + 210 pp. LCCN PN161 .T43 2000.

Talia:2003:IIC

[TT03] Domenico Talia and Paolo Trunfio. IEEE Internet computing: Peer to peer toward a synergy between P2P and Grids. IEEE Distributed Systems Online, 4(7), 2003.

ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0307/d/wp4p2p.htm.

Turner:2006:CCS

[Tur06] Fred Turner. From counterculture to cyberculture: Steward Brand, the Whole Earth network, and the rise of digital utopianism. University of Chicago Press, Chicago, IL, USA, 2006. ISBN 0-226-81741-5 (hardcover), 0-226-81742-3 (paperback). x + 327 + 16 pp. LCCN QA76.9.C66 T875 2006.

Tse:2005:FWC

[TV05] David Tse and Pramod Viswanath. Fundamentals of wireless communication. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-84527-0. xxi + 564 pp. LCCN TK5103.2 .T74 2005.

Thompson:2003:CFC

[TWW03] Timothy Thompson, Rick Weil, and Mark D. Wood. Cover feature: CPXe: Web services for Internet imaging. Computer, 36(10):54– 62, October 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/co/
                                        2003/10/rx054.htm;   http:
                                        //csdl.computer.org/dl/mags/
                                        co/2003/10/rx054.pdf;   http:
                                        //www.computer.org/computer/
                                        co2003/rx054abs.htm.

Tynan:2005:CPA

[Tyn05] Dan Tynan. Computer privacy annoyances: how to avoid the most annoying invasions of your personal and online privacy. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00775-2 (paperback). xii + 177 pp. LCCN QA76.9.A25 T96 2005.

Unger:2002:IIC

[UBM02] Herwig Unger, Thomas Bohme, and Armin Mikler, editors. Innovative Internet computing systems: second international workshop, IICS 2002, Kuhlungsborn, Germany, June 20–22, 2002: proceedings, volume 2346 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-43790-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.D5 I52 2002.

Underhill:2000:CIG

[UG00] Rod Underhill and Nat Gertler. The complete idiot’s guide to MP3: music on the Internet. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2036-1 (paperback). xiii + 294 pp. LCCN ML74.4.M6 U64 2000.

Uhl:2002:BMI

[Uhl02]  Axel Uhl.  A bandwidth

model for Internet search. In Bernstein et al. [B^{+}02], pages 538–549. ISBN 1-55860-869-9. LCCN ????

URL http://www.vldb.org/
conf/2002/S15P03.pdf.

Urbaczewski:2002:DEM

[UJ02] Andrew Urbaczewski and Leonard M. Jessup. Does electronic monitoring of employee Internet usage work? Communications of the ACM, 45 (1):80–83, January 2002. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

USOJJDP:2001:ADS

[Uni01] United States. Office of Juvenile Justice and Delinquency Prevention. Access denied: should youth access to the Internet be regulated?: a resource guide for teachers. American Bar Association, Chicago, IL, USA, 2001. ISBN 1-57073-789-4. iv + 44 pp. LCCN KF4772 .A93 2001.

USENIX:2001:PUS

[USE01] USENIX Association, editor. Proceedings of the 3rd USENIX Symposium on Internet Technologies and Systems (USITS ’01), March 26– 28, 2001, San Francisco, California, USA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-14-6. LCCN TK5105.875.I57 U96

2001. URL http://library.
ucsc.edu/science/protect/

usenix.html;   http://www.
usenix.org/publications/
library/proceedings/.

Vacca:2001:WBN

[Vac01] John R. Vacca. Wireless broadband networks handbook: 3G, LMDS and wireless Internet. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-213031-8. xxx + 960 pp. LCCN TK5103.4 .V33 2001.

Vacca:2006:GWN

[Vac06] John R. Vacca. Guide to wireless network security, volume 20 of Advances in information security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-95425-2. xxii + 848 pp. LCCN QA76.9.A25 C66

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0663/2005939009-d.html.

Vacca:2009:CIS

[Vac09] John R. Vacca, editor. Computer and information security handbook. The Morgan Kaufmann series in computer security. Elsevier, Amsterdam, The Netherlands, 2009. ISBN 0-12-374354-0. xxxii + 844 pp. LCCN QA76.9.A25 C627 2009.

Vandegrift:2001:IQC

[Van01a] Judy Vandegrift. Internet Quests: current events and world geography. Teacher Created Materials, Inc., Westminster, CA, USA, 2001. ISBN 0-7439-3444-X. 80 pp. LCCN LB1583.5 .V36 2001. Middle school/high school.

Vandegrift:2001:IQM

[Van01b] Judy Vandegrift. Internet quests: Math. Teacher Created Materials, Westminster, CA, USA, 2001. ISBN 0-7439-3445-8. 80 pp. LCCN QA139 .V36 2001.

Varshney:2003:LMM

[Var03] Upkar Varshney. Location management for mobile commerce applications in wireless Internet environment. ACM Transactions on Internet Technology (TOIT), 3(3): 236–255, August 2003. CO-DEN ???? ISSN 1533-5399.

Varghese:2005:NAI

[Var05] George Varghese. Network algorithmics: an interdisciplinary approach to designing fast networked devices. The Morgan Kaufmann series in networking. Elsevier/ Morgan Kaufmann, San Francisco, CA, USA, 2005. ISBN 0-12-088477-1. xxiv + 465 pp. LCCN TK5105.55 .V37 2005.

Vogelsang:2000:IUR

[VC00] Ingo Vogelsang and Benjamin M. Compaine. The Internet upheaval: raising questions, seeking answers in communications policy. MIT Press, Cambridge, MA, USA, 2000. ISBN 0-262-22063-6. xxix + 426 pp. LCCN HE7781 .I596 2000.

vandenBronk:2000:MIN

[vdB00] Floris van den Bronk. Management of International Networks: Cost-Effective Strategies for the New Telecom Regulations and Service. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-0739-2. 191 pp. LCCN HE7661 .B76 2000. US$69.95.

Mei:2001:IPC

[vdMdB01] Robert D. van der Mei and Frank Huebner-Szabo de Bucs, editors. Internet performance and control of network systems II: 21–22 August 2001, Denver, USA, volume 4523 of SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 2001. ISBN 0-8194-4247-X. LCCN TS510.S63 v.4523.

Vacca:2005:FJN

[VE05] John R. Vacca and Scott (Scott Raymond) Ellis. Firewalls: jumpstart for network and systems administrators. Elsevier/Digital Press, Amsterdam, The Netherlands; Boston, MA, USA, 2005. ISBN 1-55558-297-4. xxx + 418 pp. LCCN TK5105.59 .V32 2005. URL http:

//books.elsevier.com/us/
mk/us/subindex.asp?isbn=
1555582974;   http://www.
loc.gov/catdir/description/
els051/2004063211.html;

                                        http://www.loc.gov/catdir/
                                        toc/els051/2004063211.html.

Vedder:2001:EI

[Ved01] Anton Vedder. Ethics and the Internet. Intersentia, Antwerpen, Belgium, 2001. ISBN 90-5095-176-7. x + 211 pp. LCCN K4305 .E84 2001.

Verma:2001:CDN

[Ver01] Dinesh Verma. Content distribution networks: an engineering approach. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-44341-7. 200 (est.) pp. LCCN ???? US$44.95.

Vladimirov:2004:WFS

[VGM04] Andrew A. Vladimirov, Konstantin V. Gavrilenko, and Andrei A. Mikhailovsky. Wi-Foo: The Secrets of Wireless Hacking. Addison-Wes-ley, Reading, MA, USA, 2004. ISBN 0-321-20217-1. xxvii + 555 pp. LCCN TK5105.59 .V53 2004. URL http://www.

                                        awprofessional.com/title/
                                        0321202171.

Vukadinovic:2002:SSI

[VHE02] Danica Vukadinović, Polly Huang, and Thomas Erlebach. On the spectrum and structure of Internet topology graphs. Lecture Notes in Computer Science, 2346:83– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2346/23460083.htm;
http://link.springer.de/
link/service/series/0558/
papers/2346/23460083.pdf.

Vinoski:2002:ICT

[Vin02] Steve Vinoski. Internet computing: Toward integration: Middleware “Dark Matter”. IEEE Distributed Systems Online, 3(10), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0210/d/wp5midd.htm.

Vinoski:2003:IICc

[Vin03a] Steve Vinoski. IEEE Internet computing: Toward integration: Invocation styles. IEEE Distributed Systems Online, 4(8), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0308/d/w4midp.htm.

Vinoski:2003:IICb

[Vin03b] Steve Vinoski. IEEE Internet computing: Toward integration: It’s just a mapping problem. IEEE Distributed Systems Online, 4(5), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://dsonline.computer.
org/0305/d/wp3mid.htm.

Vinoski:2003:IICa

[Vin03c] Steve Vinoski. IEEE Internet computing: Toward integration: The performance presumption. IEEE Distributed Systems Online, 4(4), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0304/d/wp2midd.htm.

Viega:2002:BSS

[VM02] John Viega and Gary Mc-Graw. Building Secure Software: How to Avoid Security Problems the Right Way. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-72152-X. xxx + 493 pp. LCCN QA76.76.D47 V857 2001.

Vaughan-Nichols:2003:TNM

[VN03] Steven J. Vaughan-Nichols. Technology news: Mobile IPv6 and the future of wireless Internet access. Computer, 36(2):18–20, February 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/02/r2018.
htm;   http://csdl.computer.
org/dl/mags/co/2003/02/r2018.
pdf.

Vasseur:2004:NRP

[VPD04] Jean-Philippe Vasseur, Mario Pickavet, and Piet Demeester. Network recovery: protection and restoration of optical, SONET-SDH, IP, and MPLS. Morgan Kaufmann series in networking. Elsevier/Morgan Kaufmann, San Francisco, CA, USA, 2004. ISBN 0-12-715051-X. xx + 521 pp. LCCN TK5105.5 .V385 2004. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els051/2004558125.
html;   http://www.loc.gov/
catdir/toc/els051/2004558125.
html.

Vlamis:2001:DYB

[VS01] Anthony S. Vlamis and Bob Smith. Do you? business the Yahoo! way: secrets of the world’s most popular Internet company. Capstone, Milford, CT, USA, 2001. ISBN 1-84112-111-8. viii + 231 pp. LCCN HD9696.8.U64 Y348 2001.

VanDyk:2007:PDD

[VW07] John K. VanDyk and Matt Westgate. Pro Drupal development. Expert’s voice in open source; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-755-9. xxv + 428 pp. LCCN TK5105.888 .V363 2007.

Whittaker:2006:HBW

[WA06] James A. Whittaker and Mike Andrews. How to break Web software: functional and security testing of Web applications and Web services. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-36944-0 (paperback). ???? pp. LCCN QA76.76.T48 W485 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip064/2005034913.html.

Wallach:2000:SSM

[WAF00] Dan S. Wallach, Andrew W. Appel, and Edward W. Felten. SAFKASI: a security mechanism for languagebased systems. ACM Transactions on Software Engineering and Methodology, 9 (4):341–378, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/tosem/
                                        2000-9-4/p341-wallach/p341-
                                        wallach.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        tosem/2000-9-4/p341-wallach/

.

Wall:2001:CI

[Wal01a] David Wall. Crime and the Internet. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2001. ISBN 0-415-24428-5, 0-415-24429-3 (paperback). xii + 221 pp. LCCN HV6773 .C75 2001.

Walters:2001:EGC

[Wal01b] E. Garrison Walters. The essential guide to computing: the story of information technology. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-019469-

7. xxv + 499 pp. LCCN

QA76.6 .W335 2001.

Wang:2001:SCB

[Wan01a] Wally Wang. Steal this computer book: what they won’t tell you about the Internet. No Starch Press, San Francisco, CA, USA, second edition, 2001. ISBN 1-886411-42-5 (paperback). xviii + 462 pp. LCCN HV6773.W35 2001.

Wang:2001:IQA

[Wan01b] Zheng Wang. Internet QoS:

architectures and mechanisms for quality of service. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-608-4. xv + 239 pp. LCCN TK5105.875.I57 W373 2001.

Wang:2006:BNC

[Wan06a] Wally Wang. The book of Nero 7: CD and DVD burning made easy. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-110-7. ???? pp. LCCN TK7895.C39 W363

2006. URL http://www.loc.
gov/catdir/toc/ecip0612/
2006011705.html.

Wang:2006:SCB

[Wan06b] Wally Wang. Steal this computer book 4.0: what they won’t tell you about the Internet. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-105-0. ???? pp. LCCN TK5105.875.I57 W3727 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip065/2005037435.html.

Wang:2008:STN

[Wan08] Yun Wang. Statistical Techniques for Network Security: Modern Statistically-Based Intrusion Detection and Protection. IGI Global, Hershey, PA, USA, 2008. ISBN 1-59904-708-X (hardcover), 1-59904-710-1 (electronic). xii + 476 pp. LCCN TK5105.59. US$165.00.

Warkentin:2001:RWP

[War01] Craig Warkentin. Reshaping world politics: NGOs, the Internet, and global civil society. Rowman and Littlefield Publishers, Lanham, MD, USA, 2001. ISBN 0-7425-0971-0, 0-7425-0972-9 (paperback). xiv + 207 pp. LCCN JC337 .W37 2001.

Wasny:2000:WBR

[Was00] Garrett Wasny. World business resources.com: a directory of 8,000+ business resources on the Internet. Mc-Graw-Hill, New York, NY, USA, 2000. ISBN 0-07-136072-7. xviii + 677 pp. LCCN HF54.56 .W37 2000. This book was previously published under the title “How to Conquer the World.”.

Worthington:2002:BMP

[WB02] Shari L. S. Worthington and Walt Boyes. E-business in manufacturing: putting the Internet to work in the industrial enterprise. Instrumentation Systems and Automation Society, Research Triangle Park, NC, USA, 2002. ISBN 1-55617-758-5. xvi + 208 pp. LCCN HD9720.5 .W64 2002.

Wood:2000:WHI

[WC00] M. Sandra Wood and Janet M.

Coggan. Women’s health on the Internet. Haworth Press, New York, NY, USA, 2000. ISBN 0-7890-1300-2, 0-7890-1301-0. 153 pp. LCCN RA778.W756 2000. Copublished simultaneously as Health care on the Internet, volume 4, Numbers 2/3 2000.

Welsh:2001:SAW

[WCB01] Matt Welsh, David Culler, and Eric Brewer. SEDA: an architecture for well-conditioned, scalable Internet services. Operating Systems Review, 35(5): 230–243, December 2001. CO-DEN OSRED8. ISSN 0163-5980.

Wehner:2001:SIR

[Weh01] Christopher Wehner. Screenwriting on the Internet: researching, writing, and selling your script on the Web. Michael Wiese Productions, Studio City, CA, USA, 2001. ISBN 0-941188-36-1. xx + 235 pp. LCCN PN1996.W294 2001.

Weinstein:2000:IRR

[Wei00]  Lauren Weinstein.  Inside

risks: Risks of Internet voting. Communications of the ACM, 43(6):128, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-6/p128-weinstein/

.

Weinberg:2001:SIB

[Wei01a] Robert Weinberg. Streetwise Internet business plan: create a compelling plan for your .com business that will get it financed, and lead it to success. Adams Media Corp., Holbrook, MA, USA, 2001. ISBN 1-58062-502-9. xii + 384 pp. LCCN HF5548.32 .W46 2001.

Weiser:2001:WHN

[Wei01b] Mark Weiser. Whatever happened to the next-generation Internet? Communications of the ACM, 44(9):61– 69, September 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/articles/journals/cacm/
2001-44-9/p61-weiser/p61-
weiser.pdf;   http://www.
acm.org/pubs/citations/journals/
cacm/2001-44-9/p61-weiser/

.

Weisman:2002:EGR

[Wei02] Carl J. Weisman. The Essential Guide to RF and Wireless. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2002. ISBN 0-13-035465-

1. 332 (est.) pp. LCCN ???? US$34.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130354651.html.

Wellman:2002:DIN

[Wel02] Barry Wellman. Designing the Internet for a networked society. Communications of the ACM, 45(5):91–96, May 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wertheim:2000:PGC

[Wer00]  Margaret Wertheim.  The

pearly gates of cyberspace: a history of space from Dante to the Internet. W.W. Norton, New York, NY, USA, 2000. 335 pp. LCCN QA76.9.C66 W48 2000. Contents: ch. 1. Soul-space — ch. 2. Physical space — ch. 3. Celestial space — ch. 4. Relativistic space — ch. 5. Hyperspace — ch. 6. Cyberspace — ch. 7. Cyber soulspace — ch. 8. Cyber-utopia.

West:2000:AEI

[Wes00a] Darrell M. West. Assessing E-government: the Internet, democracy and service delivery. Inside Politics, Providence, RI, USA, 2000. 27 pp. LCCN 2001 0276. URL http:

//www.insidepolitics.org/
egovtreport00.html. Cap-

tion title. Downloaded Oct. 24, 2000.

West:2000:AGI

[Wes00b] Darrell M. West. Assessing E-government: the Internet, democracy and service delivery. Inside Politics, Providence, RI, USA, 2000. 27 pp. LCCN 2001 0276.

Westholm:2002:DGA

[Wes02] Hilmar Westholm. e-democracy goes ahead. the Internet as a tool for improving deliberative policies? Lecture Notes in Computer Science, 2456:240– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2456/24560240.htm;
http://link.springer.de/
link/service/series/0558/
papers/2456/24560240.pdf.

Weyers:2001:ECA

[Wey01] Hans-Leo Weyers. Electronic commerce: der Abschluss von Vertragen im Internet. (German) [Electronic commerce: the end of tolerance in the Internet], volume 196 of Arbeiten zur Rechtsvergleichung. Nomos, Baden-Baden, Germany, 2001. ISBN 3-7890-7091-2. 137 pp. LCCN K564.C6 E42 2001.

Warwick:2002:FIP

[WHA02] Mal Warwick, Ted Hart, and Nick Allen. Fundraising on the Internet: the e-PhilanthropyFoundation.org’s guide to success online. The Jossey-Bass nonprofit and public management series. Jossey-Bass, San Francisco, CA, USA, second edition, 2002. ISBN 0-7879-6045-4. xxvi + 291 pp. LCCN HV41.2 .F87 2002.

Weber:2000:GIR

[WHdM00] Rolf H. Weber, Reto M.

Hilty, and Rolf Auf der Maur, editors. Geschaftsplattform Internet: rechtliche und praktische Aspekte. (German) [The Internet business platform: legal and practical aspects], volume 10 of Publikationen aus dem Zentrum fur Informations- und Kommunikationsrecht der Universitat Zurich. Schulthess, Zürich, Switzerland, 2000. ISBN 3-7255-4073-X. LCCN KE3866 .G47 2000.

Windeatt:2000:I

[WHE00] Scott Windeatt, David Hardisty, and David Eastment. The Internet. Resource books for teachers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-437223-5. vii + 136 pp. LCCN LB1576.7 .W56 2000. Contents: Core Internet skills — Focus on language — Focus on language skills — Appendices: A. Glossary of Internet terms — B. The Web — C. Using computers for communication — D. Internet resources — E. Activitiy links.

Wierzbicki:2002:ICL

[Wie02]  Adam Wierzbicki.  Inter-

net cache location and design of content delivery networks. Lecture Notes in Computer Science, 2376:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2376/23760069.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2376/23760069.pdf.

Williams:2000:TII

[Wil00] Al Williams. The TINI Internet interface. Dr. Dobbs Journal, 25(10):82, 84, 86, 88, October 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.
                                        ddj.com/ftp/2000/2000_10/
                                        tini.txt.

Willis:2001:CSI

[Wil01] Peter Willis. Carrier scale IP networks: designing and operating Internet networks, volume 1 of BT communications technology series. Institution of Electrical Engineers, London, UK, 2001. ISBN 0-85296-982-1. xvi + 293 pp. LCCN TK5105.5.C37 2001.

Wilkinson:2002:NGN

[Wil02a] Neill Wilkinson. Next Generation Network Services. John Wiley and Sons, Inc., New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-48667-1. 210 (est.) pp. LCCN ???? US$65.00.

Wilson:2002:DWN

[Wil02b] Ed Wilson. Designing Windows 2000 Networks. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-066199-6. 525 pp. LCCN ???? US$49.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130661996.html.

Williams:2003:ICB

[Wil03] Al Williams. Internet connectivity in a box. Dr. Dobbs Journal, 28(1):52, 54, 56, 58, January 2003. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.
ddj.com/documents/s=7758/
ddj0301g/.

Wilson:2004:DCM

[Wil04] Connie S. Wilson, editor. Deploying Citrix MetaFrame Presentation Server 3.0 with Windows Server 2003 Terminal Services. Syngress Publishing, Inc., Rockland, MA, USA, 2004. ISBN 1-932266-50-X. xxiv + 567 pp. LCCN QA76.9.C55 D47 2004.

Wilkinson:2006:EPG

[Wil06] Julia L. Wilkinson. The eBay price guide: what sells for what (in every category!). No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-055-0. ???? pp. LCCN HF5478 .W548 2005. URL http:

//www.loc.gov/catdir/toc/
ecip0518/2005025146.html.

Wiseman:2000:IEA

[Wis00] Alan E. Wiseman. The Internet economy: access, taxes, and market structure. Brookings Institution Press, Washington, DC, USA, 2000. ISBN 0-8157-9384-7. ix + 130 pp. LCCN HF5548.32 .W57 2000. Contents: Introduction — What is the Internet?

— Pricing of access to the Internet — Pricing of goods and services on the Internet — Network effects and the Internet — Internet taxation.

Wolton:2000:IPM

[WJ00] Dominique Wolton and Olivier Jay. Internet: petit manuel de survie. (French) [Internet: Little survival manual]. Flammarion, Paris, France, 2000. ISBN 2-08-211583-6. 186 pp. LCCN HM851 .W655 2000.

Weske:2004:OOI

[WL04] Mathias Weske and Peter Liggesmeyer, editors. Object-oriented and Internetbased technologies: 5th Annual International Conference on Object-Oriented and Internet-Based Technologies, Concepts, and Applications for a Networked World, Net.ObjectDays 2004, Erfurt, Germany, September 27–30, 2004: proceedings, volume 3263 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-23201-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.64 .N632

2004.  URL http://www.
springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3263.

Waltermann:2000:POC

[WM00] Jens Waltermann and Marcel Machill. Protecting our children on the Internet: towards a new culture of responsibility. Brookings Institution Press, Washington, DC, USA, 2000. ISBN 3-89204-474-0. 512 pp. LCCN K564.C6 P76 2000. Includes the Bertelsmann Foundation’s “Memorandum on self-regulation of Internet content” and commentaries from the Internet Content Summit, held September 1999 in Munich, Germany.

Weinstein:2000:IRI

[WN00] Lauren Weinstein and Peter G.

Neumann. Inside risks: Internet risks. Communications of the ACM, 43(5):144, May 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-5/p144-weinstein/

.

Williams:2001:ICI

[WN01a] Peter Williams and David Nicholas. The Internet and the changing information environment. Managing information report. Aslib-IMI, London, UK, 2001. ISBN 0-85142-404-X. iv + 117 pp. LCCN ZA4201 .W55 2001.

Wong:2001:CEN

[WN01b] John Wong and Seok Ling Nah. China’s emerging new economy: the Internet and e-commerce. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2001. ISBN 981-02-4495-9. vii + 142 pp. LCCN HF5548.325.C5 W66 2001.

Whitaker:2006:PTN

[WN06] Andrew Whitaker and Daniel P. Newman. Penetration testing and network defense. Cisco Press networking technology series. Cisco Press, Indianapolis, IN, USA, 2006. ISBN 1-58705-208-3 (paperback). xxvii + 598 pp. LCCN TK5105.59 .W53 2006. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        fy0607/2004108262.html.

Woehr:2002:PBE

[Woe02] Jack J. Woehr. Programmer’s bookshelf: Embedded systems programming. Dr. Dobbs Journal, 27(10):86, October 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/.

Wong:2000:WBY

[Won00] Thomas Wong. 101 ways to boost your Web traffic: Internet promotion made easier. Intesync, Union City, CA, USA, 2000. ISBN 0-9638349-5-9. 318 pp. LCCN HF5415.1265 .W66 2000.

Wood:2000:HCR

[Woo00] M. Sandra Wood. Health care resources on the Internet: a guide for librarians and health care consumers. Haworth Information Press, New York, NY, USA, 2000. ISBN 0-7890-0632-4, 0-7890-0911-0 (paperback). xxi + 205 pp. LCCN W26.55 C7 H4 2000.

Worth:2002:BSI

[Wor02] Roland H. Worth. Biblical studies on the Internet: a resource guide. McFarland and Co., Jefferson, NC, USA, 2002. ISBN 0-7864-1327-1. xi + 316 pp. LCCN BS600.2 .W67 2002.

Williams:2001:CIL

[WS01] Richard D. Williams and Bruce T. Smyth. Computer and Internet liability: strategies, claims, and defenses. Aspen Law and Business, Gaithersburg, MD, USA, second edition, 2001. ISBN 0-7355-1851-3. various pp. LCCN KF390.5.C6 W544 2001. Rev. ed. of: Law of the year 2000 problem. c1999-Loose-leaf for updating.

Warschauer:2000:IET

[WSM00] Mark Warschauer, Heidi Shetzer, and Christine F. Meloni. Internet for English teaching. TESOL, Alexandria, VA, USA, 2000. ISBN 0-939791-88-9. ix + 178 pp. LCCN PE1128.A2 W37 2000.

Walrand:2000:HPC

[WV00] Jean Walrand and Pravin Varaiya. High-Performance Communication Networks. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2000. ISBN 1-55860-574-6. xxii + 693 pp. LCCN TK5105.5 .W353 2000. US$84.95.

Wu:2004:RNR

[WWD04] Richard S. L. Wu, Allan K. Y. Wong, and Tharam S. Dillon. RDCT: a novel reconfigurable dynamic cache size tuner to shorten information retrieval time over the Internet. International Journal of Computer Systems Science and Engineering, 19(6):??, November 2004. CODEN CSSEEI. ISSN 0267-6192.

Wyler:2005:ANS

[Wyl05]  Neil R. Wyler.  Aggres-

sive network self-defense. Syngress Publishing, Inc., Rockland, MA, USA, 2005. xxx + 383 pp. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://site.ebrary.com/
lib/ucsc/Doc?id=10075669.

Wittmann:2001:MCP

[WZ01a] Ralph Wittmann and Martina Zitterbart. Multicast Communication: Protocols, Programming, and Applications. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-645-9. xvii + 349 pp. LCCN TK5105.887 .W58 2001. US$47.95.

Wolf:2001:LMB

[WZ01b] Jason Wolf and Natalie Zee.

The last mile: broadband and the next Internet revolution. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-136349-1. xxii + 200 pp. LCCN TK5103.4 .W65 2001.

Watson:2000:IIM

[WZP00] Richard T. Watson, George M.

Zinkhan, and Leyland F. Pitt. Integrated Internet marketing. Communications of the ACM, 43(6):97–102, June 2000. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/cacm/
2000-43-6/p97-watson/.

Wang:2004:CSI

[WZZ04] Junfeng Wang, Mingtian Zhou, and Hongxia Zhou. Clock synchronization for Internet measurements: a clustering algorithm. Computer Networks (Amsterdam, Netherlands: 1999), 45(6): 731–741, August 21, 2004. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Young:2004:HHS

[YA04] Susan (Susan Elizabeth) Young and Dave Aitel. The hacker’s handbook: the strategy behind breaking into and defending networks. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2004. ISBN 0-8493-0888-7. xxxiv + 860 pp. LCCN TK5105.59 .Y68 2004. URL http://

www.e-streams.com/es0709/
es0709_3516.html;   http:
//www.loc.gov/catdir/enhancements/
fy0646/2003055391-d.html.

Yang:2003:ICS

[YCY03] S. H. Yang, X. Chen, and L. Yang. Integration of control system design and implementation over the Internet using the Jini technology. Software —Practice and Experience, 33 (12):1151–1175, October 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yan:2002:QAM

[YFB02] Shuqian Yan, Michalis Faloutsos, and Anindo Banerjea. QoS-aware multicast routing for the Internet: the design and evaluation of QoS-MIC. IEEE/ACM Transactions on Networking, 10(1):54– 66, February 2002.

Chao:2001:IRS

[yJC01] Sheau yueh J. Chao. Internet resources and services for international real estate information. Global guides to Internet business resources. Oryx Press, Westport, CT, USA, 2001. ISBN 1-57356-373-

0. xviii + 384 pp. LCCN HD1380.6 .C485 2001.

Yourdon:2002:MHI

[You02]  Edward Yourdon.  Man-

aging high-intensity Internet projects. Just enough series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-062110-2. xi + 226 pp. LCCN HF5548.32 .Y68 2002.

Yu:2000:TTAa

[Yu00a] Haifeng Yu. TACT: tunable availability and consistency tradeoffs for replicated Internet services. Operating Systems Review, 34(2):33, April 2000. CODEN OSRED8. ISSN 0163-5980.

Yu:2000:TTAb

[Yu00b] Haifeng Yu. TACT: tunable availability and consistency tradeoffs for replicated Internet services (poster session). Operating Systems Review, 34(2):40, April 2000. CODEN OSRED8. ISSN 0163-5980.

Yuan:2001:LPS

[Yua01] Michael Yuan. LaTeX2HTML:

Publish science to the Web. Linux Journal, 92:93–94, December 2001. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Zadok:2001:LNA

[Zad01] Erez Zadok. Linux NFS and Automounter administration. The Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2739-8. xlvi + 638 pp. LCCN QA76.76.O63 Z34 2001.

Zalech:2003:BRW

[Zal03] Wojciech Zalech. Book review: What hides behind the Internet-HTTP protocol. IEEE Distributed Systems Online, 4(8), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0308/0306/d/bks_a.htm.

Zalewski:2005:SWF

[Zal05] Michal Zalewski. Silence on the wire: a field guide to passive reconnaissance and indirect attacks. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-046-1. xxv + 281 pp. LCCN TK5105.59 .Z35 2005.

Ziccardi:2001:DAN

[ZB01] Giovanni Ziccardi and John P. Barlow. Il diritto d’autore nell’era digitale: evoluzione tecnologica e copyright: Internet, mp3, DivX;-), open source, Gnu/Linux, free software, mezzi di protezione. Diritto. Il sole 24 ore, Milano, Italy, 2001. ISBN 88-324-4459-3. xxi + 343 pp. LCCN KE1809 .Z53 2001.

Zwicky:2000:BIF

[ZCC00] Elizabeth D. Zwicky, Simon Cooper, and D. Brent Chapman. Building Internet firewalls. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2000. ISBN 1-56592-871-7. xxi + 869 pp. LCCN TK5105.59 .Z85 2000.

Zdziarski:2005:ESB

[Zdz05] Jonathan A. Zdziarski. Ending spam: Bayesian content filtering and the art of statistical language classification. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-052-6. xx + 287 pp. LCCN TK5105.743 .Z35 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip0510/2005008221.
html;   http://www.oreilly.
com/catalog/1593270526/.

Zeldman:2007:DWS

[Zel07] Jeffrey Zeldman. Designing with Web standards. Voices that matter. New Riders Publishing, Carmel, IN, USA, second edition, 2007. ISBN 0-321-38555-1 (paperback). xx + 410 pp. LCCN TK5105.888 Z46 2007. URL http:/

/www.loc.gov/catdir/toc/
fy0705/2006283016.html.

Ziviani:2005:IAM

[ZFdRD05] Artur Ziviani, Serge Fdida, José F. de Rezende, and Otto Carlos M. B. Duarte. Improving the accuracy of measurement-based geographic location of Internet hosts. Computer Networks (Amsterdam, Netherlands: 1999), 47 (4):503–523, March 15, 2005. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Zhao:2004:WSN

[ZG04] Feng Zhao and Leonidas J. Guibas. Wireless sensor networks: an information processing approach. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2004. ISBN 1-55860-914-8. xvii + 358 pp. LCCN TK7872.D48 Z43 2004. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        description/els051/2004301905.
                                        html;   http://www.loc.gov/
                                        catdir/toc/els051/2004301905.
                                        html.

Zhao:2003:SNI

[Zha03]  Feng Zhao.  Sensor Net-

works: An Information Processing Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-914-8. ???? pp. LCCN ???? US$64.95.

Zhu:2000:ZHL

[Zhu00] Yusheng Zhu. Zhongguo hu lian wang gai lan. Xin hua chu ban she, Beijing Shi, China, di 1 ban edition, 2000. ISBN 7-5011-5143-1. 221 pp. LCCN ZA4226 .Z46 2000. Includes legislation. Chinese and English.

Zimmerman:2000:MI

[Zim00] Jan Zimmerman. Marketing on the Internet. Maximum Press, Gulf Breeze, FL, USA, fourth edition, 2000. ISBN 1-885068-36-0. xxvi + 450 pp. LCCN HF6146.I58 M38 2000.

Zhang:2001:DMI

[ZLL01] David Zhang, X. Li, and Zhiyong Liu. Data management and Internet computing for image/pattern analysis, volume 11 of Kluwer international series on Asian studies in computer and information science; 11. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-7456-8. xiii + 365 pp. LCCN QA76.9.D3 Z53 2001.

Zheng:2005:SPN

[ZN05] Pei Zheng and Lionel Ni.

Smart Phone and Next-Generation Mobile Computing. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2005. ISBN 0-12-088560-3. ???? pp. LCCN ???? US$44.95.

Zhao:2001:IPT

[ZR01] J. Leon Zhao and Vincent H.

Resh. Internet publishing and transformation of knowledge processes. Communications of the ACM, 44(12):103– 109, December 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Zysman:2000:GPI

[ZW00] John Zysman and Steve Weber. Governance and politics of the Internet economy: historical transformation or ordinary politics with a new vocabulary? Working paper 141, Berkeley Roundtable on the International Economy, University of California, Berkeley, Berkeley, CA, USA, May 2000. 19 pp.

Zhang:2002:RBP

[ZY02] Yaying Zhang and Jinyuan You. An RBAC based policy enforcement coordination model in Internet environment. Lecture Notes in Computer Science, 2480:466–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2480/24800466.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2480/24800466.
                                        pdf.